A Complete Bibliography of Publications in Journal of Fish Diseases (2020–2029)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

23 April 2025 Version 1.26

Title word cross-reference

+ [LWT^{+}22]. ^{vc} [BGZW22]. _{2} [ZWB^{+}23]. _{AHPND} [CCL^{+}24]. α [KQW^{+}21]. β [LLC^{+}21, RGS^{+}21]. · [ZWB^{+}23]. Δ [DOW^{+}22]. γ [YGC^{+}20]. λ/κ [CSGB^{+}21, CSVG^{+}22]. × [AGR^{+}22, FFB^{+}21, HHW^{+}24, LPGA^{+}21, LPGW^{+}22, SLA^{+}20, VGW^{+}21, YGR^{+}24, ZSL^{+}24a].

&n [CLC^{+}22].

-carrageenin [CSGB^{+}21, CSVG^{+}22]. -defensin [LLC^{+}21]. -glucan [RGS^{+}21]. -induced [EMH^{+}22, FZC^{+}23]. -like [BGZW22]. -positive [DOW^{+}22].

. [SGR^{+}20a, WSW^{+}20].

01 [CCY^{+}23].

1} [FdSB^{+}22, ASF^{+}21, EBBIAH21, FVC^{+}21, PEN^{+}21, PPT^{+}20, YLC^{+}21, dYdRFF^{+}23]. 10 [MHA^{+}20]. 12-b [BME^{+}20]. 124 [FZC^{+}23, YCC^{+}21]. 12^{o} [LWR^{+}20]. 13326 [HHW^{+}24].

2 [FM23, FZC^{+}23, KKH^{+}24, TMM^{+}24]. 2022 [Muf23]. 22^{o} [BLJ^{+}21]. 26 [YJL^{+}23b]. 2640 [ZSL^{+}24a]. 2642 [ZSL^{+}24a]. 2nd [Muf23].

3 [PEś^{+}20, TBMP23, TR23]. 36 [KSM^{+}21]. 3D [FSS^{+}21].

4 [WCT^{+}23a].

5 [LHLS20]. 57 [AHSB^{+}22]. 57-kDa [AHSB^{+}22].

7-arylaminoisoquinolin-5 [VBL^{+}23]. 7MAnt [VBG^{+}22].

8-hydroxyl [LYZ^{+}22]. 8-quinones [VBL^{+}23].

978 [FM23]. 978-1-032-48341-2 [FM23].

A-layer [HLH^{+}23b]. A. [LH20, ZLWL23]. A2 [XQZ^{+}22]. ability [DHSM^{+}20, KKBL21, LHTW21, YPT^{+}22, ZLWL23]. abnormalities [APK^{+}23, EGRD22, FEM^{+}20, ŘB24]. Abramis [APS23]. abscessation [RZTV21]. abscesses [MAG24]. abundance [BKM21, FMK^{+}21, YK21]. acanthocephalan [CRWC22, SMP^{+}22a]. Acanthopagrus [LjY^{+}22, LZC^{+}24, ZFC^{+}23]. accurate [BFvG^{+}23]. acetic [ZVYD23]. acid [AABB23, GWL^{+}21, IKT^{+}23, LXZ^{+}21, TPSJS20a, YDM^{+}21, YY24, ZVYD23]. acid-fluorescence [IKT^{+}23]. acidophilic [CSVG^{+}22]. acids [HLH^{+}23b]. Acipenser [CVFRM^{+}23, CMLK^{+}22, GMK^{+}21, MCP^{+}22, PSH^{+}20]. acquisition [LFA20]. action [LHTW21]. active [IPC^{+}24, LYZ^{+}22]. activity [AABB23, CCV^{+}24, FWL^{+}23, GYC^{+}22, HBRN24, IFI^{+}21, LYX^{+}20, MSB^{+}21, NTB^{+}24, NSTS24, NNC^{+}23, SMD^{+}20, SZH^{+}22, SCX^{+}21, VBL^{+}23, Wel20, YJL^{+}23b, ZYZ^{+}20, ZYH^{+}23]. Acute [DNN^{+}22, SST^{+}22, NAaS^{+}22, WML^{+}21, WCT^{+}20, YDM^{+}21, ZGC^{+}22]. acyl [LRGAH21]. acyl-homoserine [LRGAH21]. adaptation [LDO^{+}21]. adapted [YY24]. adaption [LFG^{+}23]. additive [SVP^{+}24]. adhesion [GYC^{+}22, ZJF^{+}22, ZTY^{+}22]. administered [CLC^{+}22, RGT^{+}23]. Administration [LFL^{+}21, DXD^{+}21, OGR^{+}24, RKC^{+}23, ABA^{+}23]. administrative [NGC^{+}23]. adomavirus [YCL^{+}21]. Adriatic [MCP^{+}22]. adult [BKM21, HKB^{+}22, IAH21, MPA^{+}21, MFP21]. advantages [TPSJS20b]. Aeromonas [ASNA24, ASR^{+}23, BLJ^{+}21, BMO^{+}20, CZM^{+}22, CXC^{+}24, CZJ^{+}23, DCB20, DLS^{+}21, DZQ^{+}22, GPC22, HHW^{+}24, HRR^{+}24, HLH^{+}23b, JJZ^{+}22, KQW^{+}21, KB21, KKL^{+}22, LMJ^{+}24, LFL^{+}21, LH20, LFG^{+}23, MQK^{+}20, MBR^{+}22, PDM^{+}21, PGH^{+}21, PHB^{+}23, QWL^{+}21, SMD^{+}20, SYKC23, SMMM20, TSV^{+}22, VDD^{+}20, WLH^{+}20, WWW^{+}21, WQL^{+}23, XKF^{+}22, YCS^{+}21, ZJF^{+}22, ZLWL23, ZCY^{+}23, lZWC^{+}20, ZYL^{+}22, ZHY^{+}21]. aetiological [SMY^{+}22]. Aetobatus [DYN^{+}20]. affect [LEV^{+}20]. affected [AMM^{+}22, BTP^{+}24, SLPF^{+}20, TMM^{+}24, WAF^{+}21, YGR^{+}24]. affecting [PRW24, PS23, THH^{+}21]. affects [OBSH24, UMD22, WLTZ21]. affinis [JVM^{+}20, WCT^{+}23a]. Africa [AONDO24, SLPF^{+}20]. African [MDAO22]. after [ASF^{+}21, FYH^{+}21, MLHD^{+}23, RKC^{+}23, WQL^{+}23, WCT^{+}23a, XKF^{+}22]. against [ASA^{+}20, AABB23, AAZR^{+}20, BAM^{+}23, BIK^{+}21, BMK^{+}20, CZM^{+}22, CCV^{+}24, CLC^{+}22, DXD^{+}21, DZJ^{+}21, GZL^{+}21, GZW^{+}23, HHW^{+}24, HLpSC22, HWG^{+}21, HSH^{+}22, IFIZ22, JØM^{+}20, JKM^{+}20, JRZ^{+}22, KQW^{+}21, LH20, LdAPA^{+}23, LYX^{+}20, LXW^{+}20, LZZ^{+}23, LPGW^{+}22, MPS^{+}22, MKT^{+}21, MR23, NSTS24, NRKK21, PNØ^{+}21, QST^{+}21, RGS^{+}22, RGT^{+}23, REB^{+}20, RTA21, SMY^{+}24, SOE^{+}21, SGW^{+}22, SZH^{+}22, SLH^{+}24, SLA^{+}20, SCX^{+}21, TSV^{+}22, VBL^{+}23, WLK^{+}22, XFL^{+}21, XJZ^{+}22, YZM^{+}21, YGZ^{+}22, YJL^{+}23b, YYS^{+}24, YLW^{+}21, ZSL^{+}24a, ZLWL23, ZDF^{+}23, ZYH^{+}23, ZZL^{+}21, ZWB^{+}23, ZMZ^{+}22, ZSG^{+}23, dCJBG22, dAMHdOCF^{+}23]. agalactiae [HSH^{+}22, PHS^{+}20, PTK^{+}21, PSKW22, SWC21, XFL^{+}21, ZYH^{+}23]. AGD [EBA^{+}21, FSFS^{+}20]. age [LWK20, MFP21]. agent [CMM^{+}20, CDJ^{+}23, GYN^{+}24, HRR^{+}24, HSH^{+}22, JDA^{+}20, ZMZ^{+}22]. agents [AAZR^{+}20, SMY^{+}22]. agnathan [MHYW22b]. Agreement [BLJ^{+}21]. AHLQ20 [CCY^{+}23]. AHLQ20-01 [CCY^{+}23]. AHPND [BGZW22]. AHPND-causing [BGZW22]. aided [CCG24, GWF^{+}22]. AKT [ZZS^{+}24]. Alabama [BAT^{+}24]. alba [SMP^{+}22a]. albiflora [XJM^{+}21]. albino [SHS^{+}22a]. albus [LFZ^{+}22]. ALDH7A1 [LFL^{+}21]. alevin [ELB21]. algae [ZAG^{+}20]. algal [GKU21]. Algerian [BPA^{+}20]. alginate [XJZ^{+}22]. alginate-chitosan [XJZ^{+}22]. alginolyticus [AEE^{+}23, BCT^{+}20b, FYH^{+}21, LXZ^{+}21, THW^{+}22, ZTY^{+}22]. allergenic [AFKK22]. Allium [öTTS20]. alloherpesvirus [VGW^{+}21]. alloy [ZYB^{+}23]. almost [BGZW22]. along [GLS^{+}21]. alpha [FLH^{+}21]. alphavirus [ABM^{+}21, MNB^{+}24]. alpinus [MHH23]. alter [dCJBG22]. Alteration [WQL^{+}23]. Alterations [SMMM20, AEE^{+}23]. alternative [BAM^{+}23]. altivelis [THH^{+}21]. Amantadine [ZMZ^{+}22]. Amazon [dCJBG22]. amberjack [PZP^{+}20]. Ameiurus [BSS^{+}20, HCC^{+}22]. America [EGRD22, TBMP23]. American [EGRD22, OHL^{+}20, WZC^{+}24, YCL^{+}21]. Ameson [XZS^{+}23]. amnoonvirus [SLY^{+}24]. amoeba [LHTW21, VWE^{+}23]. amoebae [BTP^{+}24, JKM^{+}20]. amoebic [BHS^{+}21, CMM^{+}20, EBA^{+}21, FSFS^{+}20, WSS^{+}20]. among [CDJ^{+}23, LKV^{+}23, PSH^{+}21]. Amplicon [FMK^{+}21]. amplification [CMM^{+}20, CCG24, CDJ^{+}23, FCH^{+}22a, GWF^{+}22, HLW^{+}23, KDS^{+}21, PTK^{+}21, PYT^{+}22, PSV^{+}21, QWL^{+}21, TBK^{+}22, WML^{+}21, WCT^{+}23b, YDM^{+}21, ZHY^{+}21, ZLS^{+}21]. amplification-based [HLW^{+}23, ZLS^{+}21].

Amyloodinium [LZC^{+}24, ZHL^{+}21]. anabas [DDH^{+}22, DDH^{+}22]. anadromous [GRM^{+}23]. anaemia [AMJ^{+}21, AMB^{+}22, CGA^{+}20, DGB^{+}22, RGH^{+}22, TSS^{+}23]. anaesthetics [RCC^{+}21]. anal [HSS^{+}24]. analyses [CRP^{+}21, LHLS20, MAK^{+}23, YSW^{+}22]. Analysing [PNAS22]. Analysis [FVC^{+}21, lZWC^{+}20, ABBP24, AOT22, AEE^{+}23, AHSB^{+}22, BLL23, BFH^{+}23, CCY^{+}23, CLH^{+}21, xCqXfL^{+}21, FYH^{+}21, FLH^{+}21, GKTC^{+}21, HWG^{+}21, IHK^{+}22, KSY^{+}23, LFG^{+}23, LYF^{+}21, MIK^{+}24, MBN^{+}22, NSTS24, NNC^{+}23, NLL^{+}21, ØSH^{+}22, PTMC22, PBT^{+}23, PØN^{+}21a, QST^{+}21, RCW^{+}23, RCB^{+}22, SCAH22, SSM^{+}20, SDH22, SRLARG24, TIK^{+}22, TSC^{+}23, ÚPZ^{+}23, WZC^{+}24, WWK^{+}20, XKF^{+}22, YCS^{+}21, ZMZ^{+}23, ZGL^{+}20, ZHpT^{+}20, ZHL^{+}23, ZLZ21]. Ancyrocephalidae [IAO20]. Andrias [MJX^{+}23]. anemone [DNN^{+}22]. angasi [BWH^{+}20, BHT^{+}20]. AngHV [PEN^{+}21]. AngHV-1 [PEN^{+}21]. angiomatosis [MBF^{+}23]. angiomatosis-like [MBF^{+}23]. Anguilla [BKM21, PEN^{+}21, PTBSF22, WZC^{+}24, YCL^{+}21, ZWZ^{+}20]. anguillae [PEN^{+}21]. anguillarum [AGR^{+}22, BRWC23, HSH^{+}20, KKL^{+}22, LJHH23, RCB^{+}22, SDH22, WMS21]. anguillicaudatus [ZYZ^{+}20]. Anguillicola [BKM21, MNPS23]. Anguillicoloides [dYdRFF^{+}23]. anguillid [CCG24, dYdRFF^{+}23]. angustifolia [SCX^{+}21]. animal [AHBE^{+}20, KSY^{+}23, MWY^{+}22]. animals [NNC^{+}23]. Anisakis [MFP21]. annual [CTOH20]. anomalies [EGRD22, PDZ^{+}22]. antagonistic [ABA^{+}23]. Antarctic [FWL^{+}23, QGG^{+}21, VBG^{+}22, SDMO^{+}20]. anthelmintic [LdAPA^{+}23, MSF^{+}20]. anti [CCV^{+}24, LZZ^{+}23, ZLWL23]. anti-inflammatory [ZLWL23]. anti-parasites [LZZ^{+}23]. anti-viral [CCV^{+}24]. Antibacterial [KKL^{+}22, FWL^{+}23, HBRN24, HSH^{+}22, NSTS24, XQZ^{+}22, YJL^{+}23b, ZYZ^{+}20, ZYH^{+}23]. antibiotic [DXL^{+}20b, MWH^{+}21, QCZ^{+}22]. Antibiotics [DCB20, RCC^{+}21]. antibodies [ZDF^{+}23, ZGz^{+}22]. antibody [GZZ^{+}22, LWH^{+}21, MPS^{+}22, WWK^{+}20]. antifungal [MLL^{+}22]. antigen [DXL^{+}20a, ZSL^{+}24a]. antigen-presentation [DXL^{+}20a]. Antigenic [JXJ^{+}22]. Antimicrobial [VLSM^{+}20, ASA^{+}20, AMM^{+}22, GIS^{+}21, HLH^{+}23b, LS24, MBR^{+}22, PLS^{+}22, RBM^{+}22, XFL^{+}21, XJM^{+}21, DBDSP20]. antioxidant [CSGB^{+}21]. Antiparasitic [ZYB^{+}23, JKM^{+}20]. Antiviral [LYX^{+}20, SZH^{+}22, SCX^{+}21, YYS^{+}24]. aodp [ZJF^{+}22]. aortic [BME^{+}20]. Aphanomyces [BTN^{+}24, WBM21]. apoptosis [FZC^{+}23, HCW^{+}20, HWC^{+}21, KQW^{+}21, WKL^{+}22, ZYL^{+}22]. Aposelemis [KKC^{+}23]. Application [PTK^{+}21, ASG^{+}22, AMS^{+}24, CQG^{+}21, CCG24, GZZ^{+}22, LTW^{+}21, LjY^{+}22, MWY^{+}22, SNY^{+}21, WYR^{+}23, YTS22, ZZF^{+}23]. applications [SAAB^{+}21]. approach [WLK^{+}22]. approaches [OKL^{+}22]. Aptamer [LXW^{+}20, GYN^{+}24, WLK^{+}22]. Aptamer-based [LXW^{+}20]. aptamers [YLW^{+}21]. aptasensor [GYN^{+}24]. aquabirnavirus [dYdRFF^{+}23]. aquaculture [AHCSCI20, BTN^{+}21, CZMP21, DWK^{+}22, DBDSP20, KNG^{+}22, LPGA^{+}21, MHP^{+}20, MIK^{+}24, MPP^{+}20b, PTBSF22, QRTD^{+}22, RCG22, RGS^{+}22, SAAB^{+}21, SMD^{+}20, TPSJS20a, TPSJS20b, ÚPZ^{+}23, VWE^{+}23, WYR^{+}23, ZYB^{+}23, ZOZ^{+}20, FM23]. Aquareovirus [LKD^{+}24]. aquaria [KSY^{+}23, NLH^{+}20]. aquarium [ACGDO^{+}21, JLM^{+}20, LFSB^{+}21, SBVC21, SCS24]. aquarium-housed [ACGDO^{+}21]. aquatic [LWL^{+}24, MWY^{+}22, NNC^{+}23, RCC^{+}21, ZLS^{+}21]. aquatilis [LHLS20]. Aquimarina [MSS^{+}20]. arapaima [dCJBG22, dCJBG22, dAMHdOCF^{+}23]. Arctic [GRM^{+}23, MHH23]. areas [SPHM23]. argenteus [JZW^{+}23, ZHL^{+}21]. Argonaute [WCT^{+}23b]. Argulidae [POS^{+}20]. Argulus [POS^{+}20]. argus [HHW^{+}24]. armoured [DDH^{+}22]. arowana [LS24]. art [BONB20]. Artemia [PKF^{+}21]. arteritis [BME^{+}20]. Arthrospira [BAM^{+}23]. arugamensis [MHP^{+}20]. arylaminoisoquinolin [VBL^{+}23]. As-Cathelicidin4 [CZM^{+}22]. Ascanius [PDM^{+}21, PGH^{+}21]. ascO [ZYL^{+}22]. Asia [ZGL^{+}20]. Asian [CSDS21, CGK23, GXW^{+}23, PTW21, PPWC20, RBM^{+}22, RKC^{+}23, SSMC22, SKS^{+}24]. Aspects [ADAO22]. assay [CMM^{+}20, CQG^{+}21, CCG24, KDS^{+}21, LWH^{+}21, lLWC^{+}24, MJX^{+}23, PTP^{+}23, PS23, QWL^{+}21, SCALIAH23, SPS23, SMY^{+}22, SKR^{+}20, TBK^{+}22, TCS21, YLW^{+}21, ZZZ^{+}23]. assays [GWF^{+}22, PDM^{+}21, YTS22]. assemblages [TBB^{+}21]. assess [KH20]. Assessing [CDJ^{+}23, LAH21, PAH^{+}22, YYS^{+}22]. Assessment [BMJ^{+}21, CMN^{+}20, ASA^{+}20, LPGA^{+}21, NRY^{+}23, NLPM^{+}21]. associated [BZP^{+}22, BMJ^{+}21, BMSC21, DYN^{+}20, EBAH24, FMK^{+}21, GKU21, JØM^{+}20, JLM^{+}20, KaVDH^{+}21, KKH^{+}24, LDR^{+}22, LFP^{+}24, MWZ^{+}20, MCP^{+}22, NKS^{+}21, NPP^{+}24, PPDG^{+}22, PEN^{+}21, PTW21, PSH^{+}20, SKR^{+}20, SNT^{+}22, SLE^{+}20, SRP^{+}21, WXM^{+}22]. association [Roh22]. Astragalus [LDW^{+}23]. astrocyte [LHC^{+}22]. astrocyte-like [LHC^{+}22]. asymptomatic [KBG^{+}23]. AT-rich [YLC^{+}21]. Atlantic [ABM^{+}21, AFKK22, AMB^{+}22, BHS^{+}21, BRR^{+}23, BFH^{+}23, BMJ^{+}21, BONB20, BFC^{+}23, CWB^{+}20b, CVDO^{+}20, CVDB^{+}21, CdPM^{+}21, DHSM^{+}20, DPLL^{+}21, DMH^{+}21, EBA^{+}21, FSFS^{+}20, FCH^{+}22b, FMP^{+}21, FMH^{+}22, FEH^{+}22, HSS^{+}24, HPR^{+}22, JQT20, JBG^{+}23, JLM^{+}20, KBR^{+}22, KDR^{+}22, KKK^{+}23, LMJ^{+}24, LGP^{+}20, LNK^{+}21, LFL^{+}21, LdL20, MIR^{+}22, MBJ20, MLM^{+}20, MGG22, MBS21, MFP21, MLHD^{+}23, MR23, OBSH24, ØSH^{+}22, PSH^{+}21, PBF^{+}21, PRW24, QRTD^{+}22, RØLW23, RGH^{+}22, RTA21, SCIAH21, SRT^{+}21, SRE^{+}21, SSK23, SVO^{+}22, SST^{+}22, SKR^{+}20, SNT^{+}22, SBVC21, SdSNK^{+}22, SCD^{+}23, TBK^{+}20, TCC^{+}24, UD22, UMD22, WJPS21, WBV^{+}20, WAF^{+}21, WSS^{+}20]. atlantica [KKL^{+}22]. Atresia [CZMP21]. attachment [BONB20]. attack [JJZ^{+}22]. attenuated [ABM^{+}21, BMK^{+}20, BMSC21, LTW^{+}21, LPGW^{+}22, SMY^{+}24]. attenuation [ZJF^{+}22, ZSL^{+}24b]. atypical [DCB20, KKL^{+}22, PDM^{+}21, PGH^{+}21]. AuNP [KDS^{+}21]. aurata [AEE^{+}23, CSGB^{+}21, CSVG^{+}22, CFGCCE21, FEM^{+}20, FK21, HZFT21, KPC^{+}22, LDC^{+}22, MBN^{+}22, RGT^{+}23]. auratus [LYF^{+}21, MHA^{+}20, SCRS^{+}23, WZG^{+}20, YCC^{+}21, MLGRF^{+}20, WWW^{+}21, XKF^{+}22, YJL^{+}23b, ZLWL23]. aureus [ZSL^{+}24a, WWW^{+}21]. aurorubens [HB20]. Austrian [LFSB^{+}21]. authored [SNSP^{+}23]. Autophagy [LPL22, CZC^{+}22, JPC^{+}21, LPW^{+}20]. Availability [MHA^{+}20]. avirulent [RGS^{+}22]. axis [WQL^{+}23]. ayu [THH^{+}21].

b [BME^{+}20, AANY23, GYC^{+}22, NLL^{+}21, WCT^{+}20]. B-cell [GYC^{+}22]. bacillary [PØN^{+}21a, PØN^{+}21b]. Bacillus [CRP^{+}21]. background [CCV^{+}24, RGS^{+}22]. BacMCP [ZSG^{+}23]. bacteria [AABB23, BFC^{+}23, DWY^{+}20, MAK^{+}23, QCZ^{+}22, SRLARG24, SMMM20]. Bacterial [MAG24, Tan23, AMM^{+}22, CLC^{+}22, FFB^{+}21, GWL^{+}21, JAC^{+}22, KKBL21, NRKK21, NLL^{+}21, PRC20, PLS^{+}22, QFW^{+}23, SCIAH21, ZWZ^{+}20, ZZL^{+}21, ZLS^{+}21]. Bactericidal [ASR^{+}23, VBL^{+}23]. bacteriological [LGP^{+}20]. bacteriophage [Wel20]. bacterium [WHK^{+}23]. baculovirus [ZSG^{+}23, CCCC23]. baerii [CVFRM^{+}23]. bagarii [KNG^{+}22]. Balancing [PKF^{+}21]. ballan [PDM^{+}21, PGH^{+}21]. Baltic [BRCI23, LKV^{+}23, NAPP24, WAF^{+}21]. bamboo [MAG24]. Banggai [KSG^{+}23]. Bangladesh [DDDJ^{+}20, DDHT^{+}22]. Banyuwangi [LDR^{+}22]. Barbodes [SCS24]. barbs [CCP^{+}21, SCS24]. barramundi [GKTC^{+}21, KSD^{+}20, SZZ^{+}23]. barrier [LHC^{+}22]. Basal [SHS^{+}22a]. based [CCY^{+}23, GYN^{+}24, HLW^{+}23, LRGAH21, LXW^{+}20, LWT^{+}22, lLWC^{+}24, NAaS^{+}22, NNC^{+}23, NKZ^{+}20, OKL^{+}22, PTP^{+}23, SSMC22, TIK^{+}22, THW^{+}22, WLK^{+}22, WRMT24, YTS22, ZMZ^{+}23, ZZF^{+}23, ZZL^{+}21, ZLS^{+}21, GZL^{+}21, HPR^{+}22, ZLZ21]. Basibranchial [PRW24]. Basin [SCV^{+}20]. basR [SGW^{+}22]. bass [BBG21, CYX^{+}22, CCY^{+}23, CSDS21, FFB^{+}21, GWF^{+}22, GZW^{+}23, JXJ^{+}22, JRZ^{+}22, JSZ^{+}23, LDM^{+}20, PPWC20, RBM^{+}22, SSMC22, WBM21, WYR^{+}23, YK21, ZVB^{+}21, SGW^{+}22]. bath [AHBE^{+}20, HWM^{+}24, PGH^{+}21]. baths [dAMHdOCF^{+}23]. Bathygobius [MRM21]. batomorph [EGRD22]. Bay [PSH^{+}21, VPW^{+}21]. Bayesian [LGP^{+}20]. be [HLpSC22, KKK^{+}23]. bean [RGS^{+}21]. belly [GKTC^{+}21]. belonging [ZVYD23]. benign [SBB21]. Berg [SCRS^{+}23]. bergylta [PDM^{+}21, PGH^{+}21]. betaine [CLC^{+}22]. betanodavirus [JVM^{+}20, KPC^{+}22]. Betta [SBB21]. between [BLJ^{+}21, KSG^{+}23, KSY^{+}23, PTMC22, PAH^{+}22, PNØ^{+}21, PØN^{+}21b, PKF^{+}21, SCIAH21, SPM^{+}20, WSS^{+}20]. bicirrhosum [LS24]. bicuspidata [MLL^{+}22]. big [GKTC^{+}21]. biliary [SCD^{+}23]. bind [LHTW21]. binding [GYC^{+}22, MLGRF^{+}20, WKL^{+}22, TASP^{+}20]. biochemical [WSS^{+}20]. biofilm [DCB20, FFB^{+}21, HLH^{+}23b, LFA20, LXZ^{+}21, OCS^{+}23, ORV^{+}24, SVP^{+}24, VRC^{+}22]. biofilms [PLS^{+}22]. biofloc [HBRN24, PDL^{+}23, WGE^{+}24]. Biokos [WHK^{+}23]. biology [KKH^{+}24, PVCG20]. biomarker [WAF^{+}21]. biomarkers [ŘB24]. biosecurity [LDR^{+}22]. biosensor [THW^{+}22]. biotypes [FEMMI^{+}23]. Biotyping [RCG22]. birnavirus [LYC^{+}20]. bivalent [LTW^{+}21]. bivalve [MCA^{+}20]. Bivalvulida [KRS^{+}20]. black [FMK^{+}21, WSW^{+}20, HSG^{+}21]. bladder [MNPS23]. Bleeker [SAKC20]. Bloch [CGK23, GKU21, GKTC^{+}21, KSD^{+}20, PTW21]. blochii [ASG^{+}22, AMS^{+}24, YTH^{+}24]. blood [LHC^{+}22, MMTD^{+}21]. blooded [SFKB^{+}20]. blue [AGR^{+}22, LPGA^{+}21, LPGW^{+}22, SLA^{+}20, VGW^{+}21, ZHY^{+}21, ZAG^{+}20]. bluefin [WBV^{+}20]. Boca [FM23]. bodies [SPHM23]. body [FK21, PØN^{+}21b, TMM^{+}24]. Bohemian [SMD^{+}20]. Bonamia [BWH^{+}20, BHT^{+}20]. Bonaparte [MCP^{+}22]. Book [RM20]. Boomker [ADAO22]. Bosnia [GVB^{+}21]. Bostrychus [DWY^{+}20]. both [YYS^{+}24]. botryosa [CMLK^{+}22, GMK^{+}21]. Boxshall [GLS^{+}21]. brackish [TPSJS20b]. brain [AMS^{+}24, DHSK^{+}21, JPC^{+}21, LHC^{+}22, LZK^{+}23, MVW^{+}20]. brama [APS23]. branchial [BME^{+}20]. brasiliensis [LXW^{+}21]. Bravo [GLS^{+}21]. Brazilian [TCC^{+}24]. bream [APS23, ENS^{+}23, IHK^{+}22, KPC^{+}22, LjY^{+}22, PDZ^{+}22, REB^{+}20]. breeding [PNØ^{+}21]. brevirostris [RZTV21]. brine [ZVYD23]. British [JDA^{+}20, MBJ20]. broodfish [WZG^{+}20]. broodstock [LdL20, NCA^{+}20, SEL^{+}20]. brook [BMJ^{+}21, SCV^{+}20, SEL^{+}20]. Broussonet [TCC^{+}24]. brown [ASF^{+}21, BSS^{+}20, CDY^{+}20, FSW^{+}21, LNK^{+}21, LFSB^{+}21, SPKS^{+}20, WBB^{+}21]. brown-marbled [CDY^{+}20]. Brunswick [PSH^{+}21]. brush [GXW^{+}23]. brush-clawed [GXW^{+}23]. bryosalmonae [LKV^{+}23, OSM22a, SSM^{+}21, WBB^{+}21]. Bucephalidae [KNG^{+}22]. bulbosa [AFKK22]. bullfrog [TSC^{+}23]. bullhead [BSS^{+}20, FSW^{+}21]. bumper [SBVC21]. Bunodeopsis [DNN^{+}22]. burbot [BMO^{+}20]. Burchell [ADAO22, MDAO22]. burden [LNK^{+}21]. buttnerae [LdAPA^{+}23, SMP^{+}22a].

C [BLJ^{+}21, HHW^{+}24, LWR^{+}20, MWY^{+}20, WWW^{+}21, ZYZ^{+}20]. C-13326 [HHW^{+}24]. C-terminal [WWW^{+}21]. C-type [ZYZ^{+}20]. C. [LYF^{+}21, YCC^{+}21]. C6 [KSM^{+}21]. C6/ [KSM^{+}21]. CadC [GWL^{+}21]. cadherin [REB^{+}20]. caecum [ZDF^{+}23]. cage [NHG^{+}23, NKZ^{+}20, PNAS22, SNT^{+}22, WJPS21]. cage-farmed [SNT^{+}22]. cage-level [WJPS21]. caged [YK21]. cages [BONB20]. calcarifer [CSDS21, CGK23, GKTC^{+}21, GKAS23, KSD^{+}20, PTW21, PPWC20, RBM^{+}22, RKC^{+}23, SZZ^{+}23, SSMC22, SKS^{+}24]. calculi [DBC^{+}22]. California [GMK^{+}21]. californiensis [HRR^{+}24, LHH^{+}22, NKS^{+}21]. Caligus [BFC^{+}23, GLS^{+}21, JMA^{+}22, MCA^{+}20]. callainos [SHS^{+}22a]. campbellii [CCL^{+}24]. Canada [JDA^{+}20, JBG^{+}23, JLM^{+}20, MBJ20, RGH^{+}22, VPW^{+}21]. candidates [SGW^{+}22]. cap [CZJ^{+}23]. capacities [IFIZ22]. capsid [GYN^{+}24]. capsular [RCS^{+}22]. captive [DVR^{+}20, HCC^{+}22, IKT^{+}23, RZTV21]. capture [GYN^{+}24, MCA^{+}20]. Carangidae [SBVC21]. Carassius [KKH^{+}24, MLGRF^{+}20, SCRS^{+}23, WZG^{+}20, WWW^{+}21, WQL^{+}23, XKF^{+}22, YJL^{+}23b, ZLWL23]. carbohydrate [WWW^{+}21]. carcinoma [SHS^{+}22a, SAKC20]. card [PTK^{+}21]. cardiac [NKö^{+}21, PRW24]. cardinalfish [KSG^{+}23]. cardiomyopathy [CdPM^{+}21, FMP^{+}21, FMH^{+}22]. care [CMM^{+}20, MAG24]. caridean [CZM^{+}22]. carnivorous [SVC^{+}21]. Carnobacterium [RKL^{+}20, SMP^{+}22b]. Carolina [BTN^{+}24]. Carp [MPPF21, MPA^{+}21, MPP^{+}20b, PPDG^{+}22, ABBP24, BAS^{+}22, BPA^{+}21, DZJ^{+}21, FAG^{+}22, GZL^{+}21, JJZ^{+}22, KKH^{+}24, LHLS20, LLJ^{+}21, MPP20a, MBF^{+}23, MHA^{+}20, PEś^{+}20, RCW^{+}23, SMC^{+}22, SGR^{+}20a, STB^{+}21, SMMM20, SCRS^{+}23, ÚPZ^{+}23, WQL^{+}23, XKF^{+}22, YGC^{+}20, YJL^{+}23b, YJL^{+}23a, YLW^{+}21, ZZL^{+}21, ZOZ^{+}20]. carp-derived [ABBP24]. Carpathian [SCV^{+}20]. carpio [BAS^{+}22, MBF^{+}23, STB^{+}21, JJZ^{+}22, MPPF21, MPP20a, MPA^{+}21, SMC^{+}22, SGR^{+}20a]. carrageenin [CSGB^{+}21, CSVG^{+}22]. Carrasius [MHA^{+}20]. carrier [LTW^{+}21, SM21]. Case [AGH^{+}23, JSW^{+}21, PNØ^{+}21, SGU^{+}21, BTN^{+}24, DVR^{+}20, KRS^{+}20, MLHD^{+}23, PPDG^{+}22, PEś^{+}20]. cases [KSG^{+}23]. caspases [DWY^{+}20]. cataract [DPA^{+}24]. categorization [BLJ^{+}21]. catesbeiana [TSC^{+}23]. catfish [AGR^{+}22, DXD^{+}21, DDH^{+}22, FSW^{+}21, HCC^{+}22, KNG^{+}22, LKD^{+}24, LPGA^{+}21, LPGW^{+}22, MDAO22, OGR^{+}24, PNØ^{+}21, PØN^{+}21a, PØN^{+}21b, SOE^{+}21, SGW^{+}22, SLA^{+}20, VGW^{+}21, YGR^{+}24, YLC^{+}21, ZHL^{+}22, PRC20]. Cathelicidin4 [CZM^{+}22]. Cathelicidins [SLA^{+}20]. catus [HCC^{+}22]. Caudal [FEM^{+}20, LYF^{+}21, YCC^{+}21]. caudata [NK22]. caught [KRS^{+}20, VDC20]. causative [CMM^{+}20, HRR^{+}24]. cause [KKBL21, SPKS^{+}20, WBM21, WCT^{+}20]. caused [AONDO24, BTN^{+}24, CSVG^{+}22, DWK^{+}22, DNN^{+}22, GKTC^{+}21, HHJ^{+}21, IKT^{+}23, KKC^{+}23, LWR^{+}20, MBS21, PAS^{+}21, PØN^{+}21a, PØN^{+}21b, PSKW22, TMM^{+}24, ZVB^{+}21, dRRRPSR^{+}21]. causes [BHT^{+}20, MLHD^{+}23, TSS^{+}23, ZSL^{+}24b]. causing [BGZW22, SPKS^{+}20, XZS^{+}23]. ccBAX [YCC^{+}21, FZC^{+}23]. CD4 [ASF^{+}21]. CD4-1 [ASF^{+}21]. CD40 [XFL^{+}21]. CDC [VRC^{+}22]. cDNA [LZX^{+}23]. cell [ABBP24, ASG^{+}22, ASA^{+}20, AMS^{+}24, BBS^{+}20, DWK^{+}22, DGB^{+}22, FSS^{+}21, FAG^{+}22, FSW^{+}21, GSJ^{+}20, GPL^{+}22, GYC^{+}22, HCW^{+}20, JZW^{+}23, JVM^{+}20, KSM^{+}21, LHC^{+}22, LjY^{+}22, LZC^{+}24, LZK^{+}23, LPW^{+}20, LXZ^{+}21, LFZ^{+}22, MWY^{+}22, NRC^{+}23, PBF^{+}21, SHS^{+}22a, SMG^{+}23, WKL^{+}22, YPT^{+}22, YGC^{+}20, YCC^{+}21, ZFC^{+}23, ZSL^{+}24b]. cell-binding [GYC^{+}22]. cell-wall [HCW^{+}20]. Cells [Muf23, ABBP24, CSVG^{+}22, EBBIAH21, FVC^{+}21, FSW^{+}21, HCW^{+}20, HWC^{+}21, JPC^{+}21, LYF^{+}21, NRKK21, PHS^{+}20, SBD21, TASP^{+}20, VBG^{+}22]. cellular [CZC^{+}22, WWK^{+}20]. Center [SSM^{+}20]. central [DDH^{+}22, VLSM^{+}20, DNN^{+}22]. centres [PZP^{+}20]. cepa [öTTS20]. cephalus [RBP^{+}20a]. cerebralis [AWF22, KRS^{+}20]. cereus [CRP^{+}21]. Cestoda [APS23, HSS^{+}24, KBMD21]. cestode [IFI^{+}21]. cestodes [IFIZ22, KSI^{+}20, KBMD21]. cf [KB21]. CFP [MHA^{+}20]. CFP-10 [MHA^{+}20]. chain [SPS23, SMY^{+}22]. challenge [BMO^{+}20, EBA^{+}21, HSM^{+}22, IAH21, LWK20, LWL^{+}24, LGJ21, MLL^{+}22, MIR^{+}22, PGH^{+}21, PNØ^{+}21, PØN^{+}21b, QFW^{+}23, XKF^{+}22]. challenged [DLS^{+}21, KKL^{+}22, SZZ^{+}23, ZVYD23]. change [LNK^{+}21, RSPB22]. change-driven [LNK^{+}21]. changes [BFH^{+}23, DXL^{+}20b, FEH^{+}22, KAS^{+}24, SST^{+}22, TMM^{+}24, XXX^{+}24]. Channa [HHW^{+}24, PHB^{+}23]. Channel [ZHL^{+}22, AGR^{+}22, FSW^{+}21, KNG^{+}22, LKD^{+}24, LPGA^{+}21, LPGW^{+}22, OGR^{+}24, SLA^{+}20, VGW^{+}21, YLC^{+}21]. char [MHH23, SGU^{+}21]. character [ZHpT^{+}20]. characteristics [FAG^{+}22, LWL^{+}24]. Characterization [BGZW22, CLS^{+}22, FMP^{+}21, FMH^{+}22, LXW^{+}21, LH20, LDC^{+}22, MSS^{+}20, NLL^{+}21, WHC^{+}21, Wel20, WWK^{+}20, ASG^{+}22, AMS^{+}24, AANY23, AMM^{+}22, CYX^{+}22, CVDO^{+}20, CXC^{+}24, CCCC23, CdPM^{+}21, DWY^{+}20, EBELIAH20, ESEI20, ELS^{+}23, FLH^{+}21, GYN^{+}24, GPL^{+}22, HAK^{+}20, JZW^{+}23, KNG^{+}22, KSG^{+}23, KQW^{+}21, KaVDH^{+}21, LCTC20, LFZ^{+}21, LjY^{+}22, LHLS20, LZK^{+}23, LLR^{+}20, MLL^{+}22, MPMB^{+}20, MWY^{+}22, MIC^{+}24, NRY^{+}23, NLPM^{+}21, NGC^{+}23, NKZ^{+}20, PCWC20, QFW^{+}23, RRK^{+}23, RBP^{+}20b, STB^{+}21, VIB^{+}21, WSW^{+}20, WLH^{+}20, YLW^{+}21, ZFC^{+}23, ZDF^{+}23, ZGz^{+}22]. Charleston [BTN^{+}24]. charr [GRM^{+}23]. Cheat [WBM21]. Chelon [HZFT21]. chelonae [JW22]. chemicals [ŘB24]. Chile [DMH^{+}21, GLS^{+}21, LGP^{+}20, MIC^{+}24, VIB^{+}21]. Chilean [AHCSCI20, AHTCDI20, GIS^{+}21, IVR^{+}23, IMAH21, MSZV^{+}23, SCIAH21]. chilensis [IAH21, MCA^{+}20, SCIAH21]. Chiloscyllium [MAG24]. chimeric [MPS^{+}22]. China [DXL^{+}20b, LWR^{+}20, LLR^{+}20, QCZ^{+}22, WSW^{+}20, WLH^{+}20, WXM^{+}22, XZS^{+}23]. Chindongo [SHS^{+}22a]. Chinese [CXC^{+}24, CRP^{+}21, LFZ^{+}22, MWY^{+}22, SCS24]. Chinook [JSW^{+}21, LFP^{+}24, NKS^{+}21, NPP^{+}24, NLPM^{+}21, PPT^{+}20, SST^{+}22]. Chionodraco [FWL^{+}23]. chip [HLW^{+}23, ZLS^{+}21]. chitosan [ASA^{+}20, XJZ^{+}22]. chloride [dAMHdOCF^{+}23]. chlorine [dAMHdOCF^{+}23]. chloroquine [LL21]. Chloroscombrus [SBVC21]. Chlorpyrifos [YGC^{+}20]. chondrocranial [RZTV21]. chromatic [ZZZ^{+}23]. chromatography [EJH^{+}23]. Chromatophoroma [SBB21]. chronic [LFP^{+}24, SEL^{+}20, ZBL^{+}23]. Chryseobacterium [BMK^{+}20, XXX^{+}24]. chrysophrii [RGT^{+}23]. chrysops [FFB^{+}21]. chrysurus [SBVC21]. chuatsi [GPL^{+}22, HWG^{+}21, LZX^{+}23, MWY^{+}22]. Cichlidae [IAO20]. Cichlidogyrus [IAO20]. Ciliata [YY24]. ciliate [HHZ^{+}23, JØM^{+}20, TZL^{+}22]. Ciliophora [LLR^{+}20]. circadian [CCV^{+}24]. circulation [DDDJ^{+}20]. cis} [YLC^{+}21]. cis}-element [YLC^{+}21]. citratus [POS^{+}20]. citric [YY24]. CK108 [CLS^{+}22]. clade [ZVYD23]. cladophorae [ACGDO^{+}21]. clams [LYC^{+}20]. Clarias [ADAO22, MDAO22]. clarkii [BAT^{+}24, HSG^{+}21, RFN^{+}23, SCX^{+}21, ZWX^{+}22, ZLZ21]. class [FLH^{+}21, QFW^{+}23]. clawed [GXW^{+}23]. Climate [LNK^{+}21, RSPB22]. Clinical [MBF^{+}23, TMM^{+}24, CRP^{+}21, LCTC20, SHS^{+}22a, SKS^{+}24, JBG^{+}23]. clinically [CGK23]. ClO [ZWB^{+}23]. clonal [RGS^{+}22, SSP^{+}23]. cloned [LZX^{+}23]. cloning [FLH^{+}21, YSW^{+}22]. Clonorchis [ZGL^{+}20]. Clostridium [BZP^{+}22]. Clupea [HSH^{+}20]. clupeaformis [HKB^{+}22]. cluster [ÚPZ^{+}23]. CMS [FMP^{+}21]. Co [FEMMI^{+}23, IVR^{+}23, ALDMLEM20, CWB^{+}20b, XKF^{+}22]. Co-existence [FEMMI^{+}23]. co-expression [XKF^{+}22]. co-infection [CWB^{+}20b]. co-infections [ALDMLEM20]. Co-occurrence [IVR^{+}23]. coagulation [GKAS23]. coast [BPA^{+}20, JBG^{+}23, TCC^{+}24]. coastal [MFP21]. coastlines [KPC^{+}22]. coated [MCM^{+}23]. cobalt [SHS^{+}22a]. cobalt-zebra [SHS^{+}22a]. Coccomyxa [ZAG^{+}20]. Coccomyxa-like [ZAG^{+}20]. cocosensis [MRM21]. cod [AFKK22, BRCI23, WXM^{+}22]. cohabitation [BHT^{+}20]. coho [TMM^{+}24, KKBL21, PPT^{+}20]. cohort [ØSH^{+}22]. coinfection [BFC^{+}23, KB21]. coioides [FYH^{+}21, HWZ^{+}21, LCYL20, NGC^{+}23, ZWB^{+}23]. Cold [SZZ^{+}23, CZC^{+}22, SNT^{+}22]. coldwater [KKBL21, SEL^{+}20]. colistin [QCZ^{+}22]. colistin-resistant [QCZ^{+}22]. collected [AMM^{+}22]. collection [JBG^{+}23]. colloidal [HZFT21]. colonization [ZTY^{+}22]. colonizing [VWE^{+}23]. Colorado [LHH^{+}22]. coloration [LWT^{+}22]. coloration-based [LWT^{+}22]. colorimetric [KDS^{+}21, PTK^{+}21, PSV^{+}21]. Colossoma [LdAPA^{+}23]. colour [CCV^{+}24]. Columbia [JDA^{+}20, MBJ20]. column [TBK^{+}20]. columnare [BMO^{+}20, ELB21, FFB^{+}21, LFA20]. combats [LXZ^{+}21]. Combined [FCH^{+}22b, CCG24, FCH^{+}22a, KDS^{+}21, PTP^{+}23, PSV^{+}21, WCT^{+}23b]. commercial [PBF^{+}21]. common [FAG^{+}22, GZL^{+}21, JJZ^{+}22, MPP20a, MRM21, MPA^{+}21, PEś^{+}20, SMC^{+}22, VDC20, VDD^{+}20, YTH^{+}24, YGC^{+}20, ZZL^{+}21, ZOZ^{+}20]. commonly [JW22]. communities [KSI^{+}20]. community [MWZ^{+}20, TPSJS20a]. Comparative [BRCI23, BFH^{+}23, BRWC23, IFIZ22, LFG^{+}23, MAK^{+}23, QST^{+}21, SCAH22, WZC^{+}24, YCS^{+}21, LGJ21, YZM^{+}21]. compared [MSF^{+}20]. Comparison [HHZ^{+}23, LJK^{+}24, PPWC20, SCIAH21, SSM^{+}21, TBKY20, BMSC21, FSFS^{+}20, RKL^{+}20]. compatible [NAaS^{+}22]. complex [DXL^{+}20a, FLH^{+}21, QFW^{+}23, RGS^{+}22]. complexes [SSP^{+}23]. compliance [MMTD^{+}21]. component [DZQ^{+}22, FEM^{+}20, HSH^{+}22]. Composition [KSI^{+}20, BTN^{+}21, MWZ^{+}20, SGR^{+}20b]. compound [HLpSC22]. compounds [JKM^{+}20]. Comprehensive [AEE^{+}23]. concentration [EJH^{+}23, IAH22, POS^{+}20]. concentrations [BHS^{+}21, OCS^{+}23]. Concurrent [LCCC22, PHB^{+}23]. condition [MPP20a, MNPS23]. conditions [AHSB^{+}22, BFH^{+}23, HSH^{+}20, PEś^{+}20, QGG^{+}21, WBB^{+}21]. Confirmatory [IPC^{+}24]. conger [IAH21, SCIAH21]. congolensis [MDAO22]. conjoined [YWZ^{+}21]. Consequences [PPT^{+}20, ALDMLEM20]. conservation [PTBSF22, SPHM23]. consistent [KMH^{+}21]. constituent [SCX^{+}21]. constraints [Tan23]. Construction [ZWZ^{+}20]. contaminated [FSW^{+}21]. Contracaecum [BRCI23, VKK^{+}22]. contribute [Din21, WWW^{+}21]. Contribution [CZJ^{+}23, HWZ^{+}21]. control [HZC^{+}24, MSF^{+}20, SAAB^{+}21, SZH^{+}22]. controller [MCA^{+}20]. conventional [ÚPZ^{+}23]. cooking [CMN^{+}20]. Copaifera [MSF^{+}20]. copepod [HRR^{+}24, LHH^{+}22, NKS^{+}21, TBB^{+}21, YK21]. Copepoda [AONDO24]. copepodids [DPLL^{+}21]. copper [ZYB^{+}23, ZMZ^{+}23]. cord [BBS^{+}20, GPL^{+}22]. Core [KSY^{+}23]. Coregonus [HKB^{+}22]. corni [ACGDO^{+}21]. corpuscles [SAKC20]. correlates [VFR^{+}23]. correlation [GSS^{+}21, PNØ^{+}21]. correlations [PØN^{+}21b]. Cost [DMH^{+}21]. Cost-effectiveness [DMH^{+}21]. could [HLpSC22]. coumarin [LZZ^{+}23]. countries [TNB^{+}22]. coupled [GLS^{+}21, PTK^{+}21]. Cover [Ano21a, Ano21b, Ano21c, Ano22a, CWB^{+}20a, HLH^{+}23a, MHYW22a, SHS^{+}22b]. covert [WLTZ21]. CPA [PSV^{+}21]. crab [Din21, GXW^{+}23, LLR^{+}20, MSS^{+}20, XZS^{+}23]. crassum [HSS^{+}24]. crassus [BKM21, MNPS23, dYdRFF^{+}23]. crayfish [BBD^{+}22, BAT^{+}24, DBPH21, SCX^{+}21, SFKB^{+}20, ZWX^{+}22, ZLZ21]. CRC [FM23, RM20]. CRediT [SNSP^{+}23]. CrERV [LFZ^{+}22]. Crete [KKC^{+}23]. CRISPR [NAaS^{+}22, SSMC22]. CRISPR-based [NAaS^{+}22, SSMC22]. Critical [SYKC23, AHTCDI20, DBPH21]. croaker [LWR^{+}20, LLC^{+}21, LCW^{+}24, SJL^{+}23]. Croatia [BBD^{+}22, ZOZ^{+}20, ZVB^{+}21]. crocea [LWR^{+}20, LLC^{+}21, SJL^{+}23]. Cross [BMK^{+}20, LPGW^{+}22, MBJ20, PSV^{+}21, TBK^{+}20]. cross-priming [PSV^{+}21]. Cross-protection [BMK^{+}20]. Cross-protective [LPGW^{+}22]. Cross-sectional [MBJ20]. cross-stitch [TBK^{+}20]. crossroads [PTBSF22]. crowding [DPLL^{+}21]. Crucial [DYPN22]. crucian [LHLS20, MHA^{+}20, WQL^{+}23, XKF^{+}22, YJL^{+}23b]. Crustacea [POS^{+}20]. crustaceans [JBP24]. cryopreservation [TKB^{+}21]. Cryptocaryon [xCqXfL^{+}21, DXL^{+}20a, TZL^{+}22, WHK^{+}23, XJM^{+}21, YY24, ZYB^{+}23, ZMZ^{+}23, ZWB^{+}23]. Cryptocotyle [DMB^{+}21]. Cryptosporidium [SKS^{+}24]. Ctenopharyngodon [YJL^{+}23a]. culture [BFvG^{+}23, DGB^{+}22, FSS^{+}21, GSJ^{+}20, HBRN24, LGP^{+}20, MHA^{+}20, PTMC22, SZH^{+}22, SJK^{+}21, WGE^{+}24]. cultured [AEE^{+}23, BLL23, BMO^{+}20, CRWC22, EBA^{+}21, GMK^{+}21, LJK^{+}24, LWR^{+}20, LH20, MWH^{+}21, MAK^{+}23, NK22, NRY^{+}23, NHG^{+}23, NKZ^{+}20, PHS^{+}20, PSH^{+}20, SJK^{+}21, SRLARG24, SMP^{+}20, SWC21, SKS^{+}24, WSW^{+}20, XZS^{+}23, YTH^{+}24, YCL^{+}21]. cultures [NKö^{+}21]. Current [MPPF21]. curvature [LFP^{+}24]. curved [TBK^{+}20]. Cutaneous [HCC^{+}22, MBF^{+}23]. cutthroat [RFN^{+}23]. Cuvier [HB20]. cuvieri [WQL^{+}23]. cycle [CTOH20, JMA^{+}22, MBJ20, NKö^{+}21]. cycloancistrium [dCJBG22, dAMHdOCF^{+}23]. Cyclopterus [DPA^{+}24, KKL^{+}22, LdL20, SVO^{+}22]. CyHV [ZZZ^{+}23, FZC^{+}23, KKH^{+}24, PEś^{+}20, TBMP23, TR23]. CyHV-2 [FZC^{+}23, KKH^{+}24]. CyHV-3 [ZZZ^{+}23, PEś^{+}20, TBMP23, TR23]. Cymbopogon [POS^{+}20]. Cynoglossus [ZDZZ22]. Cyprinid [CLF^{+}22, FZC^{+}23, BAS^{+}22, BDK^{+}21, GZZ^{+}22, KKH^{+}24, LYF^{+}21, MLGRF^{+}20, UCB^{+}21, YCC^{+}21, ZGz^{+}22, ZZZ^{+}23, TBMP23]. Cyprinidae [SMMM20]. Cypriniformes [SMMM20]. Cyprinus [BAS^{+}22, JJZ^{+}22, MPPF21, MPP20a, MBF^{+}23, MPA^{+}21, SMC^{+}22, SGR^{+}20a, STB^{+}21]. Cytochrome [FSW^{+}21]. cytotoxic [EBBIAH21, EBAH24, FVC^{+}21, SMD^{+}20]. Cytotoxicity [PHS^{+}20, ASA^{+}20, GWL^{+}21]. Czech [MPP^{+}20b].

dab [VDC20, VDD^{+}20]. Dactylogyrus [LZZ^{+}23]. daily [PNAS22]. Dam [KKC^{+}23]. damage [LXZ^{+}21, SMP^{+}20]. damaged [SZZ^{+}23]. damselae [PCWC20, ESEI20, WSW^{+}20]. Dana [BBD^{+}22, LHH^{+}22]. danfordi [SCV^{+}20]. danger [DBDSP20]. Danio [APK^{+}23, AHBE^{+}20, FdSB^{+}22, JW22, KSS^{+}20, KMH^{+}21, KH20, MSB^{+}21, MVW^{+}20, PKF^{+}21, SSM^{+}20, SKAS^{+}22, SLK^{+}23, WLTZ21]. Danish [DvGJK^{+}21]. DAS-ELISA [LWH^{+}21]. data [DBPH21, Li23, SSM^{+}20]. davidianus [MJX^{+}23]. Dawestrema [dCJBG22, dAMHdOCF^{+}23]. dctP [ZTY^{+}22]. death [DNN^{+}22]. decapod [GXL^{+}24, JBP24]. Decapoda [XZS^{+}23]. decipiens [NAPP24]. decreased [YGR^{+}24, YY24]. deep [OSM22a]. defence [VBG^{+}22]. defensin [HHW^{+}24, LLC^{+}21]. defensin-like [HHW^{+}24]. definitions [JSW^{+}21]. deformities [MRM21]. deformity [PSKW22, TBK^{+}20]. degenerative [MPP20a]. delivered [NRKK21]. delivery [MMTD^{+}21, VRS^{+}24]. delousing [WJPS21]. demonstrates [WYR^{+}23]. density [ORV^{+}24, PSED^{+}20, SMP^{+}20, VSODB21]. dentex [MAK^{+}23]. deoxynivalenol [LL21]. dependent [YGC^{+}20]. depends [MPF^{+}24]. deployable [SSMC22]. derivatives [LZZ^{+}23]. derived [ABBP24, DYPN22, JVM^{+}20, LjY^{+}22, LHTW21, LFZ^{+}22, RKL^{+}20]. dermatitis [EMH^{+}22]. dermocystid [SCV^{+}20]. Dermocystidiaceae [SGR^{+}20a]. Dermocystidium [SMC^{+}22, SGR^{+}20a]. Description [DOW^{+}22, CVFRM^{+}23, DWK^{+}22, SLY^{+}24, ZVB^{+}21]. descriptions [ACGDO^{+}21]. Descriptive [RGH^{+}22, ØSH^{+}22]. Desmozoon [HPR^{+}22]. destructive [CSDS21, LdL20]. detachment [CVDB^{+}21]. detect [HPR^{+}22, SKR^{+}20, ZZF^{+}23]. detected [AMJ^{+}21, DDHT^{+}22, IKT^{+}23, LJK^{+}24]. detecting [lLWC^{+}24]. Detection [CSDS21, LWT^{+}22, MMW^{+}22, NLH^{+}20, CQG^{+}21, CCG24, CDJ^{+}23, CVFRM^{+}23, DMB^{+}21, FCH^{+}22a, GYN^{+}24, GWF^{+}22, GXL^{+}24, HSS^{+}24, HHZ^{+}23, HLW^{+}23, JAC^{+}22, JSZ^{+}23, KDS^{+}21, KPC^{+}22, KDR^{+}22, LDR^{+}22, LWH^{+}21, LCW^{+}24, MJX^{+}23, NAaS^{+}22, OKL^{+}22, PEN^{+}21, PYT^{+}22, PDZ^{+}22, PS23, PSV^{+}21, QWL^{+}21, SNY^{+}21, SKAS^{+}22, SLK^{+}23, SSM^{+}21, SPS23, SMY^{+}22, SM21, SCRS^{+}23, SSMC22, TZL^{+}22, THW^{+}22, TCS21, WCT^{+}23b, WCT^{+}20, WRMT24, YTS22, YLW^{+}21, ZGC^{+}22, ZZZ^{+}23, ZHY^{+}21, ZLS^{+}21]. determinants [SCIAH21]. Determination [TIK^{+}22, CTOH20]. determine [GIS^{+}21]. determines [KKBL21]. Development [ASG^{+}22, AMS^{+}24, CQG^{+}21, CCG24, FSS^{+}21, GSJ^{+}20, HZFT21, JSZ^{+}23, LWH^{+}21, LGJ21, MWY^{+}22, MJX^{+}23, NK22, PDM^{+}21, PSV^{+}21, QWL^{+}21, SNY^{+}21, SCALIAH23, SMY^{+}22, SMP^{+}22b, TBK^{+}22, TASP^{+}20, TCS21, WCT^{+}20, YTS22, ZFC^{+}23, ZGz^{+}22, ZZF^{+}23, ZZZ^{+}23, APK^{+}23, DPA^{+}24, FFB^{+}21, LFA20, LH20, MWZ^{+}20, PNØ^{+}21, PKF^{+}21, TCC^{+}24, VDD^{+}20, YLW^{+}21, lZWC^{+}20]. deviations [FK21]. Devriesia [PAS^{+}21]. diagnosed [KRS^{+}20]. diagnoses [PTBSF22]. Diagnosis [JBG^{+}23, SLY^{+}24, DSD^{+}22, FSFS^{+}20, HZFT21, JAC^{+}22, MHA^{+}20, PTK^{+}21, PTP^{+}23, WML^{+}21, YDM^{+}21]. Diagnostic [SHS^{+}22a, FMH^{+}22, JBG^{+}23, LGP^{+}20, LJK^{+}24, PDM^{+}21, SSM^{+}20, SSM^{+}21]. diagnostics [CMM^{+}20, ZVB^{+}21]. dicentrarchi [NLR21, AHSCEBI23, EBAH24, IAH21, IMAH21, IAH22, SCIAH21]. Dicentrarchus [ASNA24, BBS^{+}20, xCqXfL^{+}21, LDM^{+}20, MBN^{+}22, YK21, ZVB^{+}21]. Dicistro [HSG^{+}21]. Dicistro-like [HSG^{+}21]. did [EBA^{+}21]. Dietary [LKP^{+}22, YY24, DPA^{+}24, FCH^{+}22b, RGS^{+}21]. diets [APK^{+}23, YGR^{+}24]. Differences [JW22, SM21, SPM^{+}20, YZM^{+}21]. different [AFKK22, AHSCEBI23, BAM^{+}23, FSFS^{+}20, FEH^{+}22, HBRN24, JLH^{+}21, MWH^{+}21, MHP^{+}20, NTB^{+}24, PCWC20, SLPF^{+}20, UD22, WJPS21, YCS^{+}21]. Differential [FFB^{+}21, RCC^{+}21, SMG^{+}23, NK22]. differentiation [PDM^{+}21]. differently [WWW^{+}21]. Diffuse [BRR^{+}23]. diffusion [BLJ^{+}21]. Digenea [KNG^{+}22]. digest [LHTW21]. Digestion [DMB^{+}21]. digestive [KSI^{+}20, KH20, SFKB^{+}20]. digital [HHZ^{+}23, JSZ^{+}23, MMW^{+}22, MJX^{+}23, SPS23, ZGC^{+}22, ZZF^{+}23]. Dimethomorph [NTB^{+}24]. Diphyllobothriidea [KBMD21]. diploproa [PAS^{+}21]. dipstick [CCG24, PSV^{+}21]. dipsticks [FCH^{+}22a]. directions [DBPH21]. disc [BLJ^{+}21]. Discocotyle [LFSB^{+}21]. discoidalis [POS^{+}20]. discovered [HSG^{+}21, MHYW22b]. Discovery [Li23]. Disease [TBMP23, öTTS20, ABA^{+}23, BIK^{+}21, BHS^{+}21, BTP^{+}24, CMM^{+}20, CSDS21, CGK23, CTP^{+}23, DWK^{+}22, EBA^{+}21, EMH^{+}22, FSFS^{+}20, GKTC^{+}21, GKAS23, HSG^{+}21, JAC^{+}22, KSD^{+}20, KKBL21, KRS^{+}20, KB21, LMJ^{+}24, LDR^{+}22, LNK^{+}21, LWR^{+}20, LdL20, MPP^{+}20b, NAaS^{+}22, NKö^{+}21, PSV^{+}21, RSPB22, RTA21, SAAB^{+}21, SZH^{+}22, SSMC22, SPM^{+}20, TSS^{+}23, ÚPZ^{+}23, WSW^{+}20, WML^{+}21, WCT^{+}20, WSS^{+}20, YGR^{+}24, YDM^{+}21, YCL^{+}21, YJL^{+}23a, ZGC^{+}22, ZDZZ22, ZVB^{+}21, NRKK21, WAF^{+}21]. diseased [BCT^{+}20b, CXC^{+}24, LS24, MLL^{+}22, MAK^{+}23, MSS^{+}20, MIC^{+}24, MBR^{+}22, RKL^{+}20, SLY^{+}24, WZC^{+}24]. Diseases [RM20, CLC^{+}22, KSS^{+}20, PTBSF22, Tan23, YTH^{+}24, SNSP^{+}23, NSP20]. disinfectant [HZC^{+}24]. disinfection [TPSJS20a, TPSJS20b]. dismutase [HWC^{+}21]. disorders [Din21]. dispersal [LDM^{+}20, PSH^{+}21]. displaying [LMJ^{+}24, WSW^{+}20]. Disruption [ZSL^{+}24b]. Dissecting [DHSK^{+}21]. Disseminated [GKAS23]. distantly [TNB^{+}22]. distinct [NGC^{+}23, RCG22]. distinctive [BRWC23]. Distribution [RBM^{+}22, KRS^{+}20, LKV^{+}23, PVCG20, WWK^{+}20]. distributions [WJPS21]. districts [DDDJ^{+}20]. Disturbance [BBD^{+}22]. DIV1 [GXL^{+}24]. diverse [NLPM^{+}21]. Diversity [MIK^{+}24, MBR^{+}22, AHTCDI20, BPA^{+}21, KaVDH^{+}21, PYD^{+}22, PCWC20]. DNA [CDJ^{+}23, HPR^{+}22, LDW^{+}23, NLH^{+}20, YGZ^{+}22, ZZL^{+}21]. DNA-based [HPR^{+}22]. DNAzyme [THW^{+}22]. does [LRGAH21, dCJBG22]. Dollfustrema [KNG^{+}22]. dolomieu [WBM21]. Dolops [POS^{+}20]. domain [TASP^{+}20]. domains [WWW^{+}21]. domestic [TIK^{+}22]. dominant [DZJ^{+}21]. Don [Now21]. dorsal [BZQV21]. dosages [RKC^{+}23]. dose [WSS^{+}20]. double [LWH^{+}21]. double-antibody [LWH^{+}21]. Douëllou [IAO20]. downregulates [ENS^{+}23]. dracunculid [DYN^{+}20]. Draft [KSD^{+}20]. driven [KDHS^{+}23, LNK^{+}21]. drives [DXL^{+}20b]. drop [CSDS21, KSD^{+}20, PSV^{+}21, SSMC22]. droplet [HHZ^{+}23, JSZ^{+}23, MJX^{+}23, SPS23, ZGC^{+}22, ZZF^{+}23]. drug [DLS^{+}21, MMTD^{+}21]. drugs [LXW^{+}20]. drum [dRRRPSR^{+}21]. Dual [AWF22, ZLS^{+}21]. dual-sample [ZLS^{+}21]. due [GMK^{+}21]. dumerili [PZP^{+}20]. Duplex [ZGC^{+}22, SMP^{+}22b, WML^{+}21]. duration [BBG21]. during [CVDB^{+}21, xCqXfL^{+}21, GOG^{+}21, GMK^{+}21, IVR^{+}23, JQT20, MBJ20, TCC^{+}24, ZZS^{+}24]. dye [LWT^{+}22, ZHY^{+}21]. Dyková [SCRS^{+}23]. dynamics [BWH^{+}20, LNK^{+}21, LCYL20, LYC^{+}20, SRT^{+}21, SJK^{+}21]. dysfunction [YGC^{+}20]. dysfunctional [PZP^{+}20]. dysgalactiae [MWY^{+}20, ZHpT^{+}20].

E-cadherin [REB^{+}20]. eagle [DYN^{+}20]. Early [CVDO^{+}20, SLK^{+}23, FMH^{+}22, PDZ^{+}22]. earthen [HBRN24, NKZ^{+}20, QCZ^{+}22, XZS^{+}23]. east [JBG^{+}23, ZGL^{+}20]. eastern [BTN^{+}24, KRS^{+}20, LLR^{+}20, KB21, SPKS^{+}20]. eBook [FM23]. ecology [TBMP23]. economic [NTB^{+}24, Tan23]. ectoparasites [CSDS21]. ectoparasitic [UD22, UMD22]. edema [MPPF21, MPA^{+}21, PPDG^{+}22, ZOZ^{+}20]. Edited [FM23, RM20]. Edition [Man23, Muf23]. eDNA [MMW^{+}22]. edulis [CTOH20]. Edwardsiella [AGR^{+}22, BRWC23, CLS^{+}22, HWM^{+}24, LTW^{+}21, LPGA^{+}21, LPGW^{+}22, MPS^{+}22, NLPM^{+}21, NHG^{+}23, OGR^{+}24, PNØ^{+}21, PØN^{+}21a, PØN^{+}21b, PRC20, RCB^{+}22, Roh22, SOE^{+}21, SGW^{+}22, SLH^{+}24, THH^{+}21, WCT^{+}23a, YGR^{+}24, YMC^{+}21]. edwardsiellosis [SOE^{+}21, SGW^{+}22]. Edwarsiella [SLA^{+}20]. Eel [DHF^{+}22, BFvG^{+}23, BKM21, IAH21, LFZ^{+}22, MNPS23, PEN^{+}21, SCIAH21, WZC^{+}24, dYdRFF^{+}23]. eels [DAW^{+}22, YCL^{+}21, ZWZ^{+}20]. Effect [DPLL^{+}21, JJZ^{+}22, LWK20, ORV^{+}24, RTA21, SMY^{+}24, TPSJS20a, VSODB21, WBB^{+}21, WHK^{+}23, ZMZ^{+}23, CLH^{+}21, DHSM^{+}20, EBAH24, HHW^{+}24, LCYL20, NTB^{+}24, PLS^{+}22, QST^{+}21, SMP^{+}22a, UD22, XJZ^{+}22, YYS^{+}22, ZYB^{+}23]. effective [HZC^{+}24, MMTD^{+}21, NRC^{+}23, dCJBG22]. effectiveness [DMH^{+}21]. Effects [BIK^{+}21, CSGB^{+}21, MWH^{+}21, MVW^{+}20, MNPS23, REB^{+}20, WCT^{+}23a, ABA^{+}23, AAZR^{+}20, CLC^{+}22, DvGJK^{+}21, EBBIAH21, FFB^{+}21, FCH^{+}22b, IAO20, LDW^{+}23, MDAO22, RCC^{+}21, SVP^{+}24, SJK^{+}21, SKS^{+}24, YZM^{+}21, YJL^{+}23b, YYS^{+}24]. Efficacy [MKT^{+}21, RBP^{+}20a, TSV^{+}22, dAMHdOCF^{+}23, ASA^{+}20, DXD^{+}21, JBB^{+}24, JXJ^{+}22, LJHH23, LZZ^{+}23, LPGW^{+}22, MSF^{+}20, RGS^{+}21, SMY^{+}24, WSS^{+}20, ZLWL23]. efficiently [LYZ^{+}22]. efforts [LKD^{+}24]. egg [SLE^{+}20]. egg-associated [SLE^{+}20]. Eggs [öTTS20, ELB21, FdSB^{+}22, ÖTG22]. EIBS [TMM^{+}24]. eight [KSG^{+}23]. Eimeria [SKS^{+}24]. elasmobranch [AGH^{+}23]. electron [ADAO22, AONDO24, MDAO22, RCS^{+}22, SMMM20]. electrophoresis [RCS^{+}22]. Eleginops [QGG^{+}21, SDMO^{+}20]. element [YLC^{+}21]. elevated [FCH^{+}22b]. elicited [MPS^{+}22]. ELISA [LWH^{+}21]. Elizabethkingia [TSC^{+}23]. Elucidation [MLGRF^{+}20]. Elumalai [FM23]. elute [PTK^{+}21]. Emergence [LFSB^{+}21, SSP^{+}23]. emerging [AANY23, CMLK^{+}22, DWK^{+}22, IMAH21, WAF^{+}21]. enables [BRWC23]. encapsulated [XJZ^{+}22]. Encapsulation [BKM21]. encephalopathy [BPA^{+}20, ZMZ^{+}22]. encoded [CLF^{+}22, MLGRF^{+}20]. encoding [DZJ^{+}21, FVC^{+}21]. endemic [BPA^{+}20, TBMP23]. endoparasitic [MMTD^{+}21]. endopeptidase [WHC^{+}21]. endothelial [SBD21]. energy [FCH^{+}22b]. enhance [DZJ^{+}21, GWL^{+}21, JXJ^{+}22, PDL^{+}23, SLA^{+}20]. enhancement [YJL^{+}23b]. enhances [CZM^{+}22, LKP^{+}22, XJZ^{+}22]. enhancing [LDW^{+}23, WGE^{+}24, ZLWL23]. enrichment [PTP^{+}23]. Enteritis [PTW21, EMH^{+}22]. Enterobacter [LS24]. Enterocytozoon [LCCC22, SAB^{+}22, WML^{+}21, ZGC^{+}22]. Enteromyxum [LJK^{+}24, PSED^{+}20, REB^{+}20, SJK^{+}21]. Enterospora [lLWC^{+}24]. entry [LDM^{+}20]. environment [CZC^{+}22, HHZ^{+}23]. Environmental [AEE^{+}23, THH^{+}21, Tan23, KSY^{+}23, LDR^{+}22, NLH^{+}20, SPS23, WRMT24, YPT^{+}22]. environments [Roh22]. EnvZ [DZQ^{+}22]. EnvZ/OmpR [DZQ^{+}22]. enzymatic [NNC^{+}23]. enzyme [LWH^{+}21]. enzyme-linked [LWH^{+}21]. enzymes [FVC^{+}21]. Eoacanthocephala [SMP^{+}22a]. epidemic [MSZV^{+}23]. epidemiological [IHK^{+}22, NNC^{+}23]. Epidemiology [OHL^{+}20, SPHM23, RGH^{+}22, ÚPZ^{+}23]. epidermal [ASR^{+}23, HBRN24]. epidermis [SMMM20]. Epigallocatechin [ZWX^{+}22]. Epigallocatechin-3-gallate [ZWX^{+}22]. Epilithonimonas [MIC^{+}24]. epinepheli [lLWC^{+}24]. Epinephelus [BPA^{+}20, CDY^{+}20, FYH^{+}21, HWZ^{+}21, LCYL20, NGC^{+}23, ZWB^{+}23]. Episode [MCP^{+}22]. Epistylis [KB21]. epithelial [LPW^{+}20, SMG^{+}23]. Epitheliocystis [DDH^{+}22, QRTD^{+}22]. epitheliotropic [HCC^{+}22]. epitope [DZJ^{+}21, JXJ^{+}22, MPS^{+}22]. epitopes [GYC^{+}22]. Epizootic [SBVC21, BTN^{+}24, SPHM23, WBM21]. erectus [BME^{+}20, EMH^{+}22]. Ergasilidae [AONDO24]. Eriocheir [MLL^{+}22, Din21]. erm [AANY23]. Erysipelothrix [CCP^{+}21]. erythrocyte [ŘB24]. erythrocytic [TMM^{+}24]. erythropterus [CRWC22]. esculin [SZH^{+}22]. Essential [ZYL^{+}22, ÖTG22, POS^{+}20, SMP^{+}22a, TR23, ZHL^{+}22]. established [JVM^{+}20]. Establishing [PRC20]. Establishment [GPL^{+}22, GWF^{+}22, GXL^{+}24, JZW^{+}23, LHC^{+}22, LjY^{+}22, LZC^{+}24, LZK^{+}23, lLWC^{+}24, YDM^{+}21]. Estimates [PNØ^{+}21]. Estimating [PSH^{+}21, WJPS21]. estimation [LGP^{+}20, YK21]. Eubothrium [HSS^{+}24, IFI^{+}21]. Eudontomyzon [SCV^{+}20]. European [BKM21, BBS^{+}20, DAW^{+}22, DHF^{+}22, LDM^{+}20, MNPS23, PSPP^{+}22, SVO^{+}22, SMP^{+}20, ZVB^{+}21, dYdRFF^{+}23]. EUS [BTN^{+}24, SPHM23]. evaluating [JRZ^{+}22]. Evaluation [FEH^{+}22, HBRN24, IAH22, MCM^{+}23, SGW^{+}22, SCX^{+}21, VBL^{+}23, CLS^{+}22, GWF^{+}22, HZFT21, JKM^{+}20, RGT^{+}23, SLPF^{+}20, SFKB^{+}20, FSS^{+}21]. EVE [dYdRFF^{+}23]. events [KKH^{+}24]. EVEX [dYdRFF^{+}23]. Evidence [EBBIAH21, KBG^{+}23, AHSCEBI23, HZC^{+}24, LRGAH21, MLM^{+}20, WZG^{+}20, dRRRPSR^{+}21]. evolution [LGH^{+}21]. Examples [RCC^{+}21]. excelsa [BIK^{+}21]. exclusion [EJH^{+}23]. exclusive [RCG22]. exhibits [YYS^{+}24]. existence [EBBIAH21, FEMMI^{+}23]. existing [DBPH21]. exitiosa [BWH^{+}20, BHT^{+}20]. Exogenous [HLH^{+}23b, LGH^{+}21]. Exophiala [ACGDO^{+}21]. exophthalmia [HB20]. exoskeletal [DYPN22]. exoskeletal-derived [DYPN22]. Expansion [AHCSCI20]. Experimental [BHS^{+}21, CTOH20, HZC^{+}24, IAH21, KBMD21, SVO^{+}22, SFKB^{+}20, BMO^{+}20, DvGJK^{+}21, LWK20, LPL22, PSED^{+}20, PPWC20, QGG^{+}21]. experimentally [DHSK^{+}21, ESEI20, KKL^{+}22]. explains [GKAS23]. exposed [KBG^{+}23]. Exposure [CZC^{+}22, ASF^{+}21, BHS^{+}21, MBS21, NGC^{+}23, PSED^{+}20, SST^{+}22]. Expression [FYH^{+}21, YCC^{+}21, ASG^{+}22, AMS^{+}24, AFKK22, BMSC21, FLH^{+}21, FZC^{+}23, JPC^{+}21, JJZ^{+}22, KBR^{+}22, OBSH24, WGE^{+}24, XKF^{+}22, YMC^{+}21, ZWZ^{+}20]. expressional [YSW^{+}22]. extracellular [EBBIAH21, TR23]. extract [BIK^{+}21, NSTS24]. extracted [WHK^{+}23]. extracts [LYX^{+}20]. extragonadal [PK24]. Extreme [SMC^{+}22]. Eye [DvGJK^{+}21]. eyed [ELB21].

f [LYX^{+}20]. facilitating [LDO^{+}21]. Facilities [NSP20, KSS^{+}20]. factor [BCT^{+}20b, GPC22, SCAH22]. Factors [JLM^{+}20, AEE^{+}23, BRWC23, CCY^{+}23, FVC^{+}21, LDR^{+}22, THH^{+}21, WMS21, lZWC^{+}20]. failure [NCA^{+}20]. family [LGH^{+}21]. Fan [JBP24]. fancy [SLY^{+}24]. far [SRN20]. farm [IVR^{+}23, PTMC22, YDM^{+}21]. farmed [AMJ^{+}21, BRR^{+}23, BFH^{+}23, CGK23, DNN^{+}22, ELS^{+}23, FEMMI^{+}23, FEH^{+}22, GVB^{+}21, GLS^{+}21, HSS^{+}24, IMAH21, JSW^{+}21, KAS^{+}24, KDR^{+}22, KKK^{+}23, LGP^{+}20, LCTC20, LLR^{+}20, MBJ20, MIC^{+}24, MBS21, OSM^{+}22b, ØSH^{+}22, PEN^{+}21, PYD^{+}22, PTK^{+}21, PAH^{+}22, PVCG20, PRHW21, PHB^{+}23, RBM^{+}22, RBP^{+}20a, RØLW23, RTA21, SNT^{+}22, SRP^{+}21, TBK^{+}20, VIB^{+}21, ZVB^{+}21, dRRRPSR^{+}21]. farming [PNAS22, SSP^{+}23]. farms [AHTCDI20, BTP^{+}24, BWH^{+}20, CTP^{+}23, DOW^{+}22, DDDJ^{+}20, DMH^{+}21, GIS^{+}21, MSZV^{+}23, PSPP^{+}22, PBF^{+}21, QCZ^{+}22, RGH^{+}22]. fasciatus [ENS^{+}23]. Fast [BFvG^{+}23, THW^{+}22]. fathead [HCW^{+}20, HWC^{+}21, LL21]. fatty [HLH^{+}23b]. features [BRWC23, VFR^{+}23, XJM^{+}21]. feed [LL21, MCM^{+}23, RKC^{+}23, SVP^{+}24]. Feeding [BBG21, MCA^{+}20, SMP^{+}20]. feeds [BAM^{+}23, HBRN24]. female [HHW^{+}24, PZP^{+}20, SLA^{+}20]. females [FdSB^{+}22]. Fenneropenaeus [ZHL^{+}23]. ferric [WZC^{+}24]. fertilization [FdSB^{+}22]. FHM [HWC^{+}21]. Fiala [SCRS^{+}23]. Field [SPM^{+}20, AMB^{+}22, CQG^{+}21, LFZ^{+}22, RCS^{+}22, RGT^{+}23, SSMC22, DvGJK^{+}21]. field-deployable [SSMC22]. fighting [SBB21]. film [QCZ^{+}22]. film-lined [QCZ^{+}22]. filter [MCA^{+}20]. filter-feeding [MCA^{+}20]. filtrate [MHA^{+}20]. Fin [SMP^{+}20, BZQV21, FEM^{+}20, LYF^{+}21, YCC^{+}21]. findings [CVFRM^{+}23, JBG^{+}23]. Finland [ELS^{+}23]. finnmarkense [SNT^{+}22, NLR21, SNT^{+}22]. First [BAS^{+}22, BLL23, CRWC22, CVFRM^{+}23, CCL^{+}24, DGB^{+}22, GYC^{+}22, HKB^{+}22, LHLS20, NAPP24, NRY^{+}23, PEN^{+}21, RBP^{+}20b, SSK23, SAB^{+}22, VIB^{+}21, VDC20, XZS^{+}23, YWZ^{+}21, ZVB^{+}21, ZAG^{+}20, dRRRPSR^{+}21, ACGDO^{+}21, KPC^{+}22, KRS^{+}20, SCRS^{+}23, ZDF^{+}23, GVB^{+}21]. Fish [NSP20, SNSP^{+}23, ASA^{+}20, AANY23, ASR^{+}23, ACGDO^{+}21, AHSCEBI23, BFH^{+}23, BCT^{+}20b, CSE21, DBDSP20, DCB20, DvGJK^{+}21, DMB^{+}21, EBELIAH20, ELS^{+}23, FSS^{+}21, GPC22, GPL^{+}22, HWG^{+}21, IFI^{+}21, IFIZ22, JZW^{+}23, KKC^{+}23, KBG^{+}23, KaVDH^{+}21, KBMD21, LAH21, LjY^{+}22, Li23, LZC^{+}24, LZK^{+}23, LH20, LHTW21, MMTD^{+}21, MWY^{+}20, MPMB^{+}20, MRM21, MQK^{+}20, MIK^{+}24, MCM^{+}23, MHYW22b, MCA^{+}20, NLPM^{+}21, PBT^{+}23, PAH^{+}22, PCWC20, PSED^{+}20, PVCG20, QGG^{+}21, RCS^{+}22, ŘB24, RKL^{+}20, SCALIAH23, SBS^{+}22, SKO21, SSM^{+}21, SBD21, SPS23, SBB21, SPHM23, SBVC21, Tan23, VBG^{+}22, YPT^{+}22, YJL^{+}23b, ZGL^{+}20, ZCY^{+}23, ZDZZ22, dCJBG22, dRRRPSR^{+}21, FM23, Muf23, RM20]. fish-to-fish [PSED^{+}20]. fisheries [CZMP21]. fishes [CZMP21, DNN^{+}22, DŽR^{+}21, MBR^{+}22, SVC^{+}21]. five [ASG^{+}22]. fjord [LEV^{+}20]. FL [FM23]. flagellar [CZJ^{+}23]. flagellin [CLS^{+}22]. flavescens [ZYH^{+}23]. Flavobacterium [AHTCDI20, AHBE^{+}20, AWF22, BMK^{+}20, BMO^{+}20, BMJ^{+}21, BMSC21, CVFRM^{+}23, ELB21, FFB^{+}21, HKB^{+}22, HSM^{+}22, IVR^{+}23, KKBL21, LFA20, LPL22, MIR^{+}22, NHG^{+}23, VFR^{+}23]. flgJ [LDW^{+}23]. Florfenicol [IMAH21, OGR^{+}24, OCS^{+}23, ORV^{+}24, RCC^{+}21, RKC^{+}23]. flounder [CLC^{+}22, LFZ^{+}21, SJK^{+}21]. flow [CCG24, FCH^{+}22a, LWT^{+}22, PSV^{+}21, QRTD^{+}22, ZZZ^{+}23]. flow-through [QRTD^{+}22]. fluctuations [AEE^{+}23]. fluke [DvGJK^{+}21, RGT^{+}23]. fluorescence [IKT^{+}23, TBB^{+}21]. fluorogenic [ZLS^{+}21]. fluviatilis [KSI^{+}20, SMP^{+}20]. focal [BFH^{+}23]. focusing [SCIAH21]. follicles [CZMP21]. following [CDY^{+}20, LWK20, LWL^{+}24, OGR^{+}24, WJPS21]. fontinalis [BMJ^{+}21, SEL^{+}20]. food [BTN^{+}21]. formalin [LJHH23, MKT^{+}21, RBP^{+}20a, ZLY^{+}22]. formalin-inactivated [LJHH23, RBP^{+}20a]. formalin-killed [MKT^{+}21, ZLY^{+}22]. formation [DCB20, HLH^{+}23b, OCS^{+}23, PLS^{+}22, SVP^{+}24]. formed [ORV^{+}24]. forms [VRC^{+}22]. Forsskål [CDY^{+}20]. fortis [EMH^{+}22]. Fouquet [PYD^{+}22]. four [NGC^{+}23, RGH^{+}22, SLK^{+}23]. France [BPA^{+}21]. Francis [RM20]. Francisella [CQG^{+}21, PPWC20]. free [FdSB^{+}22, LWT^{+}22]. frequencies [ŘB24]. freshwater [APS23, ABA^{+}23, ASR^{+}23, DBPH21, DvGJK^{+}21, GOG^{+}21, HRR^{+}24, JKM^{+}20, KaVDH^{+}21, MWH^{+}21, MIK^{+}24, MBR^{+}22, PBF^{+}21, QRTD^{+}22, SVC^{+}21, SAB^{+}22, TAK^{+}21, ZWX^{+}22]. fry [MIR^{+}22]. FTA [PTK^{+}21]. fugu [LJK^{+}24]. function [GYC^{+}22, TASP^{+}20, VBG^{+}22, WHC^{+}21, XKF^{+}22, YLC^{+}21, ZZS^{+}24].

Functional [KQW^{+}21, JXJ^{+}22, LLC^{+}21, NLL^{+}21, QFW^{+}23, YSW^{+}22]. functions [FMK^{+}21, WWW^{+}21]. Fundy [PSH^{+}21, VPW^{+}21]. Fungal [öTTS20]. fungus [MIK^{+}24]. fungus-like [MIK^{+}24]. furcatus [AGR^{+}22, SLA^{+}20, YGR^{+}24]. furiosus [WCT^{+}23b]. Further [PSPP^{+}22, ABBP24]. furunculosis [HRR^{+}24, LH20]. Fusarium [BME^{+}20]. fuscoguttatus [CDY^{+}20]. Fustvatnet [MHH23]. future [DBPH21].

G. [MHH23]. Gadus [AFKK22]. gain [AKR^{+}21]. galectin [NLL^{+}21, WWW^{+}21]. galectin-9 [WWW^{+}21]. galectin-related [NLL^{+}21]. gallate [ZWX^{+}22]. galU [SJL^{+}23]. Gambusia [BTN^{+}24, JVM^{+}20, WCT^{+}23a]. gangliocytoma [SAKC20]. Ganoderma [LDW^{+}23]. GapA [GYC^{+}22]. gariepinus [ADAO22, MDAO22]. Garlic [öTTS20]. garvieae [AANY23, AOT22, BLL23, BHEM21, KBG^{+}23, NRY^{+}23, RBP^{+}20a, RCS^{+}22]. gastroenteropathy [BZP^{+}22]. gastrointestinal [AABB23, KSI^{+}20]. gel [RCS^{+}22]. Gene [AFKK22, KBR^{+}22, ASG^{+}22, AMS^{+}24, BMSC21, CGA^{+}20, CRP^{+}21, CZJ^{+}23, FSW^{+}21, FMK^{+}21, FLH^{+}21, GZW^{+}23, HWZ^{+}21, LFZ^{+}21, OBSH24, SJL^{+}23, SDMO^{+}20, TAK^{+}21, TCS21, WGE^{+}24, YZL^{+}23, YMC^{+}21, ZJF^{+}22, ZTY^{+}22, ZSL^{+}24b]. generated [SMC^{+}22]. Generation [GZZ^{+}22, YLW^{+}21, HAK^{+}20]. genes [CLS^{+}22, DXL^{+}20a, ENS^{+}23, FVC^{+}21, JJZ^{+}22, LFA20, RRK^{+}23, RKL^{+}20, ZHL^{+}22]. Genetic [AMM^{+}22, BPA^{+}21, GLS^{+}21, KSG^{+}23, KaVDH^{+}21, LPGA^{+}21, NLPM^{+}21, PYD^{+}22, PØN^{+}21a, PØN^{+}21b, WMS21, ELS^{+}23, FEM^{+}20, MLL^{+}22, MR23, PNØ^{+}21, Roh22, RBP^{+}20b]. Genetically [ABM^{+}21, TNB^{+}22]. Genital [IKT^{+}23]. genogroup [ABBP24]. genogroups [TBKY20]. Genome [CLH^{+}21, RKL^{+}20, CCY^{+}23, HAK^{+}20, IHK^{+}22, KSD^{+}20, MAK^{+}23, Roh22, SCIAH21, SKR^{+}20]. Genome-wide [CLH^{+}21, RKL^{+}20, Roh22]. Genomic [RRK^{+}23, VFR^{+}23, AANY23, CCCC23, LS24, LHLS20, LFG^{+}23, MIC^{+}24, SCAH22, YCS^{+}21]. Genomics [KDHS^{+}23, BRWC23]. Genomics-driven [KDHS^{+}23]. genomovar [SNT^{+}22]. genotoxic [ŘB24]. Genotype [CRP^{+}21, KPC^{+}22, LCCC22, NNC^{+}23, PPT^{+}20, SWC21]. genotypes [KKBL21]. Genotypic [PCWC20, TSC^{+}23, NKZ^{+}20, RCB^{+}22]. Genotyping [RCS^{+}22, ASNA24, DDTG^{+}21, SNY^{+}21, SSP^{+}23]. genus [BPA^{+}20, IFIZ22, WHK^{+}23]. Genypterus [IAH21, SCIAH21]. geographic [KRS^{+}20]. geographical [RCS^{+}22, SLPF^{+}20]. germ [BBS^{+}20]. GGIV [GSJ^{+}20]. Ghats [VKK^{+}22]. ghost [ZZL^{+}21, ZLY^{+}22]. ghost-based [ZZL^{+}21]. Giant [PHB^{+}23, ABA^{+}23, GSJ^{+}20, PEN^{+}21, SRP^{+}21]. gibel [SCRS^{+}23]. gibelio [KKH^{+}24, LYF^{+}21, SCRS^{+}23, WZG^{+}20, YCC^{+}21]. gigas [dCJBG22, dAMHdOCF^{+}23]. Gill [JKM^{+}20, BHS^{+}21, BTP^{+}24, CMM^{+}20, CVDB^{+}21, CTP^{+}23, EBA^{+}21, FSFS^{+}20, LZC^{+}24, LHTW21, LPW^{+}20, ØSH^{+}22, PPDG^{+}22, RGT^{+}23, WSW^{+}20, WSS^{+}20, XJM^{+}21, YCL^{+}21]. gill-related [ØSH^{+}22]. gills [xCqXfL^{+}21, IAO20, PBF^{+}21, VWE^{+}23]. gilthead [CSGB^{+}21, CSVG^{+}22, CFGCCE21, KPC^{+}22, LDC^{+}22, RGT^{+}23, REB^{+}20, AEE^{+}23, FEM^{+}20, FK21]. ginbuna [MHA^{+}20]. ginsengisoli [MIC^{+}24]. glochidiosis [CVDO^{+}20]. glucan [RGS^{+}21]. glutamyl [WHC^{+}21]. Gobiidae [MRM21]. gold [GYN^{+}24, HZFT21, KDS^{+}21, SCS24]. gold-nanoparticle-based [GYN^{+}24]. golden [HZFT21, YTH^{+}24]. goldfish [MLGRF^{+}20, SMY^{+}24, ZLWL23]. Gonadal [SCS24, STB^{+}21]. goramy [GSJ^{+}20, SRP^{+}21]. gourami [GSJ^{+}20, SAKC20, SRP^{+}21]. gracilis [SMP^{+}22a]. Gram [PLS^{+}22]. Gram-negative [PLS^{+}22]. granulocytes [CSVG^{+}22]. granulomatous [ZBL^{+}23]. graphene [THW^{+}22]. grass [LLJ^{+}21, YJL^{+}23a, YLW^{+}21]. Great [HKB^{+}22, SPHM23]. greater [PZP^{+}20]. Greece [KKC^{+}23]. Greek [BIK^{+}21]. green [QST^{+}21]. grey [HZFT21, RBP^{+}20a]. gross [GSS^{+}21]. Group [MWY^{+}20, RM20, CRP^{+}21, XQZ^{+}22]. grouper [CDY^{+}20, DXL^{+}20a, FYH^{+}21, LDW^{+}23, LYX^{+}20, lLWC^{+}24, LCYL20, NGC^{+}23, ZWB^{+}23]. groupers [BPA^{+}20, MHP^{+}20]. grow [ABBP24]. grown [AHSB^{+}22]. growth [FFB^{+}21, FCH^{+}22b, LKP^{+}22, LAH21, PØN^{+}21b, SVP^{+}24, SMP^{+}20]. Guichenot [IAH21]. Gulf [ZAG^{+}20]. guppies [SLY^{+}24]. guppy [Li23]. gut [IFIZ22, WQL^{+}23]. Gyrodactylus [MHH23].

habitat [LEV^{+}20]. haematocrit [YGR^{+}24]. Haematology [MBN^{+}22, CGK23]. haematopoietic [SMY^{+}24]. haemocyte [ZLZ21]. haemorrhagic [AKR^{+}21, AAZR^{+}20, ASF^{+}21, BAM^{+}23, CCV^{+}24, DXD^{+}21, HAK^{+}20, LL21, TBK^{+}22, UCB^{+}21, YCL^{+}21]. hainanensis [MSS^{+}20]. halioticida [CQG^{+}21]. hamatus [FWL^{+}23]. Hamilton [APK^{+}23, KSS^{+}20, KMH^{+}21, SSM^{+}20, SKAS^{+}22, SLK^{+}23]. handling [DPLL^{+}21]. haplotype [BME^{+}20]. Harbor [BTN^{+}24]. harbouring [AANY23, RKL^{+}20]. hard [LYC^{+}20]. Hardback [RM20]. harvest [PØN^{+}21b]. harvested [JBG^{+}23]. Harveyi [ZVYD23, CDY^{+}20, DXL^{+}20b, GZW^{+}23, LDW^{+}23, NSTS24, Roh22, SZZ^{+}23, WZC^{+}24, ZDZZ22]. harveyi-infected [ZDZZ22]. hatcheries [DDDJ^{+}20]. hatchery [MSS^{+}20]. HCHO [ZWB^{+}23]. HDPE [QCZ^{+}22]. head [EBAH24, SDH22]. headwaters [BTN^{+}24]. healing [CFGCCE21]. Health [FM23, NSP20, BTN^{+}21, JZW^{+}23, KSS^{+}20, KH20, MIK^{+}24, MNPS23]. healthy [CGK23]. heart [FCH^{+}22b, SKR^{+}20]. heat [MKT^{+}21]. heat-killed [MKT^{+}21]. heavy [KKC^{+}23]. hemagglutinin [CGA^{+}20]. Hematodinium [GXW^{+}23]. Hematologic [KAS^{+}24]. hematopoietic [UCB^{+}21]. heme [WKL^{+}22]. heme-binding [WKL^{+}22]. Hemigrapsus [GXW^{+}23]. hepatic [PDG^{+}24]. hepatitis [KAS^{+}24]. hepatopancreatic [LCCC22, NAaS^{+}22, WML^{+}21, WCT^{+}20, YDM^{+}21, ZGC^{+}22]. hepatopancreatitis [BBD^{+}22]. hepatopenaei [LCCC22, SAB^{+}22, WML^{+}21, ZGC^{+}22]. Hepatospora [Din21]. heritable [MGG22]. Hermetia [BAM^{+}23]. herpesviral [GKAS23, SMY^{+}24]. herpesvirus [AAZR^{+}20, BAS^{+}22, BDK^{+}21, CLF^{+}22, CCG24, EJH^{+}23, FZC^{+}23, GZZ^{+}22, KKH^{+}24, LYF^{+}21, MLGRF^{+}20, TBMP23, ÚPZ^{+}23, UCB^{+}21, WWK^{+}20, YCC^{+}21, YLC^{+}21, ZGz^{+}22, ZZZ^{+}23, ZOZ^{+}20, dYdRFF^{+}23, PEN^{+}21]. herring [HSH^{+}20]. Herzegovina [GVB^{+}21]. heterogeneity [SCAH22]. heterogeneous [IVR^{+}23]. heterologous [LPGW^{+}22, NLPM^{+}21]. hif [JJZ^{+}22]. High [MMTD^{+}21, SdSNK^{+}22, GLS^{+}21, LXW^{+}20, MMW^{+}22, ORV^{+}24, RBM^{+}22, SZZ^{+}23, SPKS^{+}20, WLK^{+}22]. high-density [ORV^{+}24]. high-throughput [LXW^{+}20]. highlands [VLSM^{+}20]. highlights [RGS^{+}21]. highly [CGA^{+}20, LJHH23, SGU^{+}21, SCRS^{+}23]. hillstream [VKK^{+}22]. hint [ZDZZ22]. Hippocampus [BME^{+}20, EMH^{+}22]. histochemical [SGR^{+}20b]. histochemistry [KH20]. histocompatibility [DXL^{+}20a, FLH^{+}21, QFW^{+}23]. Histological [STB^{+}21, FEH^{+}22, SLPF^{+}20, ZWB^{+}23]. Histology [Muf23, SKAS^{+}22]. Histopathological [DWK^{+}22, BLL23, DSD^{+}22, ESEI20, SST^{+}22, SHS^{+}22a, SMMM20, TMM^{+}24, GSS^{+}21]. Histopathology [DŽR^{+}21, BRCI23, MHP^{+}20, MBJ20]. hly [TSV^{+}22]. Hofer [KRS^{+}20]. hogansi [WBV^{+}20]. holbrooki [BTN^{+}24]. home [SCS24]. homogeneity [GLS^{+}21]. homolog [LLC^{+}21, WHC^{+}21, ZSL^{+}24b]. homoserine [LRGAH21]. honghuensis [WZG^{+}20]. honokiol [JRZ^{+}22, YYS^{+}24]. Hook [LYX^{+}20]. Horizontal [QGG^{+}21]. Host [KKBL21, APS23, BDK^{+}21, IFIZ22, LKV^{+}23, SSP^{+}23, TBMP23]. Host-specific [KKBL21, SSP^{+}23]. hosts [AHCSCI20, BONB20, LNK^{+}21, NLPM^{+}21]. housed [ACGDO^{+}21]. HPR [CGA^{+}20, DOW^{+}22]. HPR0 [DGB^{+}22]. HPR0-like [DGB^{+}22]. human [RSPB22]. humicola [LWK20]. Hungary [SCV^{+}20, SVC^{+}21, SMC^{+}22, SCRS^{+}23]. HY9901 [ZTY^{+}22]. hybrid [AGR^{+}22, BFH^{+}23, FFB^{+}21, HHW^{+}24, KDS^{+}21, LPGA^{+}21, SLA^{+}20, VGW^{+}21, YGR^{+}24, ZSL^{+}24a]. hybridization [GKTC^{+}21, HPR^{+}22, IKT^{+}23]. hybrids [LPGW^{+}22]. hydriforme [OHL^{+}20]. hydrochloride [LdAPA^{+}23]. Hydrogen [VPW^{+}21, WSS^{+}20]. hydrophila [ASR^{+}23, CZM^{+}22, DLS^{+}21, DZQ^{+}22, JJZ^{+}22, KQW^{+}21, KB21, MQK^{+}20, QWL^{+}21, SMMM20, TSV^{+}22, WWW^{+}21, WQL^{+}23, XKF^{+}22, ZLWL23, ZCY^{+}23, lZWC^{+}20]. hydroxyl [LYZ^{+}22]. hydroxynaphthol [ZHY^{+}21]. hyperostosis [TCC^{+}24]. hyperthermal [PEś^{+}20]. hypophthalmus [DXD^{+}21, PNØ^{+}21, PØN^{+}21a, PØN^{+}21b, PRC20].

I. [YGR^{+}24]. icefish [FWL^{+}23]. Ichthyobodo [KKC^{+}23]. Ichthyophonus [SdSNK^{+}22]. Ichthyophthirius [HHZ^{+}23, JØM^{+}20, PYD^{+}22, QST^{+}21]. ictaluri [LPGW^{+}22, MPS^{+}22, NHG^{+}23, OGR^{+}24, PNØ^{+}21, PØN^{+}21a, PØN^{+}21b, PRC20, SLA^{+}20, THH^{+}21, YGR^{+}24]. ictaluri-induced [THH^{+}21]. Ictalurid [YLC^{+}21]. Ictalurus [AGR^{+}22, KNG^{+}22, LKD^{+}24, OGR^{+}24, SLA^{+}20, YGR^{+}24]. ideal [PTBSF22]. idellus [YJL^{+}23a]. identical [BGZW22, BFH^{+}23].

Identification [BTP^{+}24, CYX^{+}22, CCCC23, DAW^{+}22, EBELIAH20, ELS^{+}23, FWL^{+}23, HCW^{+}20, HWC^{+}21, LTW^{+}21, LFZ^{+}21, WBM21, WKL^{+}22, ZSL^{+}24a, BLL23, BFvG^{+}23, BRWC23, CXC^{+}24, DMB^{+}21, GYC^{+}22, LLC^{+}21, NK22, NRY^{+}23, RRK^{+}23, SCALIAH23, SSK23, TBB^{+}21, XXX^{+}24, YSW^{+}22, ZDF^{+}23, ZHpT^{+}20, QFW^{+}23]. identified [KSG^{+}23]. identifies [WLK^{+}22]. Identifying [ZZS^{+}24, KPC^{+}22]. Idiopathic [BBD^{+}22]. IgZ [FYH^{+}21]. IHHNV [IPC^{+}24]. II [AANY23, DXL^{+}20a, QFW^{+}23, SLE^{+}20, ZWZ^{+}20]. IIα[QFW^{+}23].Illicium[LYX^{+}20].illucens[BAM^{+}23].Illuminating[TBB^{+}21].Image[Ano22a, CWB^{+}20a, HLH^{+}23a, MHYW22a, SHS^{+}22b, Ano21a, Ano21b, Ano21c].imaging[LFP^{+}24, SHS^{+}22a].imbalance[YGC^{+}20].imbalance − dependent[YGC^{+}20].Immersion[EBA^{+}21, MIR^{+}22, YZM^{+}21, ZZL^{+}21, BMK^{+}20, HWM^{+}24, IAH21, LGJ21].Immune[BFC^{+}23, KBG^{+}23, LDW^{+}23, ZHL^{+}21, ZLY^{+}22, ABA^{+}23, BIK^{+}21, CSGB^{+}21, CWB^{+}20b, CZM^{+}22, DZJ^{+}21, GZW^{+}23, HWZ^{+}21, HWG^{+}21, KQW^{+}21, LKP^{+}22, LWL^{+}24, MPPF21, NLL^{+}21, SLH^{+}24, SDH22, UMD22, WMS21, WQL^{+}23, XKF^{+}22, XJZ^{+}22, YYS^{+}22, YJL^{+}23b, ZSL^{+}24a].Immune − enhancing[LDW^{+}23].immunity[CLC^{+}22, DYPN22, VBG^{+}22, WWW^{+}21, WGE^{+}24, YGZ^{+}22, ZZL^{+}21].immunization[ABM^{+}21, GZL^{+}21, ZZL^{+}21].immuno[ZZZ^{+}23].immuno − chromatic[ZZZ^{+}23].immunoassay[LWT^{+}22].immunochromatographic[WCT^{+}20].immunochromatography[HZFT21].immunogenic[lZWC^{+}20].immunogenicity[VGW^{+}21, ZWZ^{+}20].immunohistochemical[STB^{+}21, SMMM20].immunohistochemistry[IPC^{+}24].immunology[LFL^{+}21].immunomodulator[VBG^{+}22].immunosorbent[LWH^{+}21].Impact[BTN^{+}21, CWB^{+}20b, LL21, LYC^{+}20, WGE^{+}24].impacts[LAH21, RSPB22, UMD22, VPW^{+}21].Implication[CSVG^{+}22, ABBP24].implications[CZMP21, MLHD^{+}23].importance[KPC^{+}22].important[PS23, RKL^{+}20].improve[DLS^{+}21, MSF^{+}20].Improved[WRMT24, SAAB^{+}21].Improvements[GPC22].Improving[PTP^{+}23].in − situ[GKTC^{+}21].inactivated[LJHH23, NRKK21, RBP^{+}20a].inactivates[TR23].Inactivation[WHL^{+}20].including[AAZR^{+}20, RSPB22].inclusion[TMM^{+}24].inclusions[KMH^{+}21].increase[EBA^{+}21, KDHS^{+}23, OCS^{+}23].Increased[PZP^{+}20, AMB^{+}22, FMK^{+}21, VFR^{+}23, YCS^{+}21].increases[ZVYD23].Increasing[MFP21].India[BAS^{+}22, SRP^{+}21, VKK^{+}22].indicate[RCS^{+}22].indicates[LFL^{+}21, RCB^{+}22, SCAH22].indicating[SPM^{+}20].indicators[CGK23, MNPS23].indigenous[AONDO24].Indole[ZVYD23].Indole − 3 − acetic[ZVYD23].indologenes[XXX^{+}24].Indonesia[LDR^{+}22].induce[NRKK21].induced[ADAO22, BFC^{+}23, EBA^{+}21, EMH^{+}22, FK21, FZC^{+}23, HB20, JPC^{+}21, SZZ^{+}23, THH^{+}21].induces[AHSB^{+}22, BCT^{+}20a, BHS^{+}21, HCW^{+}20, HWC^{+}21, SST^{+}22, XFL^{+}21, YGC^{+}20].inducing[ZYL^{+}22, ZZL^{+}21].infect[DHSM^{+}20, YPT^{+}22].infected[AKR^{+}21, BRCI23, DXL^{+}20a, DHSK^{+}21, ESEI20, FdSB^{+}22, FMK^{+}21, KSD^{+}20, LFL^{+}21, SMMM20, TBKY20, VFR^{+}23, WMS21, YYS^{+}22, ZDZZ22, ZLZ21, CMN^{+}20].infecting[GLS^{+}21, HWM^{+}24, PYD^{+}22, RGT^{+}23, SCRS^{+}23].Infection[BWH^{+}20, SCV^{+}20, SJK^{+}21, APS23, AFKK22, AHBE^{+}20, BLL23, BCT^{+}20a, CWB^{+}20b, CCP^{+}21, CTOH20, CLH^{+}21, xCqXfL^{+}21, CDY^{+}20, DSD^{+}22, ENS^{+}23, FSS^{+}21, FAG^{+}22, FYH^{+}21, GZW^{+}23, HLpSC22, HWG^{+}21, HSH^{+}22, IPC^{+}24, IAH21, JPC^{+}21, JBG^{+}23, JRZ^{+}22, KPC^{+}22, KKH^{+}24, LCCC22, LYZ^{+}22, LZC^{+}24, LYX^{+}20, LPW^{+}20, LPL22, LFZ^{+}22, LLJ^{+}21, LYF^{+}21, MIR^{+}22, MWY^{+}20, MSB^{+}21, MVW^{+}20, MGG22, NKS^{+}21, NHG^{+}23, NLL^{+}21, OGR^{+}24, OSM22a, PPDG^{+}22, PTW21, PSKW22, PPWC20, REB^{+}20, SRT^{+}21, SVO^{+}22, SKAS^{+}22, SMC^{+}22, SLH^{+}24, SLA^{+}20, SGR^{+}20a, SDH22, SKS^{+}24, TMM^{+}24, VSODB21, WLTZ21, WML^{+}21, WQL^{+}23, WCT^{+}23a, WHK^{+}23, WGE^{+}24, XJM^{+}21, XZS^{+}23, YZM^{+}21, YCC^{+}21, YLW^{+}21, ZSL^{+}24a, ZZS^{+}24, ZGL^{+}20, ZHL^{+}21, ZLWL23, ZYH^{+}23, ZHL^{+}23, ZAG^{+}20].infections[ALDMLEM20, DAW^{+}22, DHF^{+}22, DŽR^{+}21, LCTC20, MNB^{+}24, MMTD^{+}21, ØSH^{+}22, PHB^{+}23, PPT^{+}20, SVC^{+}21, SdSNK^{+}22, SRE^{+}21].Infectious[AMJ^{+}21, LDR^{+}22, SRP^{+}21, AMB^{+}22, CGA^{+}20, DGB^{+}22, EJH^{+}23, IVR^{+}23, JDA^{+}20, KSG^{+}23, KSM^{+}21, MPMB^{+}20, RGH^{+}22, SAB^{+}22, TBKY20, UCB^{+}21, ZSG^{+}23].infectivity[LAH21, PRC20].infects[GXW^{+}23].infestation[MHP^{+}20, YK21].inflammation[CSE21, CSVG^{+}22, ENS^{+}23, LFP^{+}24, SKR^{+}20, YGC^{+}20, ZBL^{+}23].inflammation − related[ENS^{+}23].inflammatory[CWB^{+}20b, LFZ^{+}21, RGS^{+}21, ZLWL23].inflicting[AONDO24].Influence[CCV^{+}24, OGR^{+}24, BAM^{+}23, CTOH20, FAG^{+}22, HLH^{+}23b].Information[Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano22m, Ano23a, Ano23b, Ano23c, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano24a, Ano24b, Ano24c, Ano24d, Ano24e, Ano24f, Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano21n, Ano21o].ingredient[HLpSC22].ingredients[WLK^{+}22].Inhibition[WYR^{+}23].inhibitor[IFI^{+}21].inhibitory[IAH22, IFIZ22].inhibits[LYZ^{+}22, TR23, ZWX^{+}22].iniae[MCP^{+}22, PDL^{+}23, PTP^{+}23, PTW21, PSH^{+}20, PHB^{+}23, TCS21].initiative[NSP20].injected[SDMO^{+}20].injection[CSGB^{+}21, CSVG^{+}22].Innate[CLC^{+}22, DYPN22, LKP^{+}22, VBG^{+}22].insight[IHK^{+}22, QST^{+}21].Insights[SGR^{+}20b, LS24, LFG^{+}23, NKö^{+}21, LHTW21].Integrated[LYF^{+}21].Integration[HWZ^{+}21].intended[DAW^{+}22].intensities[MFP21, UD22].intensity[MHP^{+}20, WBV^{+}20].intensive[ÚPZ^{+}23].intensively[SMP^{+}20].interaction[VRS^{+}24, WSS^{+}20].interest[NTB^{+}24].Intermediary[SDMO^{+}20].intermediate[APS23].intermedius[LZZ^{+}23].internal[YK21].International[SSM^{+}20].intertidal[BWH^{+}20, MRM21].intervals[CGK23, FEH^{+}22].Intestinal[NPP^{+}24, HHJ^{+}21, MMTD^{+}21, OGR^{+}24, RGS^{+}21, REB^{+}20, SSM^{+}20].intestinalis[APS23].intestine[XQZ^{+}22].intra[CGA^{+}20, LKV^{+}23].intra − host[LKV^{+}23].intra − segmental[CGA^{+}20].intracellular[BFC^{+}23, DZQ^{+}22, TR23, XFL^{+}21, ZCY^{+}23].intramuscular[MLHD^{+}23].Intranuclear[KMH^{+}21].intraperitoneal[CDY^{+}20, PPWC20].intraperitoneally[AKR^{+}21].intravascular[GKAS23].introduced[AONDO24].Introduction[NSP20, CTP^{+}23, SNSP^{+}23].Intussusception[HHJ^{+}21].invadans[BTN^{+}24, WBM21].invasion[BBD^{+}22, ZSL^{+}24b].invasive[BHEM21, MNPS23].invertebrate[PEś^{+}20].investigate[LGJ21].Investigation[LDM^{+}20, PK24, CTP^{+}23, KSD^{+}20, SMMM20].investigations[DvGJK^{+}21, MBF^{+}23].Invited[NSP20].involved[GZW^{+}23, LFZ^{+}21, NLL^{+}21, WKL^{+}22].involvement[YZL^{+}23].Iodobacter[KaVDH^{+}21, VBG^{+}22].IPNV[TIK^{+}22, TBKY20].Ireland[MBS21].iridescent[GXL^{+}24, LGH^{+}21].irido[SBB21].irido − melanocytoma[SBB21].Iridoviridae[SVO^{+}22].iridovirus[ENS^{+}23, GSJ^{+}20, IHK^{+}22, LYX^{+}20, LCW^{+}24, MJX^{+}23].Iron[YGR^{+}24, AHSB^{+}22, AHSCEBI23, BMSC21, LFA20].iron − limited[AHSB^{+}22].irradiation[TPSJS20b].irritans[xCqXfL^{+}21, DXL^{+}20a, TZL^{+}22, WHK^{+}23, XJM^{+}21, YY24, ZYB^{+}23, ZMZ^{+}23, ZWB^{+}23].ISAV[CGA^{+}20, DOW^{+}22, DGB^{+}22].ISAV − HPR[DOW^{+}22].Isavirus[TASP^{+}20].ISBN[FM23, RM20].ISKNV[ZSG^{+}23].isolate[LWL^{+}24, WLH^{+}20, YJL^{+}23b].isolated[AANY23, BCT^{+}20b, BHEM21, BRWC23, CYX^{+}22, CCL^{+}24, ELS^{+}23, HRR^{+}24, KPC^{+}22, LMJ^{+}24, LKD^{+}24, LXW^{+}21, LS24, LH20, LDC^{+}22, MWH^{+}21, MLL^{+}22, MAK^{+}23, MSS^{+}20, MIC^{+}24, MBR^{+}22, NK22, NRY^{+}23, NNC^{+}23, NKZ^{+}20, PBT^{+}23, PCWC20, QCZ^{+}22, RCW^{+}23, RRK^{+}23, SCIAH21, SMD^{+}20, SVO^{+}22, SWC21, TSC^{+}23, WZC^{+}24].isolates[AMM^{+}22, AHTCDI20, BLJ^{+}21, BMO^{+}20, CRP^{+}21, DCB20, IVR^{+}23, IMAH21, Li23, LPGA^{+}21, LPGW^{+}22, MPMB^{+}20, NLPM^{+}21, NGC^{+}23, RCB^{+}22, RCS^{+}22, TIK^{+}22, VIB^{+}21].Isolation[BMO^{+}20, CXC^{+}24, LCTC20, LLR^{+}20, XXX^{+}24, YCL^{+}21, ZHpT^{+}20, BZP^{+}22, DGB^{+}22, HKB^{+}22, MCP^{+}22, SLY^{+}24, WSW^{+}20].isothermal[CMM^{+}20, HLW^{+}23, KDS^{+}21, PTK^{+}21, PYT^{+}22, YDM^{+}21, ZHY^{+}21, ZLS^{+}21].Issue[Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano21n, Ano21o, Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano22m, Ano23a, Ano23b, Ano23c, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano24a, Ano24b, Ano24c, Ano24d, Ano24e, Ano24f].Italy[BTP^{+}24, CTP^{+}23, MPA^{+}21, MCP^{+}22].iTRAQ[ZLZ21].iTRAQ − based[ZLZ21].IVb[LPW^{+}20, LL21].

jack [MAK^{+}23]. jandaei [CXC^{+}24]. Japan [AANY23, IHK^{+}22]. Japanese [LFZ^{+}21, ZWZ^{+}20]. japonica [ZWZ^{+}20]. japonicus [AONDO24, Cow20]. jaw [SVP^{+}21]. jellyfish [MBS21]. Journal [SNSP^{+}23, NSP20]. juniper [BIK^{+}21]. Juniperus [BIK^{+}21]. juvenile [BZP^{+}22, CGK23, CVFRM^{+}23, JLM^{+}20, KKK^{+}23, LdAPA^{+}23, MKT^{+}21, NKS^{+}21, OSM^{+}22b, PGH^{+}21, PTW21, PSKW22, SKS^{+}24]. juveniles [PDZ^{+}22, YGR^{+}24].

Kamchatka [SGU^{+}21]. Kariba [BHEM21]. kasB [ZSL^{+}24b]. kauderni [KSG^{+}23]. Kavango [SPHM23]. KCL [LHLS20]. KCL-5 [LHLS20]. kDa [AHSB^{+}22, SCAH22]. key [JMA^{+}22, PSED^{+}20, QST^{+}21]. KHV [BDK^{+}21, LWH^{+}21]. kidney [EBAH24, JAC^{+}22, KSG^{+}23, LNK^{+}21, LFZ^{+}22, RSPB22, SDH22, SRP^{+}21, SPM^{+}20, ZSG^{+}23]. kill [KKC^{+}23]. killed [MKT^{+}21, ZLY^{+}22]. Kim [FM23]. Kinetics [ASF^{+}21, KBMD21]. kisutch [KKBL21, PPT^{+}20]. Klebsiella [LWL^{+}24]. knifejaw [LZK^{+}23]. knowledge [MPPF21]. Koi [ZOZ^{+}20, AAZR^{+}20, BAS^{+}22, BDK^{+}21, EJH^{+}23, MBF^{+}23, SGR^{+}20a, STB^{+}21, ÚPZ^{+}23, WWK^{+}20, BAS^{+}22, MBF^{+}23, SMC^{+}22, SGR^{+}20a, STB^{+}21]. Kokanee [LHH^{+}22]. Korea [CCL^{+}24, HAK^{+}20, LJK^{+}24, LH20, NK22]. Krabbe [MFP21]. Krøyer [DPLL^{+}21]. kroyeri [YK21]. KT [CLF^{+}22]. Kudoa [MMW^{+}22]. kuruma [Cow20].

L [SVO^{+}22, DPA^{+}24, MPP20a, PRW24]. L. [BRR^{+}23, CdPM^{+}21, DPLL^{+}21, DHSK^{+}21, FSFS^{+}20, FCH^{+}22b, FEM^{+}20, FK21, FMP^{+}21, FMH^{+}22, FEH^{+}22, JBG^{+}23, KSI^{+}20, KDR^{+}22, LGP^{+}20, LPL22, MIR^{+}22, MNB^{+}24, MBJ20, MLM^{+}20, MFP21, MLHD^{+}23, MHH23, ØSH^{+}22, PSH^{+}20, RØLW23, SMC^{+}22, SMP^{+}20, SCD^{+}23, TBK^{+}20, WAF^{+}21]. L3 [AFKK22]. label [LWT^{+}22]. label-free [LWT^{+}22]. Labeo [SMMM20, TSV^{+}22]. Laboratory [NKS^{+}21, DŽR^{+}21, HSH^{+}20, JBG^{+}23, KMH^{+}21, WBB^{+}21]. laboratory-reared [DŽR^{+}21]. Labrador [RGH^{+}22]. labrax [ASNA24, BBS^{+}20, xCqXfL^{+}21, LDM^{+}20, MBN^{+}22, YK21, ZVB^{+}21]. Labrus [PDM^{+}21, PGH^{+}21]. lac [LWT^{+}22]. Lacepede [ASG^{+}22]. Lack [MLM^{+}20]. lactate [TCS21]. lactic [AABB23]. lactococcosis [SMY^{+}22]. Lactococcus [AANY23, AOT22, BLL23, BHEM21, KBG^{+}23, MAK^{+}23, NRY^{+}23, RBP^{+}20a, RCS^{+}22]. lactones [LRGAH21]. Lake [DDHT^{+}22, TNB^{+}22, DDDJ^{+}20, DDTG^{+}21, HKB^{+}22, KDS^{+}21, Li23, MHH23, SPS23, SRN20, TSS^{+}23, YYS^{+}22, BSS^{+}20, BHEM21]. lakes [OSM22a, TBMP23, HKB^{+}22]. Lakshmi [FM23]. LAMP [KDS^{+}21, PYT^{+}22]. lamprey [SCV^{+}20]. langsdorfii [MHA^{+}20]. laparoscopy [LdL20]. large [LHH^{+}22, LWR^{+}20, LLC^{+}21, LCW^{+}24, OSM22a, SJL^{+}23, SRP^{+}21, dCJBG22]. large-scale [SRP^{+}21]. largemouth [BBG21, CYX^{+}22, CCY^{+}23, GWF^{+}22, JXJ^{+}22, JRZ^{+}22, JSZ^{+}23, PPWC20, WYR^{+}23]. Larimichthys [LWR^{+}20, LLC^{+}21, SJL^{+}23]. larvae [DBC^{+}22, LDC^{+}22, LCYL20, MSB^{+}21, MPF^{+}24, MSS^{+}20, MFP21]. larval [MWZ^{+}20]. Lateolabrax [GZW^{+}23, HWM^{+}24]. lateral [CCG24, FCH^{+}22a, LWT^{+}22, PSV^{+}21, ZZZ^{+}23]. lateritius [YJL^{+}23b]. Lates [CSDS21, CGK23, GKTC^{+}21, GKAS23, KSD^{+}20, PTW21, PPWC20, RBM^{+}22, RKC^{+}23, SZZ^{+}23, SSMC22, SKS^{+}24]. latitudinal [GLS^{+}21]. Latris [DBC^{+}22]. latus [LjY^{+}22, LZC^{+}24, ZFC^{+}23]. laurel [ÖTG22]. Laurus [ÖTG22]. layer [HLH^{+}23b]. leaching [MCM^{+}23]. lead [HLpSC22]. LEAP2 [FWL^{+}23]. learned [FAG^{+}22]. lecanii [ACGDO^{+}21]. lecanii-corni [ACGDO^{+}21]. lectin [DYPN22, ZYZ^{+}20]. leech [BBG21]. leei [LJK^{+}24, PSED^{+}20, SJK^{+}21]. leerii [SAKC20]. leg [RRK^{+}23]. Leguminosae [MSF^{+}20]. leiomyoma [NCA^{+}20]. lemon [RZTV21]. leniusculus [BBD^{+}22]. lentisuturalis [SCRS^{+}23]. Lepeophtheirus [CWB^{+}20b, DHSM^{+}20, DPLL^{+}21, PSH^{+}21, UD22, UMD22, VPW^{+}21]. lepeophtherii [HPR^{+}22]. Lernanthropus [YK21]. lesion [JSW^{+}21, KMH^{+}21]. Lesions [SCD^{+}23, DŽR^{+}21, GSS^{+}21, KaVDH^{+}21, NPP^{+}24, SGU^{+}21, VDC20]. Lessons [FAG^{+}22]. Lethal [POS^{+}20, CMM^{+}20, HSS^{+}24, JAC^{+}22]. leucomaenis [SGU^{+}21]. Levamisole [LdAPA^{+}23]. level [CTOH20, FCH^{+}22b, JJZ^{+}22, WJPS21]. levels [YGR^{+}24]. LFD [LCW^{+}24]. Lg [AOT22]. Lg-Per [AOT22]. lice [BONB20, GRM^{+}23, MCA^{+}20, PSH^{+}21, TBB^{+}21, UD22, UMD22]. life [JMA^{+}22, MPF^{+}24]. Light [MDAO22, ADAO22, WZG^{+}20]. Lightner [Now21]. Ligula [APS23]. like [BGZW22, CLH^{+}21, DGB^{+}22, HHW^{+}24, HSG^{+}21, JBB^{+}24, LHC^{+}22, Li23, MBF^{+}23, MIK^{+}24, SGU^{+}21, ZAG^{+}20]. Limanda [VDC20, VDD^{+}20]. limited [AHSB^{+}22]. limnosediminis [KaVDH^{+}21]. Limpopo [SPHM23]. line [AMS^{+}24, FAG^{+}22, FdSB^{+}22, FSW^{+}21, GPL^{+}22, JZW^{+}23, JVM^{+}20, KSM^{+}21, LHC^{+}22, LjY^{+}22, LZC^{+}24, LZK^{+}23, LPW^{+}20, LFZ^{+}22, MWY^{+}22, YPT^{+}22, ZFC^{+}23]. lineage [MSZV^{+}23]. lineages [RCG22, SPM^{+}20]. lineata [DBC^{+}22]. lined [QCZ^{+}22]. lines [ASG^{+}22, ASA^{+}20, FSS^{+}21, FAG^{+}22, JW22, SLK^{+}23, SMG^{+}23]. linked [BFvG^{+}23, KKK^{+}23, LWH^{+}21]. linking [XKF^{+}22]. links [MR23, SSM^{+}20]. Linnaeus [BBS^{+}20, HCC^{+}22, LdL20, ZVB^{+}21, dRRRPSR^{+}21]. Lipid [Din21]. lipids [NRC^{+}23]. Lipoid [EMH^{+}22]. lipomas [PDG^{+}24]. lipopeptide [WHK^{+}23]. Lippia [SMP^{+}22a]. literature [DVR^{+}20]. Lithuanian [MIK^{+}24]. Litopenaeus [BTN^{+}21, CMN^{+}20, CCCC23, HLpSC22, LDR^{+}22, LKP^{+}22, LGH^{+}21, TPSJS20a, TPSJS20b]. live [BMK^{+}20, LPGW^{+}22, SMY^{+}24, SOE^{+}21, lZWC^{+}20]. live-attenuated [BMK^{+}20, LPGW^{+}22]. liver [CSGB^{+}21, CZC^{+}22, EMH^{+}22, JZW^{+}23, LjY^{+}22, OBSH24, PZP^{+}20, SDH22, WQL^{+}23, ZBL^{+}23]. livers [BRCI23]. loach [ZYZ^{+}20]. loaches [VKK^{+}22]. load [MR23, TBKY20]. Localization [IFI^{+}21, DHSK^{+}21, FYH^{+}21]. location [BONB20]. locations [SLPF^{+}20]. loci [JØM^{+}20]. LongAmp [IPC^{+}24]. Longicollum [CRWC22]. longitudinal [DMH^{+}21]. loop [CMM^{+}20, HLW^{+}23, KDS^{+}21, PTK^{+}21, PYT^{+}22, ZHY^{+}21, ZLS^{+}21]. loop-mediated [CMM^{+}20, HLW^{+}23, KDS^{+}21, PTK^{+}21, PYT^{+}22, ZHY^{+}21, ZLS^{+}21]. lordosis [FK21]. lordosis-induced [FK21]. loss [RCG22]. lota [BMO^{+}20, BMO^{+}20]. louse [BONB20, BFC^{+}23, GLS^{+}21, MPF^{+}24]. low [BLL23, BHS^{+}21, BHEM21, LDO^{+}21]. low-temperature [LDO^{+}21]. lucidum [LDW^{+}23]. Lufenuron [OBSH24]. lugubris [BBG21]. lumpfish [DPA^{+}24, KKL^{+}22, LdL20, SVO^{+}22]. lumpsucker [SRE^{+}21]. lumpus [DPA^{+}24, KKL^{+}22, LdL20, SVO^{+}22]. lusoria [LYC^{+}20]. Lutjanus [CRWC22, GKU21]. LuxR [MQK^{+}20]. LuxR-type [MQK^{+}20]. lymphocyte [YGC^{+}20]. lymphocytes [ASF^{+}21]. Lymphoid [AGH^{+}23, KMH^{+}21]. lymphosarcoma [HCC^{+}22].

M. [FFB^{+}21]. mackerel [SSK23, SdSNK^{+}22]. maclovinus [QGG^{+}21, SDMO^{+}20]. Macrobrachium [ABA^{+}23, DYPN22, GYN^{+}24, MWZ^{+}20, NRKK21, TAK^{+}21, ZZS^{+}24]. Macrogyrodactylus [MDAO22]. macrophages [CSVG^{+}22, EBAH24]. macropomum [LdAPA^{+}23]. maculate [HHW^{+}24]. maculatus [GZW^{+}23, HWM^{+}24]. maculosa [BMO^{+}20]. Magnetic [LFP^{+}24]. magnifica [SMD^{+}20]. Magnolia [JRZ^{+}22]. magnolol [ZLWL23]. main [SPHM23]. Maine [ZAG^{+}20]. Major [QFW^{+}23, ASA^{+}20, DXL^{+}20a, FLH^{+}21, GPC22, SCX^{+}21, ZDF^{+}23]. makes [TAK^{+}21]. malachite [QST^{+}21]. MALDI [BFvG^{+}23]. MALDI-TOF [BFvG^{+}23]. male [HHW^{+}24, SBB21, SLA^{+}20]. males [FdSB^{+}22]. Malignant [BSS^{+}20]. maltaromaticum [RKL^{+}20, SMP^{+}22b]. Man [NRKK21]. managed [MAG24, SBVC21]. Management [FM23, NSP20, ABA^{+}23, KSS^{+}20, PTMC22]. mandarin [GPL^{+}22, HWG^{+}21]. mandarinfish [ZDF^{+}23, ZFC^{+}23]. mannose [GZL^{+}21, JXJ^{+}22]. marble [SVP^{+}21]. marbled [CDY^{+}20, LWK20]. Margaritifera [CVDO^{+}20, CVDB^{+}21]. Marine [Man23, DNN^{+}22, LEV^{+}20, MBJ20]. marinum [IKT^{+}23, KSY^{+}23]. maritimum [JBB^{+}24, VIB^{+}21]. Mark [OSM^{+}22b, GVB^{+}21, GSS^{+}21, VDC20]. Marlborough [JSW^{+}21]. marmorata [PEN^{+}21]. marmoratus [LWK20, SVP^{+}21]. masoucida [WLH^{+}20]. Mass [KKH^{+}24, WXM^{+}22, XZS^{+}23]. materials [FSFS^{+}20, ORV^{+}24]. mature [HHJ^{+}21]. maximus [SDH22, WMS21, ZLY^{+}22]. may [AMB^{+}22, KKK^{+}23, HSG^{+}21]. Maylandia [SHS^{+}22a]. measures [DAW^{+}22, KDHS^{+}23]. mechanism [LHTW21, NRC^{+}23, SDH22, VBG^{+}22, XKF^{+}22]. mechanisms [AHSB^{+}22, DPA^{+}24, DZQ^{+}22, KBMD21, LJHH23, ZCY^{+}23]. mediated [CMM^{+}20, HLW^{+}23, KDS^{+}21, PTK^{+}21, PYT^{+}22, ZJF^{+}22, ZHY^{+}21, ZLS^{+}21]. mediates [FZC^{+}23, ZBL^{+}23]. medicated [MCM^{+}23, RKC^{+}23]. medicinal [LXW^{+}20, ZYH^{+}23]. Medicine [RM20]. mediterraneus [BCT^{+}20a]. Megalobrama [ZHpT^{+}20]. megalocytiviruses [SNY^{+}21]. Mekong [PNØ^{+}21, PØN^{+}21a, PØN^{+}21b]. melanization [BRR^{+}23, MLHD^{+}23]. melanized [BFH^{+}23]. melanocytoma [SBB21]. melanoma [BSS^{+}20]. melanomacrophage [PZP^{+}20]. melastigma [YWZ^{+}21]. membrane [BCT^{+}20b, EBELIAH20, EBAH24, HLH^{+}23b, LXZ^{+}21, ZWZ^{+}20]. Memphremagog [BSS^{+}20]. Meningitis [DYN^{+}20]. Meningoencephalitis [RZTV21]. meningoseptica [TSC^{+}23]. Meretrix [LYC^{+}20]. merguiensis [ZHL^{+}23]. meristic [MRM21]. merus [LWL^{+}24]. Mesanophrys [LLR^{+}20]. mesh [ZYB^{+}23]. Mesomycetozoea [SGR^{+}20a]. meta [MBN^{+}22, PTMC22, ZGL^{+}20]. meta-analysis [MBN^{+}22, PTMC22, ZGL^{+}20]. metabolic [SDMO^{+}20]. metabolism [BMSC21, Din21]. metabolite [LFL^{+}21]. metabolome [ZMZ^{+}23]. metacercariae [DMB^{+}21, KNG^{+}22]. metagenomic [KSD^{+}20]. Metastatic [BZQV21, SVP^{+}21]. Method [MSZV^{+}23, TKB^{+}21, DMB^{+}21, HPR^{+}22, JSZ^{+}23, LGP^{+}20, NK22, PYT^{+}22, SMP^{+}20, TZL^{+}22, WCT^{+}23b, YDM^{+}21, ZGC^{+}22, ZZF^{+}23]. methods [ABA^{+}23, FMH^{+}22, JBG^{+}23, LJK^{+}24, PTBSF22, SSM^{+}21]. Metschnikowia [MLL^{+}22]. Mexico [SRLARG24, dRRRPSR^{+}21]. MIC [BLJ^{+}21, GIS^{+}21]. Microbial [DBPH21, KSI^{+}20, MWZ^{+}20, TPSJS20a]. Microbiome [Man23]. microbiota [OGR^{+}24]. Microcotyle [NK22]. microcystin [SST^{+}22]. microdiet [PKF^{+}21]. microfluidic [HLW^{+}23, ZLS^{+}21]. micropeltes [PHB^{+}23]. Micropterus [BBG21, CYX^{+}22, CCY^{+}23, FCH^{+}22a, HZC^{+}24, JRZ^{+}22, LYZ^{+}22, PPWC20, WBM21, XJZ^{+}22, YGZ^{+}22, YYS^{+}24, ZBL^{+}23]. Microradiography [PDZ^{+}22]. MicroRNA [JPC^{+}21, CLF^{+}22]. microRNAs [JPC^{+}21, LDO^{+}21]. microscopic [SMMM20]. microscopy [ADAO22, AONDO24, MDAO22, RCS^{+}22, WZG^{+}20]. microsphere [XJZ^{+}22]. microsporidia [lLWC^{+}24, XZS^{+}23]. migrating [MNB^{+}24]. migration [LEV^{+}20]. milky [MLL^{+}22]. Mimachlamys [BCT^{+}20a]. mimiviruses [RBP^{+}20b]. minigenome [TASP^{+}20]. minimum [IAH22]. MinION(R) [BRWC23]. minnow [HCW^{+}20, HWC^{+}21, LL21]. miR [CLF^{+}22, FZC^{+}23, YCC^{+}21]. miR-124 [FZC^{+}23, YCC^{+}21]. miR-KT-635 [CLF^{+}22]. miRNAs [LYF^{+}21]. Misgurnus [ZYZ^{+}20]. Mississippi [LPGA^{+}21]. Mitigating [RSPB22]. mitochondrial [XKF^{+}22]. Mixed [SBB21, SKS^{+}24, BBS^{+}20]. MLVA [SSP^{+}23]. model [AHBE^{+}20, CSE21, CLS^{+}22, KH20, LHC^{+}22, LXW^{+}20, LGJ21, MIR^{+}22, MSB^{+}21, NLPM^{+}21, PGH^{+}21]. Modelling [LCYL20, BAM^{+}23]. models [MLL^{+}22, PRC20]. modification [CGA^{+}20, GZL^{+}21, JXJ^{+}22]. Modifications [AKR^{+}21]. modified [ABM^{+}21, DZJ^{+}21]. modulate [DCB20]. modulation [LPW^{+}20, LPL22, SDMO^{+}20, VSOB22]. Molecular [DWY^{+}20, ESEI20, FLH^{+}21, KNG^{+}22, LLC^{+}21, LJHH23, MPF^{+}24, OKL^{+}22, WZG^{+}20, YSW^{+}22, KKH^{+}24, LCTC20, MIK^{+}24, NRY^{+}23, PCWC20, ÚPZ^{+}23]. mollusc [BCT^{+}20a]. Monitoring [dYdRFF^{+}23, KPC^{+}22, ZDZZ22]. monoclonal [GZZ^{+}22, WWK^{+}20, ZDF^{+}23, ZGz^{+}22]. monocytes [CSVG^{+}22]. monocytes/macrophages [CSVG^{+}22]. monodon [FMK^{+}21, LDR^{+}22]. Monogenea [IAO20, LFSB^{+}21, MHH23]. monogenean [MSF^{+}20, dAMHdOCF^{+}23]. monolayers [ABBP24]. Monopterus [LFZ^{+}22]. morbidity [WBB^{+}21, XZS^{+}23]. morhua [AFKK22]. moribund [LDC^{+}22]. Moritella [CWB^{+}20b, SSP^{+}23]. Morone [FFB^{+}21]. Morphological [EGRD22, KNG^{+}22, MIK^{+}24]. Morphology [KB21, FCH^{+}22b, NTB^{+}24, PRW24]. morphometric [GLS^{+}21]. morphometrics [FK21]. Morphopathological [CVDO^{+}20]. Morphopathology [CVDB^{+}21]. mortalities [PEś^{+}20, PEN^{+}21, WXM^{+}22]. Mortality [GOG^{+}21, MBS21, AMM^{+}22, BMJ^{+}21, BHT^{+}20, GKAS23, JQT20, KKH^{+}24, MPA^{+}21, MCP^{+}22, NKS^{+}21, NPP^{+}24, ØSH^{+}22, PNAS22, SZZ^{+}23, SPKS^{+}20, SRP^{+}21, THH^{+}21, TBKY20, WJPS21, WLTZ21]. mortem [FEH^{+}22]. mosquito [KSM^{+}21]. mosquitofish [BTN^{+}24, WCT^{+}23a, JVM^{+}20]. mossambicus [SLPF^{+}20]. motifs [TIK^{+}22]. motility [GWL^{+}21, RCG22]. mottled [PEN^{+}21]. mould [CMLK^{+}22]. Mourilyan [Cow20]. Mozambique [SLPF^{+}20]. mRNA [LYF^{+}21]. MSRV [LYZ^{+}22, YYS^{+}24]. Mucin [KH20, LHTW21]. Mucosal [CDY^{+}20, HSM^{+}22]. mucous [PBF^{+}21]. Mucus [ZDZZ22, ASR^{+}23, CSVG^{+}22, FFB^{+}21, HBRN24, HSM^{+}22, LFA20, LAH21, LHTW21]. mucus-secreting [CSVG^{+}22]. mud [MSS^{+}20, XZS^{+}23]. Mugil [RBP^{+}20a]. mullet [HZFT21, RBP^{+}20a]. multi [DLS^{+}21, MPS^{+}22, SNSP^{+}23, WAF^{+}21]. multi-authored [SNSP^{+}23]. multi-biomarker [WAF^{+}21]. multi-drug-resistant [DLS^{+}21]. multi-epitope [MPS^{+}22]. multidrug [HHW^{+}24]. multidrug-resistant [HHW^{+}24]. multifiliis [HHZ^{+}23, JØM^{+}20, PYD^{+}22, QST^{+}21]. multilocus [KKBL21, QCZ^{+}22]. Multiple [LMJ^{+}24, YLC^{+}21, DDDJ^{+}20, ZLS^{+}21]. multiplex [PS23]. Murray [WXM^{+}22]. muscle [BFH^{+}23, SKR^{+}20, SCRS^{+}23]. muscle-infecting [SCRS^{+}23]. musculature [BRR^{+}23]. mussel [LKD^{+}24]. mussels [CTOH20, ZAG^{+}20]. mutant [FdSB^{+}22]. mutants [SOE^{+}21, SGW^{+}22]. Mycobacterial [DŽR^{+}21]. Mycobacteriosis [RØLW23, IKT^{+}23, MHA^{+}20, SSK23, SBS^{+}22]. Mycobacterium [IKT^{+}23, JW22, KSY^{+}23, MHA^{+}20]. Mycoplasma [SVC^{+}21]. mykiss [AMJ^{+}21, AWF22, BAM^{+}23, BIK^{+}21, BZQV21, BTP^{+}24, ELB21, FEMMI^{+}23, GSS^{+}21, HSM^{+}22, HHJ^{+}21, JØM^{+}20, JKM^{+}20, KBG^{+}23, LAH21, LFSB^{+}21, LPL22, MIC^{+}24, ÖTG22, PYD^{+}22, PPT^{+}20, QGG^{+}21, SM21, TBKY20, VIB^{+}21, VWE^{+}23, öTTS20]. myocardial [PRHW21]. myocarditis [MLM^{+}20]. myonecrosis [KSM^{+}21, SAB^{+}22]. myositis [PSH^{+}20]. Mysterious [SPKS^{+}20]. Mytilus [CTOH20, MCA^{+}20, ZAG^{+}20]. Myxobolidae [KRS^{+}20]. Myxobolus [AWF22, KRS^{+}20, SCRS^{+}23, WZG^{+}20]. myxoid [BZQV21]. Myxosporea [SJK^{+}21]. myxospores [SGR^{+}20b]. Myxozoa [MMW^{+}22, REB^{+}20, SJK^{+}21, WBB^{+}21]. myxozoan [LKV^{+}23, SGR^{+}20b]. Myzobdella [BBG21].

N [WWW^{+}21]. N- [WWW^{+}21]. naccarii [MCP^{+}22]. naïve [EBA^{+}21]. Nano [SAAB^{+}21]. Nano-scale [SAAB^{+}21]. nanobioparticles [MMTD^{+}21]. nanobubble [DLS^{+}21]. Nanoemulsions [MSF^{+}20]. nanoformulation [MSF^{+}20]. nanoparticle [GYN^{+}24, KDS^{+}21]. nanoparticles [ASA^{+}20, VRS^{+}24]. Nanopore [BRWC23, DDTG^{+}21]. narinari [DYN^{+}20]. naringenin [SCX^{+}21]. nasal [WBV^{+}20]. Nasicola [WBV^{+}20]. natamycin [CMLK^{+}22]. Natural [HLpSC22, HSH^{+}22, WLTZ21, YYS^{+}24, WHK^{+}23, WLK^{+}22]. naturally [ESEI20, GXW^{+}23, JLM^{+}20]. nature [ALDMLEM20]. NB2011 [JLH^{+}21]. nebulosus [BSS^{+}20]. necrosis [CYX^{+}22, HZFT21, KSG^{+}23, LCYL20, MPMB^{+}20, NAaS^{+}22, PØN^{+}21a, PØN^{+}21b, SMY^{+}24, SRP^{+}21, TBKY20, UCB^{+}21, VSODB21, VSOB22, WML^{+}21, WXM^{+}22, WCT^{+}20, YDM^{+}21, YCL^{+}21, ZGC^{+}22, ZSG^{+}23, JBP24]. necrotizing [BBD^{+}22]. Negaprion [RZTV21]. negative [PLS^{+}22]. nematode [BRCI23, NLH^{+}20, VKK^{+}22]. nematodes [DYN^{+}20]. Neodevriesia [ACGDO^{+}21]. Neoechinorhynchidae [SMP^{+}22a]. Neoechinorhynchus [LdAPA^{+}23, SMP^{+}22a]. Neoergasilus [AONDO24]. Neoparamoeba [BHS^{+}21, CMM^{+}20, EBA^{+}21, LHTW21, WRMT24]. neoplasia [AGH^{+}23, DVR^{+}20, HHJ^{+}21, SBVC21, SCS24]. neoplasms [SSM^{+}20]. neoplastic [DŽR^{+}21]. Neotropical [dCJBG22]. Nephrocalcinosis [KKK^{+}23, KBR^{+}22, KDR^{+}22]. nerka [LHH^{+}22]. nerve [BZQV21]. Nervous [VSOB22, CYX^{+}22, HZFT21, LCYL20, VSODB21, WXM^{+}22]. Netherlands [KKH^{+}24]. netpen [JLM^{+}20]. netpen- [JLM^{+}20]. nets [ORV^{+}24]. network [XKF^{+}22]. neurophilia [FdSB^{+}22, MVW^{+}20, SKAS^{+}22]. neutralizing [Wel20]. neutrophil [MSB^{+}21]. Newfoundland [RGH^{+}22]. newly [JVM^{+}20]. next [HAK^{+}20]. next-generation [HAK^{+}20]. NGD [BTP^{+}24]. NGS [AOT22]. Nibea [XJM^{+}21]. Nie [SCRS^{+}23]. nigripulchritudo [RRK^{+}23]. Nile [AMM^{+}22, BHEM21, DLS^{+}21, DHSK^{+}21, DSD^{+}22, ESEI20, HBRN24, KDS^{+}21, LCTC20, MPS^{+}22, MKT^{+}21, NLL^{+}21, PDL^{+}23, PDG^{+}24, PPWC20, RCC^{+}21, SRLARG24, XFL^{+}21]. niloticus [ZSL^{+}24a, AMM^{+}22, BHEM21, DDH^{+}22, DLS^{+}21, DHSK^{+}21, DSD^{+}22, ESEI20, HBRN24, KAS^{+}24, LCTC20, LHC^{+}22, MKT^{+}21, PDL^{+}23, PDG^{+}24, PTK^{+}21, PPWC20, RCC^{+}21, XFL^{+}21, YSW^{+}22, YY24]. nine [NGC^{+}23]. nipponense [DYPN22]. NlpC [HCW^{+}20]. NlpC/P60 [HCW^{+}20]. nobilis [ÖTG22]. Nocardia [CCY^{+}23, GYC^{+}22, HCW^{+}20, HWC^{+}21, JRZ^{+}22, NRC^{+}23, NNC^{+}23, NGC^{+}23, WHC^{+}21, WKL^{+}22, ZBL^{+}23, ZSL^{+}24b, dRRRPSR^{+}21]. Nocardiosis [YTH^{+}24, NGC^{+}23, dRRRPSR^{+}21]. NOD2 [GZW^{+}23]. nodavirus [GYN^{+}24, WLTZ21]. Nodular [CTP^{+}23, BTP^{+}24]. nodules [LWR^{+}20]. Nolandella [EBA^{+}21]. Non [CMM^{+}20, HSS^{+}24, JAC^{+}22, BDK^{+}21, CSDS21, DDHT^{+}22, HLH^{+}23b, LHTW21, LdL20, MIC^{+}24]. non-A-layer [HLH^{+}23b]. non-cyprinid [BDK^{+}21]. non-destructive [CSDS21, LdL20]. non-fish-derived [LHTW21]. Non-lethal [CMM^{+}20, HSS^{+}24, JAC^{+}22]. non-pathogenic [MIC^{+}24]. non-tilapine [DDHT^{+}22]. Nonstructural [LLJ^{+}21]. normal [PKF^{+}21]. North [SVO^{+}22, AFKK22, DHF^{+}22, TBMP23]. Northeast [SdSNK^{+}22]. Northeastern [BTP^{+}24, CTP^{+}23]. northern [KNG^{+}22, MHH23, MCP^{+}22, NKZ^{+}20]. Norway [DOW^{+}22, GOG^{+}21, JQT20, MNB^{+}24, MFP21, MHH23, PBF^{+}21, WJPS21].

Norwegian [AMJ^{+}21, DWK^{+}22, LEV^{+}20, OSM22a, RØLW23, RCG22, RGS^{+}22, SSP^{+}23, SNT^{+}22, TBK^{+}20]. Nothobranchius [DŽR^{+}21]. notothenioid [QGG^{+}21, SDMO^{+}20]. Novel [HPR^{+}22, LKD^{+}24, MHYW22b, PGH^{+}21, PRHW21, ASG^{+}22, BMK^{+}20, DYPN22, GKTC^{+}21, HHW^{+}24, HSG^{+}21, LCCC22, LXW^{+}21, Li23, LDC^{+}22, MPP20a, PTP^{+}23, SSP^{+}23, SLY^{+}24, WSW^{+}20, Wel20, YCL^{+}21, ZWZ^{+}20]. NP [TASP^{+}20]. NS26 [LLJ^{+}21]. nucleic [IKT^{+}23, YDM^{+}21]. nucleoprotein [AKR^{+}21, TASP^{+}20]. Nucleospora [KMH^{+}21]. nucleotide [KSY^{+}23, SNY^{+}21]. number [YY24]. numbers [PBF^{+}21]. nylon [ORV^{+}24, VRC^{+}22].

O. [PPT^{+}20]. O1 [Wel20]. O1-specific [Wel20]. O1a [FEMMI^{+}23]. O1b [AABB23]. obscure [SLH^{+}24]. obscurus [FLH^{+}21, KQW^{+}21, QFW^{+}23, SLH^{+}24]. observation [YWZ^{+}21]. Observations [OSM^{+}22b]. observed [KAS^{+}24, VDC20]. Occurrence [MNB^{+}24, MHH23, PDG^{+}24, BAS^{+}22, IVR^{+}23, JDA^{+}20, Roh22, SAB^{+}22]. occurring [JLM^{+}20]. ocellatum [LZC^{+}24, ZHL^{+}21]. ocellatus [dRRRPSR^{+}21]. Ochroconis [LWK20]. ocular [SGR^{+}20a]. oedema [ABBP24, BPA^{+}21, MPP^{+}20b]. officinalis [JRZ^{+}22]. OH [ZWB^{+}23]. oil [POS^{+}20, TR23]. oils [ÖTG22, SMP^{+}22a]. oleoresin [MSF^{+}20]. olivaceus [CLC^{+}22, LFZ^{+}21, SJK^{+}21]. olive [CLC^{+}22, SJK^{+}21]. OMP [TSV^{+}22]. OmpA [BCT^{+}20b]. OmpR [DZQ^{+}22]. OmpU [ZWZ^{+}20]. OmpU-A-II [ZWZ^{+}20]. OMV [EBAH24]. on-site [YDM^{+}21]. Oncorhynchus [AMJ^{+}21, AWF22, BAM^{+}23, BIK^{+}21, BZQV21, BTP^{+}24, ELB21, FEMMI^{+}23, GSS^{+}21, HSM^{+}22, HHJ^{+}21, JØM^{+}20, JKM^{+}20, KBG^{+}23, KKBL21, LHH^{+}22, LAH21, LFSB^{+}21, LPL22, LFP^{+}24, MIC^{+}24, NKS^{+}21, NPP^{+}24, NLPM^{+}21, ÖTG22, PYD^{+}22, PPT^{+}20, QGG^{+}21, RFN^{+}23, SST^{+}22, SM21, TBKY20, VIB^{+}21, VWE^{+}23, öTTS20]. Onion [öTTS20]. onites [ÖTG22]. ontogenic [TCC^{+}24]. oomycete [NTB^{+}24]. Oomycota [MIK^{+}24]. Oplegnathus [ENS^{+}23, LZK^{+}23]. Opportunities [SAAB^{+}21]. optimal [FSFS^{+}20, RKC^{+}23]. Optimization [DMB^{+}21, GZL^{+}21]. optimized [CDJ^{+}23, TZL^{+}22]. oral [DXD^{+}21, MMTD^{+}21, RKC^{+}23, ZLWL23]. orally [CLC^{+}22, NRKK21, RGT^{+}23]. orange [DXL^{+}20a, NGC^{+}23, ZWB^{+}23]. orange-spotted [DXL^{+}20a, NGC^{+}23, ZWB^{+}23]. Orchitophryidae [LLR^{+}20]. ordalii [EBELIAH20, HSH^{+}20]. oregano [ÖTG22]. Oreochromis [AMM^{+}22, BHEM21, DDH^{+}22, DLS^{+}21, DHSK^{+}21, DSD^{+}22, ESEI20, HBRN24, KAS^{+}24, LCTC20, LHC^{+}22, MKT^{+}21, NKZ^{+}20, PDL^{+}23, PDG^{+}24, PTK^{+}21, PSKW22, PPWC20, RCC^{+}21, SLPF^{+}20, SWC21, XFL^{+}21, YSW^{+}22, YY24, ZSL^{+}24a, NHG^{+}23]. ORF121 [GZZ^{+}22]. ORF149 [MLGRF^{+}20]. ORF23 [CLF^{+}22].

ORF25 [ZHL^{+}22]. ORF3 [YLC^{+}21]. ORF63 [ZHL^{+}22]. ORF72 [WWK^{+}20]. organism [JBB^{+}24]. organisms [MIK^{+}24]. Organs [Muf23]. oriental [DYPN22]. orientalis [PPWC20]. Origanum [ÖTG22]. origin [CTOH20]. originating [VPW^{+}21]. ornamental [SBS^{+}22]. orthomyxovirus [SRT^{+}21]. orthoreovirus [MNB^{+}24, MBJ20, SKR^{+}20, TMM^{+}24, PVCG20, PPT^{+}20, WHL^{+}20]. orthoreovirus-1 [MNB^{+}24, SKR^{+}20]. Oryzias [YWZ^{+}21]. osculatum [BRCI23]. osmoregulatory [KKK^{+}23]. Osphronemus [GSJ^{+}20, SRP^{+}21]. Osteichthyes [SBVC21]. osteoblastic [SVP^{+}21]. Osteoglossum [LS24]. osteosarcoma [SVP^{+}21]. Ostrea [BWH^{+}20, BHT^{+}20]. other [TNB^{+}22, YTH^{+}24]. outbreak [CdPM^{+}21, FMP^{+}21, IVR^{+}23, LWR^{+}20]. Outbreaks [PSH^{+}20, AMB^{+}22, GVB^{+}21, MPP^{+}20b, PTMC22, SZH^{+}22, ÚPZ^{+}23]. outcomes [SKS^{+}24]. Outer [BCT^{+}20b, EBELIAH20, EBAH24, ZWZ^{+}20]. Ovarian [BBS^{+}20, CZMP21]. ovary [FSW^{+}21]. overcome [LHTW21]. overproduction [AHSB^{+}22]. Ovicidal [SMP^{+}22a]. Oxford [BRWC23]. Oxford-Nanopore [BRWC23]. oxidase [FSW^{+}21]. oxidative [BCT^{+}20a, SYKC23]. oxide [THW^{+}22]. oxide-DNAzyme-based [THW^{+}22]. oxytetracycline [IMAH21, MCM^{+}23]. ozonation [TPSJS20b]. Ozone [DLS^{+}21].

P60 [HCW^{+}20]. Pacifastacus [BBD^{+}22]. Pacific [BTN^{+}21, HSH^{+}20, LXW^{+}21, TPSJS20a, TPSJS20b, WSW^{+}20, WGE^{+}24]. Paddlefish [OHL^{+}20]. Paenibacillus [YJL^{+}23a]. paeoniflorin [HLpSC22]. pages [RM20]. pagrosomi [CRWC22]. pallasii [HSH^{+}20]. palm [NSTS24]. Pampus [JZW^{+}23, ZHL^{+}21]. pan [SCAH22]. pan-genomic [SCAH22]. pancreas [DSD^{+}22, NKö^{+}21, RTA21]. pancreatic [MPMB^{+}20, TBKY20]. Pangasianodon [DXD^{+}21, PNØ^{+}21, PØN^{+}21a, PØN^{+}21b, PRC20]. pansteatitis [SLPF^{+}20]. pansteatitis-affected [SLPF^{+}20]. paper [SNSP^{+}23]. papilloma [SGU^{+}21]. papilloma-like [SGU^{+}21]. parahaemolyticus [GWL^{+}21, LWT^{+}22, MWH^{+}21, NSTS24, WCT^{+}20, WGE^{+}24, ZSL^{+}24a, ZGC^{+}22]. Paralichthys [CLC^{+}22, LFZ^{+}21, SJK^{+}21]. paramamosain [XZS^{+}23]. parameter [AEE^{+}23]. parameters [CMLK^{+}22, JMA^{+}22, MMTD^{+}21, PSED^{+}20, SMP^{+}20, WMS21, dCJBG22]. Paramoeba [TKB^{+}21]. parasite [BKM21, CRWC22, DWK^{+}22, LNK^{+}21, LKV^{+}23, MNPS23, SCV^{+}20, SCRS^{+}23, WBV^{+}20, dAMHdOCF^{+}23]. parasites [LZZ^{+}23, MSF^{+}20, MCA^{+}20, NPP^{+}24, VKK^{+}22]. Parasitic [YK21, BRCI23, CVDB^{+}21, HRR^{+}24, HHZ^{+}23, JØM^{+}20, LHH^{+}22, MPF^{+}24, NKS^{+}21, NLH^{+}20, TZL^{+}22]. parasitica [ELS^{+}23, FSS^{+}21, MGG22, öTTS20]. parasitism [KKC^{+}23]. parasitizing [KSI^{+}20, KBMD21]. parauberis [Roh22]. partial [MPS^{+}22]. particular [RCB^{+}22]. partners [MLGRF^{+}20]. parvovirus [LCCC22, PYT^{+}22, YTS22, ZZF^{+}23, DSD^{+}22]. Pasteurella [KKL^{+}22, SRE^{+}21]. pathobiome [PPDG^{+}22]. pathogen [AANY23, AHSCEBI23, CLH^{+}21, EBELIAH20, FdSB^{+}22, GPC22, IMAH21, LAH21, LXW^{+}21, PTMC22, SCALIAH23, XFL^{+}21]. pathogen-free [FdSB^{+}22]. Pathogenesis [CCP^{+}21, Din21, FMH^{+}22, LHC^{+}22]. Pathogenic [SMD^{+}20, SRLARG24, WLH^{+}20, CMLK^{+}22, DCB20, DLS^{+}21, LHLS20, MQK^{+}20, MIC^{+}24, PCWC20, SCRS^{+}23, VLSM^{+}20, ZZS^{+}24, ZCY^{+}23]. Pathogenicity [NGC^{+}23, CCY^{+}23, CTOH20, CZJ^{+}23, CCL^{+}24, Cow20, HWZ^{+}21, JLH^{+}21, LS24, NRY^{+}23, PPWC20, SWC21, YZL^{+}23, ZTY^{+}22]. pathogens [ASA^{+}20, DBPH21, HLW^{+}23, LEV^{+}20, SKO21, YJL^{+}23b, ZWZ^{+}20, ZLS^{+}21]. Pathological [IAO20, CVFRM^{+}23, MBF^{+}23, SMC^{+}22, XXX^{+}24]. Pathology [AGR^{+}22, AONDO24, ADAO22, GKU21, GKAS23, KB21, MBS21, ØSH^{+}22, PRHW21, SPKS^{+}20, SKS^{+}24, WBV^{+}20]. pathway [CCV^{+}24, DXL^{+}20a, HWM^{+}24, OBSH24, WCT^{+}23a, ZZS^{+}24]. pathways [LFL^{+}21]. pattern [LDM^{+}20]. Patterns [OSM22a, TCC^{+}24, GOG^{+}21, IMAH21, KBR^{+}22, PNAS22]. PCR [CQG^{+}21, GKTC^{+}21, GXL^{+}24, HSS^{+}24, HHZ^{+}23, IPC^{+}24, JSZ^{+}23, MMW^{+}22, MJX^{+}23, NK22, NLR21, PS23, SCALIAH23, SKAS^{+}22, SRLARG24, TCS21, YTS22, ZGC^{+}22, ZZF^{+}23]. PCR-based [ZZF^{+}23]. PCR-RFLP [NK22]. pearl [SAKC20]. Pectinatella [SMD^{+}20]. Pediococcus [LKP^{+}22]. PEG [DZJ^{+}21]. PEG-modified [DZJ^{+}21]. pelagic [MCA^{+}20]. Pelodiscus [CRP^{+}21, XXX^{+}24]. penaei [CCCC23]. Penaeus [BLL23, Cow20, FMK^{+}21, LDR^{+}22, LCCC22, LXW^{+}21, LDO^{+}21, MWH^{+}21, QCZ^{+}22, RRK^{+}23, SAB^{+}22, WSW^{+}20]. pentosaceus [LKP^{+}22]. peptidase [HCW^{+}20]. peptide [HHW^{+}24, HLH^{+}23b, IKT^{+}23]. peptides [PLS^{+}22]. peracetic [TPSJS20a]. Perca [KSI^{+}20, SMP^{+}20]. perch [JPC^{+}21, KSI^{+}20, MWY^{+}22, PSPP^{+}22, SMP^{+}20]. percid [PSPP^{+}22]. Perciformes [MRM21]. Perciformes-Gobiidae [MRM21]. perezi [GXW^{+}23]. performance [FCH^{+}22b, PØN^{+}21b]. performed [BLJ^{+}21]. performing [GIS^{+}21]. perfringens [BZP^{+}22]. peri [LFP^{+}24]. peri-vertebral [LFP^{+}24]. Perkinsus [BCT^{+}20a]. permeability [HLH^{+}23b]. permease [TCS21]. peroxide [VPW^{+}21, WSS^{+}20]. Perry [BME^{+}20, EMH^{+}22]. persist [DHSM^{+}20]. Persistence [YPT^{+}22]. Peru [FEMMI^{+}23, VLSM^{+}20]. perurans [BHS^{+}21, CMM^{+}20, EBA^{+}21, LHTW21, TKB^{+}21, WRMT24]. perurans-induced [EBA^{+}21]. pet [DBDSP20]. petauri [AOT22]. Peters [SLPF^{+}20]. petterae [ADAO22]. PFGE [RCS^{+}22].

Phaeohyphomycosis [GMK^{+}21, ACGDO^{+}21, PAS^{+}21]. phage [DXD^{+}21]. pharmacokinetic [CMLK^{+}22]. Pharmacokinetics [RKC^{+}23, RCC^{+}21]. phase [GOG^{+}21, JQT20]. phases [FMH^{+}22]. Phenotype [SWC21, WZC^{+}24]. Phenotypic [LS24, LWL^{+}24, MLL^{+}22, MIC^{+}24, NKZ^{+}20, RCB^{+}22, AHSCEBI23, GPC22, GLS^{+}21, KaVDH^{+}21, TSC^{+}23]. phenotyping [RCS^{+}22]. philander [IAO20]. Philippines [LCTC20]. phoP [SOE^{+}21]. phoQ [SOE^{+}21]. phospholipase [XQZ^{+}22]. phospholipid [HLH^{+}23b]. Photobacterium [ESEI20, PCWC20, WSW^{+}20]. Phylogenetic [MPMB^{+}20, PBT^{+}23, RCW^{+}23, GKTC^{+}21, HAK^{+}20, TIK^{+}22]. phylogenetics [KB21]. Physalia [MBS21]. physalis [MBS21]. physiological [LDO^{+}21, ZWB^{+}23, dCJBG22]. physiology [SGR^{+}20b]. Phytochemical [NSTS24]. phytogenic [SVP^{+}24]. PI3K [ZZS^{+}24]. PI3K-AKT [ZZS^{+}24]. picornavirus [LDC^{+}22]. pilchard [SRT^{+}21]. Pimephales [LL21]. pindani [SHS^{+}22a]. Pinpointing [VDD^{+}20]. piperine [WYR^{+}23]. PirAB [BGZW22]. pirarucu [dAMHdOCF^{+}23]. piRNAs [ZDZZ22]. piscicida [AGR^{+}22, BRWC23, CLS^{+}22, HWM^{+}24, LTW^{+}21, LPGA^{+}21, LPGW^{+}22, NLPM^{+}21, PCWC20, Roh22, SOE^{+}21, SGW^{+}22, YMC^{+}21]. piscidin [PLS^{+}22]. Piscine [SBS^{+}22, MNB^{+}24, MBJ20, MLM^{+}20, SMY^{+}22, TMM^{+}24, PVCG20, PPT^{+}20, SKR^{+}20, WHL^{+}20]. Piscirickettsia [BFC^{+}23, DMH^{+}21, FVC^{+}21, LGP^{+}20, LAH21, LRGAH21, LGJ21, OCS^{+}23, ORV^{+}24, QGG^{+}21, SVP^{+}24, SDMO^{+}20, VBL^{+}23, VRC^{+}22, ZVB^{+}21]. piscirickettsiosis [JLM^{+}20]. piscisicarius [CCP^{+}21]. piscium [SCALIAH23]. pkh2101 [AANY23]. plagiosum [MAG24]. plague [SFKB^{+}20]. planktonic [TBB^{+}21]. plant [LXW^{+}20, ZYH^{+}23]. plasmid [AANY23]. plasticity [GLS^{+}21]. platensis [BAM^{+}23]. platform [SSMC22]. platforms [NAaS^{+}22]. play [XQZ^{+}22]. plecoglossicida [HWZ^{+}21, JLH^{+}21, LWR^{+}20, SJL^{+}23, YZL^{+}23]. Plecoglossus [THH^{+}21]. plerocercoids [APS23]. plumbagin [HSH^{+}22]. PMCV [MR23]. pneumoniae [LWL^{+}24]. Poaceae [POS^{+}20]. pocket [TASP^{+}20]. Poecilia [Li23, PK24, SLY^{+}24]. point [CMM^{+}20, SZH^{+}22]. point-of-care [CMM^{+}20]. Polish [MPMB^{+}20]. polyculture [DDHT^{+}22]. polyethylene [ORV^{+}24]. polymerase [FCH^{+}22a, QWL^{+}21, SPS23, SMY^{+}22, TBK^{+}22, WML^{+}21, WCT^{+}23b]. polymorphic [CGA^{+}20]. polymorphism [KSY^{+}23, SNY^{+}21]. polymyxa [YJL^{+}23a]. Polyopisthocotylea [LFSB^{+}21]. Polypodium [OHL^{+}20]. polysaccharides [LDW^{+}23]. polyunsaturated [HLH^{+}23b]. pomfret [JZW^{+}23, ZHL^{+}21]. pompano [ASG^{+}22, NRY^{+}23, YTH^{+}24]. POMV [SRT^{+}21]. pond [HBRN24, XZS^{+}23]. pond-cultured [XZS^{+}23]. ponds [MWH^{+}21, NKZ^{+}20, QCZ^{+}22]. population [MPP20a, PSED^{+}20, SBVC21, dYdRFF^{+}23]. populations [BAM^{+}23, GRM^{+}23]. Portunidae [XZS^{+}23]. Portunus [LLR^{+}20, XZS^{+}23]. positive [DOW^{+}22, WGE^{+}24]. possesses [HHW^{+}24]. possible [BTN^{+}21, SPKS^{+}20, MLHD^{+}23]. post [FEH^{+}22, MNB^{+}24]. post-mortem [FEH^{+}22]. post-smolt [MNB^{+}24]. Potential [ABBP24, CCY^{+}23, LDO^{+}21, ABM^{+}21, AFKK22, AABB23, BCT^{+}20b, BHEM21, BRWC23, CLS^{+}22, GYN^{+}24, GRM^{+}23, HSH^{+}22, LFL^{+}21, MCA^{+}20, PEś^{+}20, QST^{+}21, SOE^{+}21, SFKB^{+}20, WYR^{+}23, XQZ^{+}22, ZSL^{+}24a]. potentially [YGR^{+}24]. pp [FM23]. prawn [ABA^{+}23, DYPN22, TAK^{+}21]. praziquantel [RGT^{+}23]. prebiotic [OGR^{+}24]. predators [SFKB^{+}20]. predicting [FK21]. Preetham [FM23]. preference [KKBL21]. preferences [BONB20]. pregelatinized [MCM^{+}23]. Preliminary [JLH^{+}21]. presence [PAH^{+}22]. presentation [DXL^{+}20a, MBF^{+}23]. presmolt [PBF^{+}21]. prespawn [NPP^{+}24]. Press [FM23, RM20]. pressure [YK21]. Prevalence [MHP^{+}20, QRTD^{+}22, WBV^{+}20, ZGL^{+}20, CTP^{+}23, QCZ^{+}22, RBM^{+}22, SRLARG24, SdSNK^{+}22]. prevalent [SGU^{+}21]. Prevention [ÖTG22, öTTS20, HB20, SZH^{+}22]. primarily [SNT^{+}22]. primary [ABBP24, AMB^{+}22, GSJ^{+}20, NKö^{+}21, SGR^{+}20a]. prime [DSD^{+}22]. priming [PSV^{+}21]. principal [LLJ^{+}21, SCAH22]. probe [KDS^{+}21, lLWC^{+}24, PTP^{+}23, YTS22]. probe-based [lLWC^{+}24, PTP^{+}23, YTS22]. probiotic [AABB23, OGR^{+}24, VRS^{+}24]. Probiotics [PDL^{+}23, ABA^{+}23, YYS^{+}22]. Procambarus [BAT^{+}24, HSG^{+}21, SCX^{+}21, ZWX^{+}22, ZLZ21]. procedure [GZL^{+}21]. process [CFGCCE21, SNSP^{+}23]. processing [TZL^{+}22]. Procypris [LWL^{+}24]. produce [EBAH24]. product [YYS^{+}24]. Production [ZDF^{+}23, AHSCEBI23, GOG^{+}21, JQT20, LDR^{+}22, LHTW21, MBJ20, SAAB^{+}21]. profile [SMD^{+}20]. profiles [MBR^{+}22, ZDZZ22]. profiling [AMM^{+}22, HBRN24, JPC^{+}21, KBG^{+}23]. progeny [WZG^{+}20]. Progression [SKAS^{+}22]. proliferation [LHH^{+}22]. proliferative [LNK^{+}21, RSPB22, SPM^{+}20]. promelas [LL21]. promising [ZMZ^{+}22]. promoter [YLC^{+}21]. promotes [LKP^{+}22]. prop [FdSB^{+}22]. propagating [MSB^{+}21]. propagation [GSJ^{+}20, KSM^{+}21]. properties [SGR^{+}20b]. prophenoloxidase [TAK^{+}21]. prophylactic [KDHS^{+}23, ZZL^{+}21]. proPO [TAK^{+}21]. Proposed [GIS^{+}21]. protease [AAZR^{+}20]. protect [XJM^{+}21]. protection [BMK^{+}20, MPS^{+}22, NRKK21, SLA^{+}20]. Protective [DXD^{+}21, YGZ^{+}22, CLC^{+}22, HHW^{+}24, LPGW^{+}22, NRC^{+}23, SMY^{+}24, ZLWL23, ZZL^{+}21]. protein [AHSB^{+}22, BCT^{+}20b, GYN^{+}24, HBRN24, LGH^{+}21, LLJ^{+}21, LYF^{+}21, MPS^{+}22, MHA^{+}20, MLGRF^{+}20, NLL^{+}21, PHS^{+}20, SCAH22, SYKC23, WKL^{+}22, WWK^{+}20, ZWZ^{+}20]. protein-10 [MHA^{+}20]. proteinase [IFI^{+}21]. proteinases [IFIZ22]. proteins [BGZW22, CdPM^{+}21, FVC^{+}21, HSM^{+}22, NRKK21, QST^{+}21, ZLZ21]. Proteocephalus [IFIZ22]. proteolytic [FVC^{+}21]. proteolyticus [ZHL^{+}23]. Proteomic [AHSB^{+}22, CdPM^{+}21, LHLS20, QST^{+}21, ZLZ21]. protocol [CDJ^{+}23, GIS^{+}21]. provided [IHK^{+}22]. provides [LFG^{+}23, QST^{+}21]. Prudhoe [MDAO22]. Prussian [KKH^{+}24]. PRV [PPT^{+}20, TMM^{+}24]. PRV-1 [PPT^{+}20]. PRV-2 [TMM^{+}24]. Pseudocapillaria [NLH^{+}20, SSM^{+}20, SLK^{+}23]. Pseudocaranx [MAK^{+}23]. Pseudocrenilabrus [IAO20]. Pseudoloma [FdSB^{+}22, MVW^{+}20, SKAS^{+}22]. Pseudomonas [HWZ^{+}21, JLH^{+}21, LWR^{+}20, SJL^{+}23, VFR^{+}23, WHK^{+}23, YZL^{+}23]. Pseudoparamoeba [EBA^{+}21]. pseudoshottsii [MHA^{+}20]. Pseudoterranova [AFKK22, NAPP24]. psychrophilum [AHTCDI20, AHBE^{+}20, AWF22, BMK^{+}20, BMJ^{+}21, BMSC21, CVFRM^{+}23, HKB^{+}22, HSM^{+}22, IVR^{+}23, KKBL21, LPL22, MIR^{+}22, VFR^{+}23]. psychrophilum-associated [BMJ^{+}21]. Pterapogon [KSG^{+}23]. Pterygoplichthys [DDH^{+}22]. public [SBVC21]. PUFAs [HLH^{+}23b]. puffer [LJK^{+}24, SLH^{+}24]. pufferfish [FLH^{+}21, KQW^{+}21, QFW^{+}23]. pulse [RCS^{+}22]. pulse-field [RCS^{+}22]. punctatus [AGR^{+}22, KNG^{+}22, LKD^{+}24, LZK^{+}23, OGR^{+}24, SLA^{+}20, YGR^{+}24]. Puno [FEMMI^{+}23]. Puntigrus [CCP^{+}21]. Purification [EJH^{+}23]. purposes [PTBSF22]. putative [GKU21, MLGRF^{+}20]. putatively [RGS^{+}22]. PVN02 [DXD^{+}21]. pyloric [ZDF^{+}23]. Pyrococcus [WCT^{+}23b]. pyroptosis [ZBL^{+}23].

qPCR [AMJ^{+}21, DMH^{+}21, GWF^{+}22, JAC^{+}22, LGP^{+}20, lLWC^{+}24, PTP^{+}23, RGS^{+}22, SLK^{+}23, SKR^{+}20, SLE^{+}20, SMP^{+}22b, WRMT24]. QTL [JØM^{+}20]. quality [PTMC22]. Quantification [BMSC21, JSZ^{+}23, MPF^{+}24, OKL^{+}22, TCS21]. Quantifying [JMA^{+}22, LYC^{+}20]. Quantitative [JØM^{+}20, NLR21, TZL^{+}22, CQG^{+}21, HHZ^{+}23, SMY^{+}22, YTS22, ZLZ21]. Quebec [BSS^{+}20]. Quinoline [LYZ^{+}22]. quinones [VBL^{+}23]. quorum [LRGAH21, WGE^{+}24].

rA7 [LFL^{+}21]. Radiological [KDR^{+}22]. Rafinesque [LL21]. Rahnella [LHLS20]. Rainbow [öTTS20, AMJ^{+}21, AKR^{+}21, AWF22, BAM^{+}23, BIK^{+}21, BZQV21, BTP^{+}24, CTP^{+}23, EBAH24, ELB21, FEMMI^{+}23, GVB^{+}21, HSM^{+}22, HHJ^{+}21, JØM^{+}20, JKM^{+}20, KBG^{+}23, LFSB^{+}21, LJHH23, LPW^{+}20, LPL22, MIC^{+}24, OSM^{+}22b, ÖTG22, PYD^{+}22, PPT^{+}20, QGG^{+}21, SM21, SMG^{+}23, TBKY20, VIB^{+}21, VWE^{+}23, VFR^{+}23, YZM^{+}21]. Rana [TSC^{+}23]. ranavirus [GWF^{+}22, JSZ^{+}23, SVO^{+}22, ZFC^{+}23, ZDF^{+}23]. range [GLS^{+}21, NNC^{+}23, TBMP23]. Rapid [BHT^{+}20, CGA^{+}20, DDTG^{+}21, FCH^{+}22a, LHH^{+}22, LCW^{+}24, ZHY^{+}21, CCG24, HZFT21, PTK^{+}21, PYT^{+}22, PDZ^{+}22, QWL^{+}21, SSMC22, TBB^{+}21, WCT^{+}20, YLW^{+}21]. rare [TBB^{+}21]. RAS [OSM^{+}22b]. rates [NKS^{+}21]. Raton [FM23]. ray [DYN^{+}20, EGRD22]. reaction [SPS23, SMY^{+}22]. Reactor [VRC^{+}22]. Real [SKR^{+}20, GWF^{+}22, GXL^{+}24, HSS^{+}24, HHZ^{+}23, LCW^{+}24, MSB^{+}21, MMW^{+}22, QWL^{+}21, SRLARG24, TCS21]. Real-time [SKR^{+}20, GWF^{+}22, GXL^{+}24, HSS^{+}24, HHZ^{+}23, LCW^{+}24, MSB^{+}21, MMW^{+}22, QWL^{+}21, SRLARG24, TCS21]. reared [BFH^{+}23, DŽR^{+}21, FK21, JLM^{+}20, MHP^{+}20, MCP^{+}22, PDL^{+}23, QRTD^{+}22, SAB^{+}22]. rearing [CLH^{+}21, FCH^{+}22b, VSODB21, YPT^{+}22]. reasons [YCS^{+}21]. reassortant [KPC^{+}22]. recA [OKL^{+}22]. receptor [YGC^{+}20]. receptors [CLH^{+}21]. recirculating [BTN^{+}21]. recirculation [QRTD^{+}22, TPSJS20a, TPSJS20b]. recognition [WWW^{+}21]. Recombinant [NRKK21, LFL^{+}21, TSV^{+}22, XJZ^{+}22, ZSG^{+}23]. recombinase [CCG24, FCH^{+}22a, GWF^{+}22, QWL^{+}21, TBK^{+}22, WML^{+}21, WCT^{+}23b]. recombinase-aided [CCG24, GWF^{+}22]. recombination [CGA^{+}20]. record [NAPP24]. recovered [LCTC20, SLY^{+}24, VIB^{+}21]. Recovery [LZX^{+}23, CVDB^{+}21]. Red [BFH^{+}23, ENS^{+}23, GVB^{+}21, GSS^{+}21, BRR^{+}23, BAT^{+}24, CRWC22, GKU21, IHK^{+}22, KDS^{+}21, KB21, LMJ^{+}24, PSKW22, SCIAH21, VDC20, XKF^{+}22, dRRRPSR^{+}21, OSM^{+}22b, WAF^{+}21]. Rederivation [FdSB^{+}22]. redox [YGC^{+}20]. reduced [BTN^{+}21]. reduces [BKM21]. reducing [WGE^{+}24]. reduction [LXZ^{+}21]. reference [CGK23, MBN^{+}22, RCB^{+}22]. Regan [SCV^{+}20, SBB21]. region [CGA^{+}20, SMD^{+}20]. registrations [PNAS22]. regulate [FZC^{+}23]. regulated [CCV^{+}24]. regulates [CLF^{+}22, GWL^{+}21]. regulating [LFZ^{+}21, YMC^{+}21, ZTY^{+}22]. regulation [BMSC21, LFL^{+}21, REB^{+}20, XKF^{+}22, YCC^{+}21]. regulator [LLJ^{+}21, WZC^{+}24]. regulators [MQK^{+}20]. related [CTOH20, EBBIAH21, ENS^{+}23, JMA^{+}22, MPA^{+}21, NLL^{+}21, ØSH^{+}22, RRK^{+}23, RKL^{+}20, TNB^{+}22]. relationship [PTMC22]. relative [SMP^{+}20]. Remembering [Now21]. Renibacterium [AHSB^{+}22, EBBIAH21, GIS^{+}21, JAC^{+}22, JBG^{+}23, MSZV^{+}23, PAH^{+}22, RFN^{+}23, SCAH22]. reovirus [LLJ^{+}21, YLW^{+}21]. replicated [PNØ^{+}21]. replication [CLF^{+}22, FZC^{+}23, LLJ^{+}21, NKö^{+}21, TR23, WYR^{+}23, ZHL^{+}22, ZWX^{+}22]. report [BAS^{+}22, CRWC22, CCL^{+}24, DGB^{+}22, GVB^{+}21, LHLS20, PPDG^{+}22, SAB^{+}22, VIB^{+}21, VDC20, XZS^{+}23, ZAG^{+}20]. represses [OBSH24]. reproductively [PZP^{+}20]. required [SJL^{+}23]. rerio [APK^{+}23, AHBE^{+}20, FdSB^{+}22, JW22, KSS^{+}20, KMH^{+}21, KH20, MSB^{+}21, MVW^{+}20, PKF^{+}21, SSM^{+}20, SKAS^{+}22, SLK^{+}23, WLTZ21]. Research [NSP20, DBPH21, JW22, JLM^{+}20, KSS^{+}20, NLH^{+}20, WSW^{+}20]. resembling [VDC20]. reservoir [BDK^{+}21, LHH^{+}22]. reservoirs [KB21, PEś^{+}20, ŘB24]. Resistance [AANY23, MGG22, AWF22, BAM^{+}23, BIK^{+}21, CZM^{+}22, DBDSP20, DXL^{+}20b, HLH^{+}23b, JØM^{+}20, KDHS^{+}23, LS24, LPL22, MWH^{+}21, MR23, NRC^{+}23, OGR^{+}24, PDL^{+}23, PØN^{+}21a, QCZ^{+}22, RBM^{+}22, YGR^{+}24, YJL^{+}23a]. resistant [DLS^{+}21, HHW^{+}24, QCZ^{+}22]. resolution [CSVG^{+}22]. resonance [LFP^{+}24]. Resource [SSM^{+}20]. response [ABA^{+}23, BFC^{+}23, CLH^{+}21, CZM^{+}22, DZJ^{+}21, GWL^{+}21, GZW^{+}23, HWZ^{+}21, HSM^{+}22, HWG^{+}21, KQW^{+}21, LKP^{+}22, LFZ^{+}21, LWL^{+}24, LGH^{+}21, MPPF21, MPS^{+}22, NLL^{+}21, QFW^{+}23, SLH^{+}24, SMG^{+}23, SDH22, SDMO^{+}20, UMD22, WQL^{+}23, WSS^{+}20, XFL^{+}21, YYS^{+}22, ZSL^{+}24a, ZHL^{+}21, ZHL^{+}23, ZWB^{+}23]. responses [BDK^{+}21, BIK^{+}21, CWB^{+}20b, CDY^{+}20, DWY^{+}20, RGS^{+}21, XQZ^{+}22, ZLY^{+}22]. restoration [LKD^{+}24]. restriction [LL21]. reticulata [Li23, MSF^{+}20, SLY^{+}24]. retinopathy [BPA^{+}20, ZMZ^{+}22]. retrieved [AHTCDI20, FEMMI^{+}23, KSD^{+}20]. Retrobulbar [SAKC20]. Retrospective [ACGDO^{+}21, SSM^{+}20, DBDSP20, DVR^{+}20]. returning [MFP21]. reveal [XKF^{+}22]. revealed [AOT22, SGR^{+}20b, WZC^{+}24, LHLS20]. reveals [AHSB^{+}22, DDDJ^{+}20, FSW^{+}21, FMK^{+}21, HWZ^{+}21, JPC^{+}21, KSY^{+}23, RCG22, SDH22, SSP^{+}23, XJM^{+}21, YCS^{+}21]. Reverse [KDS^{+}21]. reversible [SST^{+}22]. Review [CSE21, FM23, JDA^{+}20, KSS^{+}20, Man23, Muf23, NSP20, RM20, ALDMLEM20, DVR^{+}20, DBPH21, EGRD22, JBP24, MBN^{+}22, PTMC22, SKO21, SBD21, VRS^{+}24, ZGL^{+}20]. RFLP [NK22]. RGM2144 [VFR^{+}23]. RGNNV [JPC^{+}21, KPC^{+}22, LDM^{+}20, WLK^{+}22]. RGNNV/SJNNV [KPC^{+}22]. rhabdovirus [FCH^{+}22a, HZC^{+}24, HWG^{+}21, LYZ^{+}22, LFZ^{+}22, LZX^{+}23, PSPP^{+}22, XJZ^{+}22, YGZ^{+}22, YYS^{+}24, dYdRFF^{+}23]. Rhine [DHF^{+}22]. Rhomboplites [HB20]. rhythm [CCV^{+}24]. rice [LFZ^{+}22]. rice-field [LFZ^{+}22]. rich [YLC^{+}21]. Richardson [GMK^{+}21]. rickettsia [JBB^{+}24]. rickettsia-like [JBB^{+}24]. Rise [SNSP^{+}23]. risk [AEE^{+}23, AMB^{+}22, CMN^{+}20, MIK^{+}24, ŘB24]. risks [BTN^{+}21, CTP^{+}23]. Risso [PZP^{+}20]. River [SCV^{+}20, WBM21, DYPN22, NKZ^{+}20]. river-based [NKZ^{+}20]. riverine [THH^{+}21]. rivers [DHF^{+}22, LMJ^{+}24]. RMS [GVB^{+}21]. RNA [HWZ^{+}21, TASP^{+}20, WRMT24]. RNA-based [WRMT24]. RNA-binding [TASP^{+}20]. RNA-seq [HWZ^{+}21]. RNAi [HWZ^{+}21]. rock [ENS^{+}23]. rockfish [LWK20, NK22, PAS^{+}21]. rogercresseyi [BFC^{+}23, GLS^{+}21, JMA^{+}22, MCA^{+}20]. roggenkampii [LS24]. rohita [SMMM20, TSV^{+}22]. Rohu [TSV^{+}22]. Role [DXL^{+}20a, MQK^{+}20, SLH^{+}24, ASNA24, ASR^{+}23, APK^{+}23, DPA^{+}24, DZQ^{+}22, VDD^{+}20, WZC^{+}24, YMC^{+}21, ZTY^{+}22, ZCY^{+}23, ZYL^{+}22]. roles [DYPN22, JPC^{+}21, JLH^{+}21, LDO^{+}21, SYKC23, XQZ^{+}22]. rosenbergii [ABA^{+}23, GYN^{+}24, MWZ^{+}20, NRKK21, TAK^{+}21, ZZS^{+}24]. Rosette [CDJ^{+}23]. rostral [KH20]. rostrata [WZC^{+}24, YCL^{+}21]. round [EGRD22]. routes [LDM^{+}20, NGC^{+}23]. RPA [LCW^{+}24]. RPA-LFD [LCW^{+}24]. RSIV [SNY^{+}21]. RSIV-type [SNY^{+}21]. rstA [ZCY^{+}23]. RT [AMJ^{+}21, KDS^{+}21, SKR^{+}20]. RT-LAMP [KDS^{+}21]. RT-qPCR [AMJ^{+}21, SKR^{+}20]. RTGILL [SMG^{+}23, LPW^{+}20]. RTGILL-W1 [SMG^{+}23, LPW^{+}20]. RTGUTGC [SMG^{+}23]. rubripes [xCqXfL^{+}21, LJK^{+}24]. ruckeri [AABB23, BLJ^{+}21, BAM^{+}23, BIK^{+}21, FEMMI^{+}23, RCG22, RGS^{+}22, SM21, VLSM^{+}20, Wel20, YZM^{+}21]. Rudolphi [MFP21]. rugosum [IFI^{+}21]. running [RSPB22]. Russia [SGU^{+}21].

S.A [RM20]. S1 [SKR^{+}20]. S3 [YJL^{+}23a]. safety [BTN^{+}21, MMTD^{+}21]. sagittata [LFSB^{+}21]. salar [AMB^{+}22, BRR^{+}23, BFH^{+}23, BMJ^{+}21, CWB^{+}20b, CdPM^{+}21, DPLL^{+}21, DMH^{+}21, FSFS^{+}20, FCH^{+}22b, FMP^{+}21, FMH^{+}22, FEH^{+}22, GLS^{+}21, HSS^{+}24, JQT20, JBG^{+}23, KBR^{+}22, KDR^{+}22, KKK^{+}23, LMJ^{+}24, LGP^{+}20, LAH21, MIR^{+}22, MNB^{+}24, MBJ20, MLM^{+}20, MBS21, MFP21, MLHD^{+}23, MR23, ØSH^{+}22, PBF^{+}21, PRW24, RØLW23, SCIAH21, SRT^{+}21, SST^{+}22, SNT^{+}22, SCD^{+}23, TBK^{+}20, UD22, UMD22, WJPS21, WAF^{+}21]. salaris [MHH23]. salinity [BLL23, BTN^{+}21, MWH^{+}21, VSOB22]. Salmincola [HRR^{+}24, LHH^{+}22, NKS^{+}21]. Salmo [AMB^{+}22, BRR^{+}23, BFH^{+}23, BMJ^{+}21, CWB^{+}20b, CdPM^{+}21, DPLL^{+}21, DMH^{+}21, FSFS^{+}20, FCH^{+}22b, FMP^{+}21, FMH^{+}22, FEH^{+}22, GLS^{+}21, HSS^{+}24, JQT20, JBG^{+}23, KBR^{+}22, KDR^{+}22, KKK^{+}23, LMJ^{+}24, LGP^{+}20, LKV^{+}23, LEV^{+}20, LAH21, LFSB^{+}21, MIR^{+}22, MNB^{+}24, MBJ20, MLM^{+}20, MBS21, MFP21, MLHD^{+}23, MR23, ØSH^{+}22, PBF^{+}21, PRW24, RØLW23, SCIAH21, SRT^{+}21, SST^{+}22, SVP^{+}21, SNT^{+}22, SCD^{+}23, TBK^{+}20, UD22, UMD22, WJPS21, WBB^{+}21, WAF^{+}21]. salmoides [BBG21, CYX^{+}22, CCY^{+}23, FCH^{+}22a, HZC^{+}24, JRZ^{+}22, LYZ^{+}22, PPWC20, XJZ^{+}22, YGZ^{+}22, YYS^{+}24, ZBL^{+}23]. Salmon [GRM^{+}23, SST^{+}22, ABM^{+}21, AMJ^{+}21, AMB^{+}22, AHTCDI20, BHS^{+}21, BRR^{+}23, BFH^{+}23, BMJ^{+}21, BONB20, BFC^{+}23, CGA^{+}20, CWB^{+}20b, CVDO^{+}20, CVDB^{+}21, CdPM^{+}21, DHSM^{+}20, DOW^{+}22, DPLL^{+}21, DMH^{+}21, DGB^{+}22, EBA^{+}21, FSFS^{+}20, FCH^{+}22b, FMP^{+}21, FMH^{+}22, FEH^{+}22, GLS^{+}21, GIS^{+}21, HSS^{+}24, HPR^{+}22, JQT20, JBG^{+}23, JSW^{+}21, JLM^{+}20, KBR^{+}22, KDR^{+}22, KKK^{+}23, KKBL21, LMJ^{+}24, LGP^{+}20, LNK^{+}21, LHH^{+}22, LFL^{+}21, LGJ21, LFP^{+}24, MIR^{+}22, MNB^{+}24, MBJ20, MLM^{+}20, MGG22, MBS21, MFP21, MLHD^{+}23, MR23, MSZV^{+}23, NKS^{+}21, NPP^{+}24, NLPM^{+}21, NKö^{+}21, OBSH24, ØSH^{+}22, PSH^{+}21, PBF^{+}21, PNAS22, PRW24, PPT^{+}20, QRTD^{+}22, RØLW23, RGH^{+}22, RTA21, SCIAH21, SRT^{+}21, SRE^{+}21, SKR^{+}20, SSP^{+}23, SNT^{+}22, SCD^{+}23, TMM^{+}24, TBB^{+}21, TASP^{+}20, TBK^{+}20, UD22, UMD22, WJPS21, WAF^{+}21, WSS^{+}20]. Salmon-lice [GRM^{+}23]. salmonicida [GPC22, WLH^{+}20, BLJ^{+}21, DCB20, GPC22, HRR^{+}24, HLH^{+}23b, KKL^{+}22, LFL^{+}21, LH20, LFG^{+}23, PDM^{+}21, PGH^{+}21, VDD^{+}20, ZHY^{+}21]. salmonid [ABM^{+}21, AHCSCI20, DWK^{+}22, HRR^{+}24, LH20, MNB^{+}24, MPMB^{+}20, ORV^{+}24, OSM22a, SSM^{+}21, SPM^{+}20]. salmonids [GOG^{+}21, IMAH21, JDA^{+}20, PRHW21, RSPB22]. salmoninarum [AHSB^{+}22, EBBIAH21, GIS^{+}21, JAC^{+}22, JBG^{+}23, MSZV^{+}23, PAH^{+}22, RFN^{+}23, SCAH22, SEL^{+}20, SLE^{+}20, SMP^{+}22b]. salmonis [BFC^{+}23, CWB^{+}20b, DHSM^{+}20, DPLL^{+}21, DMH^{+}21, FVC^{+}21, LGP^{+}20, LAH21, LRGAH21, LGJ21, MHH23, OCS^{+}23, ORV^{+}24, PSH^{+}21, QGG^{+}21, SVP^{+}24, SDMO^{+}20, UD22, UMD22, VBL^{+}23, VPW^{+}21, VRC^{+}22, ZVB^{+}21]. Salvelinus [BMJ^{+}21, MHH23, SGU^{+}21, SEL^{+}20]. same [IVR^{+}23]. sample [JBG^{+}23, TZL^{+}22, ZLS^{+}21]. samples [CQG^{+}21, CSDS21, SPS23, UCB^{+}21]. sampling [JAC^{+}22]. sandwich [LWH^{+}21]. Saprolegnia [ELS^{+}23, FSS^{+}21, MGG22, öTTS20]. Saprolegniasis [ÖTG22]. sativum [öTTS20]. Sauvage [PRC20]. saxatilis [FFB^{+}21]. scale [CSDS21, PSV^{+}21, SAAB^{+}21, SSMC22, SRP^{+}21, KSD^{+}20]. Scandinavian [LMJ^{+}24]. scanning [ADAO22, AONDO24, MDAO22, SMMM20]. Scavenger [SBD21]. Schlegel [CLC^{+}22, SJK^{+}21, THH^{+}21]. schubertii [HHW^{+}24]. Sciaenops [dRRRPSR^{+}21]. scientific [JW22]. Scomber [SSK23, SdSNK^{+}22]. scombrus [SSK23, SdSNK^{+}22]. scophthalmi [OKL^{+}22]. Scophthalmus [SDH22, WMS21, ZLY^{+}22]. scoring [GSS^{+}21]. SCPCs [NKö^{+}21]. Screening [AABB23, JRZ^{+}22, YJL^{+}23a, JXJ^{+}22, KPC^{+}22, LXW^{+}20, LdL20, RGS^{+}22]. Scylla [MSS^{+}20, XZS^{+}23]. SDDV [KSD^{+}20, PSV^{+}21]. sea [BONB20, BFC^{+}23, CSDS21, DNN^{+}22, ENS^{+}23, GLS^{+}21, GZW^{+}23, IHK^{+}22, JPC^{+}21, KPC^{+}22, LDM^{+}20, LKV^{+}23, LEV^{+}20, LjY^{+}22, MPF^{+}24, MFP21, MCA^{+}20, ØSH^{+}22, PSH^{+}21, PDZ^{+}22, RBM^{+}22, REB^{+}20, SNT^{+}22, SSMC22, YK21, ZVB^{+}21, DXL^{+}20b, NAPP24, WAF^{+}21]. sea-farmed [ØSH^{+}22]. seabass [BBS^{+}20, CGK23, NCA^{+}20, PTW21, PPWC20, RKC^{+}23, SKS^{+}24]. seabream [AEE^{+}23, CSGB^{+}21, CSVG^{+}22, CFGCCE21, FEM^{+}20, FK21, LZC^{+}24, LDC^{+}22, RGT^{+}23, ZFC^{+}23]. seafood [LWT^{+}22]. seahorse [XQZ^{+}22]. search [SPKS^{+}20]. seasonal [AEE^{+}23]. seasonality [MHH23]. Seawater [SRT^{+}21, JQT20, MWH^{+}21, MMW^{+}22, TZL^{+}22, TPSJS20a, TBK^{+}20, YY24]. seawater-adapted [YY24]. seawater-farmed [TBK^{+}20]. Sebastes [PAS^{+}21]. sebastis [NK22]. Sebastiscus [LWK20]. Second [Man23]. secondary [JJZ^{+}22]. secreted [HWC^{+}21, MHA^{+}20]. secreting [CSVG^{+}22]. secretion [KBMD21]. Secretory [XQZ^{+}22, PHS^{+}20, WKL^{+}22]. sectional [MBJ20]. seedling [SZH^{+}22]. segment [SKR^{+}20]. segmental [CGA^{+}20]. Selection [GYN^{+}24]. semi [CQG^{+}21]. semi-quantitative [CQG^{+}21]. semifasciolatus [SCS24]. semilaevis [ZDZZ22]. senegalensis [VSODB21]. Senegalese [VSODB21]. senescence [NPP^{+}24]. sensing [LRGAH21, WGE^{+}24]. sensitive [JSZ^{+}23, LCW^{+}24, MJX^{+}23, NAaS^{+}22, SPS23, SSMC22, THW^{+}22, WCT^{+}23b, WRMT24, ZZF^{+}23]. sensitivity [CTOH20, LGP^{+}20, LJK^{+}24, VLSM^{+}20]. separate [KSG^{+}23]. septicaemia [AKR^{+}21, AAZR^{+}20, ASF^{+}21, BAM^{+}23, CCV^{+}24, DXD^{+}21, HAK^{+}20, LL21, RZTV21, UCB^{+}21]. septicemia [TBK^{+}22]. seq [HWZ^{+}21]. sequence [CGA^{+}20, KSD^{+}20, KKBL21, QCZ^{+}22, TIK^{+}22]. sequence-based [TIK^{+}22]. sequences [SCIAH21, SKR^{+}20, YLC^{+}21]. sequencing [BRWC23, DDTG^{+}21, FSW^{+}21, FMK^{+}21, HAK^{+}20]. Serbia [RCW^{+}23]. serial [SSP^{+}23]. Series [NSP20, AGH^{+}23, DVR^{+}20]. Seriola [PZP^{+}20]. seriolae [CCY^{+}23, GYC^{+}22, HCW^{+}20, HWC^{+}21, JRZ^{+}22, NRC^{+}23, NNC^{+}23, NGC^{+}23, WHC^{+}21, WKL^{+}22, ZBL^{+}23, ZSL^{+}24b, dRRRPSR^{+}21]. Serological [AHTCDI20, BDK^{+}21]. serotype [AANY23, FEMMI^{+}23, Wel20]. serotypes [MAK^{+}23]. serovar [BFvG^{+}23]. serrata [MSS^{+}20]. serum [CdPM^{+}21]. settlement [UMD22]. seven [EGRD22, WMS21]. several [GPL^{+}22, LZK^{+}23]. Severe [BZP^{+}22, KBR^{+}22, PSH^{+}20, TSS^{+}23]. severity [EBA^{+}21, JLM^{+}20]. sex [BBS^{+}20]. SGIV [LXW^{+}20]. shape [FK21]. shark [RZTV21]. sharks [IKT^{+}23, MAG24]. sharptooth [MDAO22]. sheath [BZQV21]. shell [CXC^{+}24]. Shewanella [PBT^{+}23]. SHK [EBBIAH21, FVC^{+}21]. SHK-1 [EBBIAH21, FVC^{+}21]. showed [AKR^{+}21]. shows [LFP^{+}24]. shrimp [BLL23, BGZW22, BTN^{+}21, CMN^{+}20, CZM^{+}22, Cow20, FMK^{+}21, HLW^{+}23, IPC^{+}24, LDR^{+}22, LKP^{+}22, LXW^{+}21, LGH^{+}21, NAaS^{+}22, RRK^{+}23, SZH^{+}22, TPSJS20a, TPSJS20b, WSW^{+}20, WML^{+}21, WCT^{+}23b, WGE^{+}24, YDM^{+}21, ZVYD23]. SHT [KAS^{+}24]. Siamese [SBB21]. Siberian [CVFRM^{+}23]. siderophores [AHSCEBI23]. sidoides [SMP^{+}22a]. sign [HHW^{+}24, SLA^{+}20]. signal [BBD^{+}22, LTW^{+}21, TBB^{+}21]. signalling [CCV^{+}24, WCT^{+}23a]. significantly [FMK^{+}21]. signs [LMJ^{+}24, ZAG^{+}20]. Silencing [TAK^{+}21]. silver [JZW^{+}23, LS24, ZHL^{+}21]. Simple [FK21]. simplex [MFP21].

Simultaneous [PS23, WML^{+}21, ZLS^{+}21]. sinensis [CRP^{+}21, DWY^{+}20, Din21, MLL^{+}22, XXX^{+}24, ZGL^{+}20]. single [KSY^{+}23, SNY^{+}21]. Siniperca [GPL^{+}22, HWG^{+}21, LZX^{+}23, MWY^{+}22]. site [LYZ^{+}22, YDM^{+}21]. sites [APS23, AFKK22, PSH^{+}21]. situ [GKTC^{+}21, HPR^{+}22, IKT^{+}23, ŘB24]. SJNNV [KPC^{+}22]. Skeletal [MRM21, APK^{+}23, BRR^{+}23, BFH^{+}23, PDZ^{+}22, PKF^{+}21, SKR^{+}20]. Skin [HSM^{+}22, WAF^{+}21, AONDO24, CSVG^{+}22, CFGCCE21, JSW^{+}21, KaVDH^{+}21, LMJ^{+}24, LAH21, MDAO22, MWY^{+}22, MBS21, SZZ^{+}23, SGU^{+}21, VDD^{+}20]. Slovak [MPP^{+}20b]. smallmouth [WBM21]. Smartphone [NAaS^{+}22]. Smartphone-compatible [NAaS^{+}22]. Smith [RM20]. smolt [MNB^{+}24]. snakehead [HHW^{+}24, PHB^{+}23]. snapper [CRWC22, GKU21, HB20]. snapshot [BONB20]. SNP [SNY^{+}21]. snubnose [ASG^{+}22]. sobria [LH20]. socolofi [SHS^{+}22a]. SOD [HWC^{+}21]. sodium [dAMHdOCF^{+}23]. soft [CXC^{+}24, LFP^{+}24]. soft-shell [CXC^{+}24]. softshell [CRP^{+}21]. solani [BME^{+}20]. sole [VSODB21]. Solea [VSODB21]. some [KKBL21]. Sophora [ZYH^{+}23]. sore [KB21]. Sounds [JSW^{+}21]. source [KPC^{+}22]. sources [AHSCEBI23, KSY^{+}23]. South [BTN^{+}24, KRS^{+}20, AONDO24, CCL^{+}24, DXL^{+}20b, KB21, SMD^{+}20, SLPF^{+}20, ZGL^{+}20]. South-East [ZGL^{+}20]. south-eastern [KRS^{+}20, KB21]. southern [GLS^{+}21, VLSM^{+}20, dYdRFF^{+}23]. Southwest [SRLARG24]. soya [RGS^{+}21]. sp [BMO^{+}20, DNN^{+}22, EBA^{+}21, JKM^{+}20, KKC^{+}23, KMH^{+}21, LLR^{+}20, NHG^{+}23, PTBSF22, PAS^{+}21, POS^{+}20, PSKW22, SGR^{+}20a, SdSNK^{+}22, VBG^{+}22, VFR^{+}23]. spadefish [TCC^{+}24]. Spain [dYdRFF^{+}23]. Sparicotyle [RGT^{+}23]. sparrmanii [AONDO24]. Sparus [AEE^{+}23, CSGB^{+}21, CSVG^{+}22, CFGCCE21, FEM^{+}20, FK21, KPC^{+}22, LDC^{+}22, MBN^{+}22, RGT^{+}23]. Spatial [LKV^{+}23]. Spatio [JQT20]. Spatio-temporal [JQT20]. SPDV [NKö^{+}21]. special [ŘB24]. species [AGH^{+}23, AHCSCI20, DDHT^{+}22, DVR^{+}20, GKU21, GKTC^{+}21, LCTC20, LGJ21, MPF^{+}24, NTB^{+}24, OSM22a, PEś^{+}20, PCWC20, SPM^{+}20, VWE^{+}23]. Specific [PYT^{+}22, GZZ^{+}22, KKBL21, SCALIAH23, SPS23, SM21, SSP^{+}23, Wel20, WWK^{+}20, ZGz^{+}22]. specificity [CDJ^{+}23, LGP^{+}20, LJK^{+}24]. specimens [CSVG^{+}22, EGRD22]. sperm [FdSB^{+}22]. spermatogenic [NCA^{+}20]. spheroid [FSS^{+}21]. spikes [GKAS23]. spinal [GPL^{+}22, LFP^{+}24, PSKW22]. Spinitectus [ADAO22]. spleen [ENS^{+}23, GSJ^{+}20, KSG^{+}23, SRP^{+}21, ZFC^{+}23, ZSG^{+}23]. splendens [SBB21]. splendidus [CTOH20]. splendidus-related [CTOH20]. splenic [PDG^{+}24]. splitnose [PAS^{+}21]. Spontaneous [DVR^{+}20]. sportfishes [KB21]. spot [BAT^{+}24, HLpSC22, SZH^{+}22, SCX^{+}21, SAB^{+}22, TAK^{+}21, WCT^{+}23b, ZWX^{+}22, ZLZ21]. Spotted [GZW^{+}23, DXL^{+}20a, DYN^{+}20, LZK^{+}23, NGC^{+}23, ZWB^{+}23]. spp [BTN^{+}21, BPA^{+}20, BMK^{+}20, DDH^{+}22, MBR^{+}22, NKZ^{+}20, PBT^{+}23, REB^{+}20, SMD^{+}20, SRE^{+}21, SWC21, ZAG^{+}20]. sprat [NAPP24]. Sprattus [NAPP24]. spread [IHK^{+}22, PSPP^{+}22, PAH^{+}22]. spring [DZJ^{+}21, GZL^{+}21, RCW^{+}23, ZZL^{+}21]. Sreeja [FM23]. ssDNA [GYN^{+}24]. Stability [UCB^{+}21]. stage [MPF^{+}24, ZYB^{+}23]. stages [BKM21, CVDO^{+}20, MCA^{+}20, TBB^{+}21]. standard [CDJ^{+}23]. Stannius [SAKC20]. Staphylococcus [WWW^{+}21]. starch [MCM^{+}23]. starch-coated [MCM^{+}23]. starter [APK^{+}23]. starting [SZH^{+}22]. starving [BZP^{+}22]. States [KRS^{+}20, KB21]. status [CSGB^{+}21, SM21]. step [PTP^{+}23]. steric [EJH^{+}23]. stimulated [CLH^{+}21]. stimulates [LFA20]. stimulating [ZSL^{+}24a]. stimulation [SMG^{+}23]. stitch [TBK^{+}20]. stock [ŘB24]. stocking [DAW^{+}22, SMP^{+}20]. stomach [ADAO22]. storage [FEH^{+}22]. story [SRN20]. strain [BMSC21, HLH^{+}23b, LS24, LJHH23, YZM^{+}21, ZTY^{+}22]. strains [BTP^{+}24, BMK^{+}20, BCT^{+}20b, CTOH20, KKH^{+}24, LMJ^{+}24, MWH^{+}21, RGS^{+}22, SDMO^{+}20, TNB^{+}22, YCS^{+}21]. stramenopilous [MIK^{+}24]. strategy [ABM^{+}21]. streptococcal [LCTC20]. streptococcosis [KDHS^{+}23, SEL^{+}20]. Streptococcus [HSH^{+}22, MWY^{+}20, MCP^{+}22, NKZ^{+}20, PDL^{+}23, PHS^{+}20, PTK^{+}21, PTP^{+}23, PTW21, PSKW22, PSH^{+}20, PHB^{+}23, Roh22, SWC21, TCS21, XFL^{+}21, ZYH^{+}23, ZHpT^{+}20]. Streptomyces [YJL^{+}23b]. stress [BCT^{+}20a, KKK^{+}23, LDO^{+}21, SZZ^{+}23, SYKC23, UMD22]. stressed [Roh22]. strip [HZFT21, LWT^{+}22, WCT^{+}20, ZZZ^{+}23]. striped [DXD^{+}21, DBC^{+}22, FFB^{+}21, MAK^{+}23, PNØ^{+}21, PØN^{+}21a, PØN^{+}21b, PRC20]. stromal [BBS^{+}20]. strong [FEM^{+}20, MPS^{+}22]. structurally [BGZW22]. structure [BGZW22, HLH^{+}23b, LGH^{+}21, PRW24]. structures [AONDO24]. studies [ASG^{+}22, AMS^{+}24, CDJ^{+}23, DMB^{+}21, FSS^{+}21, JLH^{+}21, KPC^{+}22, MSB^{+}21]. Study [ŘB24, WLK^{+}22, ACGDO^{+}21, AONDO24, BTN^{+}24, CFGCCE21, DBDSP20, JBB^{+}24, KBMD21, MBJ20, ØSH^{+}22, PEś^{+}20, PNØ^{+}21, Roh22, SHS^{+}22a, SPM^{+}20, TASP^{+}20, TR23, WHC^{+}21, WAF^{+}21]. studying [JZW^{+}23, LZC^{+}24]. sturgeon [CMLK^{+}22, GMK^{+}21, MCP^{+}22, PSH^{+}20, RBP^{+}20b]. sturgeons [BZP^{+}22, CVFRM^{+}23]. sub [QGG^{+}21, SDMO^{+}20]. sub-Antarctic [QGG^{+}21, SDMO^{+}20]. subclinical [CGK23, PTW21]. subcutaneous [CSGB^{+}21, CSVG^{+}22, RZTV21]. subdued [RGS^{+}21]. Subinhibitory [OCS^{+}23]. submission [SNSP^{+}23]. subsp [WSW^{+}20, GPC22, PCWC20, WLH^{+}20]. subunit [DZJ^{+}21, GZL^{+}21, JXJ^{+}22, TSV^{+}22, XJZ^{+}22]. success [UMD22]. successful [DGB^{+}22]. suggest [GKTC^{+}21]. suggests [CGA^{+}20, GLS^{+}21]. sulphate [ZMZ^{+}23]. summer [AMM^{+}22]. SUMO [OBSH24]. superoxide [HWC^{+}21]. supplementation [LKP^{+}22, OGR^{+}24, YGR^{+}24]. supply [ŘB24]. surface [VRC^{+}22]. surfactant [WHK^{+}23]. surrounding [SMD^{+}20]. surveillance [DDDJ^{+}20, DMH^{+}21]. survivability [DLS^{+}21]. Survival [ZWB^{+}23, DZQ^{+}22, MR23, PØN^{+}21b, SJL^{+}23, VFR^{+}23, ZVYD23, ZCY^{+}23]. Surviving [IFIZ22]. susceptibilities [PNØ^{+}21]. Susceptibility [JVM^{+}20, LFZ^{+}22, CMLK^{+}22, DPLL^{+}21, FFB^{+}21, GPL^{+}22, GIS^{+}21, IMAH21, IAH22, JW22, LWK20, LZK^{+}23, LGJ21, MLL^{+}22, MBR^{+}22, PØN^{+}21b, SPM^{+}20, ZFC^{+}23]. susceptible [HSH^{+}20, TAK^{+}21]. Sustainable [FM23]. swab [FSFS^{+}20]. swabs [HSS^{+}24]. swamp [BAT^{+}24]. SWCNTs [GZL^{+}21]. SWCNTs-based [GZL^{+}21]. Sweden [PAH^{+}22]. swim [MNPS23]. swimming [LLR^{+}20]. Swiss [VWE^{+}23]. switching [CGA^{+}20]. Switzerland [DBDSP20, SPKS^{+}20]. sympatric [LNK^{+}21]. symptom [SMC^{+}22]. symptomatic [KBG^{+}23]. symptoms [TMM^{+}24]. syncytial [KAS^{+}24]. syndrome [AMM^{+}22, BTN^{+}24, BAT^{+}24, CdPM^{+}21, FMP^{+}21, FMH^{+}22, GVB^{+}21, GSS^{+}21, HLpSC22, JBP24, NCA^{+}20, SPKS^{+}20, SZH^{+}22, SPHM23, SCX^{+}21, SAB^{+}22, TMM^{+}24, TAK^{+}21, VDC20, WBM21, WCT^{+}23b, ZWX^{+}22, ZLZ21, OSM^{+}22b]. syndromes [JSW^{+}21]. Synergistic [NHG^{+}23, VRS^{+}24, WQL^{+}23]. syngnathid [DVR^{+}20]. Synthesis [LZZ^{+}23]. system [DZQ^{+}22, FSS^{+}21, OSM^{+}22b, SCD^{+}23, SFKB^{+}20, TPSJS20a, TPSJS20b, TASP^{+}20]. systematic [MBN^{+}22, PTMC22, ZGL^{+}20]. systematization [DBPH21]. Systemic [PAS^{+}21, TSS^{+}23]. systems [BTN^{+}21, DDHT^{+}22, HBRN24, LDR^{+}22, LEV^{+}20, MHP^{+}20, PDL^{+}23, QRTD^{+}22, SAB^{+}22].

T [YGC^{+}20]. T-cell [YGC^{+}20]. T. [NLR21]. T3SS [LTW^{+}21]. T6SS [YZL^{+}23]. T6SSs [JLH^{+}21]. Tail [NRKK21, JBP24]. Taiwan [BRWC23, CRP^{+}21, CRWC22, CCCC23, LCCC22, NRY^{+}23, NNC^{+}23, PCWC20, RCB^{+}22, SWC21, TSC^{+}23, YTH^{+}24]. takanoi [GXW^{+}23]. Takifugu [xCqXfL^{+}21, FLH^{+}21, KQW^{+}21, LJK^{+}24, QFW^{+}23, SLH^{+}24]. tank [JLM^{+}20]. tank-reared [JLM^{+}20]. tanks [PNØ^{+}21]. Tapeworms [SKO21]. TaqMan [lLWC^{+}24, YTS22]. tarda [AGR^{+}22, RCB^{+}22, SLH^{+}24, WCT^{+}23a]. target [DSD^{+}22, ZDF^{+}23]. targeting [CLF^{+}22, GYN^{+}24, SMP^{+}22b]. targets [CLS^{+}22, QST^{+}21]. Tasmanian [FCH^{+}22b]. taxonomy [LFG^{+}23]. Taylor [RM20]. technique [PDZ^{+}22]. technology [HBRN24]. tegumental [KBMD21]. teleost [DPA^{+}24, KH20].

Temminck [CLC^{+}22, SJK^{+}21, THH^{+}21]. Temperature [UMD22, CLH^{+}21, DHSM^{+}20, FCH^{+}22b, LWR^{+}20, LDO^{+}21, LFG^{+}23, LYC^{+}20, PSED^{+}20, SMY^{+}24, SZZ^{+}23, VSOB22, WBB^{+}21, WSS^{+}20]. temperatures [FEH^{+}22]. template [CGA^{+}20]. temporal [JQT20, LNK^{+}21]. temporarily [OBSH24]. Tenacibaculosis [SNT^{+}22, AHCSCI20, IAH21]. Tenacibaculum [AHCSCI20, AHSCEBI23, EBAH24, IAH21, IMAH21, IAH22, JBB^{+}24, NLR21, SCIAH21, SCALIAH23, SNT^{+}22, VIB^{+}21]. teratomas [PK24]. terminal [WWW^{+}21]. terminalis [ZHpT^{+}20]. test [HZFT21, IPC^{+}24, PNØ^{+}21, WCT^{+}20]. tested [PØN^{+}21b]. Testicular [NCA^{+}20, MPP20a]. testing [CMLK^{+}22, GIS^{+}21]. tests [BLJ^{+}21, GPC22, IAH22, JBG^{+}23]. testudineus [DDH^{+}22]. Tetracapsuloides [LKV^{+}23, OSM22a, SSM^{+}21, WBB^{+}21]. tetrazona [CCP^{+}21]. Thailand [MBR^{+}22, NKZ^{+}20, RBM^{+}22, SKS^{+}24]. their [ASG^{+}22, ASA^{+}20, CTOH20, CLS^{+}22, DWY^{+}20, KKBL21, NNC^{+}23, RBM^{+}22, SPM^{+}20]. them [TAK^{+}21]. therapeutic [ZMZ^{+}22]. theront [YY24]. Thompson [FM23]. thou [BONB20]. threat [GRM^{+}23]. threaten [ZOZ^{+}20]. three [AGH^{+}23, DWY^{+}20, GYC^{+}22, LGJ21, MAK^{+}23, OSM22a, PS23, ZWZ^{+}20]. throughout [MWZ^{+}20]. throughput [LXW^{+}20, WLK^{+}22]. Thunnus [WBV^{+}20]. Thyme [TR23]. thynnus [WBV^{+}20]. thyrsites [MMW^{+}22]. tiamulin [IAH22]. tiger [CCP^{+}21, FMK^{+}21, LJK^{+}24]. Tilapia [SRN20, TSS^{+}23, ALDMLEM20, AMM^{+}22, BHEM21, DDH^{+}22, DDDJ^{+}20, DDHT^{+}22, DDTG^{+}21, DLS^{+}21, DHSK^{+}21, DSD^{+}22, ESEI20, HBRN24, KAS^{+}24, KDS^{+}21, KDHS^{+}23, LCTC20, LHC^{+}22, Li23, MPS^{+}22, MKT^{+}21, NHG^{+}23, NKZ^{+}20, NLL^{+}21, PDL^{+}23, PHS^{+}20, PTMC22, PDG^{+}24, PTK^{+}21, PYT^{+}22, PSKW22, PPWC20, PS23, RCC^{+}21, SLPF^{+}20, SPS23, SRLARG24, SWC21, XFL^{+}21, YTS22, YYS^{+}22, YSW^{+}22, YY24, ZSL^{+}24a, ZZF^{+}23, AONDO24, DDHT^{+}22, DHSK^{+}21, DSD^{+}22, MKT^{+}21, TNB^{+}22, YPT^{+}22]. tilapine [DDHT^{+}22]. tilapinevirus [DHSK^{+}21, MKT^{+}21, YPT^{+}22]. TiLV [TNB^{+}22, DDDJ^{+}20, DDTG^{+}21, KDS^{+}21, TNB^{+}22, TSS^{+}23]. time [FEH^{+}22, GWF^{+}22, GXL^{+}24, HSS^{+}24, HHZ^{+}23, LCW^{+}24, MSB^{+}21, MMW^{+}22, PSED^{+}20, QWL^{+}21, RKC^{+}23, SKR^{+}20, SRLARG24, TCS21]. TiPV [DSD^{+}22]. Tissue [SM21, FYH^{+}21, GPL^{+}22, LZK^{+}23, MMTD^{+}21, MSB^{+}21, SPS23, ZDF^{+}23]. Tissue-specific [SM21]. tissues [HPR^{+}22, LDM^{+}20, LFP^{+}24, NLH^{+}20]. Tisza [SCV^{+}20]. TLR2 [WCT^{+}23a]. TLR2/4 [WCT^{+}23a]. TNF [KQW^{+}21]. TNF- [KQW^{+}21]. Tobago [SBS^{+}22]. TOF [BFvG^{+}23]. tolerance [GWL^{+}21, SYKC23]. Toll [LGH^{+}21, CLH^{+}21]. Toll-like [CLH^{+}21]. tomentosa [NLH^{+}20, SSM^{+}20, SLK^{+}23]. tomont [ZYB^{+}23]. tool [CMM^{+}20, KH20]. tools [DMB^{+}21]. toxicity [BGZW22]. toxicosis [GKU21]. toxin [BGZW22, WCT^{+}20]. Trachinotus [ASG^{+}22, AMS^{+}24, YTH^{+}24]. tract [KSI^{+}20, KH20]. trade [SBS^{+}22]. TRAF6 [SLH^{+}24]. TRAIL [YSW^{+}22]. trait [JØM^{+}20, MRM21, MGG22]. traits [IHK^{+}22]. transcription [KDS^{+}21, SDMO^{+}20]. transcriptional [DWY^{+}20]. Transcriptome [xCqXfL^{+}21, HWG^{+}21, SDH22, XKF^{+}22, ZHL^{+}23, MWH^{+}21, MVW^{+}20, RGS^{+}21, WZC^{+}24, XJM^{+}21]. transcriptomes [MHYW22b]. transcriptomic [Li23]. transfer [WZG^{+}20]. transfrontier [SPHM23]. translocation [LTW^{+}21]. Transmission [YYS^{+}22, BHT^{+}20, CMN^{+}20, HZC^{+}24, KSY^{+}23, LCYL20, LYC^{+}20, MLM^{+}20, PSED^{+}20, QGG^{+}21, RFN^{+}23, SRT^{+}21, SLE^{+}20, SFKB^{+}20]. transmontanus [CMLK^{+}22, GMK^{+}21, PSH^{+}20]. Trauma [HB20]. Trauma-induced [HB20]. treated [ASR^{+}23]. treatment [KKL^{+}22, MMTD^{+}21, OBSH24, VPW^{+}21, WSS^{+}20, HB20]. Treatments [ZWB^{+}23, DLS^{+}21, PTBSF22, WJPS21]. trends [LDR^{+}22]. trh [LWT^{+}22]. trials [SVO^{+}22]. Trichlorfon [dCJBG22, dAMHdOCF^{+}23]. Trichopodus [SAKC20]. triggered [CZC^{+}22]. Trinidad [SBS^{+}22]. triplex [HLW^{+}23]. Triploid [AMB^{+}22]. trituberculatus [LLR^{+}20]. trivalent [ZWZ^{+}20]. tropism [MSB^{+}21, SLE^{+}20]. Trout [öTTS20, AMJ^{+}21, AKR^{+}21, ASF^{+}21, AWF22, BAM^{+}23, BIK^{+}21, BZQV21, BTP^{+}24, BMJ^{+}21, CTP^{+}23, EBAH24, ELB21, FEMMI^{+}23, GVB^{+}21, GSS^{+}21, HSM^{+}22, HHJ^{+}21, JØM^{+}20, JKM^{+}20, KBG^{+}23, KRS^{+}20, LNK^{+}21, LKV^{+}23, LEV^{+}20, LFSB^{+}21, LJHH23, LPW^{+}20, LPL22, MIC^{+}24, OSM^{+}22b, ÖTG22, PYD^{+}22, PPT^{+}20, QGG^{+}21, RFN^{+}23, SPKS^{+}20, SM21, SMG^{+}23, SVP^{+}21, SEL^{+}20, TBKY20, VIB^{+}21, VWE^{+}23, VFR^{+}23, WBB^{+}21, YZM^{+}21]. Trpm4 [LFZ^{+}21]. trumpeter [DBC^{+}22]. trutta [LKV^{+}23, LEV^{+}20, LFSB^{+}21, WBB^{+}21]. tshawytscha [LFP^{+}24, NKS^{+}21, NPP^{+}24, NLPM^{+}21, PPT^{+}20, SST^{+}22]. tumor [BZQV21]. tumour [BBS^{+}20]. tumours [STB^{+}21]. tuna [WBV^{+}20]. turbot [SDH22, WLH^{+}20, WMS21, REB^{+}20]. Turkey [PYD^{+}22]. Turkish [KPC^{+}22, TIK^{+}22]. Türkiye [AOT22]. turtle [CRP^{+}21]. turtles [CXC^{+}24]. twins [YWZ^{+}21]. Two [DDDJ^{+}20, CTOH20, DZQ^{+}22, FEMMI^{+}23, IKT^{+}23, KSY^{+}23, LEV^{+}20, Li23, MWH^{+}21, PK24, RCG22, SDMO^{+}20]. two-component [DZQ^{+}22]. Two-year [DDDJ^{+}20]. type [FdSB^{+}22, PDM^{+}21, SNY^{+}21, ZYZ^{+}20, MQK^{+}20]. Typha [SCX^{+}21]. typing [KKBL21, MSZV^{+}23, QCZ^{+}22].

UhpA [YMC^{+}21]. Ukraine [RBP^{+}20b]. ulcerations [VDD^{+}20]. ulcerative [BTN^{+}24, SPHM23, WBM21]. ultra [WCT^{+}23b]. ultra-sensitive [WCT^{+}23b]. Ultrasonography [CFGCCE21]. Ultrastructural [NKö^{+}21, KBMD21]. underlying [QST^{+}21]. understated [DBDSP20]. unique [TBB^{+}21, XJM^{+}21]. United [KB21, KRS^{+}20]. Unusual [APS23]. Unveiling [WGE^{+}24, MR23]. update [AHTCDI20]. updated [KRS^{+}20]. upon [CMN^{+}20, YCC^{+}21]. upregulates [LFA20]. uptake [AHSB^{+}22, WZC^{+}24]. Urinary [DBC^{+}22]. USA [FM23, ZAG^{+}20, GMK^{+}21, SEL^{+}20, WBM21]. usage [PKF^{+}21]. Use [öTTS20, FSFS^{+}20, LEV^{+}20, LdL20, PNØ^{+}21]. used [JW22, LKD^{+}24, ORV^{+}24]. useful [PDZ^{+}22]. Usefulness [DSD^{+}22]. uses [CZMP21]. using [ADAO22, AHTCDI20, BRWC23, DDTG^{+}21, DMH^{+}21, DGB^{+}22, EJH^{+}23, FCH^{+}22a, FdSB^{+}22, HSS^{+}24, IAH22, NKö^{+}21, ÖTG22, PNAS22, PTP^{+}23, PYT^{+}22, PS23, RCC^{+}21, TCS21, ÚPZ^{+}23, VRC^{+}22, WRMT24, ZLS^{+}21]. Utilisation [ABA^{+}23]. utilization [AHSCEBI23]. UV [TPSJS20b].

V [PDM^{+}21]. V. [HSH^{+}20]. vaccinated [BFC^{+}23, KBG^{+}23]. vaccination [FYH^{+}21, HSM^{+}22, LCYL20, MLHD^{+}23]. vaccine [BMK^{+}20, CLS^{+}22, DZJ^{+}21, GZL^{+}21, JBB^{+}24, LTW^{+}21, LH20, LJHH23, LDW^{+}23, LPGW^{+}22, RBP^{+}20a, SMY^{+}24, SOE^{+}21, SGW^{+}22, XJZ^{+}22, YGZ^{+}22, ZZL^{+}21, ZLY^{+}22, ZSG^{+}23]. Vaccines [FM23, JXJ^{+}22, MKT^{+}21, RTA21, TSV^{+}22, YZM^{+}21, lZWC^{+}20]. Vagococcus [SEL^{+}20, SLE^{+}20, SMP^{+}22b]. Valenciennes [SDMO^{+}20]. value [CSE21]. values [MBN^{+}22]. Vanillic [LXZ^{+}21]. vannamei [BLL23, BTN^{+}21, CMN^{+}20, CCCC23, HLpSC22, LDR^{+}22, LKP^{+}22, LCCC22, LXW^{+}21, LDO^{+}21, LGH^{+}21, MWH^{+}21, QCZ^{+}22, RRK^{+}23, SAB^{+}22, TPSJS20a, TPSJS20b, WSW^{+}20]. Vannella [JKM^{+}20]. vapA [PDM^{+}21]. var [SGR^{+}20a, XKF^{+}22]. varia [BCT^{+}20a]. variability [GLS^{+}21, LPGA^{+}21]. variance [FEM^{+}20]. variant [DGB^{+}22]. variants [RGH^{+}22]. variation [LYC^{+}20, RCB^{+}22]. Variations [PBF^{+}21, ELB21, JQT20, MRM21, RCS^{+}22, Roh22]. varies [MWZ^{+}20]. various [AAZR^{+}20, FMH^{+}22, UCB^{+}21]. VC [CCL^{+}24]. vector [ZSG^{+}23]. vermilion [HB20]. Vermont [BSS^{+}20]. Vermont/Quebec [BSS^{+}20]. Veronaea [CMLK^{+}22, GMK^{+}21]. veronii [CZJ^{+}23, PHB^{+}23, SYKC23, YCS^{+}21, ZJF^{+}22, ZLWL23, ZYL^{+}22]. versus [DŽR^{+}21, KBG^{+}23]. vertebra [JVM^{+}20]. vertebra-derived [JVM^{+}20]. vertebrae [TBK^{+}20]. Vertebral [TBK^{+}20, LFP^{+}24]. Vertebrate [LFA20]. Vertical [RFN^{+}23, MLM^{+}20]. verum [LYX^{+}20]. vesicles [EBELIAH20, EBBIAH21, EBAH24]. veterinary [NTB^{+}24]. vgrG [YZL^{+}23]. VHS [MSB^{+}21]. VHSV [LPW^{+}20]. VI [PDM^{+}21]. via [DXD^{+}21, HWM^{+}24, NGC^{+}23, RKC^{+}23, ZLWL23]. viability [DMB^{+}21, VSOB22]. Vibrio [AEE^{+}23, BTN^{+}21, BFvG^{+}23, BCT^{+}20b, CTOH20, CDY^{+}20, CCL^{+}24, DXL^{+}20b, EBELIAH20, EMH^{+}22, FMK^{+}21, FYH^{+}21, GKTC^{+}21, GWL^{+}21, GZW^{+}23, HSH^{+}20, KKL^{+}22, LXW^{+}21, LJHH23, LDW^{+}23, LXZ^{+}21, LWT^{+}22, MWH^{+}21, NSTS24, OKL^{+}22, RRK^{+}23, Roh22, SZZ^{+}23, SDH22, THW^{+}22, WZC^{+}24, WMS21, WCT^{+}20, WGE^{+}24, ZSL^{+}24a, ZGC^{+}22, ZTY^{+}22, ZDZZ22, ZHL^{+}23, ZLY^{+}22]. Vibrionaceae [RBM^{+}22]. vibrios [ZVYD23]. vibriosis [FMK^{+}21, HSH^{+}20]. vibriosis-infected [FMK^{+}21]. Vietnam [DNN^{+}22, DDH^{+}22, KNG^{+}22, PEN^{+}21, TNB^{+}22]. Violacein [VBG^{+}22]. VipB [SYKC23]. viraemia [DZJ^{+}21, GZL^{+}21, RCW^{+}23, ZZL^{+}21]. Viral [BPA^{+}20, AKR^{+}21, AAZR^{+}20, ASF^{+}21, BAM^{+}23, CLF^{+}22, CCV^{+}24, CLC^{+}22, FAG^{+}22, FZC^{+}23, GYN^{+}24, HAK^{+}20, IHK^{+}22, LL21, LLJ^{+}21, LYF^{+}21, MR23, NRKK21, SMG^{+}23, TBK^{+}22, TBKY20, TASP^{+}20, UCB^{+}21, WXM^{+}22, WWK^{+}20, ZHL^{+}22, ZLS^{+}21, ZMZ^{+}22]. Virginia [WBM21]. virological [ASG^{+}22, AMS^{+}24]. virology [LjY^{+}22]. Virucidal [AAZR^{+}20]. Virulence [ELB21, SOE^{+}21, VGW^{+}21, ASNA24, AKR^{+}21, AGR^{+}22, BCT^{+}20b, BRWC23, CCY^{+}23, CRP^{+}21, CZJ^{+}23, DXL^{+}20b, FVC^{+}21, GPC22, LJHH23, LLR^{+}20, LPGA^{+}21, NRC^{+}23, NLPM^{+}21, RRK^{+}23, RKL^{+}20, SCIAH21, SCAH22, SYKC23, TIK^{+}22, WZC^{+}24, Wel20, WGE^{+}24, YCS^{+}21, YMC^{+}21, ZJF^{+}22, lZWC^{+}20, ZYL^{+}22, ZSL^{+}24b]. virulence-neutralizing [Wel20]. virulence-related [RRK^{+}23, RKL^{+}20]. virulent [BMSC21, LJHH23]. Virus [DHF^{+}22, TNB^{+}22, ABBP24, AMJ^{+}21, AKR^{+}21, AAZR^{+}20, ASF^{+}21, BPA^{+}21, BAM^{+}23, BAT^{+}24, CYX^{+}22, CGA^{+}20, CSDS21, CCV^{+}24, Cow20, DAW^{+}22, DDDJ^{+}20, DDTG^{+}21, DGB^{+}22, DZJ^{+}21, GZL^{+}21, GXL^{+}24, HZFT21, HLpSC22, HSG^{+}21, HAK^{+}20, JXJ^{+}22, KDS^{+}21, KSG^{+}23, KSM^{+}21, KKH^{+}24, LHC^{+}22, Li23, LL21, LGH^{+}21, LCYL20, MPPF21, MPMB^{+}20, MSB^{+}21, MPA^{+}21, MPP^{+}20b, MLM^{+}20, NKö^{+}21, PPDG^{+}22, PSV^{+}21, RCW^{+}23, RGH^{+}22, SPS23, SZH^{+}22, SSMC22, SCX^{+}21, SRN20, SAB^{+}22, SRP^{+}21, TBK^{+}22, TBKY20, TAK^{+}21, TR23, TSS^{+}23, UCB^{+}21, VSODB21, VSOB22, WYR^{+}23, WCT^{+}23b, YYS^{+}22, YLC^{+}21, ZHL^{+}22, ZWX^{+}22, ZZL^{+}21, ZLZ21, ZSG^{+}23, ZOZ^{+}20, DDHT^{+}22, KSD^{+}20]. virus-associated [SRP^{+}21]. virus-infected [YYS^{+}22]. virus-like [Li23]. virus-related [MPA^{+}21]. viruses [DWY^{+}20, GPL^{+}22, LZK^{+}23, MWY^{+}22, MHYW22b, PS23]. visceral [LWR^{+}20]. viscosa [CWB^{+}20b, SSP^{+}23]. visual [CCG24, FCH^{+}22a, KDS^{+}21, PTK^{+}21, WML^{+}21, ZHY^{+}21]. visualization [PSV^{+}21]. visualized [YDM^{+}21]. vitamin [DPA^{+}24]. vitro [ASA^{+}20, AHSCEBI23, BAM^{+}23, CMLK^{+}22, FSS^{+}21, FdSB^{+}22, IAH22, JKM^{+}20, KSM^{+}21, LAH21, LHC^{+}22, LYZ^{+}22, SVP^{+}24, TR23, ZHL^{+}22, YYS^{+}24]. vivo [LYZ^{+}22, SJL^{+}23, ZYH^{+}23, YYS^{+}24]. volumes [MMW^{+}22]. VP2 [TIK^{+}22]. vulnificus [BFvG^{+}23, ZLY^{+}22].

W1 [LPW^{+}20, SMG^{+}23]. Walbaum [AMJ^{+}21, AWF22, BZQV21, BTP^{+}24, GSS^{+}21, HHJ^{+}21, LHH^{+}22, PYD^{+}22]. wall [HCW^{+}20, NRC^{+}23]. warm [SFKB^{+}20]. warm-blooded [SFKB^{+}20]. was [AOT22, DDHT^{+}22]. waste [NSTS24]. Wastes [öTTS20]. Water [PSED^{+}20, TPSJS20b, AEE^{+}23, BLL23, BTN^{+}21, HHZ^{+}23, LWR^{+}20, PTMC22, ŘB24, SMD^{+}20, SPHM23, SJK^{+}21, SNT^{+}22, UCB^{+}21, VSOB22, WBB^{+}21, YPT^{+}22]. waters [MFP21, RSPB22]. weather [ZYZ^{+}20]. Weber [IAO20]. weight [MFP21, PØN^{+}21b]. West [WBM21]. western [JLM^{+}20, MNB^{+}24, WCT^{+}23a, VKK^{+}22]. Westphalian [DHF^{+}22]. Where [BONB20]. which [HCW^{+}20, HWC^{+}21]. whirling [KRS^{+}20]. White [BAT^{+}24, BTN^{+}21, BFH^{+}23, CMLK^{+}22, FFB^{+}21, GMK^{+}21, HCC^{+}22, HLpSC22, LWR^{+}20, LXW^{+}21, PSH^{+}20, RRK^{+}23, SZH^{+}22, SCX^{+}21, SAB^{+}22, TPSJS20a, TPSJS20b, TAK^{+}21, WSW^{+}20, WQL^{+}23, WCT^{+}23b, WGE^{+}24, ZWX^{+}22, ZLZ21, NRKK21]. whitefish [HKB^{+}22]. whiteleg [BLL23, LGH^{+}21]. whitespotted [MAG24, SGU^{+}21]. Whole [HAK^{+}20, IHK^{+}22, CCY^{+}23]. Whole-genome [HAK^{+}20, IHK^{+}22, CCY^{+}23]. wide [CLH^{+}21, DDDJ^{+}20, GLS^{+}21, NNC^{+}23, RCB^{+}22, RKL^{+}20, Roh22]. wild [BPA^{+}20, BFH^{+}23, ELS^{+}23, FdSB^{+}22, HKB^{+}22, JDA^{+}20, KRS^{+}20, KKH^{+}24, LNK^{+}21, MPP20a, MPA^{+}21, PAH^{+}22, PVCG20, QGG^{+}21, SPKS^{+}20, VDC20, ZAG^{+}20]. wild-caught [KRS^{+}20, VDC20]. wild-type [FdSB^{+}22]. wingei [PK24]. Wisconsin [LKD^{+}24, SEL^{+}20]. withdrawal [RKC^{+}23]. within [DDHT^{+}22, IVR^{+}23, PSH^{+}21]. without [MSF^{+}20]. working [FAG^{+}22]. wound [AONDO24, CFGCCE21]. wound-inflicting [AONDO24]. wrasse [PDM^{+}21, PGH^{+}21]. WSSV [CMN^{+}20, TAK^{+}21, ZLZ21]. WSSV-infected [CMN^{+}20]. wuhanensis [KB21].

x [DWK^{+}22]. x-cell [DWK^{+}22]. XIIBs [XQZ^{+}22].

Yamaguti [CRWC22]. year [DDDJ^{+}20]. yeast [RGS^{+}21]. yellow [LWR^{+}20, LLC^{+}21, LCW^{+}24, SJL^{+}23]. yellowfin [LjY^{+}22, LZC^{+}24, ZFC^{+}23]. Yersinia [AABB23, BLJ^{+}21, BAM^{+}23, BIK^{+}21, FEMMI^{+}23, RCG22, RGS^{+}22, SM21, VLSM^{+}20, Wel20, YZM^{+}21]. young [FMP^{+}21].

Z1 [YJL^{+}23b]. Z1-26 [YJL^{+}23b]. Zambezi [SPHM23]. Zambia [BHEM21]. Zealand [JBB^{+}24, JSW^{+}21]. zebra [SHS^{+}22a]. Zebrafish [AHBE^{+}20, MSB^{+}21, NSP20, RGS^{+}21, APK^{+}23, CSE21, CLH^{+}21, CZC^{+}22, CCV^{+}24, CLS^{+}22, FdSB^{+}22, JW22, KSS^{+}20, KMH^{+}21, MVW^{+}20, NLH^{+}20, PKF^{+}21, SSM^{+}20, SKAS^{+}22, SLK^{+}23, SSM^{+}20, WLTZ21]. Zeylanicobdella [MHP^{+}20]. Zimbabwe [SPHM23]. znuA [HWZ^{+}21]. zoeae [MWZ^{+}20]. zoonotic [BFvG^{+}23]. zooplankton [TBB^{+}21].

References

Amin:2023:SAP

[AABB23] Muhamad Amin, Mark B. Adams, Christopher M. Burke, and Christopher J. S. Bolch. Screening and activity of potential gastrointestinal probiotic lactic acid bacteria against Yersinia ruckeri O1b. Journal of Fish Diseases, 46(4):369– 379, April 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Akmal:2023:RGC

[AANY23] Muhammad Akmal, Mami Akatsuka, Issei Nishiki, and Terutoyo Yoshida. Resistance and genomic characterization of a plasmid pkh2101 harbouring erm (B) isolated from emerging fish pathogen Lactococcus garvieae serotype II in Japan. Journal of Fish Diseases, 46(8):841–848, August 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Amtmann:2020:VEV

[AAZR^{+}20] Anette Amtmann, Ibrahim Ahmed, Petra Zahner-Rimmel, Adam Mletzko, Lisa Katharina Jordan, Martin Oberle, Helmut Wedekind, Jürgen Christian, Sven Michael Bergmann, and Anna Maria Becker. Virucidal effects of various agents — including protease — against koi herpesvirus and viral haemorrhagic septicaemia virus. Journal of Fish Diseases, 43(2):185–195, February 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ahmmed:2023:UPD

[ABA^{+}23] Mirja Kaizer Ahmmed, Shuva Bhowmik, Fatema Ahmmed, Stephen G. Giteru, Shikder Saiful Islam, Mayssa Hachem, Md. Ashraf Hussain, Osman N. Kanwugu, Dominic Agyei, and Tom Defoirdt. Utilisation of probiotics for disease management in giant freshwater prawn (Macrobrachium rosenbergii): Administration methods, antagonistic effects and immune response. Journal of Fish Diseases, 46(12):1321– 1336, December 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Abdelsalam:2024:PGC

[ABBP24] Ehdaa Eltayeb Eltigani Abdelsalam, Zuzana Bláhová, Ali Asghar Baloch, and Veronika Piačková. Potential to grow carp oedema virus (genogroup I) in monolayers of carpderived primary cells with further implication in cell analysis. Journal of Fish Diseases, 47(6):e13934:1–e13934:??, June 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Aksnes:2021:GMA

[ABM^{+}21] Ida Aksnes, Stine Braaen, Turhan Markussen, Caroline Piercey Åkesson, Stephane Villoing, and Espen Rimstad. Genetically modified attenuated salmonid alphavirus: a potential strategy for immunization of Atlantic salmon. Journal of Fish Diseases, 44(7):923–937, July 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Armwood:2021:RSP

[ACGDO^{+}21] Abigail R. Armwood, Connie F. Cañete-Gibas, Jennifer A. Dill-Okubo, Nathan P. Wiederhold, and Alvin C. Camus. Retrospective study of phaeohyphomycosis in aquariumhoused fish, with first descriptions of Exophiala lecanii-corni and Neodevriesia cladophorae in fish. Journal of Fish Diseases, 44(10):1563–1577, October 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Austin:2022:API

[ADAO22] Lucinda Austin, Quinton Marco Dos Santos, and Annemariè Avenant-Oldewage. Aspects of the pathology induced by Spinitectus petterae Boomker, 1993 in the stomach of Clarias gariepinus (Burchell, 1822) using light and scanning electron microscopy. Journal of Fish Diseases, 45(6):861–869, June 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Aly:2023:CAV

[AEE^{+}23] Salah M. Aly, Mohamed A. Elatta, Noha I. ElBanna, Mustafa A. El-Shiekh, Mahmoud Mabrok, Mahmoud S. Kelany, and Mohamed Fathi. Comprehensive analysis of Vibrio alginolyticus: Environmental risk factors in the cultured Gilthead seabream (Sparus aurata) under seasonal fluctuations and water parameter alterations. Journal of Fish Diseases, 46(12):1425–1437, December 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Alt:2022:GEA

[AFKK22] Katharina G. Alt, Barbara Feldmeyer, Judith Kochmann, and Sven Klimpel. Gene expression and allergenic potential of Pseudoterranova bulbosa L3 from different infection sites in North Atlantic cod (Gadus morhua). Journal of Fish Diseases, 45(8):1073–1086, August 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Archer:2023:CSL

[AGH^{+}23] Karen R. Archer, Tanja Groll, Richard J. Harvey, Sue M. Thornton, Mark F. Stidworthy, and Daniela Denk. Case series: Lymphoid neoplasia in three elasmobranch species. Journal of Fish Diseases, 46(3):273–279, March 2023. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Armwood:2022:PVE

[AGR^{+}22] Abigail R. Armwood, Matt J. Griffin, Bradley M. Richardson, David J. Wise, Cynthia Ware, and Alvin C. Camus. Pathology and virulence of Edwardsiella tarda, Edwardsiella piscicida, and Edwardsiella anguillarum in channel (Ictalurus punctatus), blue (Ictalurus furcatus), and channel × blue hybrid catfish. Journal of Fish Diseases, 45(11):1683–1698, November 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Avendano-Herrera:2020:ZDR

[AHBE^{+}20] Ruben Avendaño-Herrera, Isabella Benavides, Jaime A. Espina, Daniela Soto-Comte, Matías Poblete-Morales, Juan A. Valdés, Carmen G. Feijóo, and Ariel E. Reyes. Zebrafish (Danio rerio) as an animal model for bath infection by Flavobacterium psychrophilum. Journal of Fish Diseases, 43 (5):561–570, May 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Avendano-Herrera:2020:NSH

[AHCSCI20] Ruben Avendaño-Herrera, Constanza Collarte, Mónica Saldarriaga-Córdoba, and Rute Irgang. New salmonid hosts for Tenacibaculum species: Expansion of tenacibaculosis in Chilean aquaculture. Journal of Fish Diseases, 43(9):1077– 1085, September 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Avendano-Herrera:2022:PAR

[AHSB^{+}22] Ruben Avendaño-Herrera, Pablo Saldivia, Jorn Bethke, Cristian Vargas, and Mauricio Hernández. Proteomic analysis reveals Renibacterium salmoninarum grown under ironlimited conditions induces iron uptake mechanisms and overproduction of the 57-kDa protein. Journal of Fish Diseases, 45(2):289–300, February 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Avendano-Herrera:2023:VPE

[AHSCEBI23] Ruben Avendaño-Herrera, Mónica Saldarriaga-Córdoba, Macarena Echeverría-Bugueño, and Rute Irgang. In vitro phenotypic evidence for the utilization of iron from different sources and siderophores production in the fish pathogen Tenacibaculum dicentrarchi. Journal of Fish Diseases, 46 (9):1001–1012, September 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Avendano-Herrera:2020:SDF

[AHTCDI20] Ruben Avendaño-Herrera, Diana Tapia-Cammas, Eric Duchaud, and Rute Irgang. Serological diversity in Flavobacterium psychrophilum: a critical update using isolates retrieved from Chilean salmon farms. Journal of Fish Diseases, 43(8):877–888, August 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Alencar:2021:MNV

[AKR^{+}21] Anna Luiza Farias Alencar, Se Ryun Kwon, Thomas Bruun Rasmussen, Emilie Mérour, Niels Jørgen Olesen, and Argelia Cuenca. Modifications of the nucleoprotein of viral haemorrhagic septicaemia virus showed gain of virulence in intraperitoneally infected rainbow trout. Journal of Fish Diseases, 44(9):1369–1383, September 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Abdel-Latif:2020:NCC

[ALDMLEM20] Hany M. R. Abdel-Latif, Mahmoud A. O. Dawood, Simon Menanteau-Ledouble, and Mansour El-Matbouli. The nature and consequences of co-infections in tilapia: a review. Journal of Fish Diseases, 43(6):651–664, June 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Aunsmo:2022:TAS

[AMB^{+}22] Arnfinn Aunsmo, Lisbeth Martinsen, Torkjel Bruheim, Mats Martin Sekkelsten-Kindt, Ane Sandtrø, Solveig Gaasø, Stine Braaen, and Espen Rimstad. Triploid Atlantic salmon (Salmo salar) may have increased risk of primary field outbreaks of infectious salmon anaemia. Journal of Fish Diseases, 45(11):1733–1743, November 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Alarcon:2021:ISA

[AMJ^{+}21] Marta Alarcón, Torfinn Moldal, Mona Dverdal Jansen, Maria Aamelfot, Hilde Sindre, Trude M. Lyngstad, and Knut Falk. Infectious salmon anaemia virus detected by RT-qPCR in Norwegian farmed rainbow trout, Oncorhynchus mykiss (Walbaum, 1792). Journal of Fish Diseases, 44(4):479–481, April 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ali:2022:GCA

[AMM^{+}22] Shimaa E. Ali, Osama Mahana, Chadag Vishnumurthy Mohan, Jerome Delamare-Deboutteville, and Mamdouh Y. Elgendy. Genetic characterization and antimicrobial profiling of bacterial isolates collected from Nile tilapia (Oreochromis niloticus) affected by summer mortality syndrome. Journal of Fish Diseases, 45(12):1857–1871, December 2022. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ahmed:2024:DCB

[AMS^{+}24] A. Nafeez Ahmed, S. Mithra, S. Suryakodi, G. Taju, M. J. Abdul Wazith, K. Kanimozhi, V. Rajkumar, A. Badhusha, S. Abdul Majeed, and A. S. Sahul Hameed. Development and characterization of brain cell line from Trachinotus blochii and its application in virological and gene expression studies. Journal of Fish Diseases, 47(5):e13927:1–e13927:??, May 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2020:IIa

[Ano20a] Anonymous. Issue Information. Journal of Fish Diseases, 43(1):i–iv, January 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2020:IIb

[Ano20b] Anonymous. Issue Information. Journal of Fish Diseases, 43 (2):i–iv, February 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2020:IIc

[Ano20c] Anonymous. Issue Information. Journal of Fish Diseases, 43(3):i–iv, March 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2020:IId

[Ano20d] Anonymous. Issue Information. Journal of Fish Diseases, 43(4):ii–v, April 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2020:IIe

[Ano20e] Anonymous. Issue Information. Journal of Fish Diseases, 43(5):ii–v, May 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2020:IIf

[Ano20f] Anonymous. Issue Information. Journal of Fish Diseases, 43(6):i–iv, June 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2020:IIg

[Ano20g] Anonymous. Issue Information. Journal of Fish Diseases, 43(7):i–iv, July 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2020:IIh

[Ano20h] Anonymous. Issue Information. Journal of Fish Diseases, 43(8):i–iv, August 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2020:IIi

[Ano20i] Anonymous. Issue Information. Journal of Fish Diseases, 43 (9):i–iv, September 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2020:IIj

[Ano20j] Anonymous. Issue Information. Journal of Fish Diseases, 43 (10):i–iv, October 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2020:IIk

[Ano20k] Anonymous. Issue Information. Journal of Fish Diseases, 43 (11):i–iv, November 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2020:IIl

[Ano20l] Anonymous. Issue Information. Journal of Fish Diseases, 43 (12):i–iv, December 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2021:CIa

[Ano21a] Anonymous. Cover image. Journal of Fish Diseases, 44(6): i, June 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2021:CIb

[Ano21b] Anonymous. Cover image. Journal of Fish Diseases, 44(9):i, September 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2021:CIc

[Ano21c] Anonymous. Cover image. Journal of Fish Diseases, 44(10): i, October 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2021:IIa

[Ano21d] Anonymous. Issue Information. Journal of Fish Diseases, 44(1):i–iv, January 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2021:IIb

[Ano21e] Anonymous. Issue Information. Journal of Fish Diseases, 44 (2):i–iv, February 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2021:IIc

[Ano21f] Anonymous. Issue Information. Journal of Fish Diseases, 44(3):i–iv, March 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2021:IId

[Ano21g] Anonymous. Issue Information. Journal of Fish Diseases, 44(4):i–iv, April 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2021:IIe

[Ano21h] Anonymous. Issue Information. Journal of Fish Diseases, 44(5):i–iv, May 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2021:IIf

[Ano21i] Anonymous. Issue Information. Journal of Fish Diseases, 44(6):i–iv, June 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2021:IIg

[Ano21j] Anonymous. Issue Information. Journal of Fish Diseases, 44(7):i–iv, July 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2021:IIh

[Ano21k] Anonymous. Issue Information. Journal of Fish Diseases, 44(8):i–iv, August 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2021:IIi

[Ano21l] Anonymous. Issue Information. Journal of Fish Diseases, 44 (9):ii–iv, September 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2021:IIj

[Ano21m] Anonymous. Issue Information. Journal of Fish Diseases, 44 (10):ii–iv, October 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2021:IIk

[Ano21n] Anonymous. Issue Information. Journal of Fish Diseases, 44 (11):i–iv, November 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2021:IIl

[Ano21o] Anonymous. Issue Information. Journal of Fish Diseases, 44 (12):i–iv, December 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2022:CI

[Ano22a] Anonymous. Cover image. Journal of Fish Diseases, 45(1): i, January 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2022:IIa

[Ano22b] Anonymous. Issue information. Journal of Fish Diseases, 45 (1):ii–iv, January 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2022:IIb

[Ano22c] Anonymous. Issue information. Journal of Fish Diseases, 45 (2):i–iv, February 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2022:IIc

[Ano22d] Anonymous. Issue information. Journal of Fish Diseases, 45(3):i–iv, March 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2022:IId

[Ano22e] Anonymous. Issue information. Journal of Fish Diseases, 45(4):i–iv, April 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2022:IIe

[Ano22f] Anonymous. Issue information. Journal of Fish Diseases, 45(5):i–iv, May 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2022:IIf

[Ano22g] Anonymous. Issue information. Journal of Fish Diseases, 45(6):ii–iv, June 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2022:IIg

[Ano22h] Anonymous. Issue information. Journal of Fish Diseases, 45(7):i–iv, July 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2022:IIh

[Ano22i] Anonymous. Issue information. Journal of Fish Diseases, 45(8):i–iv, August 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2022:IIi

[Ano22j] Anonymous. Issue information. Journal of Fish Diseases, 45 (9):ii–iv, September 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2022:IIj

[Ano22k] Anonymous. Issue information. Journal of Fish Diseases, 45 (10):i–iv, October 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2022:IIk

[Ano22l] Anonymous. Issue information. Journal of Fish Diseases, 45 (11):i–iv, November 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2022:IIl

[Ano22m] Anonymous. Issue information. Journal of Fish Diseases, 45 (12):i–iv, December 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2023:IIa

[Ano23a] Anonymous. Issue information. Journal of Fish Diseases, 46 (1):i–iv, January 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2023:IIb

[Ano23b] Anonymous. Issue information. Journal of Fish Diseases, 46 (2):i–iv, February 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2023:IIc

[Ano23c] Anonymous. Issue information. Journal of Fish Diseases, 46(3):i–iv, March 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2023:IId

[Ano23d] Anonymous. Issue information. Journal of Fish Diseases, 46(4):i–iv, April 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2023:IIe

[Ano23e] Anonymous. Issue information. Journal of Fish Diseases, 46(5):i–iv, May 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2023:IIf

[Ano23f] Anonymous. Issue information. Journal of Fish Diseases, 46(6):i–iv, June 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2023:IIg

[Ano23g] Anonymous. Issue information. Journal of Fish Diseases, 46(7):i–iv, July 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2023:IIh

[Ano23h] Anonymous. Issue information. Journal of Fish Diseases, 46(8):i–iv, August 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2023:IIi

[Ano23i] Anonymous. Issue information. Journal of Fish Diseases, 46 (9):i–iv, September 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2023:IIj

[Ano23j] Anonymous. Issue information. Journal of Fish Diseases, 46 (10):i–iv, October 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2023:IIk

[Ano23k] Anonymous. Issue information. Journal of Fish Diseases, 46 (11):i–iv, November 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2023:IIl

[Ano23l] Anonymous. Issue information. Journal of Fish Diseases, 46 (11):i–iv, November 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2023:IIm

[Ano23m] Anonymous. Issue information. Journal of Fish Diseases, 46 (12):i–iv, December 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2024:IIa

[Ano24a] Anonymous. Issue information. Journal of Fish Diseases, 47(1):e13802:1–e13802:??, January 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2024:IIb

[Ano24b] Anonymous. Issue information. Journal of Fish Diseases, 47(2):e13803:1–e13803:??, February 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2024:IIc

[Ano24c] Anonymous. Issue information. Journal of Fish Diseases, 47(3):e13804:1–e13804:??, March 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2024:IId

[Ano24d] Anonymous. Issue information. Journal of Fish Diseases, 47 (4):e13805:1–e13805:??, April 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2024:IIe

[Ano24e] Anonymous. Issue information. Journal of Fish Diseases, 47 (5):e13806:1–e13806:??, May 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2024:IIf

[Ano24f] Anonymous. Issue information. Journal of Fish Diseases, 47 (6):e13807:1–e13807:??, June 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Avenant-Oldewage:2024:PCI

[AONDO24] Annemariè Avenant-Oldewage, Kazuya Nagasawa, Quinton M. Dos Santos, and Willie H. Oldewage. Pathology caused by introduced Neoergasilus japonicus (Copepoda: Ergasilidae) to the skin of indigenous Tilapia sparrmanii in South Africa and scanning electron microscopy study of wound-inflicting structures. Journal of Fish Diseases, 47(1): e13867:1–e13867:??, January 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Altinok:2022:NAR

[AOT22] Ilhan Altinok, Rafet Cagri Ozturk, and Mustafa Ture. NGS analysis revealed that Lactococcus garvieae Lg-Per was Lactococcus petauri in Türkiye. Journal of Fish Diseases, 45 (12):1839–1843, December 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Antinero:2023:RSD

[APK^{+}23] Ariel Antinero, Alice Printzi, Chara Kourkouta, Stefanos Fragkoulis, David Mazurais, Jose Luis Zambonino-Infante, and George Koumoundouros. The role of starter diets in the development of skeletal abnormalities in zebrafish Danio rerio (Hamilton, 1822). Journal of Fish Diseases, 46(6):697– 705, June 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Adamek:2023:USI

[APS23] Zdeněk Adámek, Miroslava Palíková, and Tomáš Scholz. Unusual sites of infection of Ligula intestinalis (Cestoda) plerocercoids in an intermediate host, freshwater bream Abramis brama. Journal of Fish Diseases, 46(7):791–793, July 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ahmed:2020:VAA

[ASA^{+}20] Fatma Ahmed, Faiza M. Soliman, Mohamed A. Adly, Hamdy A. M. Soliman, Mansour El-Matbouli, and Mona Saleh. In vitro assessment of the antimicrobial efficacy of chitosan nanoparticles against major fish pathogens and their cytotoxicity to fish cell lines. Journal of Fish Diseases, 43 (9):1049–1063, September 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ashfaq:2021:KCL

[ASF^{+}21] Hassan Ashfaq, Hatem Soliman, Sabine Fajmann, Veronika Sexl, Mansour El-Matbouli, and Mona Saleh. Kinetics of CD4-1+ lymphocytes in brown trout after exposure to viral haemorrhagic septicaemia virus. Journal of Fish Diseases, 44(10):1553–1562, October 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Abdul:2022:DCF

[ASG^{+}22] Nafeez Ahmed Abdul, Abdul Majeed Seepoo, Taju Gani, Vimal Sugumar, Suryakodi Selvam, Badhusha Allahbagash, Abdul Nazar Abdul kuthoos, Ramesh Kumar Palsamy, Paknikar Kishore M., Jyutika M. Rajwade, and Sait Sahul Hameed Azeez. Development and characterization of five novel cell lines from snubnose pompano, Trachinotus blochii (Lacepede, 1801), and their application in gene expression and virological studies. Journal of Fish Diseases, 45(1):121–139, January 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Aboyadak:2024:RAG

[ASNA24] Ibrahim Mohamed Aboyadak, Magdy Khalil Soliman, Hadeer Mohamed Nageeb, and Nadia Gabr Ali. The role of Aeromonas genotyping in virulence for Dicentrarchus labrax.

Journal of Fish Diseases, 47(2):e13878:1–e13878:??, February 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ali:2023:BRE

[ASR^{+}23] Shandana Ali, Syed Ata Ur Rahman Shah, Muhammad Rauf, Maizom Hassan, Waheed Ullah, and Farman Ullah Dawar. Bactericidal role of epidermal mucus of freshwater fish treated with Aeromonas hydrophila. Journal of Fish Diseases, 46(11):1225–1237, November 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Avila:2022:DRF

[AWF22] Brian W. Avila, Dana L. Winkelman, and Eric R. Fetherman. Dual resistance to Flavobacterium psychrophilum and Myxobolus cerebralis in rainbow trout (Oncorhynchus mykiss, Walbaum). Journal of Fish Diseases, 45(6):801– 813, June 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bauer:2023:VMI

[BAM^{+}23] Julia Bauer, Mikolaj Adamek, Anne-Carina Miebach, Jakob Gährken, Stephan Wessels, Jens Tetens, Carsten Dietz, Angela Sünder, Marek Matras, Magdalena Stachnik, Michal Reichert, and Dieter Steinhagen. In vitro modelling of the influence of alternative feeds (Hermetia illucens, Arthrospira platensis) on the resistance of different rainbow trout populations (Oncorhynchus mykiss) against the viral haemorrhagic septicaemia virus and Yersinia ruckeri. Journal of Fish Diseases, 46(11):1269–1283, November 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Badhusha:2022:FRO

[BAS^{+}22] Allahbagash Badhusha, Abdul Nafeez Ahmed, Selvam Suryakodi, Mohamed Jaffer Abdul Wazith, Sivaraj Mithra, Kumarasamy Kanimozhi, Seepoo Abdul Majeed, Gani Taju, and Azeez Sait Sahul Hameed. First report on the occurrence of cyprinid herpesvirus 3 in koi carp (Cyprinus carpio koi) in India. Journal of Fish Diseases, 45(8):1087–1098, August 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bruce:2024:WSS

[BAT^{+}24] Timothy J. Bruce, Jason W. Abernathy, Nicole Tripp, Nicholas Barnes, Courtney E. Harrison, Abdulmalik A. Oladipupo, Justin D. Krol, Allison L. Wise, Janet V. Warg, and James A. Stoeckel. White spot syndrome virus in Alabama red swamp crayfish (Procambarus clarkii). Journal of Fish Diseases, 47(2):e13873:1–e13873:??, February 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bekavac:2022:DII

[BBD^{+}22] Ana Bekavac, Ana Beck, Paula Dragičević, Zrinka Dragun, Ivana Maguire, Dušica Ivanković, Željka Fiket, Romana Gračan, and Sandra Hudina. Disturbance in invasion? Idiopathic necrotizing hepatopancreatitis in the signal crayfish Pacifastacus leniusculus (Dana, 1852) in Croatia. Journal of Fish Diseases, 45(2):261–276, February 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Branco:2021:FDM

[BBG21] Ciara E. Branco, Robert Chad Boyce, and David T. Gauthier. Feeding duration of Myzobdella lugubris leech on largemouth bass (Micropterus salmoides). Journal of Fish Diseases, 44(3):355–358, March 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Budai:2020:OMG

[BBS^{+}20] Jane Budai, Paola Beraldo, Rubina Sirri, Rosanna Panebianco, Giuseppe Sarli, and Marco Galeotti. Ovarian mixed germ cell–sex cord–stromal tumour in a European seabass, Dicentrarchus labrax (Linnaeus, 1758). Journal of Fish Diseases, 43(11):1453–1457, November 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Box:2020:PMI

[BCT^{+}20a] Antonio Box, Xavier Capó, Silvia Tejada, Antoni Sureda, Laia Mejías, and José María Valencia. Perkinsus mediterraneus infection induces oxidative stress in the mollusc Mimachlamys varia. Journal of Fish Diseases, 43(1):1–7, January 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bunpa:2020:OMP

[BCT^{+}20b] Supansa Bunpa, Nattarika Chaichana, Jade L. L. Teng, Hwei Huih Lee, Patrick C. Y. Woo, Decha Sermwittayawong, Nongyao Sawangjaroen, and Natthawan Sermwittayawong. Outer membrane protein A (OmpA) is a potential virulence factor of Vibrio alginolyticus strains isolated from diseased fish. Journal of Fish Diseases, 43(2):275–284, February 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bergmann:2021:SRK

[BDK^{+}21] Sven Michael Bergmann, Juliane Dabels, Sandro Klafack, Yeonwha Jin, Pei-Yu Lee, Arndt Christian Hofmann, Yingying Wang, Qing Wang, Yingying Li, Weiwei Zeng, Angela Lusiastuti, Shucheng Zheng, Yijuan Jin, Jolanta Kiełpińska, and Sean Monaghan. Serological responses to koi herpesvirus (KHV) in a non-cyprinid reservoir host. Journal of Fish Diseases, 44(8):1229–1236, August 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bustos:2023:IRI

[BFC^{+}23] Paulina Bustos, Carolina Figueroa, Bárbara Cádiz, Tamara Santander, Brian Dixon, José A. Gallardo, and Pablo Conejeros. Immune response induced by coinfection of the sea louse Caligus rogercresseyi and the intracellular bacteria Piscirickettsia salmonis in vaccinated Atlantic salmon. Journal of Fish Diseases, 46(12):1337–1342, December 2023. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Brimsholm:2023:RMF

[BFH^{+}23] Malin Brimsholm, Per Gunnar Fjelldal, Tom Hansen, Thomas Kenneth William Fraser, Monica Solberg, Kevin Glover, Erling Olaf Koppang, and Håvard Bjørgen. Red and melanized focal changes in white skeletal muscle in Atlantic salmon (Salmo salar): Comparative analysis of farmed, wild and hybrid fish reared under identical conditions. Journal of Fish Diseases, 46(12):1377–1389, December 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Boonstra:2023:FAI

[BFvG^{+}23] Mirjam Boonstra, Belén Fouz, Betty van Gelderen, Inger Dalsgaard, Lone Madsen, Eva Jansson, Carmen Amaro, and Olga Haenen. Fast and accurate identification by MALDI-TOF of the zoonotic serovar E of Vibrio vulnificus linked to eel culture. Journal of Fish Diseases, 46(4):445–452, April 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bao:2022:CTS

[BGZW22] Shuaiming Bao, Shuangshuang Gao, Mengmeng Zhang, and Yongjie Wang. Characterization of toxicity and structure of PirAB^{vc}-like proteins that are structurally almost identical to shrimp AHPND-causing PirAB toxin. Journal of Fish Diseases, 45(2):315–326, February 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bwalya:2021:LGI

[BHEM21] Patricia Bwalya, Bernard M. Hang’ombe, Øystein Evensen, and Stephen Mutoloki. Lactococcus garvieae isolated from Lake Kariba (Zambia) has low invasive potential in Nile tilapia (Oreochromis niloticus). Journal of Fish Diseases, 44 (6):721–727, June 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bridle:2021:EEL

[BHS^{+}21] Andrew R. Bridle, Thomas Hill, Aaron Smith, Philip Crosbie, and Barbara F. Nowak. Experimental exposure to low concentrations of Neoparamoeba perurans induces amoebic gill disease in Atlantic salmon. Journal of Fish Diseases, 44 (7):1025–1031, July 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Buss:2020:RTB

[BHT^{+}20] Jessica Jamuna Buss, James Owen Harris, Jason Elliot Tanner, Kathryn Helen Wiltshire, and Marty Robert Deveney. Rapid transmission of Bonamia exitiosa by cohabitation causes mortality in Ostrea angasi. Journal of Fish Diseases, 43(2):227–237, February 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bilen:2021:EGJ

[BIK^{+}21] Soner Bilen, Selcuk Ispir, Osman Nezih Kenanoglu, Yiğit Taştan, Kerim Güney, and Ertugrul Terzi. Effects of Greek juniper (Juniperus excelsa) extract on immune responses and disease resistance against Yersinia ruckeri in rainbow trout (Oncorhynchus mykiss). Journal of Fish Diseases, 44(6): 729–738, June 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bracamonte:2021:EAC

[BKM21] Seraina E. Bracamonte, Klaus Knopf, and Michael T. Monaghan. Encapsulation of Anguillicola crassus reduces the abundance of adult parasite stages in the European eel (Anguilla anguilla). Journal of Fish Diseases, 44(6):771–782, June 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Baron:2021:ABC

[BLJ^{+}21] Sandrine Baron, Emeline Larvor, Eric Jouy, Isabelle Kempf, Sophie Le Bouquin, Claire Chauvin, Pierre-Marie Boitard, Matthieu Jamin, Alain Le Breton, Benoit Thuillier, and Peter Smith. Agreement between the categorization of isolates of Aeromonas salmonicida and Yersinia ruckeri by disc diffusion and MIC tests performed at 22^{o}C. Journal of Fish Diseases, 44(7):979–985, July 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ballantyne:2023:FIH

[BLL23] Rolissa Ballantyne, Jai-Wei Lee, and Chun-Hung Liu. First identification and histopathological analysis of Lactococcus garvieae infection in whiteleg shrimp, Penaeus vannamei cultured in low salinity water. Journal of Fish Diseases, 46(9):929–942, September 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Brown:2020:FSH

[BME^{+}20] Tori Brown, Zach Millar, Drake Evans, Phuc H. Pham, Veronique LePage, and John S. Lumsden. Fusarium solani haplotype 12-b and aortic and branchial arteritis in Hippocampus erectus Perry. Journal of Fish Diseases, 43(2): 301–304, February 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bruce:2021:AFP

[BMJ^{+}21] Timothy J. Bruce, Jie Ma, Evan M. Jones, Brent M. Vuglar, Luke P. Oliver, Christopher Knupp, Thomas P. Loch, and Kenneth D. Cain. Assessment of Flavobacterium psychrophilum-associated mortality in Atlantic salmon (Salmo salar) and brook trout (Salvelinus fontinalis). Journal of Fish Diseases, 44(5):645–653, May 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bruce:2020:CPL

[BMK^{+}20] Timothy J. Bruce, Jie Ma, Christopher Knupp, Thomas P. Loch, Mohamed Faisal, and Kenneth D. Cain. Crossprotection of a live-attenuated Flavobacterium psychrophilum immersion vaccine against novel Flavobacterium spp. and Chryseobacterium spp. strains. Journal of Fish Diseases, 43 (8):915–928, August 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bruce:2020:IEC

[BMO^{+}20] Timothy J. Bruce, Jie Ma, Luke P. Oliver, Evan M. Jones, Benjamin R. LaFrentz, and Kenneth D. Cain. Isolation and experimental challenge of cultured burbot (Lota lota maculosa) with Flavobacterium columnare and Aeromonas sp. isolates. Journal of Fish Diseases, 43(8):839–851, August 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bruce:2021:QCG

[BMSC21] Timothy J. Bruce, Jie Ma, Ponnerassery S. Sudheesh, and Kenneth D. Cain. Quantification and comparison of gene expression associated with iron regulation and metabolism in a virulent and attenuated strain of Flavobacterium psychrophilum. Journal of Fish Diseases, 44(7):949–960, July 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bui:2020:WAT

[BONB20] Samantha Bui, Frode Oppedal, Velimir Nola, and Luke T. Barrett. Where art thou louse? A snapshot of attachment location preferences in salmon lice on Atlantic salmon hosts in sea cages. Journal of Fish Diseases, 43(6):697–706, June 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Boukedjouta:2020:VER

[BPA^{+}20] Rachid Boukedjouta, Tobia Pretto, Miriam Abbadi, Lorena Biasini, Anna Toffan, and Karim Mezali. Viral encephalopathy and retinopathy is endemic in wild groupers (genus Epinephelus spp.) of the Algerian coast. Journal of Fish Diseases, 43(7):801–812, July 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Baud:2021:GDC

[BPA^{+}21] Marine Baud, Laurane Pallandre, Fabrice Almeras, Loeiz Maillet, David Stone, and Laurent Bigarré. Genetic diversity of the carp oedema virus in France. Journal of Fish Diseases, 44(10):1531–1542, October 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Behrens:2023:CHL

[BRCI23] Jane W. Behrens, Marie Plambech Ryberg, Virginia Chondromatidou, and Tine Moesgaard Iburg. Comparative histopathology of livers from Baltic cod infected with the parasitic nematode Contracaecum osculatum. Journal of Fish Diseases, 46(6):653–662, June 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Brimsholm:2023:DMR

[BRR^{+}23] Malin Brimsholm, Line Rønning, Espen Rimstad, Erling Olaf Koppang, and Håvard Bjørgen. Diffuse melanization of the red skeletal musculature in farmed Atlantic salmon (Salmo salar L.). Journal of Fish Diseases, 46(4): 453–458, April 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Byadgi:2023:CGE

[BRWC23] Omkar Vijay Byadgi, Atiek Rahmawaty, Pei-Chi Wang, and Shih-Chu Chen. Comparative genomics of Edwardsiella anguillarum and Edwardsiella piscicida isolated in Taiwan enables the identification of distinctive features and potential virulence factors using Oxford-Nanopore MinION(R) sequencing. Journal of Fish Diseases, 46(4):287–297, April 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Blazer:2020:MMB

[BSS^{+}20] Vicki S. Blazer, Cassidy H. Shaw, Cheyenne R. Smith, Peter Emerson, and Thomas Jones. Malignant melanoma of brown bullhead (Ameiurus nebulosus) in Lake Memphremagog, Vermont/Quebec. Journal of Fish Diseases, 43(1): 91–100, January 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bauer:2021:IRW

[BTN^{+}21] Julia Bauer, Felix Teitge, Lisa Neffe, Mikolaj Adamek, Arne Jung, Christina Peppler, Dieter Steinhagen, and Verena Jung-Schroers. Impact of a reduced water salinity on the composition of Vibrio spp. in recirculating aquaculture systems for Pacific white shrimp (Litopenaeus vannamei) and its possible risks for shrimp health and food safety. Journal of Fish Diseases, 44(1):89–105, January 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bruce:2024:CSE

[BTN^{+}24] Timothy J. Bruce, Carl C. Trettin, Zachary A. Noel, Alex T. Chow, Kylie Warden, Craig Roghair, and Troy M. Farmer. A case study of epizootic ulcerative syndrome (EUS) caused by Aphanomyces invadans in eastern mosquitofish (Gambusia holbrooki) from the headwaters of Charleston Harbor, South Carolina. Journal of Fish Diseases, 47(3):e13895:1– e13895:??, March 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Brocca:2024:INA

[BTP^{+}24] Ginevra Brocca, Alessandro Truant, Hana Peckova, Martina Lisnerová, Alberto Perolo, Marialetizia Fioravanti, Ivan Fiala, Gianfranco Gabai, Francesco Quaglio, and Andrea Gustinelli. Identification of new amoebae strains in rainbow trout (Oncorhynchus mykiss, Walbaum) farms affected by nodular gill disease (NGD) in Northeastern Italy. Journal of Fish Diseases, 47(6):e13933:1–e13933:??, June 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Buss:2020:IDB

[BWH^{+}20] Jessica Jamuna Buss, Kathryn Helen Wiltshire, James Owen Harris, Jason Elliot Tanner, and Marty Robert Deveney.

Infection dynamics of Bonamia exitiosa on intertidal Ostrea angasi farms. Journal of Fish Diseases, 43(3):359–369, March 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Brocca:2022:SGA

[BZP^{+}22] Ginevra Brocca, Samuele Zamparo, Tobia Pretto, Alessandro Calore, Andrea Marsella, Romy Lucon Xiccato, Matteo Cornaggia, Luana Cortinovis, Luca Bano, Anna Toffan, Francesco Quaglio, and Ranieri Verin. Severe gastroenteropathy associated with Clostridium perfringens isolation in starving juvenile sturgeons. Journal of Fish Diseases, 45 (3):471–477, March 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Brocca:2021:MMN

[BZQV21] Ginevra Brocca, Samuele Zamparo, Francesco Quaglio, and Ranieri Verin. Metastatic myxoid nerve sheath tumor of the dorsal fin in a rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 44(11):1875–1878, November 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cheng:2023:IGC

[CCCC23] Yun-Han Cheng, Yen-Chen Chang, Chia-Yu Chang, and Hui-Wen Chang. Identification and genomic characterization of Baculovirus penaei in Litopenaeus vannamei in Taiwan. Journal of Fish Diseases, 46(6):611–617, June 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chen:2024:DAR

[CCG24] Xi Chen, Hua Chen, and Jun-Qing Ge. Development and application of a recombinase-aided amplification combined with a lateral flow dipstick assay for rapid visual detection of anguillid herpesvirus 1. Journal of Fish Diseases, 47(4): e13907:1–e13907:??, April 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Choi:2024:FRP

[CCL^{+}24] Hee-Jae Choi, Da-Yeon Choi, Ji-Hoon Lee, Jun-Hwan Kim, and Yue Jai Kang. First report and pathogenicity of Vibrio campbellii (VC_{AHPND}) isolated in South Korea. Journal of Fish Diseases, 47(5):e13928:1–e13928:??, May 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chang:2021:PEP

[CCP^{+}21] Erica K. Chang, Alvin C. Camus, Eric Pomaranski, Zeinab Yazdi, and Esteban Soto. Pathogenesis of Erysipelothrix piscisicarius infection in tiger barbs (Puntigrus tetrazona). Journal of Fish Diseases, 44(11):1681–1688, November 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cho:2024:ICB

[CCV^{+}24] Se-Hyun Cho, Se-Young Cho, Bipin Vaidya, Joseph Kwon, and Duwoon Kim. Influence of colour background on antiviral activity against viral haemorrhagic septicaemia virus in zebrafish regulated by circadian rhythm signalling pathway. Journal of Fish Diseases, 47(2):e13887:1–e13887:??, February 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cao:2023:PVF

[CCY^{+}23] Shoulin Cao, Jiaojiao Chang, Xiaozhen Yue, Jinnian Li, and Xuelan Liu. Potential virulence factors of Nocardia seriolae AHLQ20-01 based on whole-genome analysis and its pathogenicity to largemouth bass (Micropterus salmoides). Journal of Fish Diseases, 46(4):333–345, April 2023. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cherif:2023:ASR

[CDJ^{+}23] Emira Cherif, Theo Deremarque, Fabienne Justy, Rodolphe Elie Gozlan, and Marine Combe. Assessing the specificity of the Rosette agent DNA amplification: An optimized protocol for the detection of standard DNA among studies. Journal of Fish Diseases, 46(2):177–180, February 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Costa:2021:PCS

[CdPM^{+}21] Janina Z. Costa, Jorge del Pozo, Kevin McLean, Neil Inglis, Philippe Sourd, Andrei Bordeianu, and Kim D. Thompson. Proteomic characterization of serum proteins from Atlantic salmon (Salmo salar L.) from an outbreak with cardiomyopathy syndrome. Journal of Fish Diseases, 44(11):1697– 1709, November 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chieng:2020:MRB

[CDY^{+}20] Catherine Cheng Yun Chieng, Hassan Mohd Daud, Fatimah Md Yusoff, Kim Dawn Thompson, and Maha Abdullah. Mucosal responses of brown-marbled grouper Epinephelus fuscoguttatus (Forsskål, 1775) following intraperitoneal infection with Vibrio harveyi. Journal of Fish Diseases, 43(10):1249–1258, October 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ceballos-Francisco:2021:USS

[CFGCCE21] Diana Ceballos-Francisco, Nuria García-Carrillo, Alberto Cuesta, and María Ángeles Esteban. Ultrasonography study of the skin wound healing process in gilthead seabream (Sparus aurata). Journal of Fish Diseases, 44(8):1091–1100, August 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cardenas:2020:RSM

[CGA^{+}20] Matías Cárdenas, Claudia Galleguillos, Karina Acevedo, Catarina Ananias, Javiera Alarcón, Sofía Michelson, Jorge Toledo, Margarita Montoya, Claudio Meneses, Eduardo Castro-Nallar, Yesseny Vásquez-Martínez, and Marcelo Cortez-San Martín. Rapid sequence modification in the highly polymorphic region (HPR) of the hemagglutinin gene of the infectious salmon anaemia virus (ISAV) suggests intrasegmental template switching recombination. Journal of Fish Diseases, 43(12):1483–1496, December 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chew:2023:HCH

[CGK23] Xian Zhe Chew and Susan Gibson-Kueh. The haematology of clinically healthy, farmed juvenile Asian seabass (Lates calcarifer Bloch) — reference intervals, and indicators of subclinical disease. Journal of Fish Diseases, 46(10):1109– 1124, October 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Choi:2022:IIP

[CLC^{+}22] Hee Jae Choi, Ji-Hoon Lee, Da Yeon Choi, Ji Hee Kim, Jin Man Cho, Hyuk Chu Kwon, and Yue Jai Kang. Innate immunity and protective effects of orally administered betaine against viral and bacterial diseases in the olive flounder Paralichthys olivaceus (Temminck &n Schlegel). Journal of Fish Diseases, 45(11):1789–1798, November 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chen:2022:MMK

[CLF^{+}22] Qikang Chen, Yang Luo, Yan Fu, Zizhao Feng, Liqun Lu, Yousheng Jiang, and Dan Xu. microRNA (miR-KT-635) encoded by cyprinid herpesvirus 2 regulates the viral replication with targeting to the ORF23. Journal of Fish Diseases, 45(5):631–639, May 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chen:2021:GWA

[CLH^{+}21] Hong Chen, Yue Liang, Yawen Han, Tengfei Liu, and Shulin Chen. Genome-wide analysis of Toll-like receptors in zebrafish and the effect of rearing temperature on the receptors in response to stimulated pathogen infection. Journal of Fish Diseases, 44(3):337–349, March 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Choe:2022:CEP

[CLS^{+}22] Yunjeong Choe, Donghee Lee, Minji Seong, Ju Bin Yoon, Jun Hyeok Yang, Jin-Young Yang, Ki Hwan Moon, and Ho Young Kang. Characterization of Edwardsiella piscicida CK108 flagellin genes and evaluation of their potential as vaccine targets in the zebrafish model. Journal of Fish Diseases, 45(2):249–259, February 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Coleman:2022:VST

[CMLK^{+}22] Denver Coleman, Beatriz Martínez-López, Heather Knych, Susan Yun, Kirsten Kenelty, Valentina Tomasi, and Esteban Soto. In vitro susceptibility testing for the emerging pathogenic mould Veronaea botryosa and pharmacokinetic parameters of natamycin in white sturgeon (Acipenser transmontanus). Journal of Fish Diseases, 45(11):1623–1633, November 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cano:2020:NLL

[CMM^{+}20] Irene Cano, Robin McCullough, Brian Mulhearn, Susie Gunning, Ava Waine, Claire Joiner, and Richard Paley. Nonlethal loop-mediated isothermal amplification assay as a point-of-care diagnostics tool for Neoparamoeba perurans, the causative agent of amoebic gill disease. Journal of Fish Diseases, 43(7):779–790, July 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Caro:2020:ATR

[CMN^{+}20] Luis Fernando Aranguren Caro, Hung N. Mai, Linda Nunan, Joshua Lin, Brenda Noble, and Arun K. Dhar. Assessment of transmission risk in WSSV-infected shrimp Litopenaeus vannamei upon cooking. Journal of Fish Diseases, 43(4): 403–411, April 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cowley:2020:MVP

[Cow20] Jeff A. Cowley. Mourilyan virus pathogenicity in kuruma shrimp (Penaeus japonicus). Journal of Fish Diseases, 43 (11):1401–1407, November 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Charles:2021:DSQ

[CQG^{+}21] Maud Charles, Yann Quesnelle, Guillaume Glais, Suzanne Trancart, and Maryline Houssin. Development of a semiquantitative PCR assay for the detection of Francisella halioticida and its application to field samples. Journal of Fish Diseases, 44(8):1169–1177, August 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cheng:2021:GVG

[CRP^{+}21] Li-Wu Cheng, Shreesha Rao, Sayuj Poudyal, Pei-Chi Wang, and Shih-Chu Chen. Genotype and virulence gene analyses of Bacillus cereus group clinical isolates from the Chinese softshell turtle (Pelodiscus sinensis) in Taiwan. Journal of Fish Diseases, 44(10):1515–1529, October 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cheng:2022:FRA

[CRWC22] Li-Wu Cheng, Shreesha Rao, Pei-Chi Wang, and Shih-Chu Chen. First report of acanthocephalan parasite, Longicollum pagrosomi Yamaguti, 1935 in cultured red snapper (Lutjanus erythropterus) in Taiwan. Journal of Fish Diseases, 45(4): 579–593, April 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Charoenwai:2021:DSD

[CSDS21] Onanong Charoenwai, Saengchan Senapin, Ha Thanh Dong, and Molruedee Sonthi. Detection of scale drop disease virus from non-destructive samples and ectoparasites of Asian sea bass, Lates calcarifer. Journal of Fish Diseases, 44(4):461– 467, April 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Campos-Sanchez:2021:RIF

[CSE21] Jose Carlos Campos-Sánchez and María Ángeles Esteban. Review of inflammation in fish and value of the zebrafish model. Journal of Fish Diseases, 44(2):123–139, February 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Campos-Sanchez:2021:ESI

[CSGB^{+}21] José Carlos Campos-Sánchez, Francisco A. Guardiola, José María García Beltrán, Diana Ceballos-Francisco, and María Ángeles Esteban. Effects of subcutaneous injection of λ/κ-carrageenin on the immune and liver antioxidant status of gilthead seabream (Sparus aurata). Journal of Fish Diseases, 44(9):1449–1462, September 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Campos-Sanchez:2022:IMS

[CSVG^{+}22] José Carlos Campos-Sánchez, Enrica Vitarelli, Francisco A. Guardiola, Diana Ceballos-Francisco, José María García Beltrán, Antonio Ieni, and María Ángeles Esteban. Implication of mucus-secreting cells, acidophilic granulocytes and monocytes/macrophages in the resolution of skin inflammation caused by subcutaneous injection of λ/κ-carrageenin to gilthead seabream (Sparus aurata) specimens. Journal of Fish Diseases, 45(1):19–33, January 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Charles:2020:EIM

[CTOH20] Maud Charles, Suzanne Trancart, Elise Oden, and Maryline Houssin. Experimental infection of Mytilus edulis by two Vibrio splendidus-related strains: Determination of pathogenicity level of strains and influence of the origin and annual cycle of mussels on their sensitivity. Journal of Fish Diseases, 43(1):9–21, January 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cocco:2023:NGD

[CTP^{+}23] Alessia Cocco, Marica Toson, Alberto Perolo, Claudia Casarotto, Eleonora Franzago, Ginevra Brocca, Ranieri Verin, Francesco Quaglio, Manuela Dalla Pozza, and Laura Bille. Nodular gill disease in Northeastern Italy: an investigation on the prevalence of the disease and the risks of introduction in rainbow trout farms. Journal of Fish Diseases, 46(10):1021–1028, October 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Castrillo:2021:MGR

[CVDB^{+}21] Pedro A. Castrillo, Catuxa Varela-Dopico, Roberto Bermúdez, Paz Ondina, and María Isabel Quiroga. Morphopathology and gill recovery of Atlantic salmon during the parasitic detachment of Margaritifera margaritifera. Journal of Fish Diseases, 44(8):1101–1115, August 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Castrillo:2020:ESM

[CVDO^{+}20] Pedro A. Castrillo, Catuxa Varela-Dopico, Paz Ondina, María Isabel Quiroga, and Roberto Bermúdez. Early stages of Margaritifera margaritifera glochidiosis in Atlantic salmon: Morphopathological characterization. Journal of Fish Diseases, 43(1):69–80, January 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chinchilla:2023:FDF

[CVFRM^{+}23] Blanca Chinchilla, Esther Vázquez-Fernández, Agustín Rebollada-Merino, Marta Pérez-Sancho, Lucas Domínguez, and Antonio Rodríguez-Bertos. First detection of Flavobacterium psychrophilum in juvenile Siberian sturgeons (Acipenser baerii) and description of the pathological findings. Journal of Fish Diseases, 46(8):887–894, August 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Carvalho:2020:CI

[CWB^{+}20a] Laura A. Carvalho, Shona K. Whyte, Laura M. Braden, Sara L. Purcell, Anthony J. Manning, Anne Muckle, and Mark D. Fast. Cover image. Journal of Fish Diseases, 43 (4):i, April 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Carvalho:2020:ICI

[CWB^{+}20b] Laura A. Carvalho, Shona K. Whyte, Laura M. Braden, Sara L. Purcell, Anthony J. Manning, Anne Muckle, and Mark D. Fast. Impact of co-infection with Lepeophtheirus salmonis and Moritella viscosa on inflammatory and immune responses of Atlantic salmon (Salmo salar). Journal of Fish Diseases, 43(4):459–473, April 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chen:2024:IIC

[CXC^{+}24] Mengmeng Chen, Mingyang Xue, Jingtao Chen, Zidong Xiao, Xiaowei Hu, Chunjie Zhang, Nan Jiang, Yuding Fan, Yan Meng, and Yong Zhou. Isolation, identification and characterization of Aeromonas jandaei from diseased Chinese soft-shell turtles. Journal of Fish Diseases, 47(4): e13919:1–e13919:??, April 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cai:2022:ICN

[CYX^{+}22] Jia Cai, Dapeng Yu, Hongli Xia, Liqun Xia, and Yishan Lu. Identification and characterization of a nervous necrosis virus isolated from largemouth bass (Micropterus salmoides). Journal of Fish Diseases, 45(4):607–611, April 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chen:2022:EZC

[CZC^{+}22] Hong Chen, Fange Zhao, Kexing Chen, Yihan Guo, Yue Liang, Huiying Zhao, and Shulin Chen. Exposure of zebrafish to a cold environment triggered cellular autophagy in zebrafish liver. Journal of Fish Diseases, 45(7):991–1000, July 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cheng:2023:CFC

[CZJ^{+}23] Zhao Cheng, Xianliang Zhao, Xinyu Jiang, Jie Zhang, Li Li, Chao Pei, Yong Zhou, Lingbing Zeng, and Xianghui Kong. Contribution of flagellar cap gene in virulence and pathogenicity of Aeromonas veronii. Journal of Fish Diseases, 46(3):247–259, March 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chen:2022:CEI

[CZM^{+}22] Zhiqiang Chen, Jinyu Zhang, Zhihao Ming, Hao Tong, Jiahui Wu, Qiaoqiao Chen, Yintao Wang, Fangmei Luo, Yipeng Wang, and Tingting Feng. As-Cathelicidin4 enhances the immune response and resistance against Aeromonas hydrophila in caridean shrimp. Journal of Fish Diseases, 45 (5):743–754, May 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Corriero:2021:AOF

[CZMP21] Aldo Corriero, Rosa Zupa, Constantinos C. Mylonas, and Letizia Passantino. Atresia of ovarian follicles in fishes, and implications and uses in aquaculture and fisheries. Journal of Fish Diseases, 44(9):1271–1291, September 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

deAraujo:2023:ECS

[dAMHdOCF^{+}23] Pollyana Alves de Araújo, Patricia Oliveira Maciel-Honda, Thaís de Oliveira Costa-Fernandes, Gracienhe Gomes dos Santos, and Maurício Laterça Martins. Efficacy of chlorine, sodium chloride and trichlorfon baths against monogenean Dawestrema cycloancistrium parasite of pirarucu Arapaima gigas. Journal of Fish Diseases, 46(2):113–126, February 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Danne:2022:IVI

[DAW^{+}22] Linna Danne, Mikolaj Adamek, Hubert Wonnemann, Theresa Pieper, Daniel Fey, and John Hellmann. Identification of virus infections of European eels intended for stocking measures. Journal of Fish Diseases, 45(9):1259– 1266, September 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dias:2022:UCL

[DBC^{+}22] Valera L. Dias, Stephen C. Battaglene, Jennifer M. Cobcroft, Karsten Goemann, and Barbara F. Nowak. Urinary calculi in larvae of striped trumpeter, Latris lineata. Journal of Fish Diseases, 45(9):1281–1287, September 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Delalay:2020:UDA

[DBDSP20] Gary Delalay, John Andrew Berezowski, Nicolas Diserens, and Heike Schmidt-Posthaus. An understated danger: Antimicrobial resistance in aquaculture and pet fish in Switzerland, a retrospective study from 2000 to 2017. Journal of Fish Diseases, 43(10):1299–1315, October 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dragicevic:2021:MPF

[DBPH21] Paula Dragičević, Ana Bielen, Ines Petrić, and Sandra Hudina. Microbial pathogens of freshwater crayfish: a critical review and systematization of the existing data with directions for future research. Journal of Fish Diseases, 44(3): 221–247, March 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Desbois:2020:AMB

[DCB20] Andrew P. Desbois, Kira J. Cook, and Elizabeth Buba. Antibiotics modulate biofilm formation in fish pathogenic isolates of atypical Aeromonas salmonicida. Journal of Fish Diseases, 43(11):1373–1379, November 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

daCruz:2022:TEA

[dCJBG22] Matheus Gomes da Cruz, Gabriela Tomas Jerônimo, Sunique Poá Cordeiro Bentes, and Ligia Uribe Gonçalves. Trichlorfon is effective against Dawestrema cycloancistrium and does not alter the physiological parameters of arapaima (Arapaima gigas): a large Neotropical fish from the Amazon. Journal of Fish Diseases, 45(1):203–212, January 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Debnath:2020:TYS

[DDDJ^{+}20] Partho Pratim Debnath, Jerome Delamare-Deboutteville, Mona Dverdal Jansen, Kornsunee Phiwsaiya, Afsana Dalia, Md. Abir Hasan, Saengchan Senapin, Chadag Vishnumurthy Mohan, Ha Thanh Dong, and Channarong Rodkhum. Twoyear surveillance of tilapia lake virus (TiLV) reveals its wide circulation in tilapia farms and hatcheries from multiple districts of Bangladesh. Journal of Fish Diseases, 43(11):1381– 1389, November 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dang:2022:EAC

[DDH^{+}22] Mai Dang, Tran Duc Dien, Vo Thi Ha, Viet Cuong Hua, Nguyen Thi Hai Thanh, and Barbara F. Nowak. Epitheliocystis in armoured catfish (Pterygoplichthys spp.), anabas (Anabas testudineus) and tilapia (Oreochromis niloticus) in central Vietnam. Journal of Fish Diseases, 45(5):755–760, May 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Debnath:2022:TLV

[DDHT^{+}22] Partho Pratim Debnath, Nguyen Dinh-Hung, Suwimon Taengphu, Vuong Viet Nguyen, Jerome Delamare-Deboutteville, Saengchan Senapin, Chadag Vishnumurthy Mohan, Ha Thanh Dong, and Channarong Rodkhum. Tilapia Lake Virus was not detected in non-tilapine species within tilapia polyculture systems of Bangladesh. Journal of Fish Diseases, 45 (1):77–87, January 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Delamare-Deboutteville:2021:RGT

[DDTG^{+}21] Jerome Delamare-Deboutteville, Suwimon Taengphu, Han Ming Gan, Pattanapon Kayansamruaj, Partho Pratim Debnath, Andrew Barnes, Shaun Wilkinson, Minami Kawasaki, Chadag Vishnumurthy Mohan, Saengchan Senapin, and Ha Thanh Dong. Rapid genotyping of tilapia lake virus (TiLV) using Nanopore sequencing. Journal of Fish Diseases, 44(10):1491–1502, October 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ditlecadet:2022:FRS

[DGB^{+}22] D. Ditlecadet, C. Gautreau, L. Boston, R. Liston, E. Johnsen, and N. Gagné. First report of successful isolation of a HPR0-like variant of the infectious salmon anaemia virus (ISAV) using cell culture. Journal of Fish Diseases, 45 (3):479–483, March 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Danne:2022:VIE

[DHF^{+}22] Linna Danne, Lisa Horn, Anita Feldhaus, Daniel Fey, Sebastian Emde, Heike Schütze, Mikolaj Adamek, and John Hellmann. Virus infections of the European eel in North Rhine Westphalian rivers. Journal of Fish Diseases, 45(1): 69–76, January 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dinh-Hung:2021:DLT

[DHSK^{+}21] Nguyen Dinh-Hung, Pattiya Sangpo, Thanapong Kruangkum, Pattanapon Kayansamruaj, Tilladit Rung-ruangkijkrai, Saengchan Senapin, Channarong Rodkhum, and Ha Thanh Dong. Dissecting the localization of Tilapia tilapinevirus in the brain of the experimentally infected Nile tilapia, Oreochromis niloticus (L.). Journal of Fish Diseases, 44(8): 1053–1064, August 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dalvin:2020:ETA

[DHSM^{+}20] Sussie Dalvin, Lars Are Hamre, Rasmus Skern-Mauritzen, Tone Vågseth, Lars Stien, Frode Oppedal, and Samantha Bui. The effect of temperature on ability of Lepeophtheirus salmonis to infect and persist on Atlantic salmon. Journal of Fish Diseases, 43(12):1519–1529, December 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ding:2021:LMD

[Din21] Zhengfeng Ding. Lipid metabolism disorders contribute to the pathogenesis of Hepatospora eriocheir in the crab Eriocheir sinensis. Journal of Fish Diseases, 44(3):305–313, March 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dien:2021:ONT

[DLS^{+}21] Le Thanh Dien, Nguyen Vu Linh, Pattiya Sangpo, Saengchan Senapin, Sophie St-Hilaire, Channarong Rodkhum, and Ha Thanh Dong. Ozone nanobubble treatments improve survivability of Nile tilapia (Oreochromis niloticus) challenged with a pathogenic multi-drug-resistant Aeromonas hydrophila. Journal of Fish Diseases, 44(9): 1435–1447, September 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Duflot:2021:OTD

[DMB^{+}21] Maureen Duflot, Graziella Midelet, Odile Bourgau, Kurt Buchmann, and Mélanie Gay. Optimization of tools for the detection and identification of Cryptocotyle metacercariae in fish: Digestion method and viability studies. Journal of Fish Diseases, 44(11):1777–1784, November 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Delphino:2021:CEL

[DMH^{+}21] Marina K. V. C. Delphino, Fernando O. Mardones, Joaquin Neumann Heise, Alicia Gallardo, Daniel Jimenez, Andrea Peña, Marco Rozas-Serri, and Ian A. Gardner. Costeffectiveness of longitudinal surveillance for Piscirickettsia salmonis using qPCR in Atlantic salmon farms (Salmo salar) in Chile. Journal of Fish Diseases, 44(3):315–326, March 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dang:2022:ADF

[DNN^{+}22] Mai Dang, Hai Thanh T. Nguyen, Van Manh Ngo, Tran Duc Dien, Tran Ngoc Thang, Nguyen Thi Phu’o’ng Thao, Binh Thuy Dang, and Ha Thanh Dong. Acute death in farmed marine fishes caused by sea anemone (Bunodeopsis sp.) in Central Vietnam. Journal of Fish Diseases, 45(11): 1799–1803, November 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dean:2022:DIH

[DOW^{+}22] Katharine R. Dean, Victor H. S. Oliveira, Cecilia Wolff, Torfinn Moldal, and Mona Dverdal Jansen. Description of ISAV-HPRΔ-positive salmon farms in Norway in 2020. Journal of Fish Diseases, 45(1):225–229, January 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Deering:2024:RDV

[DPA^{+}24] Margret J. Deering, Hélène Paradis, Raahyma Ahmad, Adil S. Al-Mehiawi, and Robert L. Gendron. The role of dietary vitamin A in mechanisms of cataract development in the teleost lumpfish (Cyclopterus lumpus L). Journal of Fish Diseases, 47(3):e13899:1–e13899:??, March 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Delfosse:2021:EHC

[DPLL^{+}21] Cyril Delfosse, Patrick Pageat, Céline Lafont-Lecuelle, Pietro Asproni, Camille Chabaud, Alessandro Cozzi, and Cécile Bienboire-Frosini. Effect of handling and crowding on the susceptibility of Atlantic salmon (Salmo salar L.) to Lepeophtheirus salmonis (Krøyer) copepodids. Journal of Fish Diseases, 44(3):327–336, March 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

delRio-Rodriguez:2021:FEF

[dRRRPSR^{+}21] Rodolfo Enrique del Rio-Rodriguez, Jose Gustavo Ramirez-Paredes, Sonia Araceli Soto-Rodriguez, Yechiam Shapira, Mariana del Jesus Huchin-Cortes, Judith Ruiz-Hernandez, Monica Isela Gomez-Solano, and David J. Haydon. First evidence of fish nocardiosis in Mexico caused by Nocardia seriolae in farmed red drum (Sciaenops ocellatus, Linnaeus). Journal of Fish Diseases, 44(8):1117–1130, August 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dong:2022:UPP

[DSD^{+}22] Ha Thanh Dong, Pattiya Sangpo, Le Thanh Dien, Thao Thu Mai, Nguyen Vu Linh, Jorge del Pozo, Krishna R. Salin, and Saengchan Senapin. Usefulness of the pancreas as a prime target for histopathological diagnosis of Tilapia parvovirus (TiPV) infection in Nile tilapia, Oreochromis niloticus. Journal of Fish Diseases, 45(9):1323–1331, September 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Duan:2021:EFE

[DvGJK^{+}21] Yajiao Duan, Louise von Gersdorff Jørgensen, Per Walter Kania, Asma M. Karami, Azmi Al-Jubury, and Kurt Buchmann. Eye fluke effects on Danish freshwater fish: Field and experimental investigations. Journal of Fish Diseases, 44(11):1785–1798, November 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Denk:2020:SNC

[DVR^{+}20] Daniela Denk, Ranieri Verin, Lorenzo Ressel, Eleanor Lewis, and Mark F. Stidworthy. Spontaneous neoplasia in captive syngnathid species: a retrospective case series (2003–2014) and literature review. Journal of Fish Diseases, 43(8):929– 939, August 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dalum:2022:HDE

[DWK^{+}22] Alf Seljenes Dalum, Helene Wisløff, Herman Høgenes Kvinnsland, Are Nylund, and Egil Karlsbakk. Histopathological description of an emerging disease in Norwegian salmonid aquaculture caused by an x-cell parasite. Journal of Fish Diseases, 45(1):213–217, January 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ding:2020:MCT

[DWY^{+}20] Yuehan Ding, Ke Wei, Xiao Yang, Fei Jing, Bin Shen, and Jianshe Zhang. Molecular characterization of three caspases from Bostrychus sinensis and their transcriptional responses to bacteria and viruses. Journal of Fish Diseases, 43(4):431– 443, April 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dang:2021:PEP

[DXD^{+}21] Thi Hoang Oanh Dang, Tran T. T. Xuan, Le T. M. Duyen, Nga P. Le, and Hoang A. Hoang. Protective efficacy of phage PVN02 against haemorrhagic septicaemia in striped catfish Pangasianodon hypophthalmus via oral administration. Journal of Fish Diseases, 44(8):1255–1263, August 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Deng:2020:RMH

[DXL^{+}20a] Jun-Jin Deng, Shun Xu, Yan-Wei Li, Dong-Dong Xu, Ze-Quan Mo, Jia-Zhou Li, Xue-Ming Dan, and Xiao-Chun Luo. Role of major histocompatibility complex II antigenpresentation pathway genes in orange-spotted grouper infected with Cryptocaryon irritans. Journal of Fish Diseases, 43(12):1541–1552, December 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Deng:2020:WDC

[DXL^{+}20b] Yiqin Deng, Liwen Xu, Songlin Liu, Qian Wang, Zhixun Guo, Chang Chen, and Juan Feng. What drives changes in the virulence and antibiotic resistance of Vibrio harveyi in the South China Sea? Journal of Fish Diseases, 43(8): 853–862, August 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

deYbanez:2023:MAC

[dYdRFF^{+}23] M. Rocío Ruiz de Ybáñez, Laura del Río, César Flores-Flores, Pilar Muñoz, Eduardo Berriatua, Silvia Rubio, and Carlos Martínez-Carrasco. Monitoring for Anguillicoloides crassus, Anguillid herpesvirus 1, aquabirnavirus EVE and rhabdovirus EVEX in the European eel population of southern Spain. Journal of Fish Diseases, 46(4):417–431, April 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dennis:2020:MAD

[DYN^{+}20] Michelle M. Dennis, Jun Wei Alric Yeow, Brittani Nicolaci, Saundra H. Sample, and Mark A. Freeman. Meningitis associated with dracunculid nematodes in a spotted eagle ray (Aetobatus narinari). Journal of Fish Diseases, 43(6):711– 713, June 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Du:2022:CRN

[DYPN22] Juan Du, Kaidi Yue, Yanxin Peng, and Qianji Ning. Crucial roles of a novel exoskeletal-derived lectin in innate immunity of the oriental river prawn, Macrobrachium nipponense. Journal of Fish Diseases, 45(5):717–728, May 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Duan:2021:PMS

[DZJ^{+}21] Hui-Xin Duan, Zhao Zhao, Ying-Jie Jin, Zi-Long Wang, Jie-Fang Deng, Jie He, and Bin Zhu. PEG-modified subunit vaccine encoding dominant epitope to enhance immune response against spring viraemia of carp virus. Journal of Fish Diseases, 44(10):1587–1594, October 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Du:2022:RMT

[DZQ^{+}22] Ziyan Du, Mengmeng Zhang, Yingxue Qin, Lingmin Zhao, Lixing Huang, Xiaojin Xu, and Qingpi Yan. The role and mechanisms of the two-component system EnvZ/OmpR on the intracellular survival of Aeromonas hydrophila. Journal of Fish Diseases, 45(11):1609–1621, November 2022. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dykova:2021:HLR

[DŽR^{+}21] Iva Dyková, Jakub Žák, Martin Reichard, Kamila Součková, Ondřej Slabý, Vojtěch Bystrý, and Radim Blažek. Histopathology of laboratory-reared Nothobranchius fishes: Mycobacterial infections versus neoplastic lesions. Journal of Fish Diseases, 44(8):1179–1190, August 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

English:2021:ICN

[EBA^{+}21] Chloe J. English, Natasha A. Botwright, Mark B. Adams, Andrew C. Barnes, James W. Wynne, Paula C. Lima, and Mathew T. Cook. Immersion challenge of naïve Atlantic salmon with cultured Nolandella sp. and Pseudoparamoeba sp. did not increase the severity of Neoparamoeba peruransinduced amoebic gill disease (AGD). Journal of Fish Diseases, 44(2):149–160, February 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Echeverria-Bugueno:2024:TDP

[EBAH24] Macarena Echeverría-Bugueño and Ruben Avendaño-Herrera. Tenacibaculum dicentrarchi produce outer membrane vesicles (OMV) that are associated with the cytotoxic effect in rainbow trout head kidney macrophages. Journal of Fish Diseases, 47(2):e13888:1–e13888:??, February 2024.

CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Echeverria-Bugueno:2021:EEE

[EBBIAH21] Macarena Echeverría-Bugueño, Cristóbal Balada, Rute Irgang, and Ruben Avendaño-Herrera. Evidence for the existence of extracellular vesicles in Renibacterium salmoninarum and related cytotoxic effects on SHK-1 cells. Journal of Fish Diseases, 44(7):1015–1024, July 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Echeverria-Bugueno:2020:ICO

[EBELIAH20] Macarena Echeverría-Bugueño, Rodrigo Espinosa-Lemunao, Rute Irgang, and Ruben Avendaño-Herrera. Identification and characterization of outer membrane vesicles from the fish pathogen Vibrio ordalii. Journal of Fish Diseases, 43 (5):621–629, May 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ehemann:2022:MAS

[EGRD22] Nicolás Roberto Ehemann, Francisco Javier García-Rodríguez, and José De La Cruz-Agüero. Morphological abnormalities in seven American round ray specimens: a review of America’s batomorph anomalies. Journal of Fish Diseases, 45(3): 395–409, March 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Eilts:2023:PCI

[EJH^{+}23] Friederike Eilts, Lisa K. Jordan, Yasmina M. J. Harsy, Sven M. Bergmann, Anna M. Becker, and Michael W. Wolff. Purification and concentration of infectious koi herpesvirus using steric exclusion chromatography. Journal of Fish Diseases, 46(8):873–886, August 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Evenhuis:2021:VVF

[ELB21] Jason P. Evenhuis, Ryan Lipscomb, and Clayton Birkett. Virulence variations of Flavobacterium columnare in rainbow trout (Oncorhynchus mykiss) eyed eggs and alevin. Journal of Fish Diseases, 44(5):533–539, May 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Engblom:2023:IGC

[ELS^{+}23] Christine Engblom, Lotta Landor, Conny Sjöqvist, Tiina Korkea-aho, Satu Viljamaa-Dirks, Lars Paulin, and Tom Wiklund. Identification and genetic characterization of Saprolegnia parasitica, isolated from farmed and wild fish in Finland. Journal of Fish Diseases, 46(8):849–860, August 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Evans:2022:LLD

[EMH^{+}22] Drayke Evans, Zachary Millar, Danielle Harding, Phuc H. Pham, Véronique LePage, and John S. Lumsden. Lipoid liver disease in Hippocampus erectus Perry with Vibrio fortis -induced dermatitis and enteritis. Journal of Fish Diseases, 45(8):1225–1229, August 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Eurlaphan:2023:RSB

[ENS^{+}23] Chalermkwan Eurlaphan, Reiko Nozaki, Motohiko Sano, Keiichiro Koiwai, Ikuo Hirono, and Hidehiro Kondo. Red sea bream iridovirus infection downregulates inflammationrelated genes in the spleen of rock bream (Oplegnathus fasciatus). Journal of Fish Diseases, 46(12):1403–1411, December 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

El-Son:2020:MHC

[ESEI20] Mai A. M. El-Son, Samia Elbahnaswy, and Iman Ibrahim. Molecular and histopathological characterization of Photobacterium damselae in naturally and experimentally infected Nile tilapia (Oreochromis niloticus). Journal of Fish Diseases, 43(12):1505–1517, December 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Felten:2022:IVI

[FAG^{+}22] Martin Felten, Mikolaj Adamek, Marina Gebert, Sebastian Rakers, and Dieter Steinhagen. The influence of viral infection on cell line characteristics: Lessons learned from working with new cell lines from common carp. Journal of Fish Diseases, 45(11):1767–1780, November 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Feng:2022:RVD

[FCH^{+}22a] Zizhao Feng, Xin Chu, Minzhen Han, Chenwei Yu, Yousheng Jiang, Hao Wang, Liqun Lu, and Dan Xu. Rapid visual detection of Micropterus salmoides rhabdovirus using recombinase polymerase amplification combined with lateral flow dipsticks. Journal of Fish Diseases, 45(3):461–469, March 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Foddai:2022:CEE

[FCH^{+}22b] Marco Foddai, Chris G. Carter, Pollyanna E. Hilder, Harley Gurr, and Nicole Ruff. Combined effects of elevated rearing temperature and dietary energy level on heart morphology and growth performance of Tasmanian Atlantic salmon (Salmo salar L.). Journal of Fish Diseases, 45(2):301–313, February 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fernandes:2022:RML

[FdSB^{+}22] Bianca H. Ventura Fernandes, Caroline Caetano da Silva, Debora Bissegato, Michael L. Kent, and Luciani R. Carvalho. Rederivation of a mutant line (prop 1 ) of zebrafish Danio rerio infected with Pseudoloma neurophilia using in vitro fertilization with eggs from pathogen-free wild-type (AB) females and sperm from prop 1 males. Journal of Fish Diseases, 45(1):35–39, January 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Furnesvik:2022:EHP

[FEH^{+}22] Lisa Furnesvik, Toni Erkinharju, Miroslava Hansen, Muhammad Naveed Yousaf, and Tore Seternes. Evaluation of histological post-mortem changes in farmed Atlantic salmon (Salmo salar L.) at different time intervals and storage temperatures. Journal of Fish Diseases, 45(10):1571–1580, October 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fragkoulis:2020:CFA

[FEM^{+}20] Stefanos Fragkoulis, Irene Economou, Gabriel Moukas, George Koumoundouros, and Costas Batargias. Caudal fin abnormalities in Gilthead seabream (Sparus aurata L.) have a strong genetic variance component. Journal of Fish Diseases, 43(7):825–828, July 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fernandez-Espinel:2023:CET

[FEMMI^{+}23] Carla Fernandez-Espinel, Marco Medina-Morillo, Rute Irgang, Giovanna Sotil, Henry Araya-León, Violeta Flores-Dominick, Jesús L. Romalde, Ruben Avendaño-Herrera, and Jefferson Yunis-Aguinaga. Co-existence of two Yersinia ruckeri biotypes and serotype O1a retrieved from rainbow trout (Oncorhynchus mykiss) farmed in Puno, Peru. Journal of Fish Diseases, 46(2):157–163, February 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Farmer:2021:DSW

[FFB^{+}21] Bradley D. Farmer, Sidney Adam Fuller, Benjamin H. Beck, Jason W. Abernathy, Miles D. Lange, and Carl D. Webster. Differential susceptibility of white bass (Morone chrysops), striped bass (Morone saxatilis) and hybrid striped bass (M. chrysops × M. saxatilis) to Flavobacterium columnare and effects of mucus on bacterial growth and biofilm development. Journal of Fish Diseases, 44(2):161–169, February 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fragkoulis:2021:SMP

[FK21] Stefanos Fragkoulis and George Koumoundouros. Simple morphometrics for predicting lordosis-induced deviations of body shape in reared Gilthead seabream (Sparus aurata L.). Journal of Fish Diseases, 44(8):1265–1267, August 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fu:2021:MCC

[FLH^{+}21] Shengli Fu, Haisu Liu, Yu Huang, Junru Wang, Kun Qian, Mingmei Ding, and Jianmin Ye. Molecular cloning, characterization and expression analysis of major histocompatibility complex class I alpha gene of pufferfish (Takifugu obscurus). Journal of Fish Diseases, 44(5):613–625, May 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Febriana:2023:RFV

[FM23] Clara Artha Febriana and Yunita Messe. Review of Fish Vaccines: Health Management for Sustainable Aquaculture. Edited by Preetham Elumalai, Kim Thompson, Sreeja Lakshmi. Boca Raton, FL, USA: CRC Press, 2023, 262 pp., eBook 108.00, ISBN 978-1-032-48341-2. Journal of Fish Diseases, 46(12):1445–1447, December 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fritsvold:2022:CEP

[FMH^{+}22] Camilla Fritsvold, Aase B. Mikalsen, Øyvind Haugland, Haitham Tartor, and Hilde Sindre. Characterization of early phases of cardiomyopathy syndrome pathogenesis in Atlantic salmon (Salmo salar L.) through various diagnostic methods. Journal of Fish Diseases, 45(9):1267–1279, September 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Foysal:2021:ASR

[FMK^{+}21] Md Javed Foysal, Farhana Momtaz, A. Q. M. Robiul Kawser, Md Hazrat Ali, Topu Raihan, Muhammad A. B. Siddik, Md Mahbubur Rahman, and Alfred Tay. Amplicon sequencing reveals significantly increased Vibrio abundance and associated gene functions in vibriosis-infected black tiger shrimp (Penaeus monodon). Journal of Fish Diseases, 44 (5):591–599, May 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fritsvold:2021:COC

[FMP^{+}21] Camilla Fritsvold, Aase B. Mikalsen, Trygve T. Poppe, Torunn Taksdal, and Hilde Sindre. Characterization of an outbreak of cardiomyopathy syndrome (CMS) in young Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 44 (12):2067–2082, December 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fernandez-Senac:2020:CUD

[FSFS^{+}20] Carolina Fernandez-Senac, Sophie Fridman, Jadwiga Sokolowska, Sean J. Monaghan, Teresa Garzon, Monica Betancor, Giuseppe Paladini, Alexandra Adams, and James E. Bron. A comparison of the use of different swab materials for optimal diagnosis of amoebic gill disease (AGD) in Atlantic salmon (Salmo salar L.). Journal of Fish Diseases, 43(11): 1463–1472, November 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Faber:2021:DSC

[FSS^{+}21] Marc N. Faber, Jerry M. Sojan, Marcia Saraiva, Pieter van West, and Christopher J. Secombes. Development of a 3D spheroid cell culture system from fish cell lines for in vitro infection studies: Evaluation with Saprolegnia parasitica. Journal of Fish Diseases, 44(6):701–710, June 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ford:2021:COG

[FSW^{+}21] Lorelei Ford, Kuttichantran Subramaniam, Thomas B. Waltzek, Paul R. Bowser, and Larry Hanson. Cytochrome oxidase gene sequencing reveals channel catfish ovary cell line is contaminated with brown bullhead cells. Journal of Fish Diseases, 44(1):119–122, January 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Figueroa:2021:AGE

[FVC^{+}21] Jaime Figueroa, Daniela Villagrán, Carlos Cartes, Camila Solis, Guillermo Nourdin-Galindo, and Denise Haussmann. Analysis of genes encoding for proteolytic enzymes and cytotoxic proteins as virulence factors of Piscirickettsia salmonis in SHK-1 cells. Journal of Fish Diseases, 44(5):495–504, May 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fei:2023:IAA

[FWL^{+}23] Yueyue Fei, Qin Wang, Jigang Lu, Linyue Ouyang, Wei Li, Ruiqin Hu, and Liangbiao Chen. Identification of antibacterial activity of LEAP2 from Antarctic icefish Chionodraco hamatus. Journal of Fish Diseases, 46(9):905–916, September 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fu:2021:EAT

[FYH^{+}21] Yao-Wu Fu, Zhan-Juan Yao, Meng-Han He, Zuo-Pei Wang, Wei-Feng Chen, Miao Cui, and Qi-Zhong Zhang. Expression analysis and tissue localization of IgZ in the grouper Epinephelus coioides after Vibrio alginolyticus infection and vaccination. Journal of Fish Diseases, 44(10):1647–1655, October 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fu:2023:MME

[FZC^{+}23] Yan Fu, Junzhe Zhang, Wenjie Cheng, Xingyu Cheng, Liqun Lu, Lang Gui, Yousheng Jiang, Ye Zhang, and Dan Xu. miR-124 mediates the expression of ccBax to regulate cyprinid herpesvirus 2 (CyHV-2)-induced apoptosis and viral replication. Journal of Fish Diseases, 46(7):743–749, July 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Grandon:2021:PPP

[GIS^{+}21] Maritza Grandón, Rute Irgang, José Saavedra, Marcos Mancilla, and Ruben Avendaño-Herrera. Proposed protocol for performing MIC testing to determine the antimicrobial susceptibility of Renibacterium salmoninarum in Chilean salmon farms. Journal of Fish Diseases, 44(3):287–296, March 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gibson-Kueh:2023:PLC

[GKAS23] S. Gibson-Kueh, S. Awate, and M. Schrittwieser. The pathology of Lates calcarifer herpesviral disease — Disseminated intravascular coagulation explains mortality spikes. Journal of Fish Diseases, 46(8):861–871, August 2023. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gibson-Kueh:2021:PSH

[GKTC^{+}21] Susan Gibson-Kueh, Celestine Terence, Xian Zhe Chew, Joseph Angelo Uichanco, and Xueyan Shen. PCR, in-situ hybridization, and phylogenetic analysis suggest that ‘big belly’ disease in barramundi, Lates calcarifer (Bloch), is caused by a novel Vibrio species. Journal of Fish Diseases, 44(12): 1985–1992, December 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gibson-Kueh:2021:PAP

[GKU21] Susan Gibson-Kueh and Joseph Angelo Uichanco. The pathology associated with putative algal toxicosis in red snapper, Lutjanus species (Bloch 1790). Journal of Fish Diseases, 44(6):857–861, June 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gonzalez:2021:GHC

[GLS^{+}21] María Teresa González, Natalia Verónica Leiva, Fabiola Sepúlveda, Gladys Asencio, and Juan Antonio Baeza. Genetic homogeneity coupled with morphometric variability suggests high phenotypic plasticity in the sea louse Caligus rogercresseyi (Boxshall and Bravo, 2000), infecting farmed salmon (Salmo salar) along a wide latitudinal range in southern Chile. Journal of Fish Diseases, 44(5):633–638, May 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Groff:2021:PDV

[GMK^{+}21] Joseph M. Groff, Mai Y. Mok, Steven V. Kubiski, Adam O. Michel, Galaxia A. Cortés-Hinojosa, Barbara A. Byrne, Brian L. Wickes, E. Scott Weber I.I.I, Lori A. Campbell, and Thomas B. Waltzek. Phaeohyphomycosis due to Veronaea botryosa in cultured white sturgeon (Acipenser transmontanus Richardson) from California USA during 2006 to 2015. Journal of Fish Diseases, 44(6):793–801, June 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gaasnes:2021:MPD

[GOG^{+}21] Siri Kristine Gåsnes, Victor H. S. Oliveira, Kristine Gismervik, Ashley Ahimbisibwe, Brit Tørud, and Britt Bang Jensen. Mortality patterns during the freshwater production phase of salmonids in Norway. Journal of Fish Diseases, 44(12):2083–2096, December 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Girard:2022:IVF

[GPC22] Sarah B. Girard, Valérie E. Paquet, and Steve J. Charette. Improvements of virulence factor phenotypic tests for Aeromonas salmonicida subsp. salmonicida, a major fish pathogen. Journal of Fish Diseases, 45(1):177–184, January 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gong:2022:ECS

[GPL^{+}22] Jinpeng Gong, Xiaoyi Pan, Lingyun Lin, Yue Zhu, Jiayun Yao, Congxu Wang, Wenlin Yin, Lei Huang, Yihan Liu, Fan Chen, and Jinyu Shen. Establishment and characterization of a spinal cord tissue cell line from mandarin fish, Siniperca chuatsi and its susceptibility to several viruses. Journal of Fish Diseases, 45(10):1419–1427, October 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Grenier:2023:SLP

[GRM^{+}23] Gabrielle Grenier, Eloïse C. Rochat, Rune Muladal, Hallvard Jensen, and Rune Knudsen. Salmon-lice as a potential threat to anadromous Arctic charr populations. Journal of Fish Diseases, 46(5):465–475, May 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gardenia:2020:DPC

[GSJ^{+}20] Lila Gardenia, Sukenda Sukenda, Muhammad Zairin Junior, Angela Lusiastuti, and Alimuddin Alimuddin. Development of primary cell culture from spleen of giant gourami Osphronemus goramy for propagation of giant gourami iridovirus (GGIV). Journal of Fish Diseases, 43(8):829–838, August 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Galeotti:2021:RMSb

[GSS^{+}21] M. Galeotti, G. Sarli, R. Sirri, L. Mandrioli, P. Beraldo, P. Bronzatti, R. Giavenni, M. Orioles, G. E. Magi, and D. Volpatti. Red mark syndrome of trout (Oncorhynchus mykiss; Walbaum, 1792): Histopathological scoring and correlation with gross lesions. Journal of Fish Diseases, 44(9): 1325–1336, September 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Galeotti:2021:RMSa

[GVB^{+}21] Marco Galeotti, Donatella Volpatti, Omkar Byadgi, Paola Beraldo, Massimo Orioles, Massimo Sarti, Sara Ciulli, and Gian Enrico Magi. Red mark syndrome (RMS) in farmed rainbow trout: First report of outbreaks in Bosnia and Herzegovina. Journal of Fish Diseases, 44(5):627–631, May 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Guo:2022:EEQ

[GWF^{+}22] Yanmin Guo, Yahui Wang, Zhaobin Fan, Xianlin Zhao, Sven M. Bergmann, Hanxu Dong, Yuqi Jin, Dongli Sun, Qianyi Mai, Weiqiang Liu, and Weiwei Zeng. Establishment and evaluation of qPCR and real-time recombinaseaided amplification assays for detection of largemouth bass ranavirus. Journal of Fish Diseases, 45(7):1033–1043, July 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gu:2021:VPC

[GWL^{+}21] Dan Gu, Kangru Wang, Tianyu Lu, Lingzhi Li, and Xinan Jiao. Vibrio parahaemolyticus CadC regulates acid tolerance response to enhance bacterial motility and cytotoxicity. Journal of Fish Diseases, 44(8):1155–1168, August 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Guo:2024:ERT

[GXL^{+}24] Xiao-Meng Guo, Jing-Yi Xing, Anqi Li, Liang Qiu, Qing-Li Zhang, and Jie Huang. Establishment of a real-time PCR for the detection of decapod iridescent virus 1 (DIV1). Journal of Fish Diseases, 47(6):e13926:1–e13926:??, June 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gong:2023:HPN

[GXW^{+}23] Miao Gong, Guosi Xie, Hailiang Wang, Xinshu Li, Ang Li, Xiaoyuan Wan, Jingxian Huang, Chengyin Shi, Qingli Zhang, and Jie Huang. Hematodinium perezi naturally infects Asian brush-clawed crab (Hemigrapsus takanoi). Journal of Fish Diseases, 46(1):67–74, January 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Guo:2022:FIN

[GYC^{+}22] Jiajing Guo, Xiaozhen Yue, Jiaojiao Chang, Zhenyuan Zhang, Jinnian Li, and Xuelan Liu. First identification of Nocardia seriolae GapA adhesion function and its three Bcell epitopes with cell-binding activity. Journal of Fish Diseases, 45(12):1845–1855, December 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ghadin:2024:SCS

[GYN^{+}24] Norazli Ghadin, Nur Afiqah Md Yusof, Baharum Syarul Nataqain, Nurul Hanun Ahmad Raston, and Chen Fei Low. Selection and characterization of ssDNA aptamer targeting Macrobrachium rosenbergii nodavirus capsid protein: a potential capture agent in gold-nanoparticle-based aptasensor for viral protein detection. Journal of Fish Diseases, 47(2): e13892:1–e13892:??, February 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gong:2021:OIP

[GZL^{+}21] Yu-Ming Gong, Chen Zhang, Yang Li, Guo Chen, Gao-Xue Wang, and Bin Zhu. Optimization of immunization procedure for SWCNTs-based subunit vaccine with mannose modification against spring viraemia of carp virus in common carp. Journal of Fish Diseases, 44(12):1925–1936, December 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Guo:2023:SSB

[GZW^{+}23] Haiwei Guo, Bo Zhang, Pengfei Wang, Chao Zhao, Lulu Yan, Yunxiang Lin, and Lihua Qiu. Spotted sea bass (Lateolabrax maculatus) NOD2 gene involved in the immune response against Vibrio harveyi infection. Journal of Fish Diseases, 46(8):897–901, August 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gao:2022:GAM

[GZZ^{+}22] Wa Gao, Lupin Zhao, Yihua Zheng, Kaixuan Wu, Feiyang Xu, Hao Wang, Liqun Lu, and Yousheng Jiang. Generation and application of a monoclonal antibody specific for the ORF121 of cyprinid herpesvirus 2. Journal of Fish Diseases, 45(3):387–394, March 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hwang:2020:WGN

[HAK^{+}20] Jee Youn Hwang, Sang Jung Ahn, Mun-Gyeong Kwon, Jung Soo Seo, Seong Don Hwang, and Bo Young Jee. Wholegenome next-generation sequencing and phylogenetic characterization of viral haemorrhagic septicaemia virus in Korea. Journal of Fish Diseases, 43(5):599–607, May 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Henderson:2020:TIE

[HB20] Stacey Elizabeth Henderson and Shane Matthew Boylan. Trauma-induced exophthalmia in vermilion snapper, Rhomboplites aurorubens (Cuvier 1829): Treatment and prevention. Journal of Fish Diseases, 43(2):295–300, February 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Habib:2024:EAA

[HBRN24] Syed Sikandar Habib, Aima Iram Batool, Muhammad Fayyaz Ur Rehman, and Saira Naz. Evaluation of the antibacterial activity and protein profiling of Nile tilapia (Oreochromis niloticus) epidermal mucus under different feeds and culture systems (biofloc technology and earthen pond). Journal of Fish Diseases, 47(2):e13884:1–e13884:??, February 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Heniff:2022:CEL

[HCC^{+}22] Ashlyn C. Heniff, Laura R. Chen, Emily F. Christiansen, Craig A. Harms, Jerry M. Law, Christian Legner, and Keith E. Linder. Cutaneous epitheliotropic lymphosarcoma in a captive white catfish (Ameiurus catus Linnaeus). Journal of Fish Diseases, 45(7):971–974, July 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hou:2020:ICW

[HCW^{+}20] Suying Hou, Guoquan Chen, Wenji Wang, Liqun Xia, Zhiwen Wang, and Yishan Lu. Identification of a cell-wall peptidase (NlpC/P60) from Nocardia seriolae which induces apoptosis in fathead minnow cells. Journal of Fish Diseases, 43(5):571–581, May 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hoitsy:2021:ICI

[HHJ^{+}21] Márton Hoitsy, György Hoitsy, Csaba Jakab, Tamás Molnár, János Gál, and Ferenc Baska. Intussusception caused by intestinal neoplasia in mature rainbow trout (Oncorhynchus mykiss, Walbaum 1792). Journal of Fish Diseases, 44(7): 893–898, July 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

He:2024:NDL

[HHW^{+}24] Huanrong He, Xinzhi Huang, Caiyi Wen, Chun Liu, Biao Jiang, Yanhua Huang, Youlu Su, and Wei Li. A novel defensin-like peptide C-13326 possesses protective effect against multidrug-resistant Aeromonas schubertii in hybrid snakehead (Channa maculate [female sign] × Channa argus [male sign]). Journal of Fish Diseases, 47(4):e13922:1– e13922:??, April 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hu:2023:CDD

[HHZ^{+}23] Guangran Hu, Ke Huang, Weitian Zhou, Runqiu Wang, Weishan Zhao, Hong Zou, Wenxiang Li, Shangong Wu, Ming Li, and Guitang Wang. Comparison of droplet digital PCR and real-time quantitative PCR for quantitative detection of the parasitic ciliate Ichthyophthirius multifiliis in the water environment. Journal of Fish Diseases, 46(4):357–367, April 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Harrison:2022:FIF

[HKB^{+}22] Courtney E. Harrison, Christopher K. Knupp, Travis O. Brenden, Mark P. Ebener, and Thomas P. Loch. First isolation of Flavobacterium psychrophilum from wild adult Great Lakes lake whitefish (Coregonus clupeaformis). Journal of Fish Diseases, 45(7):1023–1032, July 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hofer:2023:CI

[HLH^{+}23a] Rachel N. Hofer, Allen Lin, Benjamin C. House, Christopher N. Purvis, Bradley J. Harris, Steven J. K. Symes, and David K. Giles. Cover image. Journal of Fish Diseases, 46(1):i, January 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hofer:2023:EPF

[HLH^{+}23b] Rachel N. Hofer, Allen Lin, Benjamin C. House, Christopher N. Purvis, Bradley J. Harris, Steven J. K. Symes, and David K. Giles. Exogenous polyunsaturated fatty acids (PUFAs) influence permeability, antimicrobial peptide resistance, biofilm formation and membrane phospholipid structure in an A-layer and non-A-layer strain of Aeromonas salmonicida. Journal of Fish Diseases, 46(1):31–45, January 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hu:2022:NIP

[HLpSC22] Yang Hu, Lei Liu, Li peng Shan, and Jiong Chen. Natural ingredient paeoniflorin could be a lead compound against white spot syndrome virus infection in Litopenaeus vannamei. Journal of Fish Diseases, 45(2):349–359, February 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hu:2023:LMI

[HLW^{+}23] Keshun Hu, Ye Li, Feng Wang, Jianying Liu, Yuanyuan Li, Qian Zhao, Xiaoye Zheng, Ningyu Zhu, Xiaoping Yu, Shaohua Fang, and Jun Huang. A loop-mediated isothermal amplification-based microfluidic chip for triplex detection of shrimp pathogens. Journal of Fish Diseases, 46(2):137–146, February 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Herrero:2022:NDB

[HPR^{+}22] Ana Herrero, Oswaldo Palenzuela, Hamish Rodger, Chris Matthews, Mar Marcos-López, James E. Bron, Mark P. Dagleish, and Kim D. Thompson. Novel DNA-based in situ hybridization method to detect Desmozoon lepeophtherii in Atlantic salmon tissues. Journal of Fish Diseases, 45(6):871– 882, June 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Herron:2024:ASC

[HRR^{+}24] Crystal L. Herron, Natassia E. Ruse, Daniel D. Rockey, Justin L. Sanders, James T. Peterson, Carl B. Schreck, and Michael L. Kent. Aeromonas salmonicida, causative agent of salmonid furunculosis, isolated from the freshwater parasitic copepod, Salmincola californiensis. Journal of Fish Diseases, 47(2):e13885:1–e13885:??, February 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Huang:2021:NDL

[HSG^{+}21] Pengdan Huang, Guoqing Shen, Jie Gong, Mengru Zhu, Yang Wang, Xiao Zhang, Kassimu Hashimu Ame, Yanan Zang, and Huaishun Shen. A novel Dicistro-like virus discovered in Procambarus clarkii with “Black May” disease. Journal of Fish Diseases, 44(6):803–811, June 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hershberger:2020:PHC

[HSH^{+}20] Paul K. Hershberger, Matthew E. T. Stinson, Betsy Hall, Ashley H. MacKenzie, Jacob L. Gregg, William Richards, and James R. Winton. Pacific herring Clupea pallasii are not susceptible to vibriosis from Vibrio anguillarum or V. ordalii under laboratory conditions. Journal of Fish Diseases, 43 (12):1607–1609, December 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Huang:2022:NCP

[HSH^{+}22] Ai-Guo Huang, Lin-Jun Su, Wei-Hao He, Fa-Li Zhang, Chao-Shuai Wei, and Ying-Hui Wang. Natural component plumbagin as a potential antibacterial agent against Streptococcus agalactiae infection. Journal of Fish Diseases, 45 (6):815–823, June 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hoare:2022:SMP

[HSM^{+}22] R. Hoare, K. Shahin, K. McLean, A. Adams, and K. D. Thompson. Skin mucus proteins of rainbow trout (Oncorhynchus mykiss) in response to mucosal vaccination and challenge with Flavobacterium psychrophilum. Journal of Fish Diseases, 45(3):491–495, March 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hansen:2024:NLD

[HSS^{+}24] Haakon Hansen, Bjørn Spilsberg, Sigmund Sevatdal, Trine Sakariassen, Christoph Hahn, Saima Nasrin Mohammad, and Egil Karlsbakk. Non-lethal detection of Eubothrium crassum (Cestoda) in farmed Atlantic salmon, Salmo salar, using anal swabs and real-time PCR. Journal of Fish Diseases, 47(5):e13918:1–e13918:??, May 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hou:2021:ISS

[HWC^{+}21] Suying Hou, Wenji Wang, Guoquan Chen, Liqun Xia, Zhiwen Wang, and Yishan Lu. Identification of a secreted superoxide dismutase (SOD) from Nocardia seriolae which induces apoptosis in fathead minnow (FHM) cells. Journal of Fish Diseases, 44(1):63–72, January 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Huang:2021:TAI

[HWG^{+}21] Ying Huang, Ruixia Wang, Tianheng Gao, Ting Wu, Qiya Zhang, Yangbai Shi, Shuyan Ding, and Zhe Zhao. Transcriptome analysis of immune response against Siniperca chuatsi rhabdovirus infection in mandarin fish Siniperca chuatsi. Journal of Fish Diseases, 44(6):675–687, June 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hu:2024:PEP

[HWM^{+}24] Jianmei Hu, Baotun Wang, Zhuang Ma, Juan Feng, Biao Jiang, and Youlu Su. The pathway of Edwardsiella piscicida infecting Lateolabrax maculatus via the immersion bath. Journal of Fish Diseases, 47(1):e13863:1–e13863:??, January 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

He:2021:IRS

[HWZ^{+}21] Le He, Luying Wang, Lingmin Zhao, Zhixia Zhuang, Xiaoru Wang, Huabin Huang, Qi Fu, Lixing Huang, Yingxue Qin, Pan Wang, and Qingpi Yan. Integration of RNAseq and RNAi reveals the contribution of znuA gene to the pathogenicity of Pseudomonas plecoglossicida and to the immune response of Epinephelus coioides. Journal of Fish Diseases, 44(11):1831–1841, November 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

He:2024:EEE

[HZC^{+}24] Runzhen He, Ningyu Zhu, Xiaoming Chen, Qianrong Liang, Gaohua Yao, Quanquan Tian, Fan Zhou, and Xueyan Ding. Experimental evidence of effective disinfectant to control the transmission of Micropterus salmoides rhabdovirus. Journal of Fish Diseases, 47(2):e13891:1–e13891:??, February 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hassantabar:2021:DEC

[HZFT21] Fatemeh Hassantabar, Mohammad J. Zorriehzahra, Farid Firouzbakhsh, and Kim D. Thompson. Development and evaluation of colloidal gold immunochromatography test strip for rapid diagnosis of nervous necrosis virus in golden grey mullet (Chelon aurata). Journal of Fish Diseases, 44 (6):783–791, June 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Irgang:2021:ETI

[IAH21] Rute Irgang and Ruben Avendaño-Herrera. Experimental tenacibaculosis infection in adult conger eel (Genypterus chilensis, Guichenot 1948) by immersion challenge with Tenacibaculum dicentrarchi. Journal of Fish Diseases, 44 (2):211–216, February 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Irgang:2022:EVS

[IAH22] Rute Irgang and Ruben Avendaño-Herrera. Evaluation of the in vitro susceptibility of Tenacibaculum dicentrarchi to tiamulin using minimum inhibitory concentration tests. Journal of Fish Diseases, 45(6):795–799, June 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Igeh:2020:PEC

[IAO20] Patience Chwe Igeh and Annemariè Avenant-Oldewage. Pathological effects of Cichlidogyrus philander Douëllou, 1993 (Monogenea, Ancyrocephalidae) on the gills of Pseudocrenilabrus philander (Weber, 1897) (Cichlidae). Journal of Fish Diseases, 43(2):177–184, February 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Izvekova:2021:LPI

[IFI^{+}21] Galina I. Izvekova, Tatyana V. Frolova, Evgeny I. Izvekov, Elena N. Kashinskaya, and Mikhail M. Solovyev. Localization of the proteinase inhibitor activity in the fish cestode Eubothrium rugosum. Journal of Fish Diseases, 44(12): 1951–1958, December 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Izvekova:2022:SFG

[IFIZ22] Galina I. Izvekova, Tatyana V. Frolova, Evgeny I. Izvekov, and Alexander E. Zhokhov. Surviving in the fish gut: Comparative inhibitory capacities against the host proteinases in cestodes of the genus Proteocephalus. Journal of Fish Diseases, 45(7):1011–1021, July 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ishihara:2022:WGA

[IHK^{+}22] Hikaru Ishihara, Shogo Harakawa, Hidemasa Kawakami, Keisuke Yoshii, Naoya Murase, Hidetoshi Yamada, Yutaka Fukuda, Reiko Nozaki, Satoshi Kawato, Keiichiro Koiwai, Ikuo Hirono, and Hidehiro Kondo. Whole-genome analysis of red sea bream iridovirus spread in 2021 in Japan provided epidemiological and viral traits insight. Journal of Fish Diseases, 45(10):1593–1597, October 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Inohana:2023:GMC

[IKT^{+}23] Mari Inohana, Takeshi Komine, Yoshiaki Tanaka, Osamu Kurata, and Shinpei Wada. Genital mycobacteriosis caused by Mycobacterium marinum detected in two captive sharks by peptide nucleic acid-fluorescence in situ hybridization. Journal of Fish Diseases, 46(1):47–59, January 2023. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Irgang:2021:FOS

[IMAH21] Rute Irgang, Marcos Mancilla, and Ruben Avendaño-Herrera. Florfenicol and oxytetracycline susceptibility patterns in Chilean isolates of Tenacibaculum dicentrarchi: an emerging pathogen for farmed salmonids. Journal of Fish Diseases, 44(7):1043–1046, July 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Imsonpang:2024:CTA

[IPC^{+}24] Supapong Imsonpang, Arnon Pudgerd, Charoonroj Chotwiwatthanakun, Jiraporn Srisala, Piyachat Sanguanrut, Chanadda Kasamechotchung, Kallaya Sritunyalucksana, Suparat Taengchaiyaphum, and Rapeepun Vanichviriyakit. Confirmatory test of active IHHNV infection in shrimp by immunohistochemistry and IHHNV–LongAmp PCR. Journal of Fish Diseases, 47(3):e13905:1–e13905:??, March 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ilardi:2023:COH

[IVR^{+}23] Pedro Ilardi, Sara Valdes, Javier Rivera, Rute Irgang, and Ruben Avendaño-Herrera. Co-occurrence of heterogeneous Flavobacterium psychrophilum isolates within the same Chilean farm and during the same infectious outbreak. Journal of Fish Diseases, 46(10):1085–1096, October 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jansson:2022:NLS

[JAC^{+}22] Eva Jansson, Anna Aspán, Arianna Comin, Maj Hjort, Tomas Jinnerot, and Charlotte Axén. Non-lethal sampling for the detection of Renibacterium salmoninarum by qPCR for diagnosis of bacterial kidney disease. Journal of Fish Diseases, 45(6):883–894, June 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jaramillo:2024:NZR

[JBB^{+}24] D. Jaramillo, B. P. Busby, M. Bestbier, P. Bennett, and Z. Waddington. New Zealand rickettsia-like organism and Tenacibaculum maritimum vaccine efficacy study. Journal of Fish Diseases, 47(2):e13883:1–e13883:??, February 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jia:2023:DRS

[JBG^{+}23] Beibei Jia, Holly Burnley, Ian A. Gardner, Matthew E. Saab, Adele Doucet, and K. Larry Hammell. Diagnosis of Renibacterium salmoninarum infection in harvested Atlantic salmon (Salmo salar L.) on the east coast of Canada: Clinical findings, sample collection methods and laboratory diagnostic tests. Journal of Fish Diseases, 46(5):575–589, May 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jones:2024:TFN

[JBP24] John Brian Jones, Cara L. Brosnahan, and Anjali Pande. Tail Fan Necrosis syndrome in decapod crustaceans: a review. Journal of Fish Diseases, 47(5):e13920:1–e13920:??, May 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jia:2020:RIA

[JDA^{+}20] Beibei Jia, Marina K. V. C. Delphino, Babafela Awosile, Tim Hewison, Patrick Whittaker, Diane Morrison, Mykolas Kamaitis, Ahmed Siah, Barry Milligan, Stewart C. Johnson, and Ian A. Gardner. Review of infectious agent occurrence in wild salmonids in British Columbia, Canada. Journal of Fish Diseases, 43(2):153–175, February 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jiang:2022:ESA

[JJZ^{+}22] Zhou Jiang, Yingying Jia, Jiangfan Zhang, Xuejun Li, and Chuanju Dong. Effect of secondary attack by Aeromonas hydrophila on the expression level of hif genes in common carp (Cyprinus carpio). Journal of Fish Diseases, 45(6):907– 917, June 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jensen:2020:GAF

[JKM^{+}20] Hannah Malene Jensen, Asma Mohammad Karami, Heidi Mathiessen, Azmi Al-Jubury, Per Walter Kania, and Kurt Buchmann. Gill amoebae from freshwater rainbow trout (Oncorhynchus mykiss): In vitro evaluation of antiparasitic compounds against Vannella sp. Journal of Fish Diseases, 43(6):665–672, June 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jin:2021:PSD

[JLH^{+}21] Jiamin Jin, Yiying Li, Mengxia Huang, Shanshan Li, and Zhijuan Mao. Preliminary studies on the different roles of T6SSs in pathogenicity of Pseudomonas plecoglossicida NB2011. Journal of Fish Diseases, 44(11):1669–1679, November 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jones:2020:FAS

[JLM^{+}20] Simon R. M. Jones, Amy Long, Christine MacWilliams, Mark Polinski, and Kyle Garver. Factors associated with severity of naturally occurring piscirickettsiosis in netpenand tank-reared juvenile Atlantic salmon at a research aquarium in western Canada. Journal of Fish Diseases, 43 (1):49–55, January 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jeong:2022:QKP

[JMA^{+}22] Jaewoon Jeong, Gregor F. McEwan, Gabriel Arriagada, Cristian Gallardo-Escárate, and Crawford W. Revie. Quantifying key parameters related to the life cycle of Caligus rogercresseyi. Journal of Fish Diseases, 45(1):219–224, January 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jaafar:2020:QTL

[JØM^{+}20] R. Jaafar, J. Ødegård, H. Mathiessen, A. M. Karami, M. H. Marana, L. von Gersdorff Jørgensen, S. Zuo, T. Nielsen, P. W. Kania, and K. Buchmann. Quantitative trait loci (QTL) associated with resistance of rainbow trout Oncorhynchus mykiss against the parasitic ciliate Ichthyophthirius multifiliis. Journal of Fish Diseases, 43(12):1591– 1602, December 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jia:2021:MEP

[JPC^{+}21] Peng Jia, Hongbo Pan, Kuopeng Cui, Kuntong Jia, and Meisheng Yi. MicroRNA expression profiling of sea perch brain cells reveals the roles of microRNAs in autophagy induced by RGNNV infection. Journal of Fish Diseases, 44 (9):1305–1314, September 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jensen:2020:STV

[JQT20] Britt Bang Jensen, Lars Qviller, and Nils Toft. Spatiotemporal variations in mortality during the seawater production phase of Atlantic salmon (Salmo salar) in Norway. Journal of Fish Diseases, 43(4):445–457, April 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jiang:2022:SEH

[JRZ^{+}22] Xin-Yuan Jiang, Zong-Yi Ren, Nian-Kun Zhang, Ke-Chen Yang, Gao-Xue Wang, and Hai-Feng Jiang. Screening and evaluating honokiol from Magnolia officinalis against Nocardia seriolae infection in largemouth bass (Micropterus salmoides). Journal of Fish Diseases, 45(11):1599–1607, November 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Johnston:2021:CDS

[JSW^{+}21] Hazel Johnston, Jane Symonds, Seumas Walker, Mark Preece, Cesar Lopez, and Barbara Nowak. Case definitions for skin lesion syndromes in Chinook salmon farmed in Marlborough Sounds, New Zealand. Journal of Fish Diseases, 44 (2):141–147, February 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jiang:2023:DDD

[JSZ^{+}23] Nan Jiang, Jinyu Shen, Yong Zhou, Wenzhi Liu, Yan Meng, Yiqun Li, Mingyang Xue, Chen Xu, and Yuding Fan. Development of a droplet digital PCR method for the sensitive detection and quantification of largemouth bass ranavirus. Journal of Fish Diseases, 46(2):91–98, February 2023. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jyotsna:2020:SBN

[JVM^{+}20] Jyotsna, Parameswaran Vijayakumar, Ravi Mani, Sudhakaran Raja, and Tharmathass Stalin Dhas. Susceptibility of betanodavirus in a newly established vertebra-derived cell line from Mosquitofish (Gambusia affinis). Journal of Fish Diseases, 43(2):263–273, February 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Janik:2022:DSM

[JW22] Andrew J Janik and Christopher M. Whipps. Differences in susceptibility to Mycobacterium chelonae in zebrafish (Danio rerio) lines commonly used in scientific research. Journal of Fish Diseases, 45(3):435–443, March 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jia:2022:AES

[JXJ^{+}22] Yi-Jun Jia, Jun-Yao Xia, Fu-Yi Jiang, Yang Li, Guo Chen, and Bin Zhu. Antigenic epitope screening and functional modification of mannose enhance the efficacy of largemouth bass virus subunit vaccines. Journal of Fish Diseases, 45 (11):1635–1643, November 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jiang:2023:ECL

[JZW^{+}23] Huan Jiang, Youyi Zhang, Xiangbing Wang, Guanlin Wang, Jiajie Zhu, Jiachu Sun, Man Zhang, Yaya Li, Shanliang Xu, Jiabao Hu, and Yajun Wang. Establishment and characterization of a liver cell line from silver pomfret (Pampus argenteus) for studying fish health. Journal of Fish Diseases, 46(11):1193–1205, November 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kabuusu:2024:HCO

[KAS^{+}24] R. M. Kabuusu, T. A. Aire, D. F. Stroup, C. N. L. Macpherson, S. Beltran, E. Reyes, and H. W. Ferguson. Hematologic changes observed in syncytial hepatitis of farmed tilapia (SHT), Oreochromis niloticus. Journal of Fish Diseases, 47(1):e13872:1–e13872:??, January 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Korkea-aho:2021:GDP

[KaVDH^{+}21] Tiina Liisa Korkea-aho, Satu Viljamaa-Dirks, Sirpa Heinikainen, Henry Kuronen, and Marja Tiirola. Genetic diversity and phenotypic characterization of Iodobacter limnosediminis associated with skin lesions in freshwater fish. Journal of Fish Diseases, 44(11):1711–1724, November 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ksepka:2021:MPP

[KB21] Steven P. Ksepka and Stephen A. Bullard. Morphology, phylogenetics and pathology of “red sore disease” (coinfection by Epistylis cf. wuhanensis and Aeromonas hydrophila) on sportfishes from reservoirs in the South-Eastern United States. Journal of Fish Diseases, 44(5):541–551, May 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Khalil:2023:IPR

[KBG^{+}23] Sarker Mohammed Ibrahim Khalil, Chiara Bulfon, Marco Galeotti, Pier Luigi Acutis, Ilhan Altinok, Charalampos Kotzamanidis, Ana Isabel Vela, Lucio Fariano, Marino Prearo, Silvia Colussi, and Donatella Volpatti. Immune profiling of rainbow trout (Oncorhynchus mykiss) exposed to Lactococcus garvieae: Evidence in asymptomatic versus symptomatic or vaccinated fish. Journal of Fish Diseases, 46 (7):731–741, July 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kutyrev:2021:ESU

[KBMD21] Ivan Alexandrovich Kutyrev, Natalia Mikhailovna Biserova, Olga Evgenievna Mazur, and Zhargal Nimaevich Dugarov. Experimental study of ultrastructural mechanisms and kinetics of tegumental secretion in cestodes parasitizing fish (Cestoda: Diphyllobothriidea). Journal of Fish Diseases, 44 (8):1237–1254, August 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Klykken:2022:GEP

[KBR^{+}22] Christine Klykken, Lauris Boissonnot, Anne Katrine Reed, Paul Whatmore, Kari Attramadal, and Rolf Erik Olsen. Gene expression patterns in Atlantic salmon (Salmo salar) with severe nephrocalcinosis. Journal of Fish Diseases, 45 (11):1645–1658, November 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kayansamruaj:2023:GDP

[KDHS^{+}23] Pattanapon Kayansamruaj, Nguyen Dinh-Hung, Prapansak Srisapoome, Uthairat Na-Nakorn, and Satid Chatchaiphan. Genomics-driven prophylactic measures to increase streptococcosis resistance in tilapia. Journal of Fish Diseases, 46 (6):597–610, June 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Klykken:2022:RDN

[KDR^{+}22] Christine Klykken, Alf Seljenes Dalum, Anne Katrine Reed, Kari Attramadal, Rolf Erik Olsen, and Lauris Boissonnot. Radiological detection of nephrocalcinosis in farmed Atlantic salmon Salmo salar L. Journal of Fish Diseases, 45(12): 1883–1888, December 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kampeera:2021:RTL

[KDS^{+}21] Jantana Kampeera, Sirintip Dangtip, Rapheephat Suvannakad, Pakapreud Khumwan, Saengchan Senapin, and Wansika Kiatpathomchai. Reverse transcription loop-mediated isothermal amplification (RT-LAMP) combined with colorimetric gold nanoparticle (AuNP) probe assay for visual detection of tilapia lake virus (TiLV) in Nile and red hybrid tilapia. Journal of Fish Diseases, 44(10):1595–1607, October 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kotze:2020:MHT

[KH20] Sanet H. Kotzé and Ann Huysseune. Mucin histochemistry as a tool to assess rostral digestive tract health in a teleost model (Danio rerio). Journal of Fish Diseases, 43(12):1603– 1606, December 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Knupp:2021:HSP

[KKBL21] Christopher Knupp, Matti Kiupel, Travis O. Brenden, and Thomas P. Loch. Host-specific preference of some Flavobacterium psychrophilum multilocus sequence typing genotypes determines their ability to cause bacterial coldwater disease in Coho salmon (Oncorhynchus kisutch). Journal of Fish Diseases, 44(5):521–531, May 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Katharios:2023:FKA

[KKC^{+}23] Pantelis Katharios, Constantina Kokkari, Maria Chiara Cascarano, Stylianos Gyparakis, Chrissi Aggelidaki, Ioanna Antonakaki, and Efthimia Vargiakaki. A fish kill at the Aposelemis Dam (Crete, Greece) caused by heavy parasitism by Ichthyobodo sp. Journal of Fish Diseases, 46(1):61–66, January 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kurobe:2024:MME

[KKH^{+}24] Tomofumi Kurobe, Jun Kurita, Olga Haenen, Michal Voorbergen-Laarman, and Takafumi Ito. Mass mortality events associated with cyprinid herpesvirus 2 (CyHV-2) infection in wild Prussian carp Carassius gibelio in The Netherlands, and molecular biology of virus strains. Journal of Fish Diseases, 47(1):e13868:1–e13868:??, January 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Klykken:2023:NJF

[KKK^{+}23] C. Klykken, E. Khan, C. Karlsen, A. K. Reed, K. J. K. Attramadal, R. E. Olsen, and L. Boissonnot. Nephrocalcinosis in juvenile farmed Atlantic salmon (Salmo salar) may be linked to osmoregulatory stress. Journal of Fish Diseases, 46(9):943–956, September 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kverme:2022:ATL

[KKL^{+}22] Karen O. Kverme, Marielle Kallekleiv, Kristina Larsen, Anita Rønneseth, Heidrun I. Wergeland, Ole B. Samuelsen, and Gyri T. Haugland. Antibacterial treatment of lumpfish (Cyclopterus lumpus) experimentally challenged with Vibrio anguillarum, atypical Aeromonas salmonicida and Pasteurella atlantica. Journal of Fish Diseases, 45(1):153–163, January 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kent:2021:IIC

[KMH^{+}21] Michael L. Kent, Katrina N. Murray, Maurine R. Hobbs, Louis M. Weiss, Sean T. Spagnoli, and Justin L. Sanders. Intranuclear inclusions consistent with a Nucleospora sp. in a lymphoid lesion in a laboratory zebrafish, Danio rerio (Hamilton 1822). Journal of Fish Diseases, 44(1):107–112, January 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kim:2022:MMC

[KNG^{+}22] Van Van Kim, Hung Manh Nguyen, Stephen E. Greiman, Ha Van Nguyen, Chinh Ngoc Nguyen, Manh Duc Vu, Truong Dinh Hoai, and Henry Madsen. Molecular and morphological characterization of Dollfustrema bagarii (Digenea: Bucephalidae) metacercariae from aquaculture channel catfish (Ictalurus punctatus) in northern Vietnam. Journal of Fish Diseases, 45(8):1165–1171, August 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kaplan:2022:FDB

[KPC^{+}22] Murat Kaplan, Kemal Pekmez, Abdurrahman Anil Cagirgan, Fatih Arslan, Buket Özkan, and Gülnur Kalaycı. The first detection of betanodavirus reassortant genotype (RGNNV/SJNNV) isolated from gilthead sea bream (Sparus aurata) in the Turkish coastlines: The importance of screening and monitoring studies for identifying the source of the infection. Journal of Fish Diseases, 45(6):783–793, June 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kong:2021:FCT

[KQW^{+}21] Linghe Kong, Kun Qian, Siwei Wu, Bingxi Li, Zheng Guo, Xiaoxue Yin, Yu Huang, Jianmin Ye, Xiao Tu, and Shengli Fu. Functional characterization of TNF-α in pufferfish (Takifugu obscurus) in immune response and apoptosis against Aeromonas hydrophila. Journal of Fish Diseases, 44(9): 1343–1353, September 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ksepka:2020:UGD

[KRS^{+}20] Steven P. Ksepka, Jacob M. Rash, Brandon L. Simcox, Doug A. Besler, Haley R. Dutton, Micah B. Warren, and Stephen A. Bullard. An updated geographic distribution of Myxobolus cerebralis (Hofer, 1903) (Bivalvulida: Myxobolidae) and the first diagnosed case of whirling disease in wildcaught trout in the south-eastern United States. Journal of Fish Diseases, 43(7):813–820, July 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kayansamruaj:2020:DGS

[KSD^{+}20] Pattanapon Kayansamruaj, Chayanit Soontara, Ha T. Dong, Kornsunee Phiwsaiya, and Saengchan Senapin. Draft genome sequence of scale drop disease virus (SDDV) retrieved from metagenomic investigation of infected barramundi, Lates calcarifer (Bloch, 1790). Journal of Fish Diseases, 43(10):1287–1298, October 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Koda:2023:GCI

[KSG^{+}23] Samantha A. Koda, Kuttichantran Subramaniam, Joseph M. Groff, Roy P. Yanong, Deborah B. Pouder, Matt Pedersen, Craig Pelton, Michael M. Garner, Nicholas B. D. Phelps, Anibal G. Armien, Michael W. Hyatt, Paul M. Hick, Joy A. Becker, Mark F. Stidworthy, and Thomas B. Waltzek. Genetic characterization of infectious spleen and kidney necrosis virus in Banggai cardinalfish Pterapogon kauderni identified from eight separate cases between 2000 and 2017. Journal of Fish Diseases, 46(7):795–802, July 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kashinskaya:2020:CMC

[KSI^{+}20] Elena N. Kashinskaya, Evgeniy P. Simonov, Galina I. Izvekova, Aleksey N. Parshukov, Karl B. Andree, and Mikhail M. Solovyev. Composition of the microbial communities in the gastrointestinal tract of perch (Perca fluviatilis L. 1758) and cestodes parasitizing the perch digestive tract. Journal of Fish Diseases, 43(1):23–38, January 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kumar:2021:VPI

[KSM^{+}21] S. Santhosh Kumar, S. Sivakumar, S. Abdul Majeed, S. Vimal, G. Taju, and A. S. Sahul Hameed. In vitro propagation of infectious myonecrosis virus in C6/ 36 mosquito cell line. Journal of Fish Diseases, 44(7):987–992, July 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kent:2020:RDH

[KSS^{+}20] M. L. Kent, J. L. Sanders, S. Spagnoli, C. E. Al-Samarrie, and K. N. Murray. Review of diseases and health management in zebrafish Danio rerio (Hamilton 1822) in research facilities. Journal of Fish Diseases, 43(6):637–650, June 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Komine:2023:CSN

[KSY^{+}23] Takeshi Komine, Saralee Srivorakul, Mitsumi Yoshida, Yoshiaki Tanaka, Yuma Sugimoto, Mari Inohana, Hanako Fukano, Yoshihiko Hoshino, Osamu Kurata, and Shinpei Wada. Core single nucleotide polymorphism analysis reveals transmission of Mycobacterium marinum between animal and environmental sources in two aquaria. Journal of Fish Diseases, 46(5):507–516, May 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Levipan:2021:AIS

[LAH21] Héctor A. Levipan and Ruben Avendaño-Herrera. Assessing the impacts of skin mucus from Salmo salar and Oncorhynchus mykiss on the growth and in vitro infectivity of the fish pathogen Piscirickettsia salmonis. Journal of Fish Diseases, 44(2):181–190, February 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lee:2022:CIN

[LCCC22] Chi-Fen Lee, Yen-Chen Chang, Hue-Ying Chiou, and Hui-Wen Chang. Concurrent infection of a novel genotype of hepatopancreatic parvovirus and Enterocytozoon hepatopenaei in Penaeus vannamei in Taiwan. Journal of Fish Diseases, 45(8):1201–1210, August 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Legario:2020:IMC

[LCTC20] Francis S. Legario, Casiano H. Choresca, Jr., Jimmy F. Turnbull, and Margaret Crumlish. Isolation and molecular characterization of streptococcal species recovered from clinical infections in farmed Nile tilapia (Oreochromis niloticus) in the Philippines. Journal of Fish Diseases, 43(11):1431–1442, November 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Liu:2024:RSD

[LCW^{+}24] Xiaoru Liu, Yong Cao, Jiayin Wang, Suyuheng Cao, Liqun Lu, and Yousheng Jiang. Rapid and sensitive detection of large yellow croaker iridovirus by real-time RPA and RPA-LFD. Journal of Fish Diseases, 47(6):e13930:1–e13930:??, June 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lu:2020:MEV

[LCYL20] Tien-Hsuan Lu, Chi-Yun Chen, Ying-Fei Yang, and Chung-Min Liao. Modelling the effect of vaccination on transmission dynamics of nervous necrosis virus in grouper larvae Epinephelus coioides. Journal of Fish Diseases, 43(10): 1155–1165, October 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

LimaMoraes:2023:LHA

[LdAPA^{+}23] José Lima Moraes, Sanny Maria de Andrade-Porto, Eduardo Akifumi Ono, Gustavo Da Silva Claudiano, and Elizabeth Gusmão Affonso. Levamisole hydrochloride as an anthelmintic against Neoechinorhynchus buttnerae in juvenile Colossoma macropomum. Journal of Fish Diseases, 46(8): 803–811, August 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Louboutin:2022:CNP

[LDC^{+}22] Lénaïg Louboutin, Nolwenn M Dheilly, Joëlle Cabon, Sara Picon Camacho, Aurélie Leroux, Pierrick Lucas, Alain Le Breton, Yannick Blanchard, and Thierry Morin. Characterization of a novel picornavirus isolated from moribund gilthead seabream (Sparus aurata) larvae. Journal of Fish Diseases, 45(5):707–716, May 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lloyd:2020:ULN

[LdL20] Richard Lloyd and Carlos Garcia de Leaniz. The use of laparoscopy for non-destructive disease screening of broodstock Atlantic lumpfish, Cyclopterus lumpus Linnaeus. Journal of Fish Diseases, 43(9):1107–1110, September 2020. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lampou:2020:IRE

[LDM^{+}20] Eirini Lampou, Chrysostomos Dovas, Maritsa Margaroni, Taxiarchis Chasalevris, Ioannis S. Pappas, Eleni Dotsika, Evdokia Karagouni, Fotini Athanassopoulou, Dimitrios Katsaras, and Konstantina Bitchava. Investigation of routes of entry and dispersal pattern of RGNNV in tissues of European sea bass, Dicentrarchus labrax. Journal of Fish Diseases, 43(11):1363–1371, November 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Liang:2021:PRM

[LDO^{+}21] QingJian Liang, WenNa Dong, MuFei Ou, ZhongHua Li, Yuan Liu, and Weina Wang. Potential roles for microRNAs in facilitating physiological adaptation to low-temperature stress in Penaeus vannamei. Journal of Fish Diseases, 44 (8):1191–1200, August 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Laurin:2022:IDD

[LDR^{+}22] Emilie Laurin, Marina K. V. C. Delphino, Raynalfie Budhy Rahardio, Lukman Hakim, Wildan Gayuh Zulfikar, Holly Burnley, Keith Larry Hammell, and Krishna Kumar Thakur. Infectious disease detection associated with trends in production, environmental and biosecurity factors for shrimp (Litopenaeus vannamei and Penaeus monodon) production systems in Banyuwangi, Indonesia. Journal of Fish Diseases, 45(8):1231–1236, August 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lin:2023:IEE

[LDW^{+}23] Guixiang Lin, Fan Da, Xiaoju Wan, Yucong Huang, Shiping Yang, Jichang Jian, and Shuanghu Cai. Immune-enhancing effects of Astragalus polysaccharides and Ganoderma lucidum polysaccharides on Vibrio harveyi flgJ DNA vaccine in grouper. Journal of Fish Diseases, 46(2):147–156, February 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lennox:2020:HPA

[LEV^{+}20] Robert J. Lennox, Sindre H. Eldøy, Knut W. Vollset, Kristi M. Miller, Shaorong Li, Karia H. Kaukinen, Trond E. Isaksen, and Jan G. Davidsen. How pathogens affect the marine habitat use and migration of sea trout (Salmo trutta) in two Norwegian fjord systems. Journal of Fish Diseases, 43 (7):729–746, July 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lange:2020:VMS

[LFA20] Miles D. Lange, Bradley D. Farmer, and Jason Abernathy. Vertebrate mucus stimulates biofilm development and upregulates iron acquisition genes in Flavobacterium columnare. Journal of Fish Diseases, 43(1):101–110, January 2020. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Long:2023:CGA

[LFG^{+}23] Meng Long, Huimin Fan, Zhen Gan, Zenghai Jiang, Shaoshuai Tang, Hongli Xia, and Yishan Lu. Comparative genomic analysis provides insights into taxonomy and temperature adaption of Aeromonas salmonicida. Journal of Fish Diseases, 46(5):545–561, May 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Li:2021:ARA

[LFL^{+}21] Xiaohao Li, Kunpeng Fan, Yafang Liu, Ying Liu, and Peng fei Liu. Administration of a recombinant ALDH7A1 (rA7) indicates potential regulation of the metabolite and immunology pathways in Atlantic salmon infected with Aeromonas salmonicida. Journal of Fish Diseases, 44(7): 961–977, July 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lovett:2024:MRI

[LFP^{+}24] Bailey A. Lovett, Elwyn C. Firth, Matthew R. Perrott, John S. Munday, Beau P. Pontre, Anna-Maria P. Lydon, Jane E. Symonds, Mark A. Preece, and Neill A. Herbert. Magnetic resonance imaging shows spinal curvature in Chinook salmon (Oncorhynchus tshawytscha) is associated with chronic inflammation of peri-vertebral soft tissues. Journal of Fish Diseases, 47(3):e13900:1–e13900:??, March 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lewisch:2021:EDS

[LFSB^{+}21] Eva Lewisch, Hans-Peter Führer, Bita Shahi-Barogh, Josef Harl, and Mansour El-Matbouli. Emergence of Discocotyle sagittata (Monogenea: Polyopisthocotylea) in rainbow trout (Oncorhynchus mykiss) and brown trout (Salmo trutta) in an Austrian aquarium. Journal of Fish Diseases, 44(10): 1643–1646, October 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Li:2021:ICG

[LFZ^{+}21] Shuo Li, Yu Feng, Tongtong Zhang, Shan Wang, and Jinsheng Sun. Identification and characterization of Trpm4 gene involved in regulating Japanese flounder (Paralichthys olivaceus) inflammatory response. Journal of Fish Diseases, 44(11):1765–1776, November 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Liu:2022:SCL

[LFZ^{+}22] Wenzhi Liu, Yuding Fan, Yong Zhou, Nan Jiang, Yiqun Li, Yan Meng, Mingyang Xue, Zhong Li, and Lingbing Zeng. Susceptibility of a cell line derived from the kidney of Chinese rice-field eel, Monopterus albus to the infection of rhabdovirus, CrERV. Journal of Fish Diseases, 45(2):361–371, February 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Liu:2021:TPF

[LGH^{+}21] Hongtao Liu, Shengtao Guo, Yugui He, Qiong Shi, Mingqiu Yang, and Xinxin You. Toll protein family structure, evolution and response of the whiteleg shrimp (Litopenaeus vannamei) to exogenous iridescent virus. Journal of Fish Diseases, 44(8):1131–1145, August 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Long:2021:DPS

[LGJ21] Amy Long, Aidan Goodall, and Simon R. M. Jones. Development of a Piscirickettsia salmonis immersion challenge model to investigate the comparative susceptibility of three salmon species. Journal of Fish Diseases, 44(1):1–9, January 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Laurin:2020:BED

[LGP^{+}20] Emilie Laurin, Ian A. Gardner, Andrea Peña, Marco Rozas-Serri, Jorge Gayosa, Joaquin Neumann Heise, and Fernando O. Mardones. Bayesian estimation of diagnostic sensitivity and specificity of a qPCR and a bacteriological culture method for Piscirickettsia salmonis in farmed Atlantic salmon (Salmo salar L.) in Chile. Journal of Fish Diseases, 43(10):1167–1175, October 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lim:2020:CAS

[LH20] Jongwon Lim and Suhee Hong. Characterization of Aeromonas salmonicida and A. sobria isolated from cultured salmonid fish in Korea and development of a vaccine against furunculosis. Journal of Fish Diseases, 43(5):609–620, May 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Li:2022:EAL

[LHC^{+}22] Yuan Li, Yu Huang, Jia Cai, Dongneng Jiang, Ji chang Jian, Yi shan Lu, and Bei Wang. Establishment of an astrocytelike cell line from the brain of tilapia (Oreochromis niloticus) for virus pathogenesis and a vitro model of the blood–brain barrier. Journal of Fish Diseases, 45(10):1451–1462, October 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lepak:2022:RPP

[LHH^{+}22] Jesse M. Lepak, Adam G. Hansen, Mevin B. Hooten, Daniel Brauch, and Estevan M. Vigil. Rapid proliferation of the parasitic copepod, Salmincola californiensis (Dana), on Kokanee salmon, Oncorhynchus nerka (Walbaum), in a large Colorado reservoir. Journal of Fish Diseases, 45(1):89–98, January 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Liang:2020:FRC

[LHLS20] Jing Liang, Xiucai Hu, Aijun Lü, and Jingfeng Sun. First report on the characterization of pathogenic Rahnella aquatilis KCL-5 from crucian carp: Revealed by genomic and proteomic analyses. Journal of Fish Diseases, 43(8):889–914, August 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lima:2021:ANP

[LHTW21] Paula C. Lima, Lauren Hartley-Tassell, and James W. Wynne. The ability of Neoparamoeba perurans to bind to and digest non-fish-derived mucin: Insights into the amoeba’s mechanism of action to overcome gill mucus production. Journal of Fish Diseases, 44(9):1355–1367, September 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Li:2023:DTN

[Li23] Meng-Syun Li. Discovery of two novel tilapia lake viruslike virus isolates in the transcriptomic data of guppy fish (Poecilia reticulata). Journal of Fish Diseases, 46(9):1015– 1019, September 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lim:2023:MMV

[LJHH23] Jongwon Lim, Yoonyoung Jang, Hyun-Ja Han, and Suhee Hong. Molecular mechanisms of the virulence and efficacy of a highly virulent Vibrio anguillarum strain and its formalininactivated vaccine in rainbow trout. Journal of Fish Diseases, 46(5):563–574, May 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lee:2024:CSS

[LJK^{+}24] Young Juhn Lee, Lyu Jin Jun, Ye Ji Kim, Ji Eun Han, Ye Jin Ko, Yeong Eun Oh, Eung Joon Lee, Jehee Lee, and Joon Bum Jeong. Comparison of specificity and sensitivity of diagnostic methods for Enteromyxum leei and Enteromyxum fugu detected from cultured tiger puffer, Takifugu rubripes in Korea. Journal of Fish Diseases, 47(1):e13865:1–e13865:??, January 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Li:2022:ECL

[LjY^{+}22] Yong Li, Peng jia, Fangzhao Yu, Wangdong Li, Can Mao, Meisheng Yi, Qunhong Gu, and Kuntong Jia. Establishment and characterization of a liver cell line, ALL, derived from yellowfin sea bream, Acanthopagrus latus, and its application to fish virology. Journal of Fish Diseases, 45(1):141–151, January 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Leis:2024:NAI

[LKD^{+}24] Eric Leis, Ryan Katona, Sara Dziki, Rebekah McCann, Isaac Standish, Eryn Opgenorth, Audrey Dikkeboom, Ailam Lim, Charlotte E. Ford, and Tony L. Goldberg. Novel Aquareovirus isolated from channel catfish (Ictalurus punctatus) used in mussel restoration efforts in Wisconsin. Journal of Fish Diseases, 47(6):e13938:1–e13938:??, June 2024. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Le:2022:DSP

[LKP^{+}22] Bao Le, Dong-Geun Kim, Nguyen Ngoc Phuoc, Nguyen Thi Hue Linh, and Seung-Hwan Yang. Dietary supplementation with Pediococcus pentosaceus enhances the innate immune response in and promotes growth of Litopenaeus vannamei shrimp. Journal of Fish Diseases, 45(9):1343–1354, September 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lauringson:2023:SIH

[LKV^{+}23] Magnus Lauringson, Siim Kahar, Taigor Veevo, Maidu Silm, Duncan Philpott, Roland Svirgsden, Mehis Rohtla, Priit Päkk, Riho Gross, Tanel Kaart, and Anti Vasemägi. Spatial and intra-host distribution of myxozoan parasite Tetracapsuloides bryosalmonae among Baltic sea trout (Salmo trutta). Journal of Fish Diseases, 46(10):1073–1083, October 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Liu:2021:IFR

[LL21] Juan-Ting Liu and John S. Lumsden. Impact of feed restriction, chloroquine and deoxynivalenol on viral haemorrhagic septicaemia virus IVb in fathead minnow Pimephales promelas Rafinesque. Journal of Fish Diseases, 44(2):217–220, February 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Li:2021:MFI

[LLC^{+}21] Kexin Li, Wanru Li, Xiaojuan Chen, Tian Luo, Yinnan Mu, and Xinhua Chen. Molecular and functional identification of a β-defensin homolog in large yellow croaker (Larimichthys crocea). Journal of Fish Diseases, 44(4):391–400, April 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lu:2021:NPN

[LLJ^{+}21] Jian-Fei Lu, Sheng Luo, Tian-Cheng Jin, Xin-Jiang Lu, and Jiong Chen. Nonstructural protein NS26 of grass carp reovirus is a principal regulator for viral replication and infection. Journal of Fish Diseases, 44(5):661–664, May 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Liu:2020:ICV

[LLR^{+}20] Xiao Liu, Yuhua Lei, Zhiming Ren, Suming Zhou, Dong Qian, Youbin Yu, Fei Yin, and Chunlin Wang. Isolation, characterization and virulence of Mesanophrys sp. (Ciliophora: Orchitophryidae) in farmed swimming crab (Portunus trituberculatus) in eastern China. Journal of Fish Diseases, 43(11):1419–1429, November 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Liu:2024:ETP

[lLWC^{+}24] Quan lin Liu, Yuan Wang, Jie Chen, Guo qing Pan, Yan-Feng Yue, Ze-Yang Zhou, and Wen hong Fang. Establishment of a TaqMan probe-based qPCR assay for detecting microsporidia Enterospora epinepheli in grouper. Journal of Fish Diseases, 47(3):e13893:1–e13893:??, March 2024. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lagadec:2024:MAS

[LMJ^{+}24] Erwan Lagadec, Even Bysveen Mjølnerød, Øyvind Marius Jensen, Heidrun Plarre, and Are Nylund. Multiple Aeromonas strains isolated from Atlantic salmon (Salmo salar) displaying red skin disease signs in Scandinavian rivers. Journal of Fish Diseases, 47(1):e13870:1–e13870:??, January 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lauringson:2021:CCD

[LNK^{+}21] Magnus Lauringson, Ilkka Nousiainen, Siim Kahar, Oksana Burimski, Riho Gross, Tanel Kaart, and Anti Vasemägi. Climate change-driven disease in sympatric hosts: Temporal dynamics of parasite burden and proliferative kidney disease in wild brown trout and Atlantic salmon. Journal of Fish Diseases, 44(6):689–699, June 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lopez-Porras:2021:GVE

[LPGA^{+}21] Adrián López-Porras, Matt J. Griffin, Abigail R. Armwood, Alvin C. Camus, Geoffrey C. Waldbieser, Cynthia Ware, Bradley Richardson, Terrence E. Greenway, Thomas Graham Rosser, Suja Aarattuthodiyil, and David J. Wise. Genetic variability of Edwardsiella piscicida isolates from Mississippi catfish aquaculture with an assessment of virulence in channel and channel × blue hybrid catfish. Journal of Fish Diseases, 44(11):1725–1751, November 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lopez-Porras:2022:CPE

[LPGW^{+}22] Adrián López-Porras, Matt J. Griffin, Cynthia Ware, Bradley M. Richardson, Terrence E. Greenway, Thomas Graham Rosser, Suja Aarattuthodiyi, and David J. Wise. Crossprotective efficacy of a live-attenuated Edwardsiella ictaluri vaccine against heterologous Edwardsiella piscicida isolates in channel and channel × blue catfish hybrids. Journal of Fish Diseases, 45(7):1001–1010, July 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Liu:2022:AMR

[LPL22] Juan-Ting Liu, Phuc H. Pham, and John S. Lumsden. Autophagy modulation in rainbow trout Oncorhynchus mykiss L. and resistance to experimental infection with Flavobacterium psychrophilum. Journal of Fish Diseases, 45(4):535– 545, April 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Liu:2020:VII

[LPW^{+}20] Juan-Ting Liu, Phuc H. Pham, Sarah K. Wootton, Niels C. Bols, and John S. Lumsden. VHSV IVb infection and autophagy modulation in the rainbow trout gill epithelial cell line RTgill-W1. Journal of Fish Diseases, 43(10):1237–1247, October 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Levipan:2021:PSD

[LRGAH21] Héctor A. Levipan, Luis Reyes-Garcia, and Ruben Avendaño-Herrera. Piscirickettsia salmonis does not evidence quorum sensing based on acyl-homoserine lactones. Journal of Fish Diseases, 44(7):1047–1051, July 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Li:2024:PGI

[LS24] Yuerui Li and Zhongshi Sun. Phenotypic and genomic insights into the pathogenicity and antimicrobial resistance of an Enterobacter roggenkampii strain isolated from diseased silver arowana (Osteoglossum bicirrhosum). Journal of Fish Diseases, 47(3):e13898:1–e13898:??, March 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Li:2021:IAT

[LTW^{+}21] Jie Li, Lei Tang, Pengmei Wang, Guiyang Li, Huaiyuan Jin, and Zhaolan Mo. Identification and application of T3SS translocation signal in Edwardsiella piscicida attenuated carrier as a bivalent vaccine. Journal of Fish Diseases, 44(5):513–520, May 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Li:2021:DDA

[LWH^{+}21] Yingying Li, Qing Wang, Feng Hu, Yingying Wang, Sven M. Bergmann, Weiwei Zeng, Jiyuan Yin, and Cunbin Shi. Development of a double-antibody sandwich enzyme-linked immunosorbent assay (DAS-ELISA) for the detection of KHV. Journal of Fish Diseases, 44(7):913–921, July 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lee:2020:EAS

[LWK20] Yih Nin Lee, Shinpei Wada, and Osamu Kurata. Effect of age on the susceptibility of marbled rockfish, Sebastiscus marmoratus to Ochroconis humicola following experimental challenge. Journal of Fish Diseases, 43(3):399–401, March 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Liao:2024:PCI

[LWL^{+}24] Wenyu Liao, Dongdong Wei, Mingzhu Liu, Lin Huang, Bingzheng Li, Yunyi Wei, Shuyu Han, Shuaishuai Huang, Qing Yu, and Pengfei Li. Phenotypic characteristics and immune response of Procypris merus following challenge with aquatic isolate of Klebsiella pneumoniae. Journal of Fish Diseases, 47(1):e13875:1–e13875:??, January 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Li:2020:OVW

[LWR^{+}20] Chengwei Li, Shenglan Wang, Qiulei Ren, Tianliang He, and Xinhua Chen. An outbreak of visceral white nodules disease caused by Pseudomonas plecoglossicida at a water temperature of 12^{o}C in cultured large yellow croaker (Larimichthys crocea) in China. Journal of Fish Diseases, 43(11):1353– 1361, November 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Liu:2022:DTV

[LWT^{+}22] Tao Liu, Meijiao Wu, Yachen Tian, Xiaoting Liu, Cheng Liu, Qingqiang Xu, and Qing Liu. Detection of trh^{+} Vibrio parahaemolyticus in seafood by lac dye coloration-based labelfree lateral flow immunoassay strip. Journal of Fish Diseases, 45(11):1699–1709, November 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Liu:2020:ABH

[LXW^{+}20] Mingzhu Liu, Hehe Xiao, Siting Wu, Qing Yu, and Pengfei Li. Aptamer-based high-throughput screening model for medicinal plant drugs against SGIV. Journal of Fish Diseases, 43(11):1479–1482, November 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Li:2021:CNS

[LXW^{+}21] Ge Li, Guosi Xie, Hailiang Wang, Xiaoyuan Wan, Xinshu Li, Chengyin Shi, Ziyan Wang, Miao Gong, Ting Li, Ping Wang, Qingli Zhang, and Jie Huang. Characterization of a novel shrimp pathogen, Vibrio brasiliensis, isolated from Pacific white shrimp, Penaeus vannamei. Journal of Fish Diseases, 44(10):1543–1552, October 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Liu:2021:VAC

[LXZ^{+}21] Huan Liu, Miao Xiao, Jing Zuo, Xiaoxian He, Ping Lu, Yingyu Li, Yanni Zhao, and Fei Xia. Vanillic acid combats Vibrio alginolyticus by cell membrane damage and biofilm reduction. Journal of Fish Diseases, 44(11):1799– 1809, November 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lu:2020:QIT

[LYC^{+}20] Tien-Hsuan Lu, Ying-Fei Yang, Chi-Yun Chen, Wei-Ming Wang, and Chung-Min Liao. Quantifying the impact of temperature variation on birnavirus transmission dynamics in hard clams Meretrix lusoria. Journal of Fish Diseases, 43 (1):57–68, January 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Luo:2021:IAV

[LYF^{+}21] Yang Luo, Lu Yu, Zizhao Feng, Qikang Chen, Liqun Lu, Qiya Zhang, and Dan Xu. Integrated analysis of viral miR-NAs, mRNA and protein in the caudal fin cells of C. auratus gibelio with cyprinid herpesvirus 2 infection. Journal of Fish Diseases, 44(4):441–460, April 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Liu:2020:AAI

[LYX^{+}20] Mingzhu Liu, Qing Yu, Hehe Xiao, Yi Yi, Hao Cheng, Dedi Fazriansyah Putra, Yaming Huang, Qin Zhang, and Pengfei Li. Antiviral activity of Illicium verum Hook. f. extracts against grouper iridovirus infection. Journal of Fish Diseases, 43(5):531–540, May 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Li:2022:QAS

[LYZ^{+}22] Bo-Yang Li, Fei Yang, Zhong-Yu Zhang, Yu-Feng Shen, Tao Wang, Liang Zhao, Jia-Cheng Qin, Fei Ling, and Gao-Xue Wang. Quinoline, with the active site of 8-hydroxyl, efficiently inhibits Micropterus salmoides rhabdovirus (MSRV) infection in vitro and in vivo. Journal of Fish Diseases, 45 (6):895–905, June 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Li:2024:EGC

[LZC^{+}24] Zhicheng Li, Jingyu Zhuang, Jiaming Chen, Jizhen Cao, Qing Han, Zhi Luo, Baotun Wang, Hebing Wang, and Anxing Li. Establishment of a gill cell line from yellowfin seabream (Acanthopagrus latus) for studying Amyloodinium ocellatum infection of fish. Journal of Fish Diseases, 47 (5):e13923:1–e13923:??, May 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Liao:2023:ECB

[LZK^{+}23] Jiaming Liao, Xin Zhang, Shaozhu Kang, Luhao Zhang, Dongzhuo Zhang, Zhuqing Xu, Qiwei Qin, and Jingguang Wei. Establishment and characterization of a brain tissue cell line from spotted knifejaw (Oplegnathus punctatus) and its susceptibility to several fish viruses. Journal of Fish Diseases, 46(7):767–777, July 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhao:2020:AVI

[lZWC^{+}20] Xian liang Zhao, Gan Wu, He Chen, Li Li, and Xiang hui Kong. Analysis of virulence and immunogenic factors in Aeromonas hydrophila: Towards the development of live vaccines. Journal of Fish Diseases, 43(7):747–755, July 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Liu:2023:RSC

[LZX^{+}23] Xiaoyu Liu, Xiaodong Zhang, Zhendong Xu, Zhiyang Huang, Junyao Zhong, Yinzhi Zhu, and Yongwei Wei. Recovery of Siniperca chuatsi rhabdovirus from cloned cDNA.

Journal of Fish Diseases, 46(2):127–135, February 2023. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Liu:2023:SAP

[LZZ^{+}23] Guang-Lu Liu, Hui Zhang, Li-Li Zhu, Xu-Dong Liu, Yan-Jie Liu, Ya hong Chen, Lei Liu, and Yang Hu. Synthesis and anti-parasites efficacy of coumarin derivatives against Dactylogyrus intermedius. Journal of Fish Diseases, 46(9): 967–976, September 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mason:2024:BAW

[MAG24] Alexandra K. Mason, Lance Adams, and Michael M. Garner. Bacterial abscesses in whitespotted bamboo sharks (Chiloscyllium plagiosum) in managed care. Journal of Fish Diseases, 47(5):e13917:1–e13917:??, May 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mahmoud:2023:CGA

[MAK^{+}23] Mahmoud Mostafa Mahmoud, Mohamed Abdelsalam, Satoshi Kawato, Shogo Harakawa, Hidemasa Kawakami, Ikuo Hirono, and Hidehiro Kondo. Comparative genome analyses of three serotypes of Lactococcus bacteria isolated from diseased cultured striped jack (Pseudocaranx dentex). Journal of Fish Diseases, 46(8):829–839, August 2023. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mansur:2023:RMM

[Man23] Moh Masnur Mansur. Review of “The Marine Microbiome, Second Edition”. Journal of Fish Diseases, 46(9):1013–1014, September 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mandrioli:2023:CAL

[MBF^{+}23] Luciana Mandrioli, Tim Barbé, Marco Bonazza Foselli, Ginevra Brocca, Ranieri Verin, Francesca Errani, Enrico Volpe, Francesco Montesi, Rubina Sirri, Maria Morini, Jane Budai, Anna Toffan, and Sara Ciulli. Cutaneous angiomatosis-like presentation in koi carp (Cyprinus carpio koi): Clinical–pathological investigations. Journal of Fish Diseases, 46(7):723–730, July 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Marty:2020:CSS

[MBJ20] Gary D. Marty, Julie Bidulka, and Tomy Joseph. Crosssectional study of histopathology and piscine orthoreovirus during a marine production cycle of farmed Atlantic salmon (Salmo salar L.) in British Columbia, Canada. Journal of Fish Diseases, 43(9):1019–1028, September 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Michail:2022:HRV

[MBN^{+}22] George Michail, Panagiotis Berillis, Christos Nakas, Morgane Henry, and Eleni Mente. Haematology reference values for Dicentrarchus labrax and Sparus aurata: a systematic review and meta-analysis. Journal of Fish Diseases, 45(10): 1549–1570, October 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mursalim:2022:DAS

[MBR^{+}22] Muhammad Fadhlullah Mursalim, Hendri Budiyansah, Hartanto Mulyo Raharjo, Partho Pratim Debnath, Rungnapa Sakulworakan, Putita Chokmangmeepisarn, Jitrapa Yindee, Patharapol Piasomboon, Sivaramasamy Elayaraja, and Channarong Rodkhum. Diversity and antimicrobial susceptibility profiles of Aeromonas spp. isolated from diseased freshwater fishes in Thailand. Journal of Fish Diseases, 45 (8):1149–1163, August 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mitchell:2021:MSP

[MBS21] Susie Olivia Mitchell, Sophie Bresnihan, and Felix Scholz. Mortality and skin pathology of farmed Atlantic salmon (Salmo salar) caused by exposure to the jellyfish Physalia physalis in Ireland. Journal of Fish Diseases, 44(11):1861– 1864, November 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Montory:2020:FFB

[MCA^{+}20] Jaime A. Montory, Oscar R. Chaparro, Andres Averbuj, Luis P. Salas-Yanquin, Joseline A. Büchner-Miranda, Paulina Gebauer, Juan P. Cumillaf, and Edgardo Cruces. The filter-feeding bivalve Mytilus chilensis capture pelagic stages of Caligus rogercresseyi: a potential controller of the sea lice fish parasites. Journal of Fish Diseases, 43(4):475– 484, April 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Marques:2023:EOL

[MCM^{+}23] A. R. Marques, T. H. Cheng, K. Y. Man, K. P. Cheng, K. B. Cheung, and S. St-Hilaire. Evaluation of oxytetracycline leaching from pregelatinized starch-coated medicated fish feed. Journal of Fish Diseases, 46(11):1183–1192, November 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mugetti:2022:EMA

[MCP^{+}22] Davide Mugetti, Silvia Colussi, Paolo Pastorino, Katia Varello, Mattia Tomasoni, Vasco Menconi, Claudio Pedron, Elena Bozzetta, Pier Luigi Acutis, and Marino Prearo. Episode of mortality associated with isolation of Streptococcus iniae in Adriatic sturgeon (Acipenser naccarii Bonaparte, 1836) reared in Northern Italy. Journal of Fish Diseases, 45(6):939–942, June 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Maduenyane:2022:LSE

[MDAO22] Mpho Maduenyane, Quinton Marco Dos Santos, and Annemariè Avenant-Oldewage. Light and scanning electron microscopy of the effects of Macrogyrodactylus congolensis (Prudhoe, 1957) on the skin of the African sharptooth catfish Clarias gariepinus (Burchell, 1822). Journal of Fish Diseases, 45(4):595–602, April 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mo:2021:IIA

[MFP21] Tor Atle Mo, Frode Fossøy, and Trygve T. Poppe. Increasing intensities of Anisakis simplex (Rudolphi, 1809 det. Krabbe, 1878) larvae with weight and sea age in returning adult Atlantic salmon, Salmo salar L., of coastal waters of Norway. Journal of Fish Diseases, 44(8):1075–1089, August 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Misk:2022:RSP

[MGG22] Ehab Misk, Serap Gonen, and Amber F. Garber. Resistance to Saprolegnia parasitica infection: a heritable trait in Atlantic salmon. Journal of Fish Diseases, 45(9):1333–1342, September 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Matsumoto:2020:ACF

[MHA^{+}20] Megumi Matsumoto, Kazuma Hayashi, Kyosuke Araki, Teruyuki Nakanishi, and Atsushi Yamamoto. Availability of culture filtrate protein-10 (CFP-10) secreted from Mycobacterium pseudoshottsii for mycobacteriosis diagnosis in ginbuna crucian carp Carrasius auratus langsdorfii. Journal of Fish Diseases, 43(1):81–89, January 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mo:2023:OSG

[MHH23] Tor Atle Mo, Haakon Hansen, and Sigurd Hytterød. Occurrence and seasonality of Gyrodactylus salaris and G. salmonis (Monogenea) on Arctic char (Salvelinus alpinus (L.)) in the Fustvatnet lake, Northern Norway. Journal of Fish Diseases, 46(4):395–403, April 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mahasri:2020:PIH

[MHP^{+}20] Gunanti Mahasri, Umi Hafidloh, Fungky P. Pratama, Dicky Rahmawan, Sri Subekti, Putri D. Wulansari, and Muhamad Amin. Prevalence, intensity and histopathology of Zeylanicobdella arugamensis infestation on groupers reared on different aquaculture systems. Journal of Fish Diseases, 43(10): 1133–1143, October 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mercer:2022:CI

[MHYW22a] Lewis K. Mercer, Emma F. Harding, Grace J. H. Yan, and Peter A. White. Cover image. Journal of Fish Diseases, 45 (6):i, June 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mercer:2022:NVD

[MHYW22b] Lewis K. Mercer, Emma F. Harding, Grace J. H. Yan, and Peter A. White. Novel viruses discovered in the transcriptomes of agnathan fish. Journal of Fish Diseases, 45(6):931– 938, June 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Miranda:2024:PGC

[MIC^{+}24] Claudio D. Miranda, Rute Irgang, Christopher Concha, Rodrigo Rojas, and Ruben Avendaño-Herrera. Phenotypic and genomic characterization of a non-pathogenic Epilithonimonas ginsengisoli isolated from diseased farmed rainbow trout (Oncorhynchus mykiss) in Chile. Journal of Fish Diseases, 47(3):e13897:1–e13897:??, March 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Markovskaja:2024:DFL

[MIK^{+}24] Svetlana Markovskaja, Reda Iršėnaitė, Audrius Kačergius, Gintarė Sauliutė, and Milda Stankevičiūtė. Diversity of fungus-like stramenopilous organisms (Oomycota) in Lithuanian freshwater aquaculture: Morphological and molecular analysis, risk to fish health. Journal of Fish Diseases, 47(3): e13903:1–e13903:??, March 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Macchia:2022:ICM

[MIR^{+}22] Valeria Macchia, Makoto Inami, Anne Ramstad, Fabian Grammes, Andrew Reeve, Thomas Moen, Jacob Seilø Torgersen, Alexandra Adams, Andrew P. Desbois, and Rowena Hoare. Immersion challenge model for Flavobacterium psychrophilum infection of Atlantic salmon (Salmo salar L.) fry. Journal of Fish Diseases, 45(11):1781–1788, November 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Meng:2023:DDD

[MJX^{+}23] Yan Meng, Nan Jiang, Yixing Xie, Ying Wei, Cheng Wang, Mingzhu Tian, Mingyang Xue, Chen Xu, Yiqun Li, Wenzhi Liu, Yuding Fan, and Yong Zhou. Development of a droplet digital PCR assay for the sensitive detection of iridovirus in Andrias davidianus. Journal of Fish Diseases, 46(11):1249– 1256, November 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mai:2021:EHK

[MKT^{+}21] Thao Thu Mai, Pattanapon Kayansamruaj, Suwimon Taengphu, Saengchan Senapin, Janina Z. Costa, Jorge del Pozo, Kim D. Thompson, Channarong Rodkhum, and Ha Thanh Dong. Efficacy of heat-killed and formalin-killed vaccines against Tilapia tilapinevirus in juvenile Nile tilapia (Oreochromis niloticus). Journal of Fish Diseases, 44(12): 2097–2109, December 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Menanteau–Ledouble:2020:EPB

[MLGRF^{+}20] Simon Menanteau-Ledouble, Michael Gotesman, Ebrahim Razzazi-Fazeli, Sven M. Bergmann, and Mansour El-Matbouli. Elucidation of putative binding partners for the protein encoded by ORF149 of cyprinid herpesvirus 3 in goldfish (Carassius auratus). Journal of Fish Diseases, 43 (6):707–710, June 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mo:2023:CMA

[MLHD^{+}23] Jakob Mo, Ada S. Løge-Hagen, Alf S. Dalum, Toni Erkinharju, Paul J. Midtlyng, and Arnfinn Aunsmo. A case of melanization after intramuscular vaccination in Atlantic salmon (Salmo salar L.) — Possible causes and implications. Journal of Fish Diseases, 46(10):1157–1161, October 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ma:2022:MBI

[MLL^{+}22] Hongli Ma, Xiaocen Lu, Jiannan Liu, Sicong Guo, Xiaoran Zhao, and Shigen Ye. Metschnikowia bicuspidata isolated from milky diseased Eriocheir sinensis: Phenotypic and genetic characterization, antifungal susceptibility and challenge models. Journal of Fish Diseases, 45(1):41–49, January 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mikalsen:2020:LEV

[MLM^{+}20] Aase B. Mikalsen, Morten Lund, Farah Manji, Mari Viken Kjønstad, Per Helge Bergtun, Gordon Ritchie, Vidar Aspehaug, Magnus Devold, and Øystein Evensen. Lack of evidence of vertical transmission of piscine myocarditis virus in Atlantic salmon (Salmo salar L.). Journal of Fish Diseases, 43(6):715–718, June 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Madrid:2021:HCE

[MMTD^{+}21] Rafael R. M. Madrid, Omar Mertins, Marcos Tavares-Dias, Anai P. Flores-Gonzales, Ana C. M. F. Patta, Carlos A. B. Ramirez, Vera L. S. Rigoni, and Patrick D. Mathews. High compliance and effective treatment of fish endoparasitic infections with oral drug delivery nanobioparticles: Safety of intestinal tissue and blood parameters. Journal of Fish Diseases, 44(11):1819–1829, November 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Marshall:2022:DKT

[MMW^{+}22] Wyth L. Marshall, Tiffany MacWilliam, Kayla Williams, Heather Reinholt, Heidi VanVliet, Danielle New, Meghan Mills, and Diane Morrison. Detection of Kudoa thyrsites (Myxozoa) eDNA by real-time and digital PCR from high seawater volumes. Journal of Fish Diseases, 45(9):1403– 1407, September 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Madhun:2024:OSA

[MNB^{+}24] Abdullah Sami Madhun, Rune Nilsen, Bjørn T. Barlaup, Ørjan Karlsen, and Egil Karlsbakk. Occurrence of salmonid alphavirus and piscine orthoreovirus-1 infections in migrating salmon (Salmo salar L.) post-smolt in western Norway. Journal of Fish Diseases, 47(1):e13874:1–e13874:??, January 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Myrenaas:2023:EIS

[MNPS23] Elin Myrenås, Joacim Näslund, John Persson, and Josefin Sundin. Effects of the invasive swim bladder parasite Anguillicola crassus on health and condition indicators in the European eel. Journal of Fish Diseases, 46(10):1029–1047, October 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Marsella:2021:CEV

[MPA^{+}21] Andrea Marsella, Tobia Pretto, Miriam Abbadi, Rosita Quartesan, Luana Cortinovis, Eleonora Fiocchi, Amedeo Manfrin, and Anna Toffan. Carp edema virus-related mortality in wild adult common carp (Cyprinus carpio) in Italy. Journal of Fish Diseases, 44(7):939–947, July 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mertz:2024:MQP

[MPF^{+}24] Nathan E. Mertz, Rachel A. Paterson, Bengt Finstad, Hege Brandsegg, Ida Pernille Øystese Andersskog, and Frode Fossøy. Molecular quantification of parasitic sea louse larvae depends on species and life stage. Journal of Fish Diseases, 47(4):e13908:1–e13908:??, April 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Maj-Paluch:2020:PCP

[MPMB^{+}20] Joanna Maj-Paluch, Marek Matras, Ewa Borzym, Magdalena Stachnik, and Michal Reichert. Phylogenetic characterization of Polish isolates of infectious pancreatic necrosis virus in salmonid fish. Journal of Fish Diseases, 43(11): 1443–1451, November 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mahmud:2020:NTD

[MPP20a] Raihan Mahmud, John Purser, and Jawahar G. Patil. A novel testicular degenerative condition in a wild population of the common carp Cyprinus carpio (L). Journal of Fish Diseases, 43(9):1065–1076, September 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Matejickova:2020:COV

[MPP^{+}20b] Kateřina Matějíčková, Łubomír Pojezdal, Dagmar Pokorová, Stanislava Reschová, Veronika Piačková, Miroslava Palíková, Tomáš Veselý, and Ivana Papežíková. Carp oedema virus disease outbreaks in Czech and Slovak aquaculture. Journal of Fish Diseases, 43(9):971–978, September 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Machat:2021:CEV

[MPPF21] Radek Machat, Lubomir Pojezdal, Veronika Piackova, and Martin Faldyna. Carp edema virus and immune response in carp (Cyprinus carpio): Current knowledge. Journal of Fish Diseases, 44(4):371–378, April 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Machimbirike:2022:MEC

[MPS^{+}22] Vimbai Irene Machimbirike, Natapol Pornputtapong, Saengchan Senapin, Eakapol Wangkahart, Prapansak Srisapoome, Pongsak Khunrae, and Triwit Rattanarojpong. A multiepitope chimeric protein elicited a strong antibody response and partial protection against Edwardsiella ictaluri in Nile tilapia. Journal of Fish Diseases, 45(1):1–18, January 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mao:2020:RLT

[MQK^{+}20] Leilei Mao, Yingxue Qin, Jianping Kang, Bin Wu, Lixing Huang, Suyun Wang, Mengmeng Zhang, Jiahui Zhang, Ruixuan Zhang, and Qingpi Yan. Role of LuxR-type regulators in fish pathogenic Aeromonas hydrophila. Journal of Fish Diseases, 43(2):215–225, February 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mogahadam:2023:VLS

[MR23] Hooman K. Mogahadam and Magnus Vikan Røsæg. From viral load to survival: Unveiling new genetic links for resistance against PMCV in Atlantic salmon (Salmo salar). Journal of Fish Diseases, 46(11):1285–1294, November 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Malard:2021:SDM

[MRM21] Lucie Malard, Cynthia Riginos, and Katrina McGuigan. Skeletal deformities and meristic trait variations are common in the intertidal fish Bathygobius cocosensis (Perciformes-Gobiidae). Journal of Fish Diseases, 44(6):665–673, June 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Marana:2021:ZDR

[MSB^{+}21] Moonika Haahr Marana, Jacob Günther Schmidt, Stéphane Biacchesi, Niels Lorenzen, and Louise von Gersdorff Jørgensen. Zebrafish (Danio rerio) larvae as a model for real-time studies of propagating VHS virus infection, tissue tropism and neutrophil activity. Journal of Fish Diseases, 44(5):563– 571, May 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Malheiros:2020:NOC

[MSF^{+}20] Dayna Filocreão Malheiros, Icaro Rodrigues Sarquis, Irlon Maciel Ferreira, Patrick Delgado Mathews, Omar Mertins, and Marcos Tavares-Dias. Nanoemulsions with oleoresin of Copaifera reticulata (Leguminosae) improve anthelmintic efficacy in the control of monogenean parasites when compared to oleoresin without nanoformulation. Journal of Fish Diseases, 43(6):687–695, June 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Midorikawa:2020:CAH

[MSS^{+}20] Yuki Midorikawa, Tomohito Shimizu, Tetsuya Sanda, Katsuyuki Hamasaki, Shigeki Dan, Mohammad Tamrin Bin Mohamad Lal, Goshi Kato, and Motohiko Sano. Characterization of Aquimarina hainanensis isolated from diseased mud crab Scylla serrata larvae in a hatchery. Journal of Fish Diseases, 43(5):541–549, May 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mora-Salas:2023:MLT

[MSZV^{+}23] Patricia Mora-Salas, Sebastián Zapararte, Pamela Villouta, Henry Araya-León, Ruben Avendaño-Herrera, Francisco Melo, and Fernando O. Mardones. Method for lineage typing of epidemic Renibacterium salmoninarum in Chilean salmon farms. Journal of Fish Diseases, 46(5):499–506, May 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mufliyati:2023:RFH

[Muf23] Lily Mufliyati. Review of Fish Histology: From Cells to Organs, 2nd Edition. 2022. Journal of Fish Diseases, 46(8): 895–896, August 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Midttun:2020:EPN

[MVW^{+}20] Helene L. E. Midttun, Marco A. Vindas, Paul J. Whatmore, Øyvind Øverli, and Ida B. Johansen. Effects of Pseudoloma neurophilia infection on the brain transcriptome in zebrafish (Danio rerio). Journal of Fish Diseases, 43(8):863–875, August 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ma:2021:EDS

[MWH^{+}21] Rongrong Ma, Yuan Wang, Liang Huang, Shu Zhao, Letian Li, Menghe Yin, and Wenhong Fang. Effects of different salinity on the transcriptome and antibiotic resistance of two Vibrio parahaemolyticus strains isolated from Penaeus vannamei cultured in seawater and freshwater ponds. Journal of Fish Diseases, 44(12):2055–2066, December 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Maekawa:2020:GCS

[MWY^{+}20] Shun Maekawa, Yi-Ting Wang, Terutoyo Yoshida, Pei-Chi Wang, and Shih-Chu Chen. Group C Streptococcus dysgalactiae infection in fish. Journal of Fish Diseases, 43(9):963– 970, September 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Meng:2022:DCS

[MWY^{+}22] Xian-Yu Meng, Zi-Hao Wang, Xue-Dong Yu, Qi-Ya Zhang, and Fei Ke. Development and characterization of a skin cell line from Chinese perch (Siniperca chuatsi) and its application in aquatic animal viruses. Journal of Fish Diseases, 45(10):1439–1449, October 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ma:2020:CMC

[MWZ^{+}20] Rongrong Ma, Yuan Wang, Shu Zhao, Menghe Yin, and Wenhong Fang. The composition of the microbial community associated with Macrobrachium rosenbergii zoeae varies throughout larval development. Journal of Fish Diseases, 43 (4):413–421, April 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Naranitus:2022:SCC

[NAaS^{+}22] Punyaporn Naranitus, Praphutson Aiamsa-at, Thanwarat Sukonta, Pimkhuan Hannanta-anan, and Thawatchai Chaijarasphong. Smartphone-compatible, CRISPR-based platforms for sensitive detection of acute hepatopancreatic necrosis disease in shrimp. Journal of Fish Diseases, 45(12): 1805–1816, December 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nadolna-Altyn:2024:FRP

[NAPP24] Katarzyna Nadolna-Ałtyn, Joanna Pawlak, and Marzenna Pachur. First record of Pseudoterranova decipiens in sprat (Sprattus sprattus) from the Baltic Sea. Journal of Fish Diseases, 47(1):e13866:1–e13866:??, January 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Natale:2020:TLS

[NCA^{+}20] Sabrina Natale, Fabiano Capparucci, Jessica Maria Abbate, Rosanna Panebianco, Roberto Puleio, and Carmelo Iaria. Testicular leiomyoma and spermatogenic failure syndrome in a seabass from broodstock. Journal of Fish Diseases, 43 (12):1563–1569, December 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nguyen:2023:PCN

[NGC^{+}23] Phuong Nguyen, Andre Giovanni, Li-Wu Cheng, Pei-Chi Wang, and Shih-Chu Chen. Pathogenicity characterization of nocardiosis in orange-spotted grouper (Epinephelus coioides) via exposure to nine distinct isolates of Nocardia seriolae and four administrative routes. Journal of Fish Diseases, 46(11):1257–1268, November 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nhinh:2023:SIE

[NHG^{+}23] Doan Thi Nhinh, Dang Thi Hoa, Nguyen Thi Huong Giang, Kim Van Van, Lua Thi Dang, Mags Crumlish, Ha Thanh Dong, and Truong Dinh Hoai. Synergistic infection of Edwardsiella ictaluri and Flavobacterium oreochromis in cage cultured tilapia (Oreochromis sp.). Journal of Fish Diseases, 46(10):1125–1136, October 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nam:2022:DPR

[NK22] U-Hwa Nam and Jeong-Ho Kim. Development of a PCR-RFLP method for differential identification of Microcotyle sebastis and Microcotyle caudata isolated from cultured rockfish in Korea. Journal of Fish Diseases, 45(3):411–419, March 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Noguera:2021:UIR

[NKö^{+}21] Patricia Noguera, Matthias Klinger, Histro örün, Bianka Grunow, and Jorge del Pozo. Ultrastructural insights into the replication cycle of salmon pancreas disease virus (SPDV) using salmon cardiac primary cultures (SCPCs). Journal of Fish Diseases, 44(12):2031–2041, December 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Neal:2021:LIR

[NKS^{+}21] Travis Neal, Michael L. Kent, Justin Sanders, Carl B. Schreck, and James T. Peterson. Laboratory infection rates and associated mortality of juvenile Chinook salmon (Oncorhynchus tshawytscha) from parasitic copepod (Salmincola californiensis). Journal of Fish Diseases, 44(9):1423–1434, September 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Niu:2020:PGC

[NKZ^{+}20] Guoyi Niu, Rutch Khattiya, Tingrui Zhang, Sukolrat Boonyayatra, and Dilok Wongsathein. Phenotypic and genotypic characterization of Streptococcus spp. isolated from tilapia (Oreochromis spp.) cultured in river-based cage and earthen ponds in Northern Thailand. Journal of Fish Diseases, 43(3):391–398, March 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Norris:2020:DPN

[NLH^{+}20] Lauren Norris, Noah Lawler, Andree Hunkapiller, Donna M. Mulrooney, Michael L. Kent, and Justin L. Sanders. Detection of the parasitic nematode, Pseudocapillaria tomentosa, in zebrafish tissues and environmental DNA in research aquaria. Journal of Fish Diseases, 43(9):1087–1095, September 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Niu:2021:CFA

[NLL^{+}21] Jinzhong Niu, Guoling Luo, Xinchao Liu, Yu Huang, Jufen Tang, Bei Wang, Yishan Lu, Jia Cai, and Jichang Jian. Characterization and functional analysis of a galectin-related protein B from Nile tilapia involved in the immune response to bacterial infection. Journal of Fish Diseases, 44(2):171– 180, February 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nguyen:2021:GCH

[NLPM^{+}21] Diem Thu Nguyen, Adrián López-Porras, David Marancik, Leighanne Hawkins, Timothy J. Welch, Barbara Denise Petty, Cynthia Ware, Matt J. Griffin, and Esteban Soto. Genetic characterization of heterologous Edwardsiella piscicida isolates from diverse fish hosts and virulence assessment in a Chinook salmon Oncorhynchus tshawytscha model. Journal of Fish Diseases, 44(12):1959–1970, December 2021. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nowlan:2021:QPT

[NLR21] Joseph P. Nowlan, John S. Lumsden, and Spencer Russell. Quantitative PCR for Tenacibaculum dicentrarchi and T. finnmarkense. Journal of Fish Diseases, 44(5):655–659, May 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nguyen:2023:EAN

[NNC^{+}23] Phuong Nguyen, Sandra Celenia Nazareth, Mei-Yun Chen, Pei-Chi Wang, and Shih-Chu Chen. An epidemiological analysis of Nocardia seriolae isolated from a wide range of aquatic animals in Taiwan, based on their genotype and enzymatic activity. Journal of Fish Diseases, 46(4):381–394, April 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nowak:2021:RDL

[Now21] Barbara Nowak. Remembering Don Lightner. Journal of Fish Diseases, 44(9):1269, September 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nervino:2024:ILP

[NPP^{+}24] Stephanie Nervino, Tamsen Polley, James T. Peterson, Carl B. Schreck, Michael L. Kent, and Julie D. Alexander. Intestinal lesions and parasites associated with senescence and prespawn mortality in Chinook salmon (Oncorhynchus tshawytscha). Journal of Fish Diseases, 47(2):e13876:1– e13876:??, February 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nazareth:2023:NSC

[NRC^{+}23] Sandra Celenia Nazareth, Shreesha Rao, Li-Wu Cheng, Pei-Chi Wang, and Shih-Chu Chen. Nocardia seriolae cell wall lipids: An effective protective mechanism in resistance and virulence. Journal of Fish Diseases, 46(4):405–416, April 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

NaveenKumar:2021:RVP

[NRKK21] Singaiah NaveenKumar, Praveen Rai, Indrani Karunasagar, and Iddya Karunasagar. Recombinant viral proteins delivered orally through inactivated bacterial cells induce protection in Macrobrachium rosenbergii (de Man) against White Tail Disease. Journal of Fish Diseases, 44(5):601–612, May 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Neupane:2023:FIM

[NRY^{+}23] Sandip Neupane, Shreesha Rao, Wei-Xiao Yan, Pei-Chi Wang, and Shih-Chu Chen. First identification, molecular characterization, and pathogenicity assessment of Lactococcus garvieae isolated from cultured pompano in Taiwan. Journal of Fish Diseases, 46(11):1295–1309, November 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nowak:2020:IIR

[NSP20] Barbara Nowak, David Speare, and Mark Powell. Introduction to the invited review series initiative of the Journal of Fish Diseases: Health management of zebrafish in research facilities. Journal of Fish Diseases, 43(6):635, June 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nasmia:2024:PAA

[NSTS24] Nasmia Nasmia, Novalina Serdiati, Akbar Marzuki Tahya, and Muhammad Safir. Phytochemical analysis and antibacterial activity of palm waste extract against Vibrio harveyi and Vibrio parahaemolyticus. Journal of Fish Diseases, 47 (6):e13924:1–e13924:??, June 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nanni:2024:DAE

[NTB^{+}24] Irene Maja Nanni, Perla Tedesco, David Baldo, Roberta Galuppi, and Marina Collina. Dimethomorph activity and its effect on morphology in different oomycete species of economic and veterinary interest. Journal of Fish Diseases, 47 (5):e13925:1–e13925:??, May 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Olsvik:2024:LTT

[OBSH24] Pål A. Olsvik, Kristine E. Brokke, Ole B. Samuelsen, and Rita Hannisdal. Lufenuron treatment temporarily represses gene expression and affects the SUMO pathway in liver of Atlantic salmon. Journal of Fish Diseases, 47(2):e13880:1– e13880:??, February 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Oliver:2023:SCF

[OCS^{+}23] Cristian Oliver, Constanza Céspedes, Natacha Santibañez, Pamela Ruiz, and Alex Romero. Subinhibitory concentrations of florfenicol increase the biofilm formation of Piscirickettsia salmonis. Journal of Fish Diseases, 46(5):591–596, May 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Older:2024:IPP

[OGR^{+}24] Caitlin E. Older, Matt J. Griffin, Bradley M. Richardson, Geoffrey C. Waldbieser, J. Grant Reifers, Penelope M. Goodman, Cynthia Ware, Delbert M. Gatlin III, David J. Wise, and Fernando Y. Yamamoto. Influence of probiotic and prebiotic supplementation on intestinal microbiota and resistance to Edwardsiella ictaluri infection in channel catfish (Ictalurus punctatus) following florfenicol administration. Journal of Fish Diseases, 47(4):e13910:1–e13910:??, April 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Okamura:2020:EPH

[OHL^{+}20] Beth Okamura, Ashlie Hartigan, Paul F. Long, Paolo Ruggeri, Katie Smith-Easter, and Jason D. Schooley. Epidemiology of Polypodium hydriforme in American Paddlefish.

Journal of Fish Diseases, 43(9):979–989, September 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Osman:2022:MAD

[OKL^{+}22] Ebrahim Osman, Nameun Kim, Yoonhang Lee, Juri Yoo, Sung-Hyun Kim, and Do-Hyung Kim. Molecular approaches for detection and quantification of Vibrio scophthalmi based on recA. Journal of Fish Diseases, 45(2):373–378, February 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Oliver:2024:EFP

[ORV^{+}24] Cristian Oliver, Pamela Ruiz, José Miguel Vidal, Carlos Carrasco, Carla Estefanía Escalona, Javier Barros, Daniela Sepúlveda, Homero Urrutia, and Alex Romero. Effect of florfenicol on Piscirickettsia salmonis biofilm formed in materials used in salmonid nets, nylon and high-density polyethylene. Journal of Fish Diseases, 47(1):e13862:1–e13862:??, January 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ostevik:2022:CSG

[ØSH^{+}22] Liv Østevik, Marit Stormoen, Hege Hellberg, Marianne Kraugerud, Farah Manji, Kai-Inge Lie, Ane Nødtvedt, Hamish Rodger, and Marta Alarcón. A cohort study of gill infections, gill pathology and gill-related mortality in seafarmed Atlantic salmon (Salmo salar L.): a descriptive analysis. Journal of Fish Diseases, 45(9):1301–1321, September 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Oredalen:2022:PTB

[OSM22a] Tone Jøran Oredalen, Mona Sæbø, and Tor Atle Mo. Patterns of Tetracapsuloides bryosalmonae infection of three salmonid species in large, deep Norwegian lakes. Journal of Fish Diseases, 45(1):185–202, January 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Orioles:2022:ORM

[OSM^{+}22b] Massimo Orioles, Elena Saccà, Mathjis Metselaar, Michela Bulfoni, Daniela Cesselli, and Marco Galeotti. Observations on Red Mark Syndrome in juvenile rainbow trout farmed in RAS system. Journal of Fish Diseases, 45(12):1889–1892, December 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ozdemir:2022:PSR

[ÖTG22] Rahmi Can Özdemir, Yiğit Taştan, and Kerim Güney. Prevention of Saprolegniasis in rainbow trout (Oncorhynchus mykiss) eggs using oregano (Origanum onites) and laurel (Laurus nobilis) essential oils. Journal of Fish Diseases, 45 (1):51–58, January 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

ozcelik:2020:UOA

[öTTS20] Halil özçelik, Yiğit Taştan, Ertugrul Terzi, and Adem Yavuz Sönmez. Use of onion (Allium cepa) and Garlic (Allium sativum) wastes for the prevention of fungal disease (Saprolegnia parasitica) on eggs of rainbow trout (Oncorhynchus mykiss). Journal of Fish Diseases, 43(10):1325–1330, October 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Persson:2022:APS

[PAH^{+}22] David B. Persson, Anna Aspán, Paulina Hysing, Eva Blomkvist, Eva Jansson, Ludvig Orsén, Hampus Hällbom, and Charlotte Axén. Assessing the presence and spread of Renibacterium salmoninarum between farmed and wild fish in Sweden. Journal of Fish Diseases, 45(5):613–621, May 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Peel:2021:SPS

[PAS^{+}21] Melanie J. Peel, Lance Adams, Brittany Stevens, Michael M. Garner, and Cheryl Maguire. Systemic phaeohyphomycosis in splitnose rockfish (Sebastes diploproa) caused by Devriesia sp. Journal of Fish Diseases, 44(5):639–644, May 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Persson:2021:VMC

[PBF^{+}21] David Persson, Håvard Bjørgen, Alexander Figenschou, Linn-Anett Hillestad, Erling Olaf Koppang, Ane Nødtvedt, and Marit Stormoen. Variations in mucous cell numbers in gills of Atlantic salmon (Salmo salar) presmolt in commercial freshwater farms in Norway. Journal of Fish Diseases, 44 (1):25–32, January 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pazdzior:2023:PAS

[PBT^{+}23] Ewa Paździor, Arkadiusz Bomba, Katarzyna Tłuścik, Katarzyna Nadolna-Ałtyn, Magdalena Podolska, Michał Reichert, Dariusz Wasyl, and Agnieszka Pekala-Safińska. Phylogenetic analysis of Shewanella spp. isolated from fish. Journal of Fish Diseases, 46(11):1163–1171, November 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pham:2020:GDM

[PCWC20] Trung Hieu Pham, Ta-Chih Cheng, Pei-Chi Wang, and Shih-Chu Chen. Genotypic diversity, and molecular and pathogenic characterization of Photobacterium damselae subsp. piscicida isolated from different fish species in Taiwan. Journal of Fish Diseases, 43(7):757–774, July 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pedrosa:2024:OHS

[PDG^{+}24] Virgínia F. Pedrosa, Pedro A. De P. Dos Santos, Luana B. Giesta, Julio C. Zemor, Marcelo H. Okamoto, Andrezza C. Chagas, Aline da C. Bezerra, Luis H. da S. Poersch, and Luis A. Romano. Occurrence of hepatic and splenic lipomas in Nile tilapia (Oreochromis niloticus). Journal of Fish Diseases, 47(1):e13869:1–e13869:??, January 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Padeniya:2023:PER

[PDL^{+}23] Uthpala Padeniya, D. Allen Davis, Mark R. Liles, Stacey A. LaFrentz, Benjamin R. LaFrentz, Craig A. Shoemaker, Benjamin H. Beck, Daniel E. Wells, and Timothy J. Bruce. Probiotics enhance resistance to Streptococcus iniae in Nile tilapia (Oreochromis niloticus) reared in biofloc systems. Journal of Fish Diseases, 46(10):1137–1149, October 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Papadopoulou:2021:DDA

[PDM^{+}21] Athina Papadopoulou, Andrew Davie, Sean J. Monaghan, Herve Migaud, and Alexandra Adams. Development of diagnostic assays for differentiation of atypical Aeromonas salmonicida vapA type V and type VI in ballan wrasse (Labrus bergylta, Ascanius). Journal of Fish Diseases, 44 (6):711–719, June 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pousis:2022:MUT

[PDZ^{+}22] Chrysovalentinos Pousis, Mariasevera Di Comite, Rosa Zupa, Letizia Passantino, Edmond Hala, and Aldo Corriero. Microradiography as a useful technique for the rapid detection of skeletal anomalies in early sea bream juveniles. Journal of Fish Diseases, 45(8):1237–1241, August 2022. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Panicz:2021:FDH

[PEN^{+}21] Remigiusz Panicz, Piotr Eljasik, Tuan Thuc Nguyen, Kim Thinh Vo Thi, and Duat Van Hoang. First detection of Herpesvirus anguillae (AngHV-1) associated with mortalities in farmed giant mottled eel (Anguilla marmorata) in Vietnam. Journal of Fish Diseases, 44(6):847–852, June 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Panicz:2020:NIS

[PEś^{+}20] Remigiusz Panicz, Piotr Eljasik, Natalia śmietana, Jacek Sadowski, and Marcin Biernaczyk. New invertebrate species as potential CyHV-3 reservoirs: a case study of common carp mortalities in hyperthermal conditions. Journal of Fish Diseases, 43(7):821–824, July 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Papadopoulou:2021:NAA

[PGH^{+}21] Athina Papadopoulou, Kathryn Garvey, Tom Hill, Jose G. Ramirez-Paredes, Sean J. Monaghan, Johanna L. Baily, Andrew Davie, Ioanna Katsiadaki, David Verner-Jeffreys, Timothy Wallis, Herve Migaud, and Alexandra Adams. Novel atypical Aeromonas salmonicida bath challenge model for juvenile ballan wrasse (Labrus bergylta, Ascanius). Journal of Fish Diseases, 44(6):823–835, June 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pulpipat:2023:CIS

[PHB^{+}23] Theeraporn Pulpipat, Taylor I. Heckman, Visanu Boonyawiwat, Pattarawit Kerddee, Sakuna Phatthanakunanan, Esteban Soto, and Win Surachetpong. Concurrent infections of Streptococcus iniae and Aeromonas veronii in farmed giant snakehead (Channa micropeltes). Journal of Fish Diseases, 46(6):629–641, June 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Palang:2020:CSA

[PHS^{+}20] Iyapa Palang, Ikuo Hirono, Saengchan Senapin, Wanna Sirimanapong, Boonsirm Withyachumnarnkul, and Rapeepun Vanichviriyakit. Cytotoxicity of Streptococcus agalactiae secretory protein on tilapia cultured cells. Journal of Fish Diseases, 43(10):1229–1236, October 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Prazdnikov:2024:IET

[PK24] Denis V. Prazdnikov and Ekaterina A. Kondakova. Investigation of extragonadal teratomas in two Poecilia wingei. Journal of Fish Diseases, 47(5):e13929:1–e13929:??, May 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Printzi:2021:BBA

[PKF^{+}21] Alice Printzi, Chara Kourkouta, Stefanos Fragkoulis, Anastasia Dimitriadi, George Geladakis, Michail Orfanakis, David Mazurais, Jose-Luis Zambonino-Infante, and George Koumoundouros. Balancing between Artemia and microdiet usage for normal skeletal development in zebrafish (Danio rerio). Journal of Fish Diseases, 44(11):1689–1696, November 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Prior:2022:EPA

[PLS^{+}22] Benjamin S. Prior, Miles D. Lange, Scott A. Salger, Benjamin J. Reading, Eric Peatman, and Benjamin H. Beck. The effect of piscidin antimicrobial peptides on the formation of Gram-negative bacterial biofilms. Journal of Fish Diseases, 45(1):99–105, January 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Persson:2022:AMP

[PNAS22] David Persson, Ane Nødtvedt, Arnfinn Aunsmo, and Marit Stormoen. Analysing mortality patterns in salmon farming using daily cage registrations. Journal of Fish Diseases, 45 (2):335–347, February 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pham:2021:CSD

[PNØ^{+}21] Khoi Dinh Pham, Sang Van Nguyen, Jørgen Ødegård, Hans Magnus Gjøen, and Gunnar Klemetsdal. Case study development of a challenge test against Edwardsiella ictaluri in Mekong striped catfish (Pangasianodon hypophthalmus), for use in breeding: Estimates of the genetic correlation between susceptibilities in replicated tanks. Journal of Fish Diseases, 44(5):553–561, May 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pham:2021:GAR

[PØN^{+}21a] Khoi Dinh Pham, Jørgen Ødegård, Sang Van Nguyen, Hans Magnus Gjøen, and Gunnar Klemetsdal. Genetic analysis of resistance in Mekong striped catfish (Pangasianodon hypophthalmus) to bacillary necrosis caused by Edwardsiella ictaluri. Journal of Fish Diseases, 44(2):201–210, February 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pham:2021:GCB

[PØN^{+}21b] Khoi Dinh Pham, Jørgen Ødegård, Sang Van Nguyen, Hans Magnus Gjøen, and Gunnar Klemetsdal. Genetic correlations between challenge tested susceptibility to bacillary necrosis, caused by Edwardsiella ictaluri, and growth performance tested survival and harvest body weight in Mekong striped catfish (Pangasianodon hypophthalmus). Journal of Fish Diseases, 44(2):191–199, February 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pereira:2020:LCC

[POS^{+}20] Elcimara Cardoso Pereira, Elzamara Castro Oliveira, Elen Monique Oliveira Sousa, Hugo Napoleão Pereira Silva, Lincoln Lima Corrêa, Rosa Helena Veras Mourão, Marcos Tavares-Dias, and Lenise Vargas Flores Silva. Lethal concentration of Cymbopogon citratus (Poaceae) essential oil for Dolops discoidalis and Argulus sp. (Crustacea: Argulidae). Journal of Fish Diseases, 43(12):1497–1504, December 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Palikova:2022:CEV

[PPDG^{+}22] Miroslava Palíková, L’ubomír Pojezdal, Lenka Dávidová-Geržová, Věra Nováková, Jiří Pikula, Ivana Papežíková, Hana Minářová, Ivana Mikulíková, and Iva Dyková. Carp edema virus infection associated gill pathobiome: a case report. Journal of Fish Diseases, 45(10):1409–1417, October 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Purcell:2020:CPO

[PPT^{+}20] Maureen K. Purcell, Rachel L. Powers, Torunn Taksdal, Doug McKenney, Carla M. Conway, Diane G. Elliott, Mark Polinski, Kyle Garver, and James Winton. Consequences of Piscine orthoreovirus genotype 1 (PRV-1) infections in Chinook salmon (Oncorhynchus tshawytscha), Coho salmon (O. kisutch) and rainbow trout (O. mykiss). Journal of Fish Diseases, 43(7):719–728, July 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Poudyal:2020:CPF

[PPWC20] Sayuj Poudyal, Theeraporn Pulpipat, Pei-Chi Wang, and Shih-Chu Chen. Comparison of the pathogenicity of Francisella orientalis in Nile tilapia (Oreochromis niloticus), Asian seabass (Lates calcarifer) and largemouth bass (Micropterus salmoides) through experimental intraperitoneal infection. Journal of Fish Diseases, 43(9):1097–1106, September 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Phuoc:2020:EBI

[PRC20] Nguyen Ngoc Phuoc, Randolph Richards, and Margaret Crumlish. Establishing bacterial infectivity models in striped Catfish Pangasianodon hypophthalmus (Sauvage) with Edwardsiella ictaluri. Journal of Fish Diseases, 43(3):371–378, March 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Poppe:2021:NMP

[PRHW21] Trygve T. Poppe, Anne Katrine Reed, Mette Hofossæter, and Helene Wisløff. Novel myocardial pathology in farmed salmonids. Journal of Fish Diseases, 44(10):1657–1660, October 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Poppe:2024:BSA

[PRW24] Trygve T. Poppe, Anne Katrine Reed, and Helene Wisløff. Basibranchial structure affecting cardiac morphology in Atlantic salmon (Salmo salar L.). Journal of Fish Diseases, 47 (4):e13911:1–e13911:??, April 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Prasartset:2023:SDT

[PS23] Tharinthon Prasartset and Win Surachetpong. Simultaneous detection of three important viruses affecting tilapia using a multiplex PCR assay. Journal of Fish Diseases, 46 (4):459–464, April 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Picard-Sanchez:2020:WTT

[PSED^{+}20] Amparo Picard-Sánchez, Itziar Estensoro, Raquel Del Pozo, Oswaldo R. Palenzuela, Maria Carla Piazzon, and Ariadna Sitjà-Bobadilla. Water temperature, time of exposure and population density are key parameters in Enteromyxum leei fish-to-fish experimental transmission. Journal of Fish Diseases, 43(4):491–502, April 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pierezan:2020:OSM

[PSH^{+}20] Felipe Pierezan, Khalid Shahin, Taylor I. Heckman, June Ang, Barbara A. Byrne, and Esteban Soto. Outbreaks of severe myositis in cultured white sturgeon (Acipenser transmontanus L.) associated with Streptococcus iniae. Journal of Fish Diseases, 43(4):485–490, April 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Parent:2021:EDL

[PSH^{+}21] Marianne I. Parent, Henrik Stryhn, K. Larry Hammell, Mark D. Fast, Jon Grant, and Raphaël Vanderstichel. Estimating the dispersal of Lepeophtheirus salmonis sea lice within and among Atlantic salmon sites of the Bay of Fundy, New Brunswick. Journal of Fish Diseases, 44(12):1971–1984, December 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Piamsomboon:2022:SAI

[PSKW22] Patharapol Piamsomboon, Sayamon Srisuwatanasagul, Kittipot Kongsonthana, and Janenuj Wongtavatchai. Streptococcus agalactiae infection caused spinal deformity in juvenile red tilapia (Oreochromis sp.). Journal of Fish Diseases, 45(4):603–606, April 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pallandre:2022:FSP

[PSPP^{+}22] Laurane Pallandre, Heike Schmidt-Posthaus, Françoise Pozet, Mélanie Rupp, Elodie Cristina, and Laurent Bigarré. Further spread of perch rhabdovirus on European percid farms. Journal of Fish Diseases, 45(12):1831–1837, December 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Prasitporn:2021:DCP

[PSV^{+}21] Terawut Prasitporn, Saengchan Senapin, Akapon Vaniksampanna, Siwaporn Longyant, and Parin Chaivisuthangkura. Development of cross-priming amplification (CPA) combined with colorimetric and lateral flow dipstick visualization for scale drop disease virus (SDDV) detection. Journal of Fish Diseases, 44(9):1411–1422, September 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Parchemin:2022:ASD

[PTBSF22] Christelle Parchemin, Nathalie Tapissier-Bontemps, Pierre Sasal, and Elisabeth Faliex. Anguilla sp. diseases diagnoses and treatments: The ideal methods at the crossroads of conservation and aquaculture purposes. Journal of Fish Diseases, 45(7):943–969, July 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pepey:2021:AFE

[PTK^{+}21] Elodie Pepey, Taukhid Taukhid, Nicolas Keck, Angela Lusiastuti, Jean-Christophe Avarre, Gusnia Sundari, Samira Sarter, and Domenico Caruso. Application of the FTA elute card coupled with visual colorimetric loop-mediated isothermal amplification for the rapid diagnosis of Streptococcus agalactiae in farmed tilapia (Oreochromis niloticus). Journal of Fish Diseases, 44(5):505–512, May 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Paredes-Trujillo:2022:SRM

[PTMC22] Amelia Paredes-Trujillo and Manuel Mendoza-Carranza. A systematic review and meta-analysis of the relationship between farm management, water quality and pathogen outbreaks in tilapia culture. Journal of Fish Diseases, 45(10): 1529–1548, October 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Phasunon:2023:IDS

[PTP^{+}23] Ramida Phasunon, Suwimon Taengphu, Wattana Panphut, Satid Chatchaiphan, Ha Thanh Dong, and Saengchan Senapin. Improving the diagnosis of Streptococcus iniae using a novel probe-based qPCR assay combined with an enrichment step. Journal of Fish Diseases, 46(12):1391–1401, December 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Piamsomboon:2021:EAS

[PTW21] Patharapol Piamsomboon, Pornpawit Tanpichai, and Janenuj Wongtavatchai. Enteritis associated with subclinical infection of Streptococcus iniae in juvenile Asian seabass Lates calcarifer (Bloch, 1790). Journal of Fish Diseases, 44 (11):1879–1882, November 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Polinski:2020:POB

[PVCG20] Mark P. Polinski, Niccoló Vendramin, Argelia Cuenca, and Kyle A. Garver. Piscine orthoreovirus: Biology and distribution in farmed and wild fish. Journal of Fish Diseases, 43(11):1331–1352, November 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pekmezci:2022:GDI

[PYD^{+}22] Gokmen Zafer Pekmezci, Alparslan Yildirim, Onder Duzlu, Emrah Simsek, Fikri Balta, Banu Yardimci, Ertan Emek Onuk, Zuhal Onder, Arif Ciloglu, Gamze Yetismis, Erdal Yilmaz, and Abdullah Inci. Genetic diversity of Ichthyophthirius multifiliis (Fouquet, 1876) infecting farmed rainbow trout (Oncorhynchus mykiss Walbaum, 1792) in Turkey. Journal of Fish Diseases, 45(8):1109–1115, August 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Phusantisampan:2022:SRD

[PYT^{+}22] Theerawut Phusantisampan, Jidapa Yamkasem, Puntanat Tattiyapong, Malinee Sriariyanun, and Win Surachetpong. Specific and rapid detection of tilapia parvovirus using loopmediated isothermal amplification (LAMP) method. Journal of Fish Diseases, 45(12):1893–1898, December 2022. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Passantino:2020:IMC

[PZP^{+}20] Letizia Passantino, Rosa Zupa, Chrysovalentinos Pousis, Constantinos C. Mylonas, Edmond Hala, Emilio Jirillo, and Aldo Corriero. Increased melanomacrophage centres in the liver of reproductively dysfunctional female greater amberjack Seriola dumerili (Risso, 1810). Journal of Fish Diseases, 43(4):503–514, April 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Que:2022:PAR

[QCZ^{+}22] Muyi Que, Weiwei Cao, Huang Zhang, Lei Shi, and Lei Ye. The prevalence, antibiotic resistance and multilocus sequence typing of colistin-resistant bacteria isolated from Penaeus vannamei farms in earthen ponds and HDPE filmlined ponds in China. Journal of Fish Diseases, 45(9):1289– 1299, September 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Qian:2023:MHC

[QFW^{+}23] Kun Qian, Shengli Fu, Junru Wang, Yuan Li, Jianan Xian, and Jianmin Ye. Major histocompatibility complex class IIαandIIβofpufferfish(Takifuguobscurus) : Identificationandfunctionalcharacterizationinresponsetobacterialchallenge.JournalofFishDiseases, 46 (10) : 1049−−1064, October2023.CODENJF IDDI.ISSN0140− 7775(print), 1365 − 2761(electronic).

Quintanilla:2021:HTP

[QGG^{+}21] Juan Carlos Quintanilla, Margarita P. González, Juan Pablo García, Paola Olmos, and Sergio Contreras-Lynch. Horizontal transmission of Piscirickettsia salmonis from the wild sub-Antarctic notothenioid fish Eleginops maclovinus to rainbow trout (Oncorhynchus mykiss) under experimental conditions. Journal of Fish Diseases, 44(7):993–1004, July 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Quezada-Rodriguez:2022:PEF

[QRTD^{+}22] Petra R. Quezada-Rodriguez, Richard S. Taylor, Jamie Downes, Fintan Egan, Samantha White, Aisling Brenan, Megan Rigby, Barbara F. Nowak, Neil M. Ruane, and James W. Wynne. Prevalence of epitheliocystis in freshwater Atlantic salmon reared in flow-through and recirculation aquaculture systems. Journal of Fish Diseases, 45(11):1721– 1731, November 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Qu:2021:CPA

[QST^{+}21] Shenye Qu, Chenguang Song, Xiaoping Tan, Gaoxue Wang, and Fei Ling. Comparative proteomic analysis provides insight into the key proteins as potential targets underlying the effect of malachite green against Ichthyophthirius multifiliis. Journal of Fish Diseases, 44(7):881–892, July 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Qu:2021:DRT

[QWL^{+}21] Yang Qu, Qing Wang, Yingying Li, Yingying Wang, Jiyuan Yin, Yan Ren, Chun Liu, Xiaofang Liu, Yahui Wang, and Weiwei Zeng. Development of a real-time recombinase polymerase amplification assay for rapid detection of Aeromonas hydrophila. Journal of Fish Diseases, 44(4):469–477, April 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rehulka:2024:SFE

[ŘB24] Jiří Řehulka and Jaroslav Bradík. Study of the frequencies of erythrocyte abnormalities as in situ biomarkers of genotoxic risk of chemicals in special fish stock in water supply reservoirs. Journal of Fish Diseases, 47(4):e13909:1–e13909:??, April 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Raharjo:2022:DVF

[RBM^{+}22] Hartanto Mulyo Raharjo, Hendri Budiyansah, Muhammad Fadhlullah Mursalim, Putita Chokmangmeepisarn, Rungnapa Sakulworakan, Sulaiman Madyod, Mariya Sewaka, Molruedee Sonthi, Partho Pratim Debnath, Sivaramasamy Elayaraja, Tilladit Rung-ruangkijkrai, Ha Thanh Dong, and Channarong Rodkhum. Distribution of Vibrionaceae in farmed Asian sea bass, Lates calcarifer in Thailand and their high prevalence of antimicrobial resistance. Journal of Fish Diseases, 45(9):1355–1371, September 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rao:2020:EFI

[RBP^{+}20a] S. Rao, O. Byadgi, T. Pulpipat, P-C. Wang, and S-C. Chen. Efficacy of a formalin-inactivated Lactococcus garvieae vaccine in farmed grey mullet (Mugil cephalus). Journal of Fish Diseases, 43(12):1579–1589, December 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rud:2020:FGC

[RBP^{+}20b] Yuriy Rud, Laurent Bigarré, Laurane Pallandre, François-Xavier Briand, and Leonid Buchatsky. First genetic characterization of sturgeon mimiviruses in Ukraine. Journal of Fish Diseases, 43(11):1391–1400, November 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rahmawaty:2022:PGA

[RCB^{+}22] Atiek Rahmawaty, Mei-Yun Chen, Omkar Vijay Byadgi, Pei-Chi Wang, and Shih-Chu Chen. Phenotypic and genotypic analysis of Edwardsiella isolates from Taiwan indicates wide variation with a particular reference to Edwardsiella tarda and Edwardsiella anguillarum. Journal of Fish Diseases, 45 (11):1659–1672, November 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rairat:2021:DEA

[RCC^{+}21] T. Rairat, Y. Chi, S-K. Chang, C-Y. Hsieh, N. Chuchird, and C-C. Chou. Differential effects of aquatic anaesthetics on the pharmacokinetics of antibiotics: Examples using florfenicol in Nile tilapia (Oreochromis niloticus). Journal of Fish Diseases, 44(10):1579–1586, October 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Riborg:2022:BRL

[RCG22] Andreas Riborg, Duncan J. Colquhoun, and Snorre Gulla. Biotyping reveals loss of motility in two distinct Yersinia ruckeri lineages exclusive to Norwegian aquaculture. Journal of Fish Diseases, 45(5):641–653, May 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rao:2022:GPL

[RCS^{+}22] Shreesha Rao, Mei-Yun Chen, Chiranan Sudpraseart, Peiry Lin, Terutoyo Yoshida, Pei-Chi Wang, and Shih-Chu Chen. Genotyping and phenotyping of Lactococcus garvieae isolates from fish by pulse-field gel electrophoresis (PFGE) and electron microscopy indicate geographical and capsular variations. Journal of Fish Diseases, 45(6):771–781, June 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Radosavljevic:2023:PAS

[RCW^{+}23] Vladimir Radosavljevic, Argelia Cuenca, Gareth Wood, Dimitrije Glisic, Jelena Maksimovic-Zoric, and David Stone. Phylogenetic analysis of spring viraemia of carp virus isolated in Serbia. Journal of Fish Diseases, 46(12):1343–1355, December 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ronza:2020:EES

[REB^{+}20] Paolo Ronza, Itziar Estensoro, Roberto Bermúdez, Ana Paula Losada, Gregorio Pérez-Cordón, Belén G. Pardo, Ariadna Sitjà-Bobadilla, and Ma. Isabel Quiroga. Effects of Enteromyxum spp. (Myxozoa) infection in the regulation of intestinal E-cadherin: Turbot against gilthead sea bream. Journal of Fish Diseases, 43(3):337–346, March 2020. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Riepe:2023:VTR

[RFN^{+}23] Tawni B. Riepe, Eric R. Fetherman, Brad Neuschwanger, Tracy Davis, Andrew Perkins, and Dana L. Winkelman. Vertical transmission of Renibacterium salmoninarum in cutthroat trout (Oncorhynchus clarkii). Journal of Fish Diseases, 46(4):309–319, April 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Romero:2022:DEV

[RGH^{+}22] João F. Romero, Ian A. Gardner, Larry Hammell, David Groman, Daryl Whelan, Nicole O’Brien, Leighanne J. Hawkins, Holly Burnley, and Krishna Thakur. Descriptive epidemiology of variants of infectious salmon anaemia virus in four Atlantic salmon farms in Newfoundland and Labrador, Canada. Journal of Fish Diseases, 45(6):919–930, June 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rehman:2021:ZIT

[RGS^{+}21] Saima Rehman, Adnan H. Gora, Prabhugouda Siriyappagouder, Sylvia Brugman, Jorge M. O. Fernandes, Jorge Dias, and Viswanath Kiron. Zebrafish intestinal transcriptome highlights subdued inflammatory responses to dietary soya bean and efficacy of yeast β-glucan. Journal of Fish Diseases, 44(10):1619–1637, October 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Riborg:2022:QSY

[RGS^{+}22] Andreas Riborg, Snorre Gulla, David Strand, Jannicke Wiik-Nielsen, Anita Rønneseth, Timothy J. Welch, Bjørn Spilsberg, and Duncan J. Colquhoun. qPCR screening for Yersinia ruckeri clonal complex 1 against a background of putatively avirulent strains in Norwegian aquaculture. Journal of Fish Diseases, 45(8):1211–1224, August 2022. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rigos:2023:FEO

[RGT^{+}23] George Rigos, Alexis Glaropoulos, Kostantinos Tzokas, Evgenia Gourzioti, Dimitra Kogiannou, and Eleni Golomazou. A field evaluation of orally administered praziquantel against the gill fluke Sparicotyle chrysophrii infecting gilthead seabream (Sparus aurata). Journal of Fish Diseases, 46(12):1439–1443, December 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rairat:2023:POD

[RKC^{+}23] Tirawat Rairat, Sunisa Kumphaphat, Niti Chuchird, Prapansak Srisapoome, Putsucha Phansawat, Arunothai Keetanon, Yi-Kai Liu, and Chi-Chung Chou. Pharmacokinetics, optimal dosages and withdrawal time of florfenicol in Asian seabass (Lates calcarifer) after oral administration via medicated feed. Journal of Fish Diseases, 46(1):75–84, January 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roh:2020:GWC

[RKL^{+}20] HyeongJin Roh, Bo Seong Kim, Mu Kun Lee, Chan-Il Park, and Do-Hyung Kim. Genome-wide comparison of Carnobacterium maltaromaticum derived from diseased fish harbouring important virulence-related genes. Journal of Fish Diseases, 43(9):1029–1037, September 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Radkhah:2020:BRF

[RM20] Ali Reza Radkhah and Esmaeil Sadeghinejad Masouleh. Book review: Fish Diseases and Medicine. 2019. Edited by Smith, S.A. CRC Press, Taylor & Francis Group. 413 pages. ISBN: 978–1–4987–2786–0 (Hardback). Journal of Fish Diseases, 43(10):1111–1113, October 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roh:2022:GWA

[Roh22] HyeongJin Roh. A genome-wide association study of the occurrence of genetic variations in Edwardsiella piscicida, Vibrio harveyi, and Streptococcus parauberis under stressed environments. Journal of Fish Diseases, 45(9):1373–1388, September 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Reed:2023:MGN

[RØLW23] William Reed, Liv Østevik, Kai-Inge Lie, and Helene Wisløff. Mycobacteriosis in Norwegian farmed Atlantic salmon (Salmo salar L.). Journal of Fish Diseases, 46(10):

1151–1155, October 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rathnapala:2023:GCI

[RRK^{+}23] Jayasundara Mudiyanselage Sajani Nisansala Rathnapala, Wafaa Ragab, Satoshi Kawato, Miho Furukawa, Reiko Nozaki, Hidehiro Kondo, and Ikuo Hirono. Genomic characterization and identification of virulence-related genes in Vibrio nigripulchritudo isolated from white leg shrimp Penaeus vannamei. Journal of Fish Diseases, 46(7):779–790, July 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ros:2022:MHI

[RSPB22] Albert Ros, Heike Schmidt-Posthaus, and Alexander Brinker. Mitigating human impacts including climate change on proliferative kidney disease in salmonids of running waters. Journal of Fish Diseases, 45(4):497–521, April 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rosaeg:2021:EVA

[RTA21] Magnus Vikan Røsæg, Ragnar Thorarinsson, and Arnfinn Aunsmo. Effect of vaccines against pancreas disease in farmed Atlantic salmon. Journal of Fish Diseases, 44(12): 1911–1924, December 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rich:2021:MCS

[RZTV21] Andrew Frederick Rich, Flavia Zendri, Susan Thornton, and Ranieri Verin. Meningoencephalitis with chondrocranial subcutaneous abscessation and septicaemia in a captive lemon shark (Negaprion brevirostris). Journal of Fish Diseases, 44 (1):113–117, January 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sabo-Attwood:2021:NSA

[SAAB^{+}21] Tara Sabo-Attwood, Onur G. Apul, Joseph H. Bisesi, Jr., Andrew S. Kane, and Navid B. Saleh. Nano-scale applications in aquaculture: Opportunities for improved production and disease control. Journal of Fish Diseases, 44(4):359–370, April 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Suryakodi:2022:FRO

[SAB^{+}22] Selvam Suryakodi, Abdul Nafeez Ahmed, Allahbagash Badhusha, Shanmugam Santhosh Kumar, Selvam Sivakumar, Seepoo Abdul Majeed, Gani Taju, Syed Rahamathulla, and Azeez Sait Sahul Hameed. First report on the occurrence of white spot syndrome virus, infectious myonecrosis virus and Enterocytozoon hepatopenaei in Penaeus vannamei reared in freshwater systems. Journal of Fish Diseases, 45(5):699– 706, May 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stilwell:2020:RGC

[SAKC20] Justin M. Stilwell, Chelsea Anderson, Spencer Kehoe, and Alvin C. Camus. Retrobulbar gangliocytoma and carcinoma of the corpuscles of Stannius in a pearl gourami, Trichopodus leerii (Bleeker, 1852). Journal of Fish Diseases, 43(11):1473– 1478, November 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shivley:2021:MCB

[SBB21] Jacob Michael Shivley, Wilson Cooper Brookshire, and Wes Arend Baumgartner. Mixed Chromatophoroma (benign irido-melanocytoma) in a male Siamese fighting fish, Betta splendens, Regan. Journal of Fish Diseases, 44(3): 351–354, March 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Seternes:2021:SEC

[SBD21] Tore Seternes, Jarl Bøgwald, and Roy A. Dalmo. Scavenger endothelial cells of fish, a review. Journal of Fish Diseases, 44(9):1385–1397, September 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Savage:2022:PMO

[SBS^{+}22] Ayanna Carla N. Phillips Savage, Lemar Blake, Rod Suepaul, O’Shane McHugh, Ray Rodgers, Calvern Thomas, Christopher Oura, and Esteban Soto. Piscine mycobacteriosis in the ornamental fish trade in Trinidad and Tobago. Journal of Fish Diseases, 45(4):547–560, April 2022. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stilwell:2021:ENM

[SBVC21] Justin M. Stilwell, Shane M. Boylan, Bryan Vorbach, and Alvin C. Camus. Epizootic neoplasia in a managed population of Atlantic bumper fish, Chloroscombrus chrysurus (Osteichthyes: Carangidae), at a public aquarium. Journal of Fish Diseases, 44(4):415–427, April 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Saldarriaga-Cordoba:2022:CPG

[SCAH22] Mónica Saldarriaga-Córdoba and Ruben Avendaño-Herrera. Comparative pan-genomic analysis of 51 Renibacterium salmoninarum indicates heterogeneity in the principal virulence factor, the 57 kDa protein. Journal of Fish Diseases, 45 (8):1173–1188, August 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Saldarriaga-Cordoba:2023:DPA

[SCALIAH23] Mónica Saldarriaga-Córdoba, Henry Araya-León, Rute Irgang, and Ruben Avendaño-Herrera. Development of a PCR assay for the specific identification of the fish pathogen Tenacibaculum piscium. Journal of Fish Diseases, 46(5): 517–526, May 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sveen:2023:LBS

[SCD^{+}23] Lene R. Sveen, Elvis Chikwati, Alf Dalum, Trond M. Kortner, Jarle Ravndal, Olai Einen, and Turid Mørkøre. Lesions of the biliary system in Atlantic salmon (Salmo salar L.). Journal of Fish Diseases, 46(3):281–285, March 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Saldarriaga-Cordoba:2021:CBG

[SCIAH21] Mónica Saldarriaga-Córdoba, Rute Irgang, and Ruben Avendaño-Herrera. Comparison between genome sequences of Chilean Tenacibaculum dicentrarchi isolated from red conger eel (Genypterus chilensis) and Atlantic salmon (Salmo salar) focusing on bacterial virulence determinants. Journal of Fish Diseases, 44(11):1843–1860, November 2021. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Suhaimi:2023:FDM

[SCRS^{+}23] Nadhirah Syafiqah Suhaimi, Graciela Colunga-Ramírez, Boglárka Sellyei, Gábor Cech, Kálmán Molnár, and Csaba Székely. The first detection of Myxobolus lentisuturalis Dyková, Fiala et Nie, 2002, a highly pathogenic muscleinfecting parasite of gibel carp (Carassius auratus gibelio Berg, 1932) in Hungary. Journal of Fish Diseases, 46(12): 1367–1376, December 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stilwell:2024:GNC

[SCS24] Justin M. Stilwell, Alvin C. Camus, and Natalie K. Stilwell. Gonadal neoplasia in Chinese gold barbs (Barbodes semifasciolatus) from a home aquarium. Journal of Fish Diseases, 47(3):e13901:1–e13901:??, March 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sellyei:2020:ICB

[SCV^{+}20] Boglárka Sellyei, Gábor Cech, Ádám Varga, Kálmán Molnár, Csaba Székely, Dóra Somogyi, Krisztián Nyeste, and László Antal. Infection of the Carpathian brook lamprey (Eudontomyzon danfordi Regan, 1911) with a dermocystid parasite in the Tisza River Basin, Hungary. Journal of Fish Diseases, 43(12):1571–1577, December 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sun:2021:EAA

[SCX^{+}21] Zhong-Chen Sun, Cheng Chen, Fei-Fan Xu, Bing-Ke Li, Jing-Lei Shen, Tao Wang, Hai-Feng Jiang, and Gao-Xue Wang. Evaluation of the antiviral activity of naringenin, a major constituent of Typha angustifolia, against white spot syndrome virus in crayfish Procambarus clarkii. Journal of Fish Diseases, 44(10):1503–1513, October 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Song:2022:TAT

[SDH22] Yuting Song, Xianzhi Dong, and Guobin Hu. Transcriptome analysis of turbot (Scophthalmus maximus) head kidney and liver reveals immune mechanism in response to Vibrio anguillarum infection. Journal of Fish Diseases, 45(7):1045– 1057, July 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Soto-Davila:2020:IMR

[SDMO^{+}20] Manuel Soto-Dávila, Danixa Martinez, Ricardo Oyarzún, Juan P. Pontigo, Carolina Vargas-Lagos, Francisco J. Morera, Julia Saravia, Fábio Zanuzzo, and Luis Vargas-Chacoff. Intermediary metabolic response and gene transcription modulation on the Sub-Antarctic notothenioid Eleginops maclovinus (Valenciennes, 1930) injected with two strains of Piscirickettsia salmonis. Journal of Fish Diseases, 43(1): 111–127, January 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Storesund:2022:HPI

[SdSNK^{+}22] Julia E. Storesund, Caroline da Silva Nylund, Egil Karlsbakk, Lucilla Giulietti, Miguel Bao, Paolo Cipriani, and Arne Levsen. High prevalence of Ichthyophonus sp. infections in Northeast Atlantic mackerel (Scomber scombrus). Journal of Fish Diseases, 45(8):1243–1249, August 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Standish:2020:VSC

[SEL^{+}20] Isaac Standish, Sara Erickson, Eric Leis, Wes Baumgartner, Thomas Loch, Christopher Knupp, Rebekah McCann, Corey Puzach, Ryan Katona, Ellen Lark, Jennifer Bailey, Jorge Buening, Carey Edwards, and Kenneth Phillips. Vagococcus salmoninarum I — a chronic coldwater streptococcosis in broodstock brook trout (Salvelinus fontinalis) in Wisconsin, USA. Journal of Fish Diseases, 43(3):305–316, March 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Svoboda:2020:EEP

[SFKB^{+}20] Jiří Svoboda, David Fischer, Eva Kozubíková-Balcarová, Andrea Šťástková, Marie Brůčková, Antonín Kouba, and Adam Petrusek. Experimental evaluation of the potential for crayfish plague transmission through the digestive system of warm-blooded predators. Journal of Fish Diseases, 43(1):129–138, January 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sirri:2020:DSM

[SGR^{+}20a] Rubina Sirri, Andrea Gustinelli, Silva Rubini, Francesco Quaglio, and Maria Letizia Fioravanti. Dermocystidium sp. (Mesomycetozoea: Dermocystidiaceae) primary ocular infection in a koi carp (Cyprinus carpio var . koi). Journal of Fish Diseases, 43(4):515–517, April 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stilwell:2020:IMC

[SGR^{+}20b] Justin M. Stilwell, Matt J. Griffin, Thomas G. Rosser, Haitham H. Mohammed, Inga F. Sidor, and Alvin C. Camus. Insights into myxozoan composition and physiology revealed by histochemical properties of myxospores. Journal of Fish Diseases, 43(5):583–597, May 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sergeenko:2021:CHP

[SGU^{+}21] Natalia V. Sergeenko, Tatyana V. Gavruseva, Elena A. Ustimenko, Elena V. Bochkova, and Tatyana V. Ryazanova. Case of highly prevalent papilloma-like skin lesions in whitespotted char (Salvelinus leucomaenis) in Kamchatka (Russia). Journal of Fish Diseases, 44(5):487–493, May 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sayed:2022:EEP

[SGW^{+}22] Mohamed Sayed, Matt Griffin, Cynthia Ware, Ozan Ozdemir, Hasan C. Tekedar, Manal Essa, Attila Karsi, Mark L. Lawrence, and Hossam Abdelhamed. Evaluation of Edwardsiella piscicida basS and basR mutants as vaccine candidates in catfish against edwardsiellosis. Journal of Fish Diseases, 45(12):1817–1829, December 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shokrpoor:2022:BCC

[SHS^{+}22a] Sara Shokrpoor, Hooman Rahmati Holasoo, Sarang Soroori, Amin Marandi, and Bahram Imantalab. Basal cell carcinoma in an albino pindani (Chindongo socolofi) and a cobaltzebra (Maylandia callainos): Diagnostic imaging, clinical and histopathological study. Journal of Fish Diseases, 45 (9):1251–1258, September 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shokrpoor:2022:CI

[SHS^{+}22b] Sara Shokrpoor, Hooman Rahmati Holasoo, Sarang Soroori, Amin Marandi, and Bahram Imantalab. Cover image. Journal of Fish Diseases, 45(9):i, September 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sohn:2021:IDE

[SJK^{+}21] Hanchang Sohn, Chang Nam Jin, Bong Jo Kang, Sang Phil Shin, and Jehee Lee. Infection dynamics of Enteromyxum leei (Myxozoa, Myxosporea) in culture water and its effects on cultured olive flounder, Paralichthys olivaceus (Temminck & Schlegel). Journal of Fish Diseases, 44(9):1475–1479, September 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shu:2023:GGR

[SJL^{+}23] Feng-Ling Shu, Li-Yun Jin, Hao Liu, Zhen Tao, Fei Yin, Jia-Song Xie, and Su-Ming Zhou. The galU gene is required for in vivo survival of Pseudomonas plecoglossicida in large yellow croaker (Larimichthys crocea). Journal of Fish Diseases, 46 (3):229–238, March 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Schuster:2022:PID

[SKAS^{+}22] Corbin J. Schuster, Taylor G. Kreul, Colleen E. Al-Samarrie, James T. Peterson, Justin L. Sanders, and Michael L. Kent. Progression of infection and detection of Pseudoloma neurophilia in zebrafish Danio rerio Hamilton by PCR and histology. Journal of Fish Diseases, 45(10):1463–1475, October 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Scholz:2021:TPF

[SKO21] Tomáš Scholz, Roman Kuchta, and Mikuláš Oros. Tapeworms as pathogens of fish: a review. Journal of Fish Diseases, 44(12):1883–1900, December 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Siah:2020:RTR

[SKR^{+}20] Ahmed Siah, Emilie Knutsen, Zina Richmond, Meghan Mills, Kathleen Frisch, James F. F. Powell, Øyvind Brevik, and Henrik Duesund. Real-time RT-qPCR assay to detect sequences in the Piscine orthoreovirus-1 genome segment S1 associated with heart and skeletal muscle inflammation in Atlantic salmon. Journal of Fish Diseases, 43(8):955–962, August 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Suyapoh:2024:MEC

[SKS^{+}24] Watcharapol Suyapoh, Narissara Keawchana, Peerapon Sornying, Sirikachorn Tangkawattana, Pokphon Khirilak, and Sasibha Jantrakajorn. Mixed Eimeria and Cryptosporidium infection and its effects on pathology and clinical outcomes in juvenile Asian seabass (Lates calcarifer) cultured in Thailand. Journal of Fish Diseases, 47(4):e13914:1– e13914:??, April 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Simora:2020:CEP

[SLA^{+}20] Rhoda Mae C. Simora, Shangjia Li, Nermeen Y. Abass, Jeffery S. Terhune, and Rex A. Dunham. Cathelicidins enhance protection of channel catfish, Ictalurus punctatus, and channel catfish [female sign] × blue catfish, Ictalurus furcatus [male sign] hybrid catfish against Edwarsiella ictaluri infection. Journal of Fish Diseases, 43(12):1553–1562, December 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Standish:2020:VSI

[SLE^{+}20] Isaac Standish, Eric Leis, Sara Erickson, Rebekah McCann, Corey Puzach, Ryan Katona, Ellen Lark, Jennifer Bailey, Eric Kleman, Jorge Buening, Carey Edwards, Thomas Loch, and Kenneth Phillips. Vagococcus salmoninarum II — qPCR, tropism and egg-associated transmission. Journal of Fish Diseases, 43(3):317–325, March 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shi:2024:RTO

[SLH^{+}24] Yan Shi, Xiaying Liang, Sufei Hu, MengMeng Wang, Yifan Wang, and Zhe Zhao. Role of TRAF6 from obscure puffer (Takifugu obscurus) in immune response against Edwardsiella tarda infection. Journal of Fish Diseases, 47(2): e13877:1–e13877:??, February 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Schuster:2023:EDP

[SLK^{+}23] Corbin J. Schuster, Connor Leong, Kristin D. Kasschau, Thomas J. Sharpton, and Michael L. Kent. Early detection of Pseudocapillaria tomentosa by qPCR in four lines of zebrafish, Danio rerio (Hamilton 1882). Journal of Fish Diseases, 46(6):619–627, June 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sara:2020:HEP

[SLPF^{+}20] Joseph R. Sara, Wilmien J. Luus-Powell, Susan B. Fogelson, Hannes Botha, Theresa C. Guillette, Willem J. Smit, Andre Hoffman, Katlego D. Kunutu, Jeremy P. Koelmel, and John A. Bowden. A histological evaluation of pansteatitisaffected Mozambique tilapia, Oreochromis mossambicus (Peters 1852), from different geographical locations in South Africa. Journal of Fish Diseases, 43(10):1185–1199, October 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Soto:2024:DID

[SLY^{+}24] Esteban Soto, Benjamin R. LaFrentz, Susan Yun, Dorothea Megarani, Eileen Henderson, Chutchai Piewbang, Amber E. Johnston, Somporn Techangamsuwan, Terry Fei Fan Ng, Janet Warg, Win Surachetpong, and Kuttichantran Subramaniam. Diagnosis, isolation and description of a novel amnoonvirus recovered from diseased fancy guppies, Poecilia reticulata. Journal of Fish Diseases, 47(6):e13937:1– e13937:??, June 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sibinga:2021:TSD

[SM21] Nathaniel A. Sibinga and Hélène Marquis. Tissue-specific differences in detection of Yersinia ruckeri carrier status in rainbow trout (Oncorhynchus mykiss). Journal of Fish Diseases, 44(12):2013–2020, December 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sellyei:2022:EPS

[SMC^{+}22] Boglárka Sellyei, Kálmán Molnár, Gábor Cech, Ádám Varga, and Csaba Székely. Extreme pathological symptom generated by Dermocystidium koi infection of common carp (Cyprinus carpio L.) in Hungary. Journal of Fish Diseases, 45(6):825–832, June 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Salmonova:2020:PPC

[SMD^{+}20] Hana Šubrtová Salmonová, Matilde Marchi, Ivo Doskočil, Tereza Kodešová, and Eva Vlková. Pathogenic profile and cytotoxic activity of Aeromonas spp. isolated from Pectinatella magnifica and surrounding water in the South Bohemian aquaculture region. Journal of Fish Diseases, 43(10): 1213–1227, October 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Simon:2023:DRR

[SMG^{+}23] Rocío Simón, Patricia Martínez, Lucía González, M. Camino Ordás, and Carolina Tafalla. Differential response of RTGUTGC and RTGILL-W1 rainbow trout epithelial cell lines to viral stimulation. Journal of Fish Diseases, 46(4): 433–443, April 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Srivastava:2020:AEC

[SMMM20] Ayan Srivastava, Arup Mistri, Swati Mittal, and Ajay Kumar Mittal. Alterations in the epidermis of the carp, Labeo rohita (Cyprinidae: Cypriniformes), infected by the bacteria, Aeromonas hydrophila: a scanning electron microscopic, histopathological and immunohistochemical investigation. Journal of Fish Diseases, 43(8):941–953, August 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stejskal:2020:FDG

[SMP^{+}20] Vlastimil Stejskal, Jan Matoušek, Markéta Prokešová, Peter Podhorec, Jiří Křišťan, Tomáš Policar, and Tatyana Gebauer. Fin damage and growth parameters relative to stocking density and feeding method in intensively cultured European perch (Perca fluviatilis L.). Journal of Fish Diseases, 43(2):253–262, February 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Serra:2022:OEE

[SMP^{+}22a] Bianca Natally Viana Serra, Maurício Laterça Martins, Scheila Anelise Pereira, Edsandra Campos Chagas, Francisco Célio Maia Chaves, Maria Inês Braga de Oliveira, and Gabriela Tomas Jerônimo. Ovicidal effect of essential oils of Lippia alba, Lippia sidoides and Lippia gracilis on the acanthocephalan Neoechinorhynchus buttnerae (Eoacanthocephala: Neoechinorhynchidae). Journal of Fish Diseases, 45(4):569–577, April 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Standish:2022:DDQ

[SMP^{+}22b] Isaac Standish, Rebekah McCann, Corey Puzach, Eric Leis, Jennifer Bailey, Sara Dziki, Ryan Katona, Ellen Lark, Carey Edwards, Brandon Keesler, Stephen Reichley, Stacy King, Christopher Knupp, Courtney Harrison, Thomas Loch, and Kenneth Phillips. Development of duplex qPCR targeting Carnobacterium maltaromaticum and Vagococcus salmoninarum. Journal of Fish Diseases, 45(5):667–677, May 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shahin:2022:DQP

[SMY^{+}22] Khalid Shahin, Kaveramma Mukkatira, Zeinab Yazdi, Christine Richey, Kevin Kwak, Taylor I. Heckman, Haitham H. Mohammed, Cesar Ortega, Ruben Avendaño-Herrera, Bill Keleher, Michael W. Hyatt, John D. Drennan, Mark Adkison, Matt J. Griffin, and Esteban Soto. Development of a quantitative polymerase chain reaction assay for detection of the aetiological agents of piscine lactococcosis. Journal of Fish Diseases, 45(6):847–859, June 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Saito:2024:ETP

[SMY^{+}24] Hiroaki Saito, Shungo Minami, Manami Yuguchi, Aiko Shitara, Hidehiro Kondo, Goshi Kato, and Motohiko Sano. Effect of temperature on the protective efficacy of a live attenuated vaccine against herpesviral haematopoietic necrosis in goldfish. Journal of Fish Diseases, 47(4):e13906:1–e13906:??, April 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Speare:2023:RMA

[SNSP^{+}23] David J. Speare, Barbara Nowak, Heike Schmidt-Posthaus, Saengchan Senapin, and Emily Chappell. Rise of the multiauthored paper & introduction of CRediT to the Journal of Fish Diseases submission process. Journal of Fish Diseases, 46(9):903, September 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Spilsberg:2022:TNA

[SNT^{+}22] Bjørn Spilsberg, Hanne K. Nilsen, Saraya Tavornpanich, Snorre Gulla, Mona Dverdal Jansen, Karin Lagesen, Duncan J. Colquhoun, and Anne-Berit Olsen. Tenacibaculosis in Norwegian Atlantic salmon (Salmo salar) cage-farmed in cold sea water is primarily associated with Tenacibaculum finnmarkense genomovar finnmarkense. Journal of Fish Diseases, 45(4):523–534, April 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sakseepipad:2021:DSN

[SNY^{+}21] Channapha Sakseepipad, Reiko Nozaki, Keisuke Yoshii, Yutaka Fukuda, Kaori Mizuno, Hidemasa Kawakami, Ikuo Hirono, and Hidehiro Kondo. Development of single nucleotide polymorphism (SNP) application for detection and genotyping of RSIV-type megalocytiviruses. Journal of Fish Diseases, 44(9):1337–1342, September 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sayed:2021:VLV

[SOE^{+}21] Mohamed Sayed, Ozan Ozdemir, Manal Essa, Alicia Olivier, Attila Karsi, Mark L. Lawrence, and Hossam Abdelhamed. Virulence and live vaccine potential of Edwardsiella piscicida phop and phoq mutants in catfish against edwardsiellosis. Journal of Fish Diseases, 44(9):1463–1474, September 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sibanda:2023:EEU

[SPHM23] Sitokozile Sibanda, Davies M. Pfukenyi, Bernard Hang’ombe, and Gift Matope. Epidemiology of epizootic ulcerative syndrome (EUS) in fish in the main water bodies of the Kavango–Zambezi and Great Limpopo transfrontier conservation areas of Zimbabwe. Journal of Fish Diseases, 46(3): 201–213, March 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Schmidt-Posthaus:2020:MSC

[SPKS^{+}20] Heike Schmidt-Posthaus, Michel Koch, Torsten Seuberlich, Christoph Birrer, Regula Hirschi, and Michael Kugler. Mysterious syndrome causing high mortality in wild brown trout in Eastern Switzerland, pathology and search for a possible cause. Journal of Fish Diseases, 43(10):1317–1324, October 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Syrova:2020:FSI

[SPM^{+}20] Eva Syrová, Miroslava Palíková, Jan Mendel, Veronika Seidlová, Ivana Papežíková, Heike Schmidt-Posthaus, Kristina Somerlíková, Hana Minářová, Lukáš Mareš, Ivana Mikulíková, Jiří Pikula, and Jan Mareš. Field study indicating susceptibility differences between salmonid species and their lineages to proliferative kidney disease. Journal of Fish Diseases, 43(10):1201–1211, October 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shahi:2023:SSD

[SPS23] Neetu Shahi, Tharinthon Prasartset, and Win Surachetpong. A specific and sensitive droplet digital polymerase chain reaction assay for the detection of tilapia lake virus in fish tissue and environmental samples. Journal of Fish Diseases, 46(9):957–966, September 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sandlund:2021:PSI

[SRE^{+}21] Nina Sandlund, Anita Rønneseth, Rebecca Marie Ellul, Stian Nylund, and Liv Sandlund. Pasteurella spp. Infections in Atlantic salmon and lumpsucker. Journal of Fish Diseases, 44(8):1201–1214, August 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Soto-Rodriguez:2024:PBP

[SRLARG24] Sonia A. Soto-Rodriguez, Francis I. Marrujo Lopez, Karla G. Aguilar-Rendon, and Rafael Hernández Guzmán. Pathogenic bacteria prevalence in cultured Nile tilapia in Southwest Mexico: a real-time PCR analysis. Journal of Fish Diseases, 47(5):e13921:1–e13921:??, May 2024. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Surachetpong:2020:TLV

[SRN20] Win Surachetpong, Sri Rajiv Kumar Roy, and Pamela Nicholson. Tilapia lake virus: the story so far. Journal of Fish Diseases, 43(10):1115–1132, October 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Swaminathan:2021:ISK

[SRP^{+}21] Thangaraj Raja Swaminathan, Nithianantham Sundar Raj, Prasannan Geetha Preena, Pravata Kumar Pradhan, Neeraj Sood, Rahul Girish Kumar, Arun Sudhagar, and Naresh Kumar Sood. Infectious spleen and kidney necrosis virusassociated large-scale mortality in farmed giant gourami, Osphronemus goramy, in India. Journal of Fish Diseases, 44 (12):2043–2053, December 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Samsing:2021:STI

[SRT^{+}21] Francisca Samsing, Megan Rigby, Hedda K. Tengesdal, Richard S. Taylor, Daniela Farias, Richard N. Morrison, Scott Godwin, Carla Giles, Jeremy Carson, Chloe J. English, Roger Chong, and James W. Wynne. Seawater transmission and infection dynamics of pilchard orthomyxovirus (POMV) in Atlantic salmon (Salmo salar). Journal of Fish Diseases, 44(1):73–88, January 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sandlund:2023:FIM

[SSK23] Nina Sandlund, Cecilie Skår, and Egil Karlsbakk. First identification of mycobacteriosis in Atlantic mackerel (Scomber scombrus). Journal of Fish Diseases, 46(5):527–533, May 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Schaaf:2020:RAZ

[SSM^{+}20] Russel M. Schaaf, Thomas J. Sharpton, Katrina N. Murray, Andrew D. Kent, and Michael L. Kent. Retrospective analysis of the Zebrafish International Resource Center diagnostic data links Pseudocapillaria tomentosa to intestinal neoplasms in zebrafish Danio rerio (Hamilton 1822). Journal of Fish Diseases, 43(11):1459–1462, November 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Seidlova:2021:CDM

[SSM^{+}21] Veronika Seidlova, Eva Syrova, Hana Minarova, Jan Zukal, Vojtech Balaz, Monika Nemcova, Ivana Papezikova, Jiri Pikula, Heike Schmidt-Posthaus, Jan Mares, and Miroslava Palikova. Comparison of diagnostic methods for Tetracapsuloides bryosalmonae detection in salmonid fish. Journal of Fish Diseases, 44(8):1147–1153, August 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sukonta:2022:CBP

[SSMC22] Thanwarat Sukonta, Saengchan Senapin, Watcharachai Meemetta, and Thawatchai Chaijarasphong. CRISPR-based platform for rapid, sensitive and field-deployable detection of scale drop disease virus in Asian sea bass (Lates calcarifer). Journal of Fish Diseases, 45(1):107–120, January 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sorgaard:2023:MGM

[SSP^{+}23] Martin Sørgaard, Karoline Sveinsson, Sonal Patel, Hanne K. Nilsen, Anne Berit Olsen, Øyvind Vaagnes, Duncan J. Colquhoun, and Snorre Gulla. MLVA genotyping of Moritella viscosa reveals serial emergence of novel, hostspecific clonal complexes in Norwegian salmon farming. Journal of Fish Diseases, 46(5):535–543, May 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shartau:2022:AME

[SST^{+}22] Ryan B. Shartau, Heindrich N. Snyman, Lenora Turcotte, Pearse McCarron, Julia C. Bradshaw, and Stewart C. Johnson. Acute microcystin exposure induces reversible histopathological changes in Chinook salmon (Oncorhynchus tshawytscha) and Atlantic salmon (Salmo salar). Journal of Fish Diseases, 45(5):729–742, May 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sirri:2021:HIC

[STB^{+}21] Rubina Sirri, Giorgia Tura, Jane Budai, Paola Beraldo, Michelangelo Fiorentino, Tim Barbé, Marco Galeotti, Giuseppe Sarli, and Luciana Mandrioli. Histological and immunohistochemical characterization of 17 gonadal tumours in koi carp (Cyprinus carpio koi). Journal of Fish Diseases, 44(3):273–285, March 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sellyei:2021:MIF

[SVC^{+}21] Boglárka Sellyei, Zsuzsanna Varga, Gábor Cech, Ádám Varga, and Csaba Székely. Mycoplasma infections in freshwater carnivorous fishes in Hungary. Journal of Fish Diseases, 44(3):297–304, March 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Scholz:2022:EIT

[SVO^{+}22] Felix Scholz, Niccolò Vendramin, Niels Jørgen Olesen, Argelia Cuenca, Tine Moesgaard Iburg, Luca Mirimin, Ian O’Connor, Neil M. Ruane, Hamish D. Rodger, and Eugene MacCarthy. Experimental infection trials with European North Atlantic ranavirus (Iridoviridae) isolated from lumpfish (Cyclopterus lumpus, l.). Journal of Fish Diseases, 45 (11):1745–1756, November 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sirri:2021:MOO

[SVP^{+}21] Rubina Sirri, Ranieri Verin, Alberto Perolo, Fernando Lunelli, Filippo Faccenda, Maria Letizia Fioravanti, and Francesco Quaglio. Metastatic osteoblastic osteosarcoma of the jaw in a marble trout (Salmo marmoratus). Journal of Fish Diseases, 44(4):483–485, April 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Santibanez:2024:VEP

[SVP^{+}24] Natacha Santibáñez, Matías Vega, Tatiana Pérez, Ricardo Enriquez, Carla Estefanía Escalona, Cristian Oliver, and Alex Romero. In vitro effects of phytogenic feed additive on Piscirickettsia salmonis growth and biofilm formation. Journal of Fish Diseases, 47(6):e13913:1–e13913:??, June 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sudpraseart:2021:PGP

[SWC21] Chiranan Sudpraseart, Pei-Chi Wang, and Shih-Chu Chen. Phenotype, genotype and pathogenicity of Streptococcus agalactiae isolated from cultured tilapia (Oreochromis spp.) in Taiwan. Journal of Fish Diseases, 44(6):747–756, June 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Song:2023:CRV

[SYKC23] Haichao Song, Bintong Yang, Yuanhuan Kang, and Wei Cong. Critical roles of VipB protein on virulence and oxidative stress tolerance in Aeromonas veronii. Journal of Fish Diseases, 46(5):487–497, May 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shan:2022:AAE

[SZH^{+}22] Li-Peng Shan, Xu Zhang, Yang Hu, Lei Liu, and Jiong Chen. Antiviral activity of esculin against white spot syndrome virus: a new starting point for prevention and control of white spot disease outbreaks in shrimp seedling culture. Journal of Fish Diseases, 45(1):59–68, January 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Samsing:2023:CTS

[SZZ^{+}23] Francisca Samsing, William Zhang, Ruth N. Zadoks, Richard Whittington, Carola Venturini, Carla Giles, Jeremy Carson, and Joy A. Becker. Cold temperature stress and damaged skin induced high mortality in barramundi (Lates calcarifer) challenged with Vibrio harveyi. Journal of Fish Diseases, 46(7):751–766, July 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Thamizhvanan:2021:SPP

[TAK^{+}21] S. Thamizhvanan, A. Nafeez Ahmed, D. Vinoth Kumar, S. Vimal, S. Abdul Majeed, G. Taju, Chris Hauton, and A. S. Sahul Hameed. Silencing of prophenoloxidase (proPO) gene in freshwater prawn, Macrobrachium rosenbergii, makes them susceptible to white spot syndrome virus (WSSV). Journal of Fish Diseases, 44(5):573–584, May 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tanjung:2023:BFD

[Tan23] Cahyani Fitrah Tanjung. Bacterial fish diseases: Environmental and economic constraints. Journal of Fish Diseases, 46(6):707–708, June 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Toro-Ascuy:2020:DIM

[TASP^{+}20] Daniela Toro-Ascuy, Alvaro Santibañez, Victor Peña, Carolina Beltran-Pavez, Luis Cottet, Cristian Molina, Margarita Montoya, Nicolas Sandoval, Yesseny Vásquez-Martínez, Carolina Mascayano, and Marcelo Cortez-San Martín. Development of an Isavirus minigenome system to study the function of the pocket RNA-binding domain of the viral nucleoprotein (NP) in salmon cells. Journal of Fish Diseases, 43(2): 197–206, February 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Thompson:2021:IPS

[TBB^{+}21] Cameron R. S. Thompson, James E. Bron, Samantha Bui, Sussie Dalvin, Mark J. Fordyce, Tomasz Furmanek, Gunnvør Á. Nordi, and Rasmus Skern-Mauritzen. Illuminating the planktonic stages of salmon lice: a unique fluorescence signal for rapid identification of a rare copepod in zooplankton assemblages. Journal of Fish Diseases, 44 (7):863–879, July 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Trangerud:2020:VCD

[TBK^{+}20] Cathrine Trangerud, Håvard Bjørgen, Erling Olaf Koppang, Randi Nygaard Grøntvedt, Hege Kippenes Skogmo, Nina Ottesen, and Agnar Kvellestad. Vertebral column deformity with curved cross-stitch vertebrae in Norwegian seawaterfarmed Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 43(3):379–389, March 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tamer:2022:DRP

[TBK^{+}22] Cuneyt Tamer, Jessica Benkaroun, Hanne Nur Kurucay, Harun Albayrak, and Manfred Weidmann. Development of a recombinase polymerase amplification assay for viral haemorrhagic septicemia virus. Journal of Fish Diseases, 45(8): 1065–1071, August 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tapia:2020:CMV

[TBKY20] David Tapia, Agustín Barría, Juan Kuznar, and José M. Yáñez. Comparison of mortality and viral load in rainbow trout (Oncorhynchus mykiss) infected with infectious pancreatic necrosis virus (IPNV) genogroups 1 and 5. Journal of Fish Diseases, 43(1):139–146, January 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tolo:2023:DEH

[TBMP23] Isaiah E. Tolo, Przemyslaw G. Bajer, Sunil K. Mor, and Nicholas B. D. Phelps. Disease ecology and host range of Cyprinid herpesvirus 3 (CyHV-3) in CyHV-3 endemic lakes of North America. Journal of Fish Diseases, 46(6):679–696, June 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tuna:2024:PHD

[TCC^{+}24] Fernando A. P. Tuna, Flávia A. A. Calixto, Woodson L. Cordeiro, Marcelline Luz, and Eliana F. M. Mesquita. Patterns of hyperostosis during the ontogenic development of Atlantic spadefish (Broussonet, 1782) in Brazilian coast. Journal of Fish Diseases, 47(2):e13882:1–e13882:??, February 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Torres-Corral:2021:DRT

[TCS21] Yolanda Torres-Corral and Ysabel Santos. Development of a real-time PCR assay for detection and quantification of Streptococcus iniae using the lactate permease gene. Journal of Fish Diseases, 44(1):53–61, January 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Takeuchi:2021:EFA

[THH^{+}21] Hisato Takeuchi, Motoyuki Hiratsuka, Kazutomo Hori, Hiroki Oinuma, Yoshiyuki Umino, Daiki Nakano, Mayu Iwadare, Ryuji Tomono, Toshihiro Imai, Hirohito Mashiko, Aki Namba, Tomohiro Takase, Sonoko Shimizu, Toshihiro Nakai, and Nobuhiro Mano. Environmental factors affecting Edwardsiella ictaluri-induced mortality of riverine ayu, Plecoglossus altivelis (Temminck & Schlegel). Journal of Fish Diseases, 44(8):1065–1074, August 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tian:2022:FSG

[THW^{+}22] Xueqing Tian, Jinfei Hu, Tong Wei, Wen Ding, Qingzhen Miao, Zhe Ning, Shihui Fan, Hangjie Wu, Jing Lu, Mingsheng Lyu, and Shujun Wang. Fast and sensitive graphene oxide-DNAzyme-based biosensor for Vibrio alginolyticus detection. Journal of Fish Diseases, 45(5):687–697, May 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tamer:2022:DVS

[TIK^{+}22] Cuneyt Tamer, Hakan Isıdan, Gulnur Kalaycı, Emre Ozan, Buket Ozkan, and Harun Albayrak. Determination of VP2 sequence-based virulence motifs and phylogenetic analysis of domestic Turkish IPNV isolates. Journal of Fish Diseases, 45(2):327–334, February 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Trosse:2021:MCP

[TKB^{+}21] Christiane Tröße, Mats Kindt, Steffen Blindheim, Linda Andersen, and Are Nylund. Method for cryopreservation of Paramoeba perurans. Journal of Fish Diseases, 44(6):739– 745, June 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Takano:2024:CSH

[TMM^{+}24] Tomokazu Takano, Satoshi Miwa, Tomomasa Matsuyama, Ikunari Kiryu, Miho Honjo, Takamitsu Sakai, Yuta Matsuura, Masatoshi Yamasaki, Akira Kumagai, and Chihaya Nakayasu. Clinical symptoms and histopathological changes in coho salmon affected by the erythrocytic inclusion body syndrome (EIBS) are caused by the infection of piscine orthoreovirus 2 (PRV-2). Journal of Fish Diseases, 47(6): e13939:1–e13939:??, June 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tran:2022:TLV

[TNB^{+}22] Triet Hanh Tran, Vy Thuy Hoang Nguyen, Hieu Chi Nguyen Bui, Yen Binh Thi Tran, Huong Thanh Thi Tran, Thao Thu Thi Le, Huong Thanh Thi Vu, and Thao Phuong Huynh Ngo. Tilapia Lake Virus (TiLV) from Vietnam is genetically distantly related to TiLV strains from other countries. Journal of Fish Diseases, 45(9):1389–1401, September 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Teitge:2020:EDP

[TPSJS20a] Felix Teitge, Christina Peppler, Dieter Steinhagen, and Verena Jung-Schroers. Effect of disinfection with peracetic acid on the microbial community of a seawater aquaculture recirculation system for Pacific white shrimp (Litopenaeus vannamei). Journal of Fish Diseases, 43(9):991–1017, September 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Teitge:2020:WDO

[TPSJS20b] Felix Teitge, Christina Peppler, Dieter Steinhagen, and Verena Jung-Schroers. Water disinfection by ozonation has advantages over UV irradiation in a brackish water recirculation aquaculture system for Pacific white shrimp (Litopenaeus vannamei). Journal of Fish Diseases, 43(10):1259– 1285, October 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Troszok:2023:TEO

[TR23] Agnieszka Troszok and Marek Roszko. Thyme essential oil inhibits intracellular replication of CyHV-3 and inactivates extracellular virus. An in vitro study. Journal of Fish Diseases, 46(6):663–677, June 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tsai:2023:GPA

[TSC^{+}23] Ming-An Tsai, Ming She See, Cheng-Hsun Chiu, Pei-Chi Wang, and Shih-Chu Chen. Genotypic and phenotypic analysis of Elizabethkingia meningoseptica in bullfrog Rana catesbeiana isolated in Taiwan. Journal of Fish Diseases, 46(11):1239–1248, November 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Turner:2023:TLV

[TSS^{+}23] Jessica Kaye Turner, Sasakorn Sakulpolwat, Suchanuch Sukdanon, Tuchakorn Lertwanakarn, Pitchaporn Waiyamitra, Chutchai Piewbang, Felipe Pierezan, Somporn Techangamsuwan, Esteban Soto, and Win Surachetpong. Tilapia lake virus (TiLV) causes severe anaemia and systemic disease in tilapia. Journal of Fish Diseases, 46(6):643–651, June 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Thirumalaikumar:2022:ERS

[TSV^{+}22] Eswaramoorthy Thirumalaikumar, Ramamoorthy Sathishkumar, Sugumar Vimal, Mariavincent Michael Babu, Ganapathi Uma, Angela Mariana Lusiastuti, and Thavasimuthu Citarasu. Efficacy of recombinant subunit OMP and hly vaccines against Aeromonas hydrophila in Rohu (Labeo rohita). Journal of Fish Diseases, 45(10):1581–1592, October 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tang:2022:QDP

[TZL^{+}22] Jia-Jia Tang, Zhi-Hong Zhong, Zhi-Cheng Li, Qing-Kai Guo, Shi-Yu Li, Yi-Xuan Guo, Biao Jiang, and An-Xing Li. Quantitative detection of parasitic ciliate Cryptocaryon irritans in seawater with an optimized sample processing method. Journal of Fish Diseases, 45(5):623–630, May 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ullrich:2021:SVH

[UCB^{+}21] Johanna Ullrich, Jürgen Christian, Sven Michael Bergmann, Martin Oberle, and Anna Maria Becker. Stability of viral haemorrhagic septicaemia virus, infectious hematopoietic necrosis virus and cyprinid herpesvirus 3 in various water samples. Journal of Fish Diseases, 44(4):379–390, April 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ugelvik:2022:EDI

[UD22] Mathias Stølen Ugelvik and Sussie Dalvin. The effect of different intensities of the ectoparasitic salmon lice (Lepeophtheirus salmonis) on Atlantic salmon (Salmo salar). Journal of Fish Diseases, 45(8):1133–1147, August 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ugelvik:2022:TAS

[UMD22] Mathias Stølen Ugelvik, Stig Mæhle, and Sussie Dalvin. Temperature affects settlement success of ectoparasitic salmon lice (Lepeophtheirus salmonis) and impacts the immune and stress response of Atlantic salmon (Salmo salar). Journal of Fish Diseases, 45(7):975–990, July 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ulehlova:2023:ACK

[ÚPZ^{+}23] Zuzana Úlehlová, Łubomír Pojezdal, Gabriela Zelenková, Stanislava Reschová, and Martin Faldyna. The analysis of a cluster of koi herpesvirus disease outbreaks in intensive carp aquaculture using molecular and conventional epidemiology. Journal of Fish Diseases, 46(6):709–713, June 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Valenzuela:2022:VAI

[VBG^{+}22] Beatriz Valenzuela, Almendra Benavides, Daniela Gutierrez, Lotsé Blamey, María Teresa Monsalves, Brenda Modak, and Jenny M. Blamey. Violacein from an Antarctic Iodobacter sp. 7MAnt and its function as immunomodulator of the defence mechanism of innate immunity in fish cells. Journal of Fish Diseases, 45(3):485–489, March 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Valenzuela:2023:EBA

[VBL^{+}23] Beatriz Valenzuela, Almendra Benavides, Francisco Leyton, Franco Moreno, Marcos Cortés, and Juana A. Ibacache. Evaluation of bactericidal activity of 7-arylaminoisoquinolin-5,8-quinones against Piscirickettsia salmonis. Journal of Fish Diseases, 46(1):85–89, January 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Vercauteren:2020:FRL

[VDC20] Maaike Vercauteren, Annemie Decostere, and Koen Chiers. First report of lesions resembling red mark syndrome observed in wild-caught common dab (Limanda limanda). Journal of Fish Diseases, 43(1):147–151, January 2020. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Vercauteren:2020:PRA

[VDD^{+}20] Maaike Vercauteren, Evelien De Swaef, Annelies Maria Declercq, Johan Aerts, Bart Ampe, Snorre Gulla, Freddy Haesebrouck, Lisa Devriese, Annemie Decostere, and Koen Chiers. Pinpointing the role of Aeromonas salmonicida in the development of skin ulcerations in common dab (Limanda limanda). Journal of Fish Diseases, 43(3):347–357, March 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Vidal:2023:GFP

[VFR^{+}23] José M. Vidal, Alexis Fonseca, Pamela Ruiz, Daniela Sepúlveda, Carlos Carrasco, Stefano Scilipoti, Javier Barros, Ariel Valenzuela, Ricardo Saavedra, Nathaly Ruiz-Tagle, and Homero Urrutia. Genomic features of Pseudomonas sp. RGM2144 correlates with increased survival of rainbow trout infected by Flavobacterium psychrophilum. Journal of Fish Diseases, 46(1):1–15, January 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Venugopalan:2021:VIB

[VGW^{+}21] Arun Venugopalan, Matt J. Griffin, David J. Wise, Danielle’ White, Lorelei Ford, Adrián López-Porras, Alvin C. Camus, and Larry A. Hanson. Virulence and immunogenicity of blue catfish alloherpesvirus in channel, blue and blue × channel hybrid catfish. Journal of Fish Diseases, 44(9):1399–1409, September 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Valdes:2021:FRC

[VIB^{+}21] Sara Valdes, Rute Irgang, María C. Barros, Pedro Ilardi, Mónica Saldarriaga-Córdoba, Javier Rivera-Bohle, Enrique Madrid, Johana Gajardo-Córdova, and Ruben Avendaño-Herrera. First report and characterization of Tenacibaculum maritimum isolates recovered from rainbow trout (Oncorhynchus mykiss) farmed in Chile. Journal of Fish Diseases, 44(10):1481–1490, October 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Verma:2022:CNP

[VKK^{+}22] Chandani R. Verma, Pradeep Kumkar, Tushar Khare, Manoj Pise, Lukáš Kalous, and Neelesh Dahanukar. Contracaecum nematode parasites in hillstream loaches of the Western Ghats, India. Journal of Fish Diseases, 45(12): 1873–1882, December 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Valle:2020:ASP

[VLSM^{+}20] Fernando Mesías Valle, Luis Llanco, Enrique Serrano-Martínez, Carmen Hurtado, Tania Rodríguez, and Peter Smith. Antimicrobial sensitivity of pathogenic Yersinia ruckeri from central and southern highlands of Peru. Journal of Fish Diseases, 43(5):631–633, May 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

VanIderstine:2021:HPT

[VPW^{+}21] Carter R. VanIderstine, Jordan D. Poley, Shona K. Whyte, Sara L. Purcell, and Mark D. Fast. Hydrogen peroxide treatment and its impacts on Lepeophtheirus salmonis originating from the Bay of Fundy, Canada. Journal of Fish Diseases, 44 (6):757–769, June 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Vidal:2022:PSF

[VRC^{+}22] José Miguel Vidal, Pamela Ruiz, Carlos Carrasco, Javier Barros, Daniela Sepúlveda, Nathaly Ruiz-Tagle, Alex Romero, Homero Urrutia, and Cristian Oliver. Piscirickettsia salmonis forms a biofilm on nylon surface using a CDC Biofilm Reactor. Journal of Fish Diseases, 45(8):1099–1107, August 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Vijayaram:2024:SIN

[VRS^{+}24] Srirengaraj Vijayaram, Hary Razafindralambo, Yun Zhang Sun, Giuseppe Piccione, Cristiana Roberta Multisanti, and Caterina Faggio. Synergistic interaction of nanoparticles and probiotic delivery: a review. Journal of Fish Diseases, 47 (5):e13916:1–e13916:??, May 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Vazquez-Salgado:2022:NNV

[VSOB22] Lucia Vázquez-Salgado, Jose G. Olveira, and Isabel Bandín. Nervous necrosis virus viability modulation by water salinity and temperature. Journal of Fish Diseases, 45(4):561–568, April 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Vazquez-Salgado:2021:ERD

[VSODB21] Lucía Vázquez-Salgado, Jose G. Olveira, Carlos P. Dopazo, and Isabel Bandín. Effect of rearing density on nervous necrosis virus infection in Senegalese sole (Solea senegalensis). Journal of Fish Diseases, 44(12):2003–2012, December 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Vannetti:2023:ASC

[VWE^{+}23] Stefania M. Vannetti, James W. Wynne, Chloe English, Christine Huynh, Ralph Knüsel, Carolina de Sales-Ribeiro, Maro Widmer, Gary Delalay, and Heike Schmidt-Posthaus. Amoeba species colonizing the gills of rainbow trout (Oncorhynchus mykiss) in Swiss aquaculture. Journal of Fish Diseases, 46(9):987–999, September 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Weichert:2021:MBS

[WAF^{+}21] Fabian G. Weichert, Charlotte Axén, Lars Förlin, Pedro A. Inostroza, Ulrike Kammann, Annikki Welling, Joachim Sturve, and Noomi Asker. A multi-biomarker study on Atlantic salmon (Salmo salar L.) affected by the emerging Red Skin Disease in the Baltic Sea. Journal of Fish Diseases, 44 (4):429–440, April 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Waldner:2021:EWT

[WBB^{+}21] Karoline Waldner, Martin Borkovec, Florian Borgwardt, Günther Unfer, and Mansour El-Matbouli. Effect of water temperature on the morbidity of Tetracapsuloides bryosalmonae (Myxozoa) to brown trout (Salmo trutta) under laboratory conditions. Journal of Fish Diseases, 44(7): 1005–1013, July 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Walsh:2021:IAI

[WBM21] Heather L. Walsh, Vicki S. Blazer, and Patricia M. Mazik. Identification of Aphanomyces invadans, the cause of epizootic ulcerative syndrome, in smallmouth bass (Micropterus dolomieu) from the Cheat River, West Virginia, USA. Journal of Fish Diseases, 44(10):1639–1641, October 2021. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Warshafsky:2020:PIP

[WBV^{+}20] Zoemma T. Warshafsky, Stephen A. Bullard, Wolfgang K. Vogelbein, Emily A. Chandler, and Molly E. Lutcavage. Prevalence, intensity and pathology of the nasal parasite Nasicola hogansi in Atlantic bluefin tuna (Thunnus thynnus). Journal of Fish Diseases, 43(3):327–335, March 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wangman:2020:DRI

[WCT^{+}20] Pradit Wangman, Parin Chaivisuthangkura, Suparat Taengchaiyaphum, Chalinan Pengsuk, Paisarn Sithigorngul, and Siwaporn Longyant. Development of a rapid immunochromatographic strip test for the detection of Vibrio parahaemolyticus toxin B that cause acute hepatopancreatic necrosis disease. Journal of Fish Diseases, 43(2):207–214, February 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:2023:ETS

[WCT^{+}23a] Chong Wang, Qingshi Chen, Manfei Tang, Tianli Wei, and Jixing Zou. Effects of TLR2/4 signalling pathway in western mosquitofish (Gambusia affinis) after Edwardsiella tarda infection. Journal of Fish Diseases, 46(4):299–307, April 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:2023:RPA

[WCT^{+}23b] Yu Wang, Yukang Chen, Yixin Tang, Yue Wang, Song Gao, Lihong Yang, and Pei Wang. A recombinase polymerase amplification and Pyrococcus furiosus Argonaute combined method for ultra-sensitive detection of white spot syndrome virus in shrimp. Journal of Fish Diseases, 46(12):1357–1365, December 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Welch:2020:CNY

[Wel20] Timothy J. Welch. Characterization of a novel Yersinia ruckeri serotype O1-specific bacteriophage with virulenceneutralizing activity. Journal of Fish Diseases, 43(2):285– 293, February 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Widanarni:2024:UPI

[WGE^{+}24] Widanarni Widanarni, Muhamad Gustilatov, Julie Ekasari, Pande Gde Sasmita Julyantoro, Diana Elizabeth Waturangi, and Sukenda Sukenda. Unveiling the positive impact of biofloc culture on Vibrio parahaemolyticus infection of Pacific white shrimp by reducing quorum sensing and virulence gene expression and enhancing immunity. Journal of Fish Diseases, 47(6):e13932:1–e13932:??, June 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:2021:CFS

[WHC^{+}21] Wenji Wang, Suying Hou, Guoquan Chen, Liqun Xia, Jianlin Chen, Zhiwen Wang, and Yishan Lu. Characterization and function study of a glutamyl endopeptidase homolog from Nocardia seriolae. Journal of Fish Diseases, 44(6):813–821, June 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Watanabe:2023:EBN

[WHK^{+}23] Yuho Watanabe, Jørgen Hansen, Maho Kotake, Ryotaro Fujii, Hiromi Matsuoka, and Tomoyoshi Yoshinaga. Effect of Biokos, a natural lipopeptide surfactant extracted from a bacterium of the Pseudomonas genus, on infection of Cryptocaryon irritans. Journal of Fish Diseases, 46(11):1311– 1319, November 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wessel:2020:IPO

[WHL^{+}20] Øystein Wessel, Elisabeth F. Hansen, Marie Løvoll, Makoto Inami, Asbjørn Husby, Gesa Kruse, Maria K. Dahle, and Espen Rimstad. Inactivation of Piscine orthoreovirus. Journal of Fish Diseases, 43(9):1039–1048, September 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Walde:2021:ECL

[WJPS21] Cecilie Sviland Walde, Britt Bang Jensen, Jostein Mulder Pettersen, and Marit Stormoen. Estimating cage-level mortality distributions following different delousing treatments of Atlantic salmon (Salmo salar) in Norway. Journal of Fish Diseases, 44(7):899–912, July 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wen:2022:ISH

[WKL^{+}22] Yiming Wen, Xu Kang, Zhiyuan Li, Liqun Xia, and Yishan Lu. Identification of a secretory heme-binding protein from Nocardia seriolae involved in cell apoptosis. Journal of Fish Diseases, 45(8):1189–1199, August 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:2020:PCA

[WLH^{+}20] Ping Wang, Jie Li, Tian Tian He, Nan Li, Zhao Lan Mo, Pin Nie, and Hai Xia Xie. Pathogenic characterization of Aeromonas salmonicida subsp. masoucida turbot isolate from China. Journal of Fish Diseases, 43(10):1145–1154, October 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wei:2022:SAB

[WLK^{+}22] Hongling Wei, Mingzhu Liu, Ke Ke, Shuangyan Xiao, Lin Huang, Qiongyu He, Changping Mo, Hai Pang, Guozhu Xiao, Pengfei Li, and Qing Yu. Study on aptamer based high throughput approach identifies natural ingredients against RGNNV. Journal of Fish Diseases, 45(11):1711–1719, November 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:2021:NIC

[WLTZ21] Chong Wang, Shuang Liu, Kathy F. J. Tang, and Qingli Zhang. Natural infection of covert mortality nodavirus affects Zebrafish (Danio rerio). Journal of Fish Diseases, 44 (9):1315–1324, September 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:2021:SVD

[WML^{+}21] Pei Wang, Chao Ma, Lei Liao, Junwei Yu, Longyu Yi, Yi Qiao, Xin Liu, Song Gao, Hui Shen, and Qunwei Lu. Simultaneous visual diagnosis of acute hepatopancreatic necrosis disease and Enterocytozoon hepatopenaei infection in shrimp with duplex recombinase polymerase amplification. Journal of Fish Diseases, 44(11):1753–1763, November 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:2021:GPS

[WMS21] Xin-An Wang, Ai-Jun Ma, and Zhi-Bin Sun. Genetic parameters of seven immune factors in turbot (Scophthalmus maximus) infected with Vibrio anguillarum. Journal of Fish Diseases, 44(3):263–271, March 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:2023:ASI

[WQL^{+}23] Fei Wang, Zi-Le Qin, Wei-Sheng Luo, Ning-Xia Xiong, Ming-Zhu Huang, Jie Ou, Sheng-Wei Luo, and Shao-Jun Liu. Alteration of synergistic immune response in gut–liver axis of white crucian carp (Carassius cuvieri) after gut infection with Aeromonas hydrophila. Journal of Fish Diseases, 46 (9):917–927, September 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wynne:2024:IED

[WRMT24] James W. Wynne, Megan L. Rigby, Ben T. Maynard, and Richard S. Taylor. Improved environmental detection of Neoparamoeba perurans using sensitive RNA-based qPCR. Journal of Fish Diseases, 47(2):e13879:1–e13879:??, February 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wynne:2020:IBT

[WSS^{+}20] James W. Wynne, Chris Stratford, Joel Slinger, Francisca Samsing, Megan Rigby, Russell McCulloch, Petra Quezada-Rodriguez, and Richard S. Taylor. The interaction between temperature and dose on the efficacy and biochemical response of Atlantic salmon to hydrogen peroxide treatment for amoebic gill disease. Journal of Fish Diseases, 43(1):39– 48, January 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:2020:NRI

[WSW^{+}20] Ziyan Wang, Chengyin Shi, Hailiang Wang, Xiaoyuan Wan, Qingli Zhang, Xiaoling Song, Ge Li, Miao Gong, Shigen Ye, Guosi Xie, and Jie Huang. A novel research on isolation and characterization of Photobacterium damselae subsp . damselae from Pacific white shrimp, Penaeus vannamei, displaying black gill disease cultured in China. Journal of Fish Diseases, 43(5):551–559, May 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wu:2020:CMA

[WWK^{+}20] Ching-Wei Wu, Ting-Yun Wu, Chih-Jung Kuo, Yi-Ping Lu, Maw-Sheng Chien, and Chienjin Huang. Characterization of the monoclonal antibody specific to the ORF72 protein of koi herpesvirus and cellular distribution analysis of the viral protein. Journal of Fish Diseases, 43(7):791–799, July 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:2021:CTC

[WWW^{+}21] Li Wang, Qiuxia Wang, Lei Wang, Shixiu Wu, Yan Yu, Yanhong Zhang, Pei Gao, Xianghui Kong, and Jinyou Ma. The N- and C-terminal carbohydrate recognition domains of galectin-9 from Carassius auratus contribute differently to its immunity functions to Aeromonas hydrophila and Staphylococcus aureus. Journal of Fish Diseases, 44(11):1865–1873, November 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:2022:MMA

[WXM^{+}22] Yifan Wang, Liwen Xu, Weixiang Ma, Heng Sun, Zengchao Huang, Shuanghu Cai, Jichang Jian, and Yucong Huang. Mass mortalities associated with viral nervous necrosis in Murray cod in China. Journal of Fish Diseases, 45(2):277– 287, February 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:2023:ILB

[WYR^{+}23] Mengmeng Wang, Bin Yang, Zongyi Ren, Jietao Liu, Chenwang Lu, Haifeng Jiang, Fei Ling, Gaoxue Wang, and Tianqiang Liu. Inhibition of the largemouth bass virus replication by piperine demonstrates potential application in aquaculture. Journal of Fish Diseases, 46(3):261–271, March 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wan:2024:CPT

[WZC^{+}24] Qijuan Wan, Shaowei Zhai, Minxia Chen, Ming Xu, and Songlin Guo. Comparative phenotype and transcriptome analysis revealed the role of ferric uptake regulator (Fur) in the virulence of Vibrio harveyi isolated from diseased American eel (Anguilla rostrata). Journal of Fish Diseases, 47 (6):e13931:1–e13931:??, June 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:2020:MLM

[WZG^{+}20] Sijia Wang, Bo Zhang, Qingxiang Guo, Yanhua Zhai, and Zemao Gu. Molecular and light microscopy evidence for the transfer of Myxobolus honghuensis from Carassius auratus gibelio broodfish to progeny. Journal of Fish Diseases, 43 (10):1177–1184, October 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cheng:2021:TAT

[xCqXfL^{+}21] Jian xin Cheng, Yu qing Xia, Ya fang Liu, Peng fei Liu, and Ying Liu. Transcriptome analysis in Takifugu rubripes and Dicentrarchus labrax gills during Cryptocaryon irritans infection. Journal of Fish Diseases, 44(3):249–262, March 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Xia:2021:CIA

[XFL^{+}21] Hongli Xia, Huimin Fan, Meng Long, Jun Cheng, Wenjie Chen, Dapeng Yu, Liqun Xia, and Yishan Lu. CD40 induces an antimicrobial response against the intracellular pathogen Streptococcus agalactiae in Nile tilapia, Oreochromis niloticus. Journal of Fish Diseases, 44(1):45–52, January 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Xie:2021:GTR

[XJM^{+}21] Xiao Xie, Yunyan Jiang, Rujiang Miao, Jiashuang Huang, Liyao Zhou, Jindong Kong, and Fei Yin. The gill transcriptome reveals unique antimicrobial features that protect Nibea albiflora from Cryptocaryon irritans infection. Journal of Fish Diseases, 44(8):1215–1227, August 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Xu:2022:RSV

[XJZ^{+}22] Fei-Fan Xu, Fu-Yi Jiang, Guo-Qing Zhou, Jun-Yao Xia, Fei Yang, and Bin Zhu. The recombinant subunit vaccine encapsulated by alginate-chitosan microsphere enhances the immune effect against Micropterus salmoides rhabdovirus. Journal of Fish Diseases, 45(11):1757–1765, November 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Xiong:2022:TAC

[XKF^{+}22] Ning-Xia Xiong, Xu-Ying Kuang, Zi-Xuan Fang, Jie Ou, Shi-Yun Li, Jia-Hui Zhao, Jin-Fang Huang, Ke-Xin Li, Rou Wang, Lan-Fen Fan, Sheng-Wei Luo, and Shao-Jun Liu. Transcriptome analysis and co-expression network reveal the mechanism linking mitochondrial function to immune regulation in red crucian carp (Carassius auratus red var) after Aeromonas hydrophila challenge. Journal of Fish Diseases, 45(10):1491–1509, October 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Xiao:2022:SPA

[XQZ^{+}22] Wanghong Xiao, Geng Qin, Yanhong Zhang, Bo Zhang, Tao Li, Zelin Chen, and Qiang Lin. Secretory phospholipase A2 group XIIBs play potential roles in intestine antibacterial responses in seahorse. Journal of Fish Diseases, 45(7):1059– 1063, July 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Xing:2024:IIC

[XXX^{+}24] Yangyang Xing, Mingyang Xue, Zidong Xiao, Xiaowei Hu, Jiale Zhai, Chunjie Zhang, Nan Jiang, Yuding Fan, Yan Meng, and Yong Zhou. Isolation and identification of Chryseobacterium indologenes and its pathological changes in Pelodiscus sinensis. Journal of Fish Diseases, 47(1): e13864:1–e13864:??, January 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Xintong:2023:FRA

[XZS^{+}23] Zhang Xintong, Xin Zaozhe, Xue Sijia, Tang Shaolin, and Jinyong Zhang. First report of Ameson portunus (microsporidia) infection in earthen pond-cultured mud crab, Scylla paramamosain (Decapoda: Portunidae) in China, causing mass morbidity. Journal of Fish Diseases, 46(6):715– 721, June 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yu:2021:ERC

[YCC^{+}21] Lu Yu, Qikang Chen, Xin Chu, Yang Luo, Zizhao Feng, Liqun Lu, Ye Zhang, and Dan Xu. Expression and regulation of ccBAX by miR-124 in the caudal fin cell of C. auratus gibelio upon cyprinid herpesvirus 2 infection. Journal of Fish Diseases, 44(6):837–845, June 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yang:2021:INA

[YCL^{+}21] Jin-Xian Yang, Xi Chen, Ying-Ying Li, Tie-Ying Song, and Jun-Qing Ge. Isolation of a novel adomavirus from cultured American eels, Anguilla rostrata, with haemorrhagic gill necrosis disease. Journal of Fish Diseases, 44(11):1811– 1818, November 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yang:2021:CGA

[YCS^{+}21] Bintong Yang, Chong Chen, Yufeng Sun, Linan Cao, Dongxing Zhang, Wuwen Sun, Lei Zhang, Guiqin Wang, Xiaofeng Shan, Yuanhuan Kang, and Aidong Qian. Comparative genomic analysis of different virulence strains reveals reasons for the increased virulence of Aeromonas veronii. Journal of Fish Diseases, 44(1):11–24, January 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yang:2021:EVI

[YDM^{+}21] Xiaohan Yang, Yu Dong, Chao Ma, Yi Qiao, Ge Jiang, Shiqi Chen, Jingquan Dong, Hui Shen, and Song Gao. Establishment of a visualized isothermal nucleic acid amplification method for on-site diagnosis of acute hepatopancreatic necrosis disease in shrimp farm. Journal of Fish Diseases, 44(9):1293–1303, September 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yang:2020:CIR

[YGC^{+}20] Jie Yang, Yafan Gong, Jingzeng Cai, Yingying Zheng, Honggui Liu, and Ziwei Zhang. Chlorpyrifos induces redox imbalance-dependent inflammation in common carp lymphocyte through dysfunction of T-cell receptor γ. Journal of Fish Diseases, 43(4):423–430, April 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yamamoto:2024:ISD

[YGR^{+}24] Fernando Y. Yamamoto, Matt J. Griffin, Bradley M. Richardson, Justin M. Stilwell, Nicholas Romano, Penelope M. Goodman, J. Grant Reifers, and David J. Wise. Iron supplementation in the diets of hybrid catfish (Ictalurus punctatus × I. furcatus) juveniles affected haematocrit levels and potentially decreased disease resistance to Edwardsiella ictaluri. Journal of Fish Diseases, 47(3):e13902:1–e13902:??, March 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yang:2022:PID

[YGZ^{+}22] Bin Yang, Zi-Rao Guo, Zhao Zhao, Tao Wang, Fei Yang, Fei Ling, Bin Zhu, and Gao-Xue Wang. Protective immunity by DNA vaccine against Micropterus salmoides rhabdovirus. Journal of Fish Diseases, 45(10):1429–1437, October 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yang:2023:SNP

[YJL^{+}23a] Shijia Yang, Duo Jin, Hui Li, Lingzhi Jiang, Jun Cui, Weitao Huang, Jie Rang, YunLong Li, and Liqiu Xia. Screening of new Paenibacillus polymyxa S3 and its disease resistance of grass carp (Ctenopharyngodon idellus). Journal of Fish Diseases, 46(1):17–29, January 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yang:2023:NIS

[YJL^{+}23b] Yahui Yang, Duo Jin, Wensu Long, Ximiao Lai, Yunjun Sun, Feng Zhai, Pan Wang, Xixun Zhou, Yibo Hu, Liqiu Xia, and Ganfeng Yi. A new isolate of Streptomyces lateritius (Z1-26) with antibacterial activity against fish pathogens and immune enhancement effects on crucian carp (Carassius auratus). Journal of Fish Diseases, 46(2):99–112, February 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yildiz:2021:PCL

[YK21] Hijran Yavuzcan Yıldız and Ahmet Şeref Korkmaz. Parasitic copepod (Lernanthropus kroyeri) on caged sea bass (Dicentrarchus labrax): an estimation of abundance and internal infestation pressure. Journal of Fish Diseases, 44(12):1901– 1909, December 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yu:2021:MRS

[YLC^{+}21] Fei Yu, Shuxin Li, Hongxun Chen, Kai Hao, Lihui Meng, Jiayue Yang, and Zhe Zhao. Multiple AT-rich sequences function as a cis-element in the ORF3 promoter in channel catfish virus (Ictalurid herpesvirus 1 ). Journal of Fish Diseases, 44(10):1609–1617, October 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yu:2021:GCA

[YLW^{+}21] Qing Yu, Mingzhu Liu, Siting Wu, Hehe Xiao, Xinling Qin, and Pengfei Li. Generation and characterization of aptamers against grass carp reovirus infection for the development of rapid detection assay. Journal of Fish Diseases, 44(1):33–44, January 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ye:2021:RUR

[YMC^{+}21] Tingqi Ye, Cuimin Mu, Jiakang Chen, Guangchen Pan, and Xuepeng Wang. The role of UhpA in regulating the virulence gene expression in Edwardsiella piscicida. Journal of Fish Diseases, 44(5):585–590, May 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yamkasem:2022:PTT

[YPT^{+}22] Jidapa Yamkasem, Tharinthon Prasartset, Puntanat Tattiyapong, Kwanrawee Sirikanchana, Skorn Mongkolsuk, Esteban Soto, and Win Surachetpong. Persistence of Tilapia tilapinevirus in fish rearing and environmental water and its ability to infect cell line. Journal of Fish Diseases, 45 (5):679–685, May 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yi:2022:MCE

[YSW^{+}22] Liyuan Yi, Caiyun Sun, Jinying Wu, Zengxin Hou, and Wensheng Li. Molecular cloning, expressional analyses and functional identification of TRAIL in tilapia, Oreochromis niloticus. Journal of Fish Diseases, 45(6):833–846, June 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yan:2024:NOC

[YTH^{+}24] Wei-Xiao Yan, Yu-Han Tseng, Wen-Rou Huang, Li-Wu Cheng, Pei-Chi Wang, and Shih-Chu Chen. Nocardiosis and other common diseases of cultured golden pompano (Trachinotus blochii) in Taiwan. Journal of Fish Diseases, 47(3): e13894:1–e13894:??, March 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yamkasem:2022:DAT

[YTS22] Jidapa Yamkasem, Puntanat Tattiyapong, and Win Surachetpong. Development and application of TaqMan probebased quantitative PCR assays for the detection of tilapia parvovirus. Journal of Fish Diseases, 45(3):379–386, March 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yu:2021:FOC

[YWZ^{+}21] Dao De Yu, Hai Yi Wu, Zhi Peng Zhang, Wen Dan Chi, Kai Kai Liu, Ming Liu, Ming Qi Wang, and Jing Jing Song. First observation of conjoined twins in Oryzias melastigma. Journal of Fish Diseases, 44(6):853–856, June 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yildiz:2024:DCA

[YY24] Hijran Yavuzcan Yildiz and Bilgenur Harmansa Yilmaz. Dietary citric acid decreased the theront number of Cryptocaryon irritans (Ciliata) in seawater-adapted tilapia (Oreochromis niloticus). Journal of Fish Diseases, 47(2): e13881:1–e13881:??, February 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yang:2022:AEP

[YYS^{+}22] Ying-Fei Yang, Jidapa Yamkasem, Win Surachetpong, Yi-Jun Lin, Shu-Han You, Tien-Hsuan Lu, Chi-Yun Chen, Wei-Min Wang, and Chung-Min Liao. Assessing the effect of probiotics on tilapia lake virus-infected tilapia: Transmission and immune response. Journal of Fish Diseases, 45(8): 1117–1132, August 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yang:2024:NPH

[YYS^{+}24] Fei Yang, Bin Yang, Kaige Song, Yingjie Jin, Gaoxue Wang, Pengfei Li, Qing Yu, and Fei Ling. Natural product honokiol exhibits antiviral effects against Micropterus salmoides rhabdovirus (MSRV) both in vitro and in vivo. Journal of Fish Diseases, 47(4):e13915:1–e13915:??, April 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yang:2023:ITG

[YZL^{+}23] Dou Yang, Lingmin Zhao, Qi Li, Lixing Huang, Yingxue Qin, Pan Wang, Chuanzhong Zhu, and Qingpi Yan. The involvement of the T6SS vgrG gene in the pathogenicity of Pseudomonas plecoglossicida. Journal of Fish Diseases, 46 (10):1097–1108, October 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yang:2021:IVA

[YZM^{+}21] He Yang, Ding Zhujin, Moonika H. Marana, Inger Dalsgaard, Jaafar Rzgar, Mathiessen Heidi, Karami M. Asma, Kania W. Per, and Buchmann Kurt. Immersion vaccines against Yersinia ruckeri infection in rainbow trout: Comparative effects of strain differences. Journal of Fish Diseases, 44(12):1937–1950, December 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zuykov:2020:FRS

[ZAG^{+}20] Michael Zuykov, Bassem Allam, Michel Gosselin, Philippe Archambault, Graeme Spiers, and Michael Schindler. First report of signs of infection by Coccomyxa-like algae in wild blue mussels, Mytilus spp., in the Gulf of Maine (USA, Maine). Journal of Fish Diseases, 43(7):775–778, July 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhou:2023:NSM

[ZBL^{+}23] Zheng-Yang Zhou, Shang-Jie Bai, Chen-Wang Lu, Bin Yang, Jun Wang, Sheng Lian, Zheng-Ran Jian, Er-Long Wang, Gao-Xue Wang, and Tao Liu. Nocardia seriolae mediates liver granulomatous chronic inflammation in Micropterus salmoides through pyroptosis. Journal of Fish Diseases, 46 (4):321–332, April 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhang:2023:RMG

[ZCY^{+}23] Xiaoxu Zhang, Liwei Chen, Qingpi Yan, Lingmin Zhao, Lixing Huang, Dongling Zhang, and Yingxue Qin. The role and mechanisms of rstA in the intracellular survival of fish pathogenic Aeromonas hydrophila. Journal of Fish Diseases, 46(8):813–827, August 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhang:2023:PCM

[ZDF^{+}23] Wenfeng Zhang, Hengwei Deng, Yuting Fu, Weixuan Fu, Shaoping Weng, Jianguo He, and Chuanfu Dong. Production and characterization of monoclonal antibodies against mandarinfish ranavirus and first identification of pyloric caecum as the major target tissue. Journal of Fish Diseases, 46 (3):189–199, March 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhao:2022:MPP

[ZDZZ22] Na Zhao, Qiuxia Deng, Chunhua Zhu, and Bo Zhang. Mucus piRNAs profiles of Vibrio harveyi-infected Cynoglossus semilaevis: a hint for fish disease monitoring. Journal of Fish Diseases, 45(1):165–175, January 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhan:2023:DCS

[ZFC^{+}23] Zhipeng Zhan, Jiajie Fu, Huiquan Chen, Hongbo Pan, Shaoping Weng, Jianguo He, and Changjun Guo. Development and characterization of a spleen cell line from yellowfin seabream Acanthopagrus latus and its susceptibility to Mandarinfish ranavirus. Journal of Fish Diseases, 46(11):1173– 1181, November 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhang:2022:DDD

[ZGC^{+}22] Huang Zhang, Han-Yue Gong, Wei-Wei Cao, Mu-Yi Que, Lei Ye, and Lei Shi. Duplex droplet digital PCR method for the detection of Enterocytozoon hepatopenaei and Vibrio parahaemolyticus acute hepatopancreatic necrosis disease. Journal of Fish Diseases, 45(6):761–769, June 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhang:2020:PCS

[ZGL^{+}20] Ying Zhang, Qing-Long Gong, Qing-Bo Lv, Yang-Yuan Qiu, Yan-Chun Wang, Hong-Yu Qiu, Xin-Rui Guo, Jun-Feng Gao, Qiao-Cheng Chang, and Chun-Ren Wang. Prevalence of Clonorchis sinensis infection in fish in South-East Asia: a systematic review and meta-analysis. Journal of Fish Diseases, 43(11):1409–1418, November 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhao:2022:DCM

[ZGz^{+}22] Lupin Zhao, Wa Gao, Yihua zheng, Liqun Lu, Qiang Li, and Yousheng Jiang. Development and characterization of monoclonal antibodies specific for cyprinid herpesvirus 2. Journal of Fish Diseases, 45(11):1673–1681, November 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhang:2021:IRS

[ZHL^{+}21] Youyi Zhang, Jiabao Hu, Yaya Li, Man Zhang, Kimran Jean Jacques, Weiwei Gu, Yibo Sun, Jiachu Sun, Yang Yang, Shanliang Xu, Yajun Wang, and Xiaojun Yan. Immune response of silver pomfret (Pampus argenteus) to Amyloodinium ocellatum infection. Journal of Fish Diseases, 44(12): 2111–2123, December 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhang:2022:CCV

[ZHL^{+}22] Xiaodong Zhang, Kai Hao, Shuxin Li, Lihui Meng, Hongxun Chen, Fucheng Wei, Fei Yu, Jing Xu, and Zhe Zhao. Channel catfish virus ORF25 and ORF63 genes are essential for viral replication in vitro. Journal of Fish Diseases, 45(5):655– 666, May 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhong:2023:TAF

[ZHL^{+}23] Yunqi Zhong, Zihao He, Xinxin Long, Danqing Hou, Xianye Hu, and Chengbo Sun. Transcriptome analysis of Fenneropenaeus merguiensis in response to Vibrio proteolyticus infection. Journal of Fish Diseases, 46(11):1207–1224, November 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhao:2020:IIC

[ZHpT^{+}20] Fei Zhao, Shan He, Ai ping Tan, Xue zhong Guo, Lan Jiang, Cui Liu-Fu, Yu ting Deng, and Rui quan Zhang. Isolation, identification and character analysis of Streptococcus dysgalactiae from Megalobrama terminalis. Journal of Fish Diseases, 43(2):239–252, February 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhou:2021:RVD

[ZHY^{+}21] Shun Zhou, Qing Huang, Mingming Yu, Ying Li, Xujia Zheng, and Yunji Xiu. Rapid visual detection of Aeromonas salmonicida by loop-mediated isothermal amplification with hydroxynaphthol blue dye. Journal of Fish Diseases, 44(12): 1993–2001, December 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhang:2022:AVV

[ZJF^{+}22] Lei Zhang, Shengnan Jin, Chao Feng, Haichao Song, Sayed Haidar Abbas Raza, Huabo Yu, Liang Zhang, Teng Chi, Yanling Qi, Dongxing Zhang, Aidong Qian, Ning Liu, and Xiaofeng Shan. Aeromonas veronii virulence and adhesion attenuation mediated by the gene aodp. Journal of Fish Diseases, 45(2):231–247, February 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhou:2021:SDM

[ZLS^{+}21] Qian-Jin Zhou, Jian-Fei Lu, Xiu-Rong Su, Jing-Lei Jin, Shang-Yang Li, Yan Zhou, Lei Wang, Xin-Bin Shao, Yao-Hua Wang, Mao-Cang Yan, Ming-Yun Li, and Jiong Chen. Simultaneous detection of multiple bacterial and viral aquatic pathogens using a fluorogenic loop-mediated isothermal amplification-based dual-sample microfluidic chip. Journal of Fish Diseases, 44(4):401–413, April 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhang:2023:OPE

[ZLWL23] Zhao Zhang, Jing Li, Gaoxue Wang, and Fei Ling. The oral protective efficacy of magnolol against Aeromonas hydrophila and A. veronii infection via enhancing anti-inflammatory ability in goldfish (Carassius auratus). Journal of Fish Diseases, 46(12):1413–1423, December 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhou:2022:IRV

[ZLY^{+}22] Shun Zhou, Ying Li, Jingyuan Yi, Xujia Zheng, Qing Huang, Lin Su, Baoshan Guo, Zongrui Yang, and Yunji Xiu. Immune responses to Vibrio vulnificus formalin-killed vaccine and ghost vaccine in Scophthalmus maximus. Journal of Fish Diseases, 45(10):1511–1527, October 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhou:2021:IBQ

[ZLZ21] Xiujuan Zhou, Yongyong Lai, and Fei Zhu. iTRAQ-based quantitative proteomic analysis of haemocyte proteins from crayfish (Procambarus clarkii) infected with white spot syndrome virus (WSSV). Journal of Fish Diseases, 44(10): 1661–1668, October 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhu:2022:APT

[ZMZ^{+}22] Song Zhu, Bo Miao, Yu-Zhou Zhang, De-Shou Wang, and Gao-Xue Wang. Amantadine, a promising therapeutic agent against viral encephalopathy and retinopathy. Journal of Fish Diseases, 45(3):451–459, March 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhan:2023:ECS

[ZMZ^{+}23] Quanjun Zhan, Liang Miao, Jing Zhao, Haojie Hu, Hao Cui, Shan Jin, Jiasong Xie, Dong Qian, and Rongrong Ma. Effect of copper sulphate on Cryptocaryon irritans based on metabolome analysis. Journal of Fish Diseases, 46(4):347– 356, April 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zrncic:2020:KHC

[ZOZ^{+}20] Snježana Zrnčić, Dražen Oraić, Ivana Giovanna Zupičić, Željko Pavlinec, Dragan Brnić, Žaklin Acinger Rogić, Ivica Sučec, Dieter Steinhagen, and Mikolaj Adamek. Koi herpesvirus and carp edema virus threaten common carp aquaculture in Croatia. Journal of Fish Diseases, 43(6):673–685, June 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhu:2023:RBV

[ZSG^{+}23] Min Zhu, Zeen Shen, Yuchao Gu, Xinyu Tong, Yaxin Zhang, Jun Pan, Yongjie Feng, Xiaolong Hu, Yujun Wang, Guangli Cao, Renyu Xue, and Chengliang Gong. A recombinant baculovirus vector vaccine (BacMCP) against the infectious spleen and kidney necrosis virus (ISKNV). Journal of Fish Diseases, 46(2):165–176, February 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zeng:2024:IPA

[ZSL^{+}24a] Qingsong Zeng, Yunxiao Sun, Peifang Lai, Mingqin Huang, Ximing Peng, Junjie Huang, Qintao Chen, Yanxu Chen, and Huaqian Wang. Identification of a potential antigen stimulating immune response against Vibrio parahaemolyticus infection in hybrid tilapia (Oreochromis aureus ˘2642 × Oreochromis niloticus ˘2640 ). Journal of Fish Diseases, 47(3): e13904:1–e13904:??, March 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhou:2024:DKH

[ZSL^{+}24b] Su-Ming Zhou, Feng-Ling Shu, Hao Liu, Dong Qian, Jia-Song Xie, Ya-Jun Wang, and Fei Yin. Disruption of a kasB homolog gene (kasB) causes attenuation of cell invasion and virulence of Nocardia seriolae. Journal of Fish Diseases, 47(3):e13896:1–e13896:??, March 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhang:2022:RDG

[ZTY^{+}22] Yilin Zhang, Huimin Tan, Shiping Yang, Yucong Huang, Shuanghu Cai, Jichang Jian, Jia Cai, and Qiwei Qin. The role of dctp gene in regulating colonization, adhesion and pathogenicity of Vibrio alginolyticus strain HY9901. Journal of Fish Diseases, 45(3):421–434, March 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zrncic:2021:FDD

[ZVB^{+}21] Snježana Zrnčić, Niccolò Vendramin, Torsten S. Boutrup, Mette Boye, Lone Madsen, Bettina Nonneman, Dragan Brnić, and Dražen Oraić. First description and diagnostics of disease caused by Piscirickettsia salmonis in farmed European sea bass (Dicentrarchus labrax Linnaeus) from Croatia. Journal of Fish Diseases, 44(7):1033–1042, July 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhang:2023:IAA

[ZVYD23] Shanshan Zhang, Jana Van Haesebroeck, Qian Yang, and Tom Defoirdt. Indole-3-acetic acid increases the survival of brine shrimp challenged with vibrios belonging to the Harveyi clade. Journal of Fish Diseases, 46(5):477–486, May 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhong:2023:TOS

[ZWB^{+}23] Ziqing Zhong, Xiping Wu, Mindong Bai, Xiaodian Huang, Qilin Zheng, and Chunxiang Ai. Treatments of orangespotted grouper (Epinephelus coioides) against Cryptocaryon irritans with ·OH, ClO_{2} or HCHO: Survival, physiological and histological response. Journal of Fish Diseases, 46(3):215–227, March 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhang:2022:EGI

[ZWX^{+}22] Yitong Zhang, Jinxuan Wen, Yao Xu, Hao Wang, Liqun Lu, Rui Song, and Jixing Zou. Epigallocatechin-3-gallate inhibits replication of white spot syndrome virus in the freshwater crayfish Procambarus clarkii. Journal of Fish Diseases, 45 (3):445–450, March 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhao:2020:CEI

[ZWZ^{+}20] Jinping Zhao, Liqun Wu, Shaowei Zhai, Peng Lin, and Songlin Guo. Construction expression and immunogenicity of a novel trivalent outer membrane protein (OmpU-A-II) from three bacterial pathogens in Japanese eels (Anguilla japonica). Journal of Fish Diseases, 43(5):519–529, May 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zahid:2023:AEC

[ZYB^{+}23] Aysha Zahid, Youbin Yu, Bushra, Suming Zhou, Xiao Xie, and Fei Yin. Antiparasitic effect of copper alloy mesh on tomont stage of Cryptocaryon irritans in aquaculture. Journal of Fish Diseases, 46(3):181–188, March 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhang:2023:VAA

[ZYH^{+}23] Fa-Li Zhang, Liu Yang, Wei-Hao He, Ling-Jun Xie, Fei Yang, Ying-Hui Wang, and Ai-Guo Huang. In vivo antibacterial activity of medicinal plant Sophora flavescens against Streptococcus agalactiae infection. Journal of Fish Diseases, 46 (9):977–986, September 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhao:2022:ERA

[ZYL^{+}22] Tong Zhao, Bintong Yang, Hongjin Li, Aidong Qian, Wei Cong, Wuwen Sun, and Yuanhuan Kang. Essential role of asco for virulence of Aeromonas veronii and inducing apoptosis. Journal of Fish Diseases, 45(10):1477–1489, October 2022. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhang:2020:CTL

[ZYZ^{+}20] Xiao-Wen Zhang, Cong-Hui Yang, Hong-Quan Zhang, Xin-Tong Pan, Ze-Yu Jin, Hong-Wei Zhang, and Xiao-Hua Xia. A C-type lectin with antibacterial activity in weather loach, Misgurnus anguillicaudatus. Journal of Fish Diseases, 43 (12):1531–1539, December 2020. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhao:2023:DAS

[ZZF^{+}23] Haikuo Zhao, Yong Zhou, Yuding Fan, Nan Jiang, Yan Meng, Yiqun Li, Mingyang Xue, Chen Xu, Weiliang Guo, and Wenzhi Liu. Development and application of a sensitive droplet digital PCR-based method to detect tilapia parvovirus. Journal of Fish Diseases, 46(3):239–245, March 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zheng:2021:IIC

[ZZL^{+}21] Yu-Ying Zheng, Chen Zhang, Yang Li, Peng-Qi Zhang, Guo Chen, Gao-Xue Wang, and Bin Zhu. Immersion immunization of common carp with bacterial ghost-based DNA vaccine inducing prophylactic protective immunity against spring viraemia of carp virus. Journal of Fish Diseases, 44 (12):2021–2029, December 2021. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhan:2024:IFP

[ZZS^{+}24] Fanbin Zhan, Shichun Zhou, Fei Shi, Qingqing Li, Li Lin, and Zhendong Qin. Identifying the function of the PI3K-AKT pathway during the pathogenic infection of Macrobrachium rosenbergii. Journal of Fish Diseases, 47(2): e13890:1–e13890:??, February 2024. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zheng:2023:DLF

[ZZZ^{+}23] Yihua Zheng, Yi Zhou, Lupin Zhao, Jiaxun Li, Liqun Lu, and Yousheng Jiang. Development of a lateral flow immunochromatic strip assay for the detection of cyprinid herpesvirus 3 (CyHV-3). Journal of Fish Diseases, 46(10): 1065–1071, October 2023. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).